### v1

# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

Abstract
Advancements in image segmentation play an integral role within the greater scope of Deep Learning-based computer vision. Furthermore, their widespread applicability in critical real-world tasks has given rise to challenges related to the reliability of such algorithms. Hence, uncertainty quantification has been extensively studied within this context, enabling expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision making. Due to the rapid adoption of Convolutional Neural Network (CNN)-based segmentation models in high-stake applications, a substantial body of research has been published on this very topic, causing its swift expansion into a distinct field. This work provides a comprehensive overview of probabilistic segmentation by discussing fundamental concepts in uncertainty that govern advancements in the field as well as the application to various tasks. We identify that quantifying aleatoric and epistemic uncertainty approximates Bayesian inference w.r.t. to either latent variables or model parameters, respectively. Moreover, literature on both uncertainties trace back to four key applications; (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) active learning. Then, a discussion follows that includes an overview of utilized datasets for each of the applications and comparison of the available methods. We also highlight challenges related to architectures, uncertainty-based active learning, standardization and benchmarking, and recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
Index Terms:
image segmentation, uncertainty quantification, probability theory
I Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest [1]. With the rapid development of Convolution Neural Networks (CNNs), Deep-Learning based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7]. Nevertheless, such methodologies rely on extensive assumptions and relaxations on the Bayesian learning paradigm, omitting crucial information on the uncertainty associated with the model predictions. This ignorance diminishes the reliability and interpretability of such models. For example, difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
Extensive efforts have been made to align modern neural network optimization with Bayesian Machine Learning [8, 9, 10, 11], such as learning parameter densities, rather than point estimates, to include a notion of epistemic uncertainty. Furthermore, explicitly modeling the output likelihood distribution enables expressing the aleatoric uncertainty. Notably, literature mention that determining the nature of uncertainty is not often straightforward. For example, Hüllermeier et al. [12] mentioned that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”. This sentiment is also shared by Kiureghian and Ditlevsen [13], noting that “In one model an addressed uncertainty may be aleatory, in another model it may be epistemic”. Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic.
The merits of uncertainty quantification have fortunately been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview. Nonetheless, most surveys take perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19]. There is a notable absence of a comprehensive overview in this field, relating theoretical foundation to the multitude of applications. Furthermore, the abundance of available works can often be overwhelming for both new-coming and seasoned researchers. This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are clarified through standardized notation. After reading this work, readers will be able to discern various forms of uncertainty with their pertinent applications to segmentation tasks. Additionally, they will have developed comprehensive understanding of challenges and unexplored avenues in the field.
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section II, where closely related surveys are also referenced. Then, the theoretical framework and notation that governs the remainder of the paper is introduced in Section III. The role of these concepts in image segmentation will be treated in Section IV and V, which includes all architectures and approaches with significant impact on the field. We then consider the applications that use these uncertainty estimates in Section VI. Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII and we conclude in Section VIII. For a brief overview of the sections, refer to Figure 1.
[…]
Figure 1: Overview of the sections.
II Background
As summarized by Minaee et al. [20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watersheds [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29]. After the application of CNNs [30], the domain of image segmentation underwent rapid developments. Notably, the Fully Convolutional Network (FCN) [3] adapted the AlexNet [31], VGG16 [32] and GoogLeNet [33] architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 [34], and the MobileNetv3 [35] have also been commonly used.
As the research progressed, increasing success has been observed with encoder-decoder models [36, 4, 37, 2]. Initially developed for the medical applications, Ronneberger et al. [2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures. In fact, reports in recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
Semantic Segmentation focuses solely on assigning a class label(s) to each pixel and is particularly suitable for amorphous or uncountable subjects of interest. In contrast, Instance Segmentation not only detects, but also delineates individual objects within the image. This form of segmentation is more applicable when identifying and outlining countable instances of objects. The third category, Panoptic Segmentation, combines both class and instance level classification [20].
III Probabilistic Image Segmentation
Let random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, where $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively.
III-A Bayesian inference
Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
$p(\bm{\theta}|\mathbf{y},\mathbf{x})=\frac{p(\mathbf{y}|\mathbf{x},\bm{\theta})p(\bm{\theta})}{p(\mathbf{y|x})},$ (1)
where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence). After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
[…]
As evident in Equation (2), both the variability in the empirical data and the inferred parameters of the model influence the variance of the predictive distribution. Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e. the variance of the model parameters. A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy defined as
$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\mathbb{E}\,\left[-\log p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})\right]$ (3)
or variance
$\mathrm{Var[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]}=\mathbb{E}\,[\,p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})^{2}\,]-\mathbb{E}\,[\,p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})\,]^{2}.$ (4)
III-B Conventional segmentation
The so-called “deterministic” segmentation networks are trained by Maximum Likelihood Estimation (MLE) as
$\bm{\theta}_{MLE}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta}),$ (5)
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the variance of the model parameters approaches zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution. This is typically done through Bayesian Maximum A Posteriori (MAP) estimation with
$\bm{\theta}_{MAP}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (6)
For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the L2-norm (also known as ridge regression or weight decay) or L1-norm of $\bm{\theta}$, respectively [41, 42]. Additionally, the output of such models can be interpreted as a probability distributiont For instance, let a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$, where we denote the input image and segmentation dimensionality as $\mathbb{R}^{C\times D}$ and $\mathbb{R}^{K\times D}$, respectively, for concise notation. Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
$p(\mathbf{Y}=k\,|\,\mathbf{x}^{*},\bm{\theta})=\frac{e^{\mathbf{a}_{k}}}{\sum_{k}e^{\mathbf{a}_{k}}},$ (7)
with channel-wise indexing over in the denominator, which is commonly known as the SoftMax activation. See Figure 2 for an illustration. While not referred to as such in common nomenclature, this is probabilistic modeling in the technical sense, and the approximated distribution can represent and localize uncertain regions. However, the implicit pixel-independence assumption
$p(\mathbf{Y}|\mathbf{X})=\prod^{K\times D}_{i}p(Y_{i}|\mathbf{X}),$ (8)
omits information on structural variation in the segmentation masks. At the very core, this limitation (See Figure 3) has driven research on probabilistic segmentation, enabling the sampling of spatially coherent segmentation masks. This challenge can be addressed either from the aleatoric or epistemic perspective, inferring Bayes theorem w.r.t. the hidden latent variables or model parameters, respectively. Both approaches have relevant specific use-cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
Figure 2: Aleatoric uncertainty quantification by modeling pixel-level outputs as parameters of a probability mass function.
(a) Likelihoods
(b) Thresholding
(c) Sampling
Figure 3: Interpretation of the likelihood function in segmentation models. Color intensity in (a) reflect the normalized confidence values that can be interpreted as probabilities. Maximum likelihood thresholding can be applied to obtain (b). However, the coherence of the segmentation suffers when one attempts to sample, as is shown in (c).
IV Aleatoric uncertainty
Aleatoric uncertainty quantification reconsiders the non-deterministic relationship between $\mathbf{x}\in\mathcal{X}$ and $\mathbf{y}\in\mathcal{Y}$, which implies that
$p(\mathbf{Y}|\mathbf{X})=\frac{p(\mathbf{Y},\mathbf{X})}{p(\mathbf{X})}\neq\delta(\mathbf{Y}-F(\mathbf{X})),$ (9)
with Dirac-delta function $\delta$, and arbitrary function $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$, and is inherently probabilistic due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, detecting substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition. The possible input-dependency drives further categorization in to either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, one can model a distribution parameterized by the output of a CNN (Section IV-A). Also, generative models have extensively been used, where the data likelihood is learned through latent variables (Section IV-B). Finally, image augmentation during test-time inference can also be applied to obtain a notion of aleatoric uncertainty (Section IV-C).
IV-A Pixel-level sampling
A valid approach employs direct pixel-level uncertainty in the annotations. As discussed before, uncertainty can be quantified in case the independence assumption holds. In this case, it is important to ensure proper calibration, which is discussed in Section IV-A1. In another method, the spatial correlation is explicitly modeled across the pixels of the segmentation mask and results in spatially coherent samples, which will be treated in Section IV-A2.
IV-A1 Independence
The normalized confidence values that result from SoftMax activation can only be interpreted as probabilities after proper validation, which is referred to as model calibration. Ideally, the empirical accuracy of a model should approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over- confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Different methods can be used to measure calibration, but often introduce their own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy and confidence bins as
$ECE=\sum_{b=1}^{B}\frac{n_{b}}{N}|\operatorname{acc}(b)-\operatorname{conf}(b)|,$ (10)
with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins. The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under- confidence. Furthermore, Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
(a) Calibrated
(b) Miscalibrated
Figure 4: Reliability diagram indicating if the model confidence accurately reflect the empirical probabilities.
Contemporary SoftMax-activated neural networks often portray a significantly incongruous reflection of the true data uncertainty because of negative log-likelihood optimization and techniques such as batch normalization, weight decay and other forms of regularization [9]. Most calibration techniques are post-hoc, i.e. they occur after training and thus require a separate validation set. For example, Temperature Scaling [9] has been applied in a pixel-wise manner for segmentation problems [43]. Nonetheless, some methods, such as Label Smoothing [44, 45] or using Focal loss [46] can directly be applied on the training data. Furthermore, overfitting has often been considered to be the cause of overconfidence [47, 48] and erroneous pixels can therefore be penalized through regularizing low-entropy outputs [49].
IV-A2 Spatial correlation
Figure 5: Depiction of Stochastic Segmentation Networks [50]. Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
. To explicitly model spatial correlation within pixel space of the likelihood distribution, Monteiro et al. [50] propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
$p(\mathbf{a}|\mathbf{x},\bm{\theta})=\mathcal{N}(\mathbf{a};\bm{\mu}=f^{\bm{\mu}}_{\bm{\theta}}(\mathbf{x}),\bm{\Sigma}=f^{\bm{\Sigma}}_{\bm{\theta}}(\mathbf{x})),$ (11)
where the covariance matrix has low-rank structure
$\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}.$ (12)
Here, $\mathbf{P}$ has dimensionality $((H\times W\times K)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ a diagonal matrix. This results in a more structured and expressive distribution, while retaining reasonable efficiency. As the SoftMax transform on this low-rank multivariate normal distribution pertains an intractable integral, Monte Carlo sampling is employed. SSNs can be augmented as an additional layer to any existing CNN (see Figure 5).
Another method uses an autoregressive approach to predict pixel $Y_{i}$ based on the preceding pixels. In this case, we can rephrase Equation (8) to
$p(\mathbf{Y}|\mathbf{X})=\prod_{i}^{K\times D}p(Y_{i}|Y_{1},...,Y_{i-1},\mathbf{X}).$ (13)
The PixelCNN remains a popular method to model such relationship due to its substantial receptive field [51, 52]. Zhang et al. [53] suggest to use this to predict a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$). Fusing the two masks is done through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples. See Figure 6 for an illustration. Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
Figure 6: An illustration of the PixelCNN-based PixelSeg [53]. ‘C’ indicat es the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation. Partially transparent elements appear during test-time sampling.
IV-B Latent-level sampling
Directly learning the conditional data distribution is a challenging task. Hence, generative models often employ a simpler latent (unobserved) variable $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to then learn the approximate joint density $p(\mathbf{Y}|\mathbf{Z},\mathbf{X})$. The marginalized distribution is obtained through decomposition
$p_{\bm{\theta},\bm{\psi}}(\mathbf{Y}|\mathbf{X})=\int p_{\bm{\theta}}(\mathbf{Y}|\mathbf{z},\mathbf{X})p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\mathrm{d}\mathbf{z},$ (14)
with parameters $\bm{\theta},\bm{\psi}$. Conditioning the latent density on the input images is not a necessity but usually preferred for smooth optimization trajectories [54]. As such, the spatial correlation is not explicitly modeled but rather induced through mapping the latent variables to segmentation masks. Notable architectures relevant to the context of this paper are Generative Adversarial Networks (Section IV-B1), Variational Autoencoders (Section IV-B2) and Denoising Diffusion Probabilistic Models (Section IV-B3).
IV-B1 Generative Adversarial Networks
A straightforward approach is to simply learn the decomposition in Equation (14) through sampling from an unconditional prior density
$p_{\mathbf{Z}}=\mathcal{N}(\bm{\mu}=\mathbf{0},\bm{\Sigma}=\bm{\sigma}\cdot\mathbf{I}),$ (15)
and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$. Goodfellow et al.  [55] show that this approach can notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$. In this way, $G_{\bm{\phi}}$ learns reconstruct realistic looking images, guided by the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient resistance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative. We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
$J_{G}=-J_{D}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]$ (16)
$-\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[\,\log(1-D_{\bm{\psi}}(G_{\bm{\phi}}(\mathbf{z},\mathbf{x})))\,].$
Figure 7: The Calibrated Adversarial Refinement network [56] based on the Generative Adversarial Network with additional loss terms.
While conditional GANs had been used for semantic segmentation before [57], Kassapis et al. [56] explicitly contextualized the architecture within aleatoric uncertainty quantification using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7). The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\mathbf{c}$, with (cross entropy) reconstruction loss
$\mathcal{L}_{rec}=-\mathbb{E}_{p_{\mathcal{D}}}[\log p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})].$ (17)
Then, the conditional refinement network $G_{\bm{\theta}}$ uses $\mathbf{c}$ together with $\mathbf{X}$ and latent samples $\mathbf{Z}_{i}\sim p_{\mathbf{Z}}$ injected at multiple decomposition scales $i$, to predict various segmentation maps. Furthermore, the refinement network is subject to adversarial objective
$\mathcal{L}_{adv}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (18)
which is argued to elicit superior structural qualities compared to relying on cross-entropy alone. At the same time, the discriminator opposes the optimization with
[…]
$\hskip 5.0pt-\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,].$
Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are then compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities denoted as
$\mathcal{L}_{cal}=\mathbb{E}_{p_{\mathcal{D}}}\operatorname{KL}[\,p_{\bm{\phi}}(\mathbf{y}|\mathbf{c},\mathbf{x})||p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})\,].$ (20)
In this way, the generator loss can be defined as
$\mathcal{L}_{G}=\mathcal{L}_{adv}+\lambda\cdot\mathcal{L}_{cal},$ (21)
with hyperparameter $\lambda\geq 0$. The purpose of the calibration network is argued to be three-fold. Namely, it (1) sets a calibration target for $\mathcal{L}_{cal}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification. The refinement network can be seen as modeling the spatial dependence across the pixels, which enables sampling coherent segmentation maps through latent variable $\mathbf{Z}$.
IV-B2 Variational Autoencoders
Techniques such as GANs rely on implicit distributions and are void of any notion of data likelihoods. An alternative approach estimates the Bayesian posterior w.r.t. the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$ obtained my maximizing conditional Evidence Lower Bound (ELBO)
$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
$=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
$\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
$\hskip 50.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,].$ (22)
This is also known as Variational Inference (VI). The first term in Equation (22) represents the reconstruction cost of the decoder subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Normal densities and amortized through neural networks parameterized by $\phi$, $\theta$ and $\psi$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
[…]
Implementing the conditional ELBO in Equation (22) can be achieved through a VAE-like architecture [10]. A few additional design choices specific for uncertainty quantification for segmentation result in the The Probabilistic U-Net (PU-Net) [58]. Firstly, the latent variable $\mathbf{z}$ is only introduced at the final layers of a U-Net conditioned on $\mathbf{X}$, where the vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolutions. When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation is performed in the U-Net, while the information $\mathbf{Z}$ provides is almost exclusively regarding the segmentation variability. Therefore, relatively smaller values of $d$ are feasible than what is commonly used in image generation tasks.
Figure 8: The Probabilistic U-Net [58] based on a conditional Variational Autoencoder [10]. Latent samples are inserted at the final stages of a U-Net through a tiling operation.
Similar to related research on the VAE [59, 60, 61, 62, 63], much work has been dedicated to improving the PU-Net. For instance, investigation on improving VI with novel parameterization of the amortized densities also provide interesting insights into model behaviour [64, 65, 66, 67, 68, 69, 70]. Furthermore, extending the architecture to multiple decomposition hierarchies [71, 72], three-dimensional modalities [73, 74, 75, 76, 77] and conditioning on the annotator [78] also resulted in substantial performance gains.
Density parameterization
Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness [63]. This phenomena has also been successful for cVAE-like models such as the PU-Net [64, 66]. NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $\mathbf{f}=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex distributions $\log p(\mathbf{x}|\mathbf{z})$. These are often denoted as $\log p(\mathbf{x})$ for simplicity and can be determined through
$\log p(\mathbf{x})=\log p_{\mathbf{z}}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (24)
where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are intermediate variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$. Equation (24) can be substituted in the conditional ELBO objective in Equation (22) to obtain
$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,]$
$\hskip 20.0pt-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (25)
where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF.
Bhat et al. [67, 68] compare this approach with other parameterizations of the latent space including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. [79] show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse [80, 59]. Their proposition considers the alternative formulation of the ELBO
$\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (26)
$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})||p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]$
$\hskip 70.0pt-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}),$
which the novel objective maximizes the contribution of the (expected) mutual information between latent and output variables. This can be rewritten to the objective
$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (27)
$\hskip 40.0pt+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]$
$\hskip 80.0pt+\beta\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})||p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$
with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence [81] and $\alpha,\beta$ being hyperparameters, which results in a more uniform latent space leading to increased model performance. Also, modeling the ELBO of the joint density has also been explored [69]. This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data. Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this counters the model collapse phenomena in latent space [70].
Multi-scale approach
Learning latent features over several hierarchical scales can provide expressive densities and interpretable features across various abstraction levels [82, 83, 84, 85, 86]. Such models commonly fall under hierarchical VAE umbrella term. Often, an additional Markov assumption of length $T$ is placed on the posterior as
$q_{\theta}(\mathbf{Z}_{1:T}|\mathbf{Y},\mathbf{X})=q_{\theta}(\mathbf{Z}_{1}|\mathbf{Y},\mathbf{X})\prod^{T}_{t=2}q_{\theta}(\mathbf{Z}_{t}|\mathbf{Z}_{t-1}).$ (28)
Consequently, the conditional ELBO is denoted as
$p(\mathbf{\mathbf{Y}|\mathbf{X}})\geq$ $\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$ (29)
$-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})||q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]$
$\hskip 20.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
This objective is implemented in the Hierarchical PU-Net [71] (HPU-Net, depicted in Figure 9). Simply stated, the architecture learns a latent representation at multiple decomposition levels of the U-Net. Furthermore, residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL-divergence rapidly approaching zero. For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective was employed, which extends on Equation (29) as
$\mathcal{L}_{GECO}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (30)
$-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})||q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]$
$\hskip 20.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$ set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick [87, 88]. PHiSeg [72] takes a similar approach to the HPU-Net. However, instead of placing the residual connections in the convolutional layers, PHiSeg uses these between the latent vectors across decomposition. Furthermore, auxiliary outputs, or deep supervision, at each decomposition scale is used to enforce disentanglement across latent variables.
Figure 9: Hierarchical Probabilistic U-Net [71] based on a hierarchical Variational Autoencoder [10, 82, 84]. Instead of a single latent code, multiple decomposition scales encode the segmentation variability, depending on the depth of the U-Net structure.
Extension to 3D
Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to potential loss of critical spatial information and subtle nuances often necessary for accurate delineation. This limitation has spurred extensive research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures. Initial works [76, 77] demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants. Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation. Viviers et al. [75] additionally augment a Normalizing Flow to the posterior density, as described in Section IV-B2. Further enhancements to the architecture include the implementation of the 3D hierarchical PU-Net [74] or an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications [73].
Conditioning on annotator
It can be relevant to model the annotators independently in cases with consistent annotator-segmentation pairs in the dataset. This can be achieved by conditioning the learned densities on the annotator itself [78, 89]. For example, features of a U-Net can be combined with samples from annotator specific Gaussian distributed posterior distributions [89]. Considering the approach from Gao et al. [78], generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ with for the $k$-th annotator. Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through image-conditional decoder as $\mathbf{y}=F(\mathbf{x},\mathbf{z}_{k})$. The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble. Additionally, consistency between the model and ground-truth distribution is additionally enforced through an optimal transport loss between the set of predictions and labels.
IV-B3 Denoising Diffusion Probabilistic Models
Recent developments in generative modeling have resulted a family of models known as Denoising Diffusion Probabilistic Models (DDPMs) [90, 91, 92]. Such models are especially renowned for their expressive power by able to encapsulate large and diverse datasets. While several perspectives can be used to introduce the the DDPMs, we build upon the earlier discussed HVAE (Section IV-B2) to maintain cohesiveness with the overall manuscript. In specific, let us introduce three additional modifications to the HVAE. Firstly, the latent dimensionality is set equal to the data dimensions, i.e. $d\,$=$\,D$. As a consequence, redundant notation of $\mathbf{Z}$ is removed and $\mathbf{Y}$ is instead subscripted with $t\in\{1,...,T\}$, indicating the encoding depth, where $\mathbf{Y}_{0}$ is the initial segmentation mask. Secondly, the encoding process (or forward process) is predefined as a linear Gaussian model such that
$q(\mathbf{y}_{T}|\mathbf{y}_{0})=p(\mathbf{y}_{0})\prod^{T}_{t=1}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$ (31)
and
$q(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\bm{\mu}=\sqrt{\alpha_{t}}\mathbf{y}_{t-1},\bm{\Sigma}=(1-\alpha_{t})\cdot\mathbf{I}),$ (32)
with noise schedule $\bm{\alpha}=\{\alpha_{t}\}^{T}_{t=1}$. Then, the decoding or reverse process can be learned through $p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$. The ELBO for this objective is defined as
$p(\mathbf{y}$ $|\mathbf{x})\geq\mathbb{E}_{q(\mathbf{y}_{1}|\mathbf{y}_{0})}[\log p_{\phi}(\mathbf{y}_{0}|\mathbf{y}_{1},\mathbf{x})]$ (33)
$+\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{y}_{t}|\mathbf{Y}_{0})}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})||p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]$
$\hskip 40.0pt+\underbrace{\mathbb{E}_{q(\mathbf{y}_{T}|\mathbf{y}_{0})}\left[\log\frac{p(\mathbf{y}_{T})}{q(\mathbf{y}_{T}|\mathbf{y}_{0})}\right]}_{\approx 0}.$
As denoted, the regularization term is assumed to be zero, since we assume that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed. With the reparameterization trick, the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$. As such, the KL-divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$ depending on the parameterization, and is in almost all instances approximated with a U-Net [2].
It has also be proposed to model Bernoulli noise instead of Gaussian [93, 94, 95, 96]. However, most methodologies vary in the conditioning of the reverse process on the input image [97, 98, 99, 100]. For instance, Wolleb et al. [97] concatenate the input image with the noised segmentation mask. Wu et al. [98] insert encoded image features to the U-Net bottleneck. Additionally, information on predictions $\mathbf{Y}_{t}$ at a time step $t$ is also provided in the intermediate layers of the conditioning encoder. This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attentive map and the inverse FFT. The procedure of applying attention on the spectral domain of the U-Net encoding has also been done with transformers in follow-up work [99]. Segdiff [100] also encode both current time step and input image, but combine the extracted features by simple summation before applying the U-Net.
Figure 10: An illustration of the Diffusion Probabilistic Models [90]. The model learns to remove noise that has gradually been added to the input image.
IV-C Test-time augmentation
An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest. For example, systematic noise, translation or rotation result in many instances realistic variations that approximately retain image semantics. Hence, data augmentation [31] has been used at test time (hence the name test-time augmentation, or TTA) for image classification to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function [101]. This technique has been applied to image segmentation as well [102, 103, 104, 105]. By randomly augmenting input images with invertible transformation $T$ as $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$. Repeatedly performing this procedure results in a set segmentations masks, which can serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
V Epistemic uncertainty
The crucial difference between epistemic and aleatoric uncertainty is that the former is related to model ignorance, while the latter reflects statistical ambiguity inherent in the data. Epistemic uncertainty can be further categorized into two distinct types [12]. The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, this is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data. Both model and approximation uncertainty contribute to the epistemic uncertainty.
Unfortunately, evaluation of the true Bayesian posterior (formulated in Equation (1)) is inhibited by the intractability of the data-likelihood in the denominator. Hence, extensive efforts have been taken to obtain viable approximations such as using Mean-Field Variational Inference [8], Markov Chain Monte Carlo (MCMC) [106], Monte-Carlo Dropout [107, 108], Model Ensembling [109], Laplace approximations [110], Stochastic Gradient MCMCs [111, 112, 113], assumed density filtering [114] and expectation propagation [115, 116]. We refer to any neural network that approximates the Bayesian posterior of the model parameters as a Bayesian Neural Network (BNN). The following sections treat methodologies that have been applied within the context of this paper, which are usually straightforward extensions of BNN networks used for conventional regression and classification tasks. Additionally, an illustration of these techniques are presented in Figure 11.
V-A Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Then, we can achieve Variational Inference (VI) w.r.t. to the parameters by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior can be written as
$\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathbf{y},\mathbf{x})\,\right]$
$=\operatorname*{arg\,min}_{\bm{\theta}}\int q(\bm{\theta}|\bm{\eta})\log\frac{q(\bm{\theta}|\bm{\eta})}{p(\mathbf{Y}|\mathbf{x},\bm{\theta})p(\mathbf{x},\bm{\theta})}d\bm{\theta}$
$=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[q(\bm{\theta}|\bm{\eta})||p(\bm{\theta})\right]-\mathbb{E}_{q(\bm{\theta}|\bm{\eta})}[\,\log p(\mathbf{y}|\mathbf{x},\bm{\theta})\,],$ (34)
where the parameter-independent terms are constant and therefore excluded from notation. A popular choice for the approximated variational posterior is a the Gaussian distribution, i.e. a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities. However, the priors can also be learned through Empirical Bayes [117]. Furthermore, backpropagation is possible with the reparameterization trick [10] and within this context, the procedure is referred to as Bayes by Backprop (BBB) [8]. During testing, a sample-based approach is utilized to approximate the posterior. Since using several parameter permutations effectively enriches the hypothesis space due to the model combining effect and can already improve performance [118, 119, 120].
V-B Monte Carlo Dropout
Dropout, a common technique used to regularize neural networks [121], can mimic sampling from an implicit parameter distribution, $q(\tilde{\bm{\theta}}|\bm{\theta},p)$, defined as
$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p)$ (35a)
$\tilde{\bm{\theta}}$ $=\bm{\theta}\odot\mathbf{n},$ (35b)
with probability $p$ and $\mathbf{n}$ operating element-wise on the parameters. Using Dropout can also be interpreted as a first-order equivalent $L_{2}$-regularization with additional transforming the input by the inverse diagonal Fisher information matrix [121]. With Monte-Carlo Dropout (MC Dropout), the random node switching is continued during testing, effectively sampling new sets of parameters. While seemingly arbitrary, it has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process [107]. In this manner, the authors showcase such method is able to provide multi-modal estimates of the model uncertainty.
As noted by Gal et al. [122], the model output variance is balanced with the weight magnitudes rather than the dropout rate $p$, which is usually optimized through grid search or simply fixed to $0.5$. Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, such that uncertainty estimates are governed by $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution [123, 87], to enable path-wise derivatives through $p$.
V-C Model Ensembling
As mentioned earlier, Monte Carlo dropout effectively optimizes over a set of sparse neural networks. This ensemble can also be designed in an explicit manner. Let us define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{N}$, with $N$ representing the number of models in the ensemble. Then, it is relatively simple to obtain $\bm{\Theta}=\{\bm{\theta}_{n}\}_{n=1}^{N}$, which can be interpreted as samples from an approximate posterior. Ensembling in only the latter parts of a neural network (typically the decoder) is referred to as M-heads, i.e. the network has multiple outputs. Often, the $N$ obtained parameters are from $N$ separate training sessions. However, it has also been shown effective to ensemble from single training session by saving the parameters at multiple stages or training with different weight initializations [124, 125, 126].
(a) Variational Inference
(b) Monte Carlo Dropout
(c) Ensembling
Figure 11: Varying techniques to sample parameters for the convolutional kernels. An approximation of the parameter density can be made (a), taking samples can be mimicked (b), or an ensemble of $N$ different configurations can explicitly be modeled (c).
A closely related concept to ensembling is known as Mixture of Experts (MoE), where each model in the ensemble (an ‘expert’) is trained on specific subsets of the data [127]. In such settings, a gating mechanism is usually applied after combining the expert hypotheses. While uncommon, incorporating all predictions can also be regarded as ensembling technique.
VI Applications
This section explores literature that employs uncertainty-based downstream tasks on segmentation models. These include estimating the segmentation mask distribution subject to observer variability (Section VI-A), model introspection (i.e. ability to self-assess, Section VI-B), improved generalization (Section VI-C) and reduced labeling costs using Active Learning (Section VI-D).
VI-A Observer variability
After observing sufficient data, the variability in the predictive distribution is often considered to be negligible and is therefore omitted. Nevertheless, this assumption becomes excessively strong in ambiguous modalities, where its consequence is often apparent with multiple varying, yet plausible annotations for a single image. Additionally, such annotations can also vary due to differences in expertise and experience of annotators. This phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator is referred to as the intra-observer variability (see Figure 12).
To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be treated as models themselves. For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with $k=1,2,...,K$. For a simple segmentation task, it can be expected that $\operatorname{Var}[p(\bm{\phi}_{k})]\rightarrow 0$. In other words, each annotator is consistent in their delineation and the intra-observer variability is low. For cases with consensus across experts, i.e. negligible inter-observer variability, the marginal converges to $\operatorname{Var}[p(\bm{\phi})]\rightarrow 0$. Asserting these two assumptions, it is valid to simply consider a point estimate of the posterior. Yet, this rarely the case in many real-life application and, as such, explicitly modeling the involved distributions becomes imperative.
Figure 12: A visualization of intra-observer variability in parameter space (left) and inter-observer variability in data space (right).
For evaluation, a common metric is to minimize the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick. This know as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [128], denoted as
$\operatorname{GED}^{2}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
$\hskip 10.0pt+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]$
$\hskip 20.0pt-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (37)
with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and some kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the 1--IoU or 1--Dice score. An alternative metric known as Hungarian Matching (HM) compares the predictions against the ground-truth labels through a cost matrix [71]. Subsequently, the unique optimal coupling between the two sets that minimizes the average cost is determined through combinatorial optimization algorithm. This can also be formally denoted as finding the permutation matrix $\mathbf{P}$ subject to the objective
$\operatorname{HM}(Y,\hat{Y})=\min_{\mathbf{P}}\frac{1}{N^{2}}\operatorname{Tr}(\mathbf{P}\mathbf{M}),$ (37)
where $M_{i,j}=k(y_{i},\hat{y}_{j})$ and $N^{2}$ represents the number of elements of the matrix.
For accurate comparison and evaluation of the problem, benchmarking is often constrained to publicly available multi-annotated data. For example, the Lung Image Database Consortium and Image Database Resource Initiative (LIDC-IDRI) [129] contains manually annotated lesions from lung patients and is one of the few datasets that has extensively been evaluated with appropriate metrics (see Table I). Multiple versions have been used, either with 15,096 (LIDCv1) or 12,816 (LIDCv2) patches. LIDCv1 is employed with a 60:20:20 split while for LIDCv2 this is 70:15:15 for train:validation:test. LIDCv2 has also been used with threefold cross validation (LIDCv2-cv), with either a 90:10 or 80:20 train:test split.
Other less commonly datasets include the CityScapes [130] dataset, which contains street view images of German cities from the perspective of a driving car, have been used to artificially create class-level label ambiguity [71, 58, 78]. Some classes are switched to an arbitrary auxiliary class with some probability $p$. Since the underlying probabilities are known, the empirical fraction of the classes from the model predictions can directly be compared to the ground-truth values of $p$. Furthermore, the QUBIQ Challenge [131] contains MRI and CT data from varying organs. Also, the retinal fundus images for glaucoma analysis (RIGA) [132] dataset contains delineations of optic cup and disc boundaries by six experienced ophthalmologists.
TABLE I: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table adapted from [94].
LIDCv1 LIDCv2
Method Year GED16 HM-IoU16 GED16 HM-IoU16
PU-Net [58] 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
HPU-Net [71] 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
PhiSeg [72] 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
SSN [50] 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
CAR [56] 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
JProb. U-Net [133] 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
PixelSeg [53] 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
MoSE [78] 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
AB [134] 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
CIMD [95] 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
CCDM [94] 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotations pairs, provided that annotators have an intrinsically associated expertise [135]. Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data can become quite burdensome. Hence, it can be much simpler to directly model the empirical stochasticity in the annotations and has been extensively explored in modalities such as Lung nodule detection in 2D [58, 72, 71, 136, 137, 70, 66, 68, 65, 50, 64], as well as 3D [76, 138, 139], Brain Tumor [140, 50, 97, 93, 95], White Matter Hyperintensities [141], Pulmonary Tumour Growth [142], Prostate [143, 73] and vascular [144], street scene [145, 71, 94], aerial imaging [100], optic cup [98, 146], abdominal multi-organ [99] and nuclei microscopy [96] segmentation. The overwhelming majority of research for this particular application has been executed with variants of conditional VAEs [58, 66, 64, 65, 70, 68, 71, 72, 137, 142, 147, 140, 148, 139, 144, 76, 73, 143, 138]. More recently, the growing popularity of DDPMs is also apparent in the field [94, 100, 97, 93, 95]. Also, GAN-based approaches have been also employed [56].
VI-B Model introspection
The uncertainties obtained from probabilistic models can provide an insight into the reliability of a model when the correlation between uncertainty and model accuracy is strong. This relationship has been formalized by Mukhoti et al. [149] through two conditional likelihoods. Firstly, the accuracy given a certain prediction, $p(\mathrm{A}|\mathrm{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\mathrm{U}|\mathrm{I})$. Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{ac}$, $u_{au}$, $u_{ic}$, $u_{iu}$, respectively. Consequently, the authors combine $p(\mathrm{A}|\mathrm{C})=\frac{n_{ac}}{n_{ac}+n_{au}}$ and $p(\mathrm{U}|\mathrm{I})=\frac{n_{ic}}{n_{ic}+n_{iu}}$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
$\mathrm{PAvPU}=\frac{n_{ac}+n_{au}}{n_{ac}+n_{ic}+n_{ic}+n_{iu}}.$ (38)
This downstream task has been applied to (video) street scene [150, 151], remote sensing [152, 153], instance segmentation of various objects [154], point-cloud indoor scenes [155] brain [156, 157, 158, 17, 159, 160, 161], Multiple Sclerosis [162], cardiac [163, 164], heart ventricle [161], prostate [161], carotid artery [165], lumbosacral [19] MRI, and Optical Coherence Tomography [166, 167], skin imaging [168, 169], lung [169] and liver CT [170] and MRI  [171], and ultrasound [124, 172]. Concrete dropout has been applied to instance segmentation of C. Elegans assays [173] and street-scenes [149]. To a lesser degree, ensembling [174, 124], Variational Inference both in 2D [175, 176] and 3D [177], M-heads (auxiliary networks) [14, 178, 179], and test-time augmentation [124] have also been used to quantify the uncertainty for quality assessment.
It can be noted that uncertainty is usually only obtained on a pixel basis, while crucial information can be present in structural statistics. Hence, the Coefficient of Variation (CV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples. Also, Roy et al. [160] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to then determine the pair-wise average overlap between all respective samples as
$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\left\{\mathrm{Dice}(\mathbf{Y}_{i}=t(\mathbf{y}),\mathbf{Y}_{j}=t(\mathbf{y}))\right\}_{i\neq j}\right].$ (39)
Finally, including information on localized uncertainty to the training objective has shown to improve generalization capabilities [180, 181, 182, 183]. Note that using uncertainty to guide model training is closely related to Active Learning, which is discussed in Section VI-D.
VI-C Model generalization
As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance due to the model combining effect. For instance, dropout layers at the deeper decomposition levels of the SegNet [4] improves model performance [150]. Literature also reports improved performance with Concrete Dropout for semantic (instance) segmentation [149, 173]. This has been observed in multiple domains including out-/indoor scene understanding [150, 149, 184, 185, 179], brain tumor MRI [157, 186, 159, 175], Optical Coherence Tomography [166], low dose Computed Tomography of lung nodules [187], colectoral polyps [188], cardiac MRI [18] and C.elegans roundworm microscopy images [173]. Ng et al. [18] benchmark multiple techniques for cardiac MRI segmentation and find that ensembling results in best performance improvement, while Bayes By Backprop [8] is more robust to noise distortions.
Furthermore, the improved generalization from ensembling has also shown to produce more calibrated outputs [161]. In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits [49]. Nonetheless, individual models in conventional ensembles receive data in an unstructured manner. However, it is also possible to assign specific subsets of the data to particular models (so-called ‘experts’) in the ensemble [127], commonly referred to as a Mixture of Experts (MoE). While MoE resembles ensembling in many ways, the approach additionally relies on a learnable gate that inherits the decision-taking logic. Pavlitskaya et al. [179] show merits of such approach has been observed in urban outdoor scene segmentation. For the optic cup segmentation, Ji et al. [135] also condition on a normalized expertness vector, where each each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net. Gao et al. [78] introduced Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach. Nevertheless, the methodology addresses aleatoric uncertainty quantification and has hence been discussed in Section IV-B2.
VI-D Active Learning
The field of active learning [189, 190] aims to reduce the costly annotation procedure by careful selection of unlabeled training samples. A wide range of methodologies exist, but since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate. In terms of metrics, the expected increase in posterior entropy, defined by
$H[\,\bm{\theta}|\mathcal{D}\,]-\mathbb{E}_{p(\mathbf{y}|\mathbf{x}^{*},\mathcal{D})}[H[\,\bm{\theta}|\mathbf{y},\mathbf{x}^{*},\mathcal{D}\,]],$ (40)
can provide a notion to describe information gain, and thus the uncertainty, from specific datapoints. Notably, minimizing the expected increase in posterior entropy is equivalent to maximizing the mutual information between the data and model parameters. This enables reformulation in terms of the output space, rather than the complex parameter space and is known as Bayesian Active Learning by Disagreement (BALD) [191], denoted as
$I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D})$ $=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]].$ (41a)
Active Learning through uncertainty quantification has been been beneficial for medical [192, 193, 194, 195, 196, 197, 198, 199] multi-view [200], remote sensing [201], street-views [202] and 3D point-cloud data [203]. These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [183] and novel temporal-ensembling methods [204] have also been used to reduce required human labeling. Some methods extend beyond pixel-level uncertainty, incorporating boundary information in the uncertainty quantification process [202, 198]. For example, Kasarla et al. [202] make use of both image and pixel-level uncertainty. Furthermore, pixels are weighed by their closeness to edges, as boundary cases are more likely to be uncertain. Similarly, Ma et al. [198] combine target and boundary-based uncertainty sampling, ensuring diversity, effective and balanced utilization of the available information.
VII Discussion
This section builds on the preceding, by highlighting the main trends and discussion points in the field. The challenges and limitations related to the methods (Section VII-A) and downstream applications (Section VII-B) are discussed. Based on these observations, recommendations for future work will be provided in Section VII-C.
VII-A Methods
The core challenge with modeling segmentation distributions entails calibrated uncertainty estimates and secondly, producing spatial coherence in its samples. While the former can be achieved with conventional CNN models, the latter requires modeling correlation across pixels of the segmentation mask. As discussed in earlier sections, the desired nature of the uncertainty estimates governs the required methodology. Hence, our discussion will follow a similar format as the preceding sections, i.e. the aleatoric and epistemic approaches will initially be treated in a separate manner.
VII-A1 Aleatoric methods
The literature overview indicates that three distinct routes can be taken to model the correlation. The first method is to model the correlation directly in pixels space. The second method entails using latent-variable modeling, where the correlation is encoded in latent space. Finally, test-time augmentation is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models. We will discuss the strengths and weaknesses of each of those approaches. A summary of this is presented in Table II.
TABLE II: Comparison between models that quantify aleatoric uncertainty.
Method Advantages Disadvantages Examples
TTA model agnostic, no additional training parameters implicit likelihoods [102, 103, 104, 105]
SSN model agnostic, explicit likelihoods, fast sampling unstable training [50]
PixelCNN explicit, exact likelihoods sequential sampling, memory intensive [53]
GAN fast sampling, flexible unstable training, poorly defined objective, implicit likelihoods [56, 57]
VAE fast sampling, flexible, interpretable latent space, ELBO mode/posterior collapse, amortization gap [58, 71, 72, 64, 66, 65, 68, 67, 139, 89, 78, 133, 70]
DDPM flexible, expressive sequential sampling [97, 98, 99, 100, 93, 94, 95, 96]
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification. Nonetheless, the shortcomings of VAEs are well-known. For example, such models suffer from inference suboptimality related to ELBO optimization [205, 59] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [65, 79, 70], which is hypothesized to be caused by excessively strong decoders [80]. This is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [71], residual connections [71, 72] or deep supervision [72] are required for generalization. A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotator styles or the exploration of low-likelihood regions. Hence, VAE-based models serve as good choice for the task given its shortcomings are sufficiently addressed.
Compared to the VAE-based models, the adoption of DDPMs is rather limited. This is regardless of the fact that it outperforms the VAE-based methods. Besides the point that DDPMs emerged much later in the field, its crucial limitation is regarding the tedious sequential inference procedure [92, 206, 207]. This shortcoming is exacerbated in supervised settings, which often validate through sampling on a separate data split. Furthermore, it can be noted that the best performing DDPM models are discrete in nature. While it is debatable whether shifting to categorical distributions is required for complex image generation [134], this observation does signals further investigation of the merits of categorical distributions in segmentation settings. This can visually be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13). Regardless, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g. blurry reconstructions of VAEs).
(a) LIDC-IDRI [129]
(b) CityScapes [130].
Figure 13: Continuous vs. categorical forward diffusion process with cosine noise scheduling [208]. Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
Implicit methods such as test-time augmentation (TTA) [102, 103, 104, 105] have received some use in literature but have quickly been surpassed by alternative methods. However, TTA does lead in the category of simplicity, requiring almost no additional mechanisms or modifications to the employed architecture. Furthermore, CAR [56] is a lesser used method and this is likely due to well-known training problems of GAN-based models [209]. Similar to inference suboptimality in VAEs, GANs similarly required additional heuristic terms in the training objective due to instability. For example, the training objective of CAR is a summation of four separate losses.
SSNs [50] directly model correlations in pixel space and deems a simple, fast and model-agnostic modeling. Notably, SSNs also suffer from training instability due to the invertibility requirement of the covariance matrix. This can often simply be circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular. In practice, nevertheless, this quick-fix solution was much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
VII-A2 Epistemic methods
In the context of segmentation, almost all discussed literature employ approximations of Variational Inference. In fact, usage of MC Dropout dominates the realm of epistemic uncertainty quantification, mainly due it being a relatively simple, cheap and straightforward approach. Nevertheless, MC dropout has been subject to substantial criticism [210, 211, 108, 122]. For example, it has been shown that MC dropout can assign 0 probability to the true posterior, which can also erroneously possesses multi-modality [210]. Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate rather than the observed data [211, 108]. These weaknesses of dropout have been used as a basis for alternative dropout techniques where the dropout rate is learned [108, 122]. Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of a functional model. If accurate uncertainty quantification is critical, MC dropout should be avoided altogether.
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling [187, 160], while in other works ensembling excels [161]. Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout when evaluating with the PAvPU metric [149]. All things considered, the preference for a particular methodology is seems to carry a strong data-dependency [124, 14]. Our recommendation is to experiment with both explicit VI and approximations such as MC Dropout and ensembling.
VII-B Applications
Up until now, theoretical insights have mostly been discussed. In this section, a deeper dive is taken into the domain-/downstream- level applications of uncertainty quantification. See Table III for an overview. Notably, most models are employed in healthcare use cases, where ambiguity frequently arises due to the trade-off between accuracy and the incisiveness of medical diagnosis systems. Furthermore, in the automotive industry, sensors often operate under constraints, and objects of interest are typically at a considerable distance which induces ambiguity in the acquired images. It is evident that in these fields, uncertainty is primarily utilized to quantify observer variability and/or to correlate uncertainty with prediction accuracy. Research on improved generalization and uncertainty-based active learning is much sparser. In the following sections, each respective downstream task is further discussed.
TABLE III: Overview of domains using uncertainty quantification for segmentation.
Observer variability Model introspection Active Learning Model generalization
Lung [50, 58, 64, 65, 66, 68, 70, 71, 72, 136, 137, 76, 138, 139, 93] [169, 175, 187] [180] [187]
Brain [50, 97, 95, 140, 93, 98, 137, 142] [161, 157, 159, 158, 175] [195] [157, 186, 159, 175]
Outdoor scenes [71, 50, 145, 100, 154] [150, 151, 155, 149, 179, 152, 153] [200, 202] [149, 179, 150, 184, 185]
Cardiovascular [163, 164, 174, 161, 176, 124, 165] [192] [212]
Prostate [143, 73, 72, 136, 137] [161] [192]
Eye [99, 98, 66] [166, 167] [166]
Skin [99] [169, 170] [193, 196]
Indoor scenes [100, 154] [150, 155] [200, 203] [150]
Microscopy [96, 100, 71] [178, 173] [173]
Others [139, 65, 98, 64, 95, 137] [170, 171, 175, 19, 172] [192, 198, 95] [188]
VII-B1 Observer variability
Varying delineation hypotheses present themselves as “noise” in the ground-truth of the data. This observer variability is often a result of viewing limitations and therefore heavily correlate to the input image. In turn, this explains the success of learning input-conditional observer variability directly through explicitly modeling the likelihood distribution, rather than to infer the underlying latent parameter distribution. Nonetheless, this approach does have its limitations. For example, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles [213].
A significant challenge in this domain is the lack of standardization, making benchmarking extremely difficult. The reasoning for this is twofold. Firstly, it is evident from Table III that a wide range of datasets are used in literature, often involving proprietary in-house data. This makes replicating the presented results impossible. Secondly, consensus on data splitting is also lacking on the publicly available datasets. For example, significant incongruity across literature is observed for the LIDC-IDRI dataset regarding data splitting and preprocessing. Similarly, there is a lack of agreement on the implementation for the GED and HM-IoU metric. Several factors, including the choice of kernel, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
Regarding the evaluation metrics, we find that the literature often places excessive emphasis on improving them. This essentially forces the algorithm to predict segmentations identical to the available ground-truth masks. This can be problematic, as it defeats the original goal to predict plausible unseen segmentations, becoming a textbook case of Goodhart’s law: ”When a measure becomes a target, it ceases to be a good measure”. Furthermore, the quality of the GED evaluations have been subject to substantial criticism as well [214, 71, 79]. Hence, practitioners should be extra vigilant and consider the importance to inquire with domain-level experts for qualitative evaluation. Because the GED and HM-IoU evaluations are dependent on on the number of segmentation masks available an alternative is to involve additional annotators per data point for more accurate evaluation, but is likely to be equally expensive. This fact additionally calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.
VII-B2 Model introspection
Literature that evaluates models by correlating model uncertainty with error is, comparatively, much more thorough and standardized. This is especially evident by the frequent use of the PAvPU metric. Furthermore, a majority of the available research use Monte Carlo Dropout and correlated this to the output entropy to the prediction accuracy. The disadvantage of relying on MC Dropout when the uncertainty is crucial, has been discussed in Section VII-A2. Therefore, we recommend incorporating explicit Variational Inference at specific points in the network to achieve more accurate uncertainty quantification. These points could be selected based on the layers that most significantly influence the output segmentation.
VII-B3 Model generalization
Epistemic uncertainty quantification methods can result in improved performance as it often serves as a surrogate for model combination. Therefore, uncertainty quantification is can readily be available in the toolbox of deep learning researchers together with other commonly used regularization techniques. In this case, simple methods such as ensembling and MC dropout are despite their criticisms relatively harmless. Furthermore, parallels can be made with MC Dropout and placing an L2-norm penalty on the model weights. While, this benefits is compelling, improved model performance can also be obtained with more computationally efficient regularizers. Therefore, improved performance should not be the end goal of uncertainty quantification, but considered as an ancillary advantage.
VII-B4 Active Learning
Active Learning, while being a challenging task, holds significant promise by potentially reducing the need for intensive labeling procedures that often require specialized expertise. Furthermore, this kind of approach hints towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate adoption of such models in sectors requiring extensive specialization. Furthermore, Active Learning can accelerate privacy-centric collaboration when combined in a federating setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local modals.
However, active learning requires a well-generalized model on little data. This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and often requires datasets of substantial size. Furthermore, many uncertainty-based approaches often extend traditional active learning for classification by simply aggregating the pixel’s metrics together. For imbalanced settings, it has been shown that it can perform even worse than random selection [215]. Therefore, additional modifications such as target, boundary or diversity awareness [198, 203], or region-based annotating [202, 216] are often required to apply Active Learning to segmentation tasks.
VII-C Future work
In this penultimate section, we will provide point-by-point recommendations for future work.
VII-C1 Exploring generative models
Generative modeling, a rapidly evolving field, has successfully been employed for quantifying observer variability. The benefits of quickly applying its developments to segmentation problems have been evident with state-of-the-art DDPMs [90], which were initially proposed for unsupervised image generation. Also, literature on unsupervised VAEs have greatly benefited the PU-Net [65, 79, 70]. Therefore, we advocate for a deeper contextualization of contemporary research within probabilistic segmentation models. In fact, any unsupervised generative model can theoretically be translated to the supervised setting through intricate conditioning and can therefore be used as a probabilistic segmentation model.
Given this flexibility, it remains unclear why Normalizing Flow (NF) remains underutilized for this task. Specifically continuous NFs, which approximate the time-dependent score function, strongly resembles DDPMs which have extensively been used for segmentation problems. Furthermore, NFs enable explicit and exact evaluation of the likelihood, which can aid in further interpretation of model predictions. Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian and inference times are therefore much faster than that of DDPMs. It should be noted, however, that the invertibility constraint greatly hinders expressivity of intermediate functions, leading to memory-intensive architectures.
VII-C2 Variational inference and beyond
A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred over explicit VI). Additionally, recent studies have shown the benefits of moving from a few large to many small experts for when using a MoE ensemble (see Section V-C for language modeling [217]. Future work should also experiment with this.
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are useful when expecting to fit multi-modal parameter distributions. To the best of our knowledge, these approaches have not been studied within the context of uncertainty quantification in segmentation.
VII-C3 Single-pass uncertainty
The multiple forward passes required in Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards deterministic uncertainty models [218, 219, 220], which only depend on a single forward pass. Mukhoti et al. [218] show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. This approach achieves faster computation, while also providing both epistemic and aleatoric uncertainty.
Along similar lines, Evidential Deep Learning also possess the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) [221]. While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems [222], and Ancha et al. [223] recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model. Unfortunately, there has not been much research beyond this.
VII-C4 Improved CNN architectures
With the introduction of vision transformers [224], CNN-based models have been challenged in the task of semantic segmentation [225, 226, 227]. Regardless the success of vision transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net remain the preferred backbone [39]. This is mainly because CNNs already possess desirable inductive biases, while transformers conversely require extensive pretraining with large datasets [228]. Nonetheless, CNNs have benefited from the recent developments in transformers. For instance, “ConvNext” takes inspiration from contemporary transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains [229]. Since many innovations in this field focus on the technique of uncertainty quantification, used backbones receive less attention and might often be outdated. Our recommendation is to improve current models with developments in general CNN-based architectures.
VII-C5 Distributing-free modeling
A distribution-free framework known as Conformal Deep Learning produces prediction sets that guarantee to contain the ground-truth with a user-defined probability. With the help of an additional calibration set, a heuristic notion of ambiguity (i.e. miscalibrated softmax outputs) is transformed to rigorous uncertainty and is especially renowned for being model-agnostic, simple and highly flexible [230]. Very recently, conformal prediction has been applied to segmentation problems [231, 232, 233, 234], thereby enjoying the aforementioned benefits of this framework and indicating an increased traction. We recommend further research in this direction to discover novel applications and to benchmark against current architectures.
VII-C6 Volumetric segmentation
In many clinical datasets, the volumetric data is in most instances sliced to patches and processed with conventional 2D-based CNN-models. However, such models are often straightforward extensions of existing 2D models (and almost exclusively VAE-based), rarely addressing novel challenges introduced with additional dimensionalities. For example, 3D extensions of the PU-Net simply remain to use similar techniques to insert latent samples in the decoding networks. Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization. Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners looking to implement models for volumetric segmentations problems.
Also, methodologies need to be developed in order to appropriately compare 2D to 3D models. For example, the GED can not be estimated volumetrically with 2D models. Furthermore, translating already computationally intensive models (e.g. based on DDPMs or PixelCNNs) to three dimensions is another huge challenge due to the memory requirement involved volumetric data, which will even increase the training load and inference time even further. Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underlines the need of more investigation related to three-dimensional uncertainty quantification, which more closely aligns with real-world clinical practice and therefore encourages faster adoption.
VIII Conclusion
Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability in their predictions. With the vast body of literature, encompassing diverse applications and modalities, the need for a comprehensive and systematic overview of the field is addressed. We present clear definitions and notation of methodologies that attempt uncertainty modeling by considering the field from a theoretical perspective and relating this to various pertinent applications. Aleatoric uncertainty can be modeled in pixel or latent space with generative models, or implicitly be expressed with test-time augmentation. Epistemic uncertainty is captured with Variational Inference on the parameter distribution, or approximated with Monte Carlo Dropout or model ensembling. Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks. Which in turn enable us to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and real-world applications.
Our future recommendation pertain aligning the field with advancements in general generative modeling and deep learning architectures. Furthermore, the adoption of deterministic uncertainty quantification methods that do no require multiple forward passes such as Conformal and Evidential Deep Learning is suggested. Especially the latter approach is interesting due to its ability to encapsulate and express both uncertainty types. Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation is a beneficial contribution to the field. Finally, due to the clinical relevancy of uncertainty in semantic segmentation, more attention to models catered to volumetric data is advised. In this manner, the review paper guides researchers on the topic of probabilistic segmentation and suggests future endeavors within the lightning-fast evolving field of Deep Learning-based Computer Vision.

### v3

@@
 Abstract
-Advancements in image segmentation play an integral role within the greater scope of Deep Learning-based computer vision.
-Furthermore, their widespread applicability in critical real-world tasks has given rise to challenges related to the reliability of such algorithms.
-Hence, uncertainty quantification has been extensively studied within this context, enabling expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision making.
+Advancements in image segmentation play an integral role within the broad scope of Deep Learning-based Computer Vision.
+Furthermore, their widespread applicability in critical real-world tasks has resulted in challenges related to the reliability of such algorithms.
+Hence, uncertainty quantification has been extensively studied within this context, enabling the expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision-making.
 Due to the rapid adoption of Convolutional Neural Network (CNN)-based segmentation models in high-stake applications, a substantial body of research has been published on this very topic, causing its swift expansion into a distinct field.
-This work provides a comprehensive overview of probabilistic segmentation by discussing fundamental concepts in uncertainty that govern advancements in the field as well as the application to various tasks.
-We identify that quantifying aleatoric and epistemic uncertainty approximates Bayesian inference w.r.t.
-to either latent variables or model parameters, respectively.
-Moreover, literature on both uncertainties trace back to four key applications; (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) active learning.
-Then, a discussion follows that includes an overview of utilized datasets for each of the applications and comparison of the available methods.
-We also highlight challenges related to architectures, uncertainty-based active learning, standardization and benchmarking, and recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
+This work provides a comprehensive overview of probabilistic segmentation, by discussing fundamental concepts of uncertainty quantification, governing advancements in the field as well as the application to various tasks.
+Moreover, literature on both types of uncertainties trace back to four key applications: (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) Active Learning.
+An extensive discussion follows that includes an overview of utilized datasets for each of the applications and evaluation of the available methods.
+We also highlight challenges related to architectures, uncertainty quantification methods, standardization and benchmarking, and finally end with recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
 Index Terms:
-image segmentation, uncertainty quantification, probability theory
+image segmentation, uncertainty quantification, probabilistic modeling, epistemic, aleatoric
 I Introduction
 Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest [1].
-With the rapid development of Convolution Neural Networks (CNNs), Deep-Learning based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7].
+With the rapid development of Convolution Neural Networks (CNNs), Deep Learning-based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7].
 Nevertheless, such methodologies rely on extensive assumptions and relaxations on the Bayesian learning paradigm, omitting crucial information on the uncertainty associated with the model predictions.
-This ignorance diminishes the reliability and interpretability of such models.
-For example, difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
+This omission diminishes the reliability and interpretability of such models.
+For example, a difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
 Extensive efforts have been made to align modern neural network optimization with Bayesian Machine Learning [8, 9, 10, 11], such as learning parameter densities, rather than point estimates, to include a notion of epistemic uncertainty.
@@
 For example, Hüllermeier et al.
-[12] mentioned that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”.
+[12] stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”.
 This sentiment is also shared by Kiureghian and Ditlevsen [13], noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic”.
 Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic.
-The merits of uncertainty quantification have fortunately been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview.
-Nonetheless, most surveys take perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19].
+Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview.
+Nonetheless, most surveys take the perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19].
 There is a notable absence of a comprehensive overview in this field, relating theoretical foundation to the multitude of applications.
-Furthermore, the abundance of available works can often be overwhelming for both new-coming and seasoned researchers.
-This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are clarified through standardized notation.
+Furthermore, the abundance of available works can be often overwhelming for both emerging and seasoned researchers.
+This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are discussed through standardized notation and clarified by mutual comparison.
 After reading this work, readers will be able to discern various forms of uncertainty with their pertinent applications to segmentation tasks.
@@
 Then, the theoretical framework and notation that governs the remainder of the paper is introduced in Section III.
-The role of these concepts in image segmentation will be treated in Section IV and V, which includes all architectures and approaches with significant impact on the field.
+The role of these concepts in image segmentation will be treated in Sections IV and V, which include all architectures and approaches with significant impact on the field.
 We then consider the applications that use these uncertainty estimates in Section VI.
-Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII and we conclude in Section VIII.
-For a brief overview of the sections, refer to Figure 1.
+Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII, which is concluded in Section VIII.
+Figure 1 illustrates a brief overview of the sections.
 […]
-Figure 1: Overview of the sections.
+Figure 1: Overview of the sections. The leafs of the presented hierarchical tree are related to subsections of the article.
 II Background
+This section highlights a selection of literature crucial to the development of current probabilistic segmentation models.
+While literature on segmentation itself is vast, this decision has resulted in a limited collection of studies.
 As summarized by Minaee et al.
-[20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watersheds [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29].
+[20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watershedding [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29].
 After the application of CNNs [30], the domain of image segmentation underwent rapid developments.
@@
 Initially developed for the medical applications, Ronneberger et al.
-[2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path to preserve high-frequency details in the encoded feature maps.
-To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures.
-In fact, reports in recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
-Semantic Segmentation focuses solely on assigning a class label(s) to each pixel and is particularly suitable for amorphous or uncountable subjects of interest.
+[2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps.
+To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain.
+In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
+Semantic segmentation focuses solely on assigning class labels to each pixel and is particularly suitable for amorphous or uncountable subjects of interest.
 In contrast, instance segmentation not only detects, but also delineates individual objects within the image.
 This form of segmentation is more applicable when identifying and outlining countable instances of objects.
-The third category, Panoptic Segmentation, combines both class and instance level classification [20].
+The third category, panoptic segmentation, combines both class and instance-level classification [20].
 III Probabilistic Image Segmentation
-Let random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, where $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image.
+This section describes the fundamentals of Bayesian inference and its relationship with conventional models for segmentation.
+The provided details introduce standard variables and notations as we proceed throughout the manuscript.
+III-A Bayesian inference
+Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image.
 The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively.
-III-A Bayesian inference
-Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
+Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e.
+posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
 $p(\bm{\theta}|\mathbf{y},\mathbf{x})=\frac{p(\mathbf{y}|\mathbf{x},\bm{\theta})p(\bm{\theta})}{p(\mathbf{y|x})},$ (1)
@@
 After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
-[…]
-As evident in Equation (2), both the variability in the empirical data and the inferred parameters of the model influence the variance of the predictive distribution.
+$p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})=\int\underbrace{p(\mathbf{Y}|\mathbf{x}^{*},\bm{\theta})}_{\text{Data}}\overbrace{p(\bm{\theta}|\mathcal{D})}^{\text{Model}}\mathrm{d}\bm{\theta}.$ (2)
+As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution.
 Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e.
 the variance of the model parameters.
-A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy defined as
-$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\mathbb{E}\,\left[-\log p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})\right]$ (3)
-or variance
-$\mathrm{Var[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]}=\mathbb{E}\,[\,p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})^{2}\,]-\mathbb{E}\,[\,p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})\,]^{2}.$ (4)
+A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy, which can similarly be decomposed as
+$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}}.$ (3)
 III-B Conventional segmentation
-The so-called “deterministic” segmentation networks are trained by Maximum Likelihood Estimation (MLE) as
-$\bm{\theta}_{MLE}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta}),$ (5)
+Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks, which are trained by Maximum Likelihood Estimation (MLE) and is specified as
+$\bm{\theta}_{\text{MLE}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta}),$ (4)
 which simplifies the training procedure by taking a point estimate of the posterior.
-This approximation improves as the training data increases and the variance of the model parameters approaches zero.
+This approximation improves as the training data increases and the model parameter variances approach zero.
 As such, MLE does not include any prior knowledge on the structure of the parameter distribution.
-This is typically done through Bayesian Maximum A Posteriori (MAP) estimation with
-$\bm{\theta}_{MAP}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (6)
-For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the L2-norm (also known as ridge regression or weight decay) or L1-norm of $\bm{\theta}$, respectively [41, 42].
-Additionally, the output of such models can be interpreted as a probability distributiont For instance, let a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$, where we denote the input image and segmentation dimensionality as $\mathbb{R}^{C\times D}$ and $\mathbb{R}^{K\times D}$, respectively, for concise notation.
+This is typically achieved through Bayesian Maximum A Posteriori (MAP) estimation with
+$\bm{\theta}_{\text{MAP}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (5)
+For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively [41, 42].
+Additionally, the output of such models can be interpreted as a probability distribution.
+For instance, consider spatial image dimensions $D$ and a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$.
 Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
 $p(\mathbf{Y}=k\,|\,\mathbf{x}^{*},\bm{\theta})=\frac{e^{\mathbf{a}_{k}}}{\sum_{k}e^{\mathbf{a}_{k}}},$ (6)
-with channel-wise indexing over in the denominator, which is commonly known as the SoftMax activation.
-See Figure 2 for an illustration.
-While not referred to as such in common nomenclature, this is probabilistic modeling in the technical sense, and the approximated distribution can represent and localize uncertain regions.
+with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2).
+This approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature.
+In fact, the approximated distribution can represent and localize uncertain regions.
 However, the implicit pixel-independence assumption
@@
 At the very core, this limitation (See Figure 3) has driven research on probabilistic segmentation, enabling the sampling of spatially coherent segmentation masks.
-This challenge can be addressed either from the aleatoric or epistemic perspective, inferring Bayes theorem w.r.t.
-the hidden latent variables or model parameters, respectively.
-Both approaches have relevant specific use-cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
+This challenge can be addressed either from the aleatoric or epistemic perspective.
+Both approaches have relevant specific use cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
 Figure 2: Aleatoric uncertainty quantification by modeling pixel-level outputs as parameters of a Probability Mass Function.
@@
 (c) Sampling
-Figure 3: Interpretation of the likelihood function in segmentation models.
-Color intensity in (a) reflect the normalized confidence values that can be interpreted as probabilities.
-Maximum likelihood thresholding can be applied to obtain (b).
-However, the coherence of the segmentation suffers when one attempts to sample, as is shown in (c).
+Figure 3: Illustration of the likelihood function in segmentation models.
+Color intensities reflect the normalized confidence values that can be interpreted as probabilities.
+(a) The continuous output results in a horizontal gradient.
+(b) Maximum likelihood thresholding can be applied.
+(c) However, the coherence of the segmentation suffers when sampling.
 IV Aleatoric uncertainty
@@
 $p(\mathbf{Y}|\mathbf{X})=\frac{p(\mathbf{Y},\mathbf{X})}{p(\mathbf{X})}\neq\delta(\mathbf{Y}-F(\mathbf{X})),$ (8)
-with Dirac-delta function $\delta$, and arbitrary function $F:\mathcal{X}\rightarrow\mathcal{Y}$.
-This relationship is characterized by the ambiguity in $\mathbf{X}$, and is inherently probabilistic due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g.
+with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$.
+This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g.
 not all cats have tails).
-Hence, detecting substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition.
-The possible input-dependency drives further categorization in to either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
+Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality.
+The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
 In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$.
-For example, one can model a distribution parameterized by the output of a CNN (Section IV-A).
-Also, generative models have extensively been used, where the data likelihood is learned through latent variables (Section IV-B).
-Finally, image augmentation during test-time inference can also be applied to obtain a notion of aleatoric uncertainty (Section IV-C).
+For example, it is possible to model a distribution parameterized by the output of a CNN (Section IV-A).
+Also, generative models have been used extensively, where the data likelihood is learned through latent variables (Section IV-B).
+Finally, image augmentation during test-time inference can be also applied to obtain a notion of aleatoric uncertainty (Section IV-C).
 IV-A Pixel-level sampling
 A valid approach employs direct pixel-level uncertainty in the annotations.
-As discussed before, uncertainty can be quantified in case the independence assumption holds.
-In this case, it is important to ensure proper calibration, which is discussed in Section IV-A1.
+In case the independence assumption holds, it is important to ensure proper calibration of the model, which is discussed in Section IV-A1.
 In another method, the spatial correlation is explicitly modeled across the pixels of the segmentation mask and results in spatially coherent samples, which will be treated in Section IV-A2.
@@
 The normalized confidence values that result from SoftMax activation can only be interpreted as probabilities after proper validation, which is referred to as model calibration.
-Ideally, the empirical accuracy of a model should approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e.
+Ideally, the empirical accuracy of a model should be approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e.
 $P(Y=k\,|\,c_{k})=c_{k}$.
@@
 Different methods can be used to measure calibration, but often introduce their own biases.
-A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy and confidence bins as
-$ECE=\sum_{b=1}^{B}\frac{n_{b}}{N}|\operatorname{acc}(b)-\operatorname{conf}(b)|,$ (10)
+A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
+$E_{\text{ECE}}=\sum_{b=1}^{B}\frac{n_{b}}{N}|\text{acc}(b)-\text{conf}(b)|,$ (9)
 with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins.
 The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence.
-Furthermore, Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
+Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
 (a) Calibrated
 (b) Miscalibrated
-Figure 4: Reliability diagram indicating if the model confidence accurately reflect the empirical probabilities.
+Figure 4: Visualization of a reliability diagram indicating if the model confidences accurately reflect the empirical probabilities.
 Contemporary SoftMax-activated neural networks often portray a significantly incongruous reflection of the true data uncertainty because of negative log-likelihood optimization and techniques such as batch normalization, weight decay and other forms of regularization [9].
@@
 For example, Temperature Scaling [9] has been applied in a pixel-wise manner for segmentation problems [43].
-Nonetheless, some methods, such as Label Smoothing [44, 45] or using Focal loss [46] can directly be applied on the training data.
-Furthermore, overfitting has often been considered to be the cause of overconfidence [47, 48] and erroneous pixels can therefore be penalized through regularizing low-entropy outputs [49].
+Nonetheless, some methods, such as Label Smoothing [44, 45] or using the Focal Loss [46] can be directly applied on the training data.
+Furthermore, over-fitting has often been considered to be the cause of over-confidence [47, 48] and erroneous pixels can therefore be penalized through regularizing low-entropy outputs [49].
 IV-A2 Spatial correlation
 Figure 5: Depiction of Stochastic Segmentation Networks [50]. Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
-.
 To explicitly model spatial correlation within pixel space of the likelihood distribution, Monteiro et al.
-[50] propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
+[50] propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
 $p(\mathbf{a}|\mathbf{x},\bm{\theta})=\mathcal{N}(\mathbf{a};\bm{\mu}=f^{\bm{\mu}}_{\bm{\theta}}(\mathbf{x}),\bm{\Sigma}=f^{\bm{\Sigma}}_{\bm{\theta}}(\mathbf{x})),$ (10)
-where the covariance matrix has low-rank structure
+where the covariance matrix has a low-rank structure
 $\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}.$ (11)
-Here, $\mathbf{P}$ has dimensionality $((H\times W\times K)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ a diagonal matrix.
+Here, $\mathbf{P}$ has dimensionality $((K\times D)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ a diagonal matrix.
 This results in a more structured and expressive distribution, while retaining reasonable efficiency.
@@
 $p(\mathbf{Y}|\mathbf{X})=\prod_{i}^{K\times D}p(Y_{i}|Y_{1},...,Y_{i-1},\mathbf{X}).$ (12)
-The PixelCNN remains a popular method to model such relationship due to its substantial receptive field [51, 52].
+A popular method to model such relationship due to its substantial receptive field is known as the PixelCNN [51, 52].
 Zhang et al.
-[53] suggest to use this to predict a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$).
-Fusing the two masks is done through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples.
-See Figure 6 for an illustration.
+[53] suggest to use this network to predict a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$).
+The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples.
+Figure 6 illustrates this concept.
 Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
-Figure 6: An illustration of the PixelCNN-based PixelSeg [53].
-‘C’ indicat es the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation.
-Partially transparent elements appear during test-time sampling.
+Figure 6: Illustration of the PixelCNN-based PixelSeg [53].
+Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation.
+Dotted elements appear during test-time sampling.
 IV-B Latent-level sampling
 Directly learning the conditional data distribution is a challenging task.
-Hence, generative models often employ a simpler latent (unobserved) variable $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to then learn the approximate joint density $p(\mathbf{Y}|\mathbf{Z},\mathbf{X})$.
-The marginalized distribution is obtained through decomposition
+Hence, generative models often employ a simpler latent (unobserved) variable $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{z}\in\mathbb{R}^{d}$, to instead learn the approximate joint density $p(\mathbf{Y}|\mathbf{Z},\mathbf{X})$.
+The marginalized distribution is obtained through integral
 $p_{\bm{\theta},\bm{\psi}}(\mathbf{Y}|\mathbf{X})=\int p_{\bm{\theta}}(\mathbf{Y}|\mathbf{z},\mathbf{X})p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\mathrm{d}\mathbf{z},$ (13)
 with parameters $\bm{\theta},\bm{\psi}$.
-Conditioning the latent density on the input images is not a necessity but usually preferred for smooth optimization trajectories [54].
-As such, the spatial correlation is not explicitly modeled but rather induced through mapping the latent variables to segmentation masks.
-Notable architectures relevant to the context of this paper are Generative Adversarial Networks (Section IV-B1), Variational Autoencoders (Section IV-B2) and Denoising Diffusion Probabilistic Models (Section IV-B3).
-IV-B1 Generative Adversarial Networks
-A straightforward approach is to simply learn the decomposition in Equation (14) through sampling from an unconditional prior density
-$p_{\mathbf{Z}}=\mathcal{N}(\bm{\mu}=\mathbf{0},\bm{\Sigma}=\bm{\sigma}\cdot\mathbf{I}),$ (15)
+Conditioning the latent density on the input images is not a necessity, but usually preferred for smooth optimization trajectories [54].
+As such, the spatial correlation is not explicitly modeled, rather induced through mapping the latent variables to segmentation masks.
+Notable generative model architectures relevant to this section are Generative Adversarial Networks (Section IV-B1), Variational Autoencoders (Section IV-B2) and Denoising Diffusion Probabilistic Models (Section IV-B3).
+IV-B1 Generative Adversarial Networks (GANs)
+A straightforward approach is to simply learn the marginalization in Equation (13) through sampling from an unconditional prior density
+$p_{\mathbf{Z}}=\mathcal{N}(\bm{\mu}=\mathbf{0},\bm{\Sigma}=\mathbf{I}),$ (14)
 and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$.
 Goodfellow et al.
-[55] show that this approach can notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$.
-In this way, $G_{\bm{\phi}}$ learns reconstruct realistic looking images, guided by the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient resistance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative.
+[55] show that this approach can be notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$.
+In this way, $G_{\bm{\phi}}$ learns to reconstruct realistically looking images through the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient guidance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative.
 We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
-$J_{G}=-J_{D}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]$ (16)
+$J_{G_{\bm{\phi}}}=-J_{D_{\bm{\psi}}}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]$ (15)
 $-\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[\,\log(1-D_{\bm{\psi}}(G_{\bm{\phi}}(\mathbf{z},\mathbf{x})))\,].$
-Figure 7: The Calibrated Adversarial Refinement network [56] based on the Generative Adversarial Network with additional loss terms.
-While conditional GANs had been used for semantic segmentation before [57], Kassapis et al.
-[56] explicitly contextualized the architecture within aleatoric uncertainty quantification using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7).
-The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\mathbf{c}$, with (cross entropy) reconstruction loss
-$\mathcal{L}_{rec}=-\mathbb{E}_{p_{\mathcal{D}}}[\log p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})].$ (17)
+Figure 7: Diagram of the Calibrated Adversarial Refinement network [56], based on the Generative Adversarial Network with additional loss terms.
+While conditional GANs have been used for semantic segmentation earlier [57], Kassapis et al.
+[56] explicitly contextualized the architecture within aleatoric uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7).
+The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\mathbf{c}$, with (cross-entropy) reconstruction loss
+$\mathcal{L}_{\text{rec}}=-\mathbb{E}_{p_{\mathcal{D}}}[\log p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})].$ (16)
 Then, the conditional refinement network $G_{\bm{\theta}}$ uses $\mathbf{c}$ together with $\mathbf{X}$ and latent samples $\mathbf{Z}_{i}\sim p_{\mathbf{Z}}$ injected at multiple decomposition scales $i$, to predict various segmentation maps.
-Furthermore, the refinement network is subject to adversarial objective
-$\mathcal{L}_{adv}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (18)
-which is argued to elicit superior structural qualities compared to relying on cross-entropy alone. At the same time, the discriminator opposes the optimization with
+Furthermore, the refinement network is subject to the adversarial objective
+$\mathcal{L}_{\text{adv}}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (17)
+which is argued to elicit superior structural qualities compared to relying solely on cross-entropy loss. At the same time, the discriminator opposes the optimization with
 […]
 $\hskip 5.0pt-\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,].$
-Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are then compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities denoted as
-$\mathcal{L}_{cal}=\mathbb{E}_{p_{\mathcal{D}}}\operatorname{KL}[\,p_{\bm{\phi}}(\mathbf{y}|\mathbf{c},\mathbf{x})||p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})\,].$ (20)
+Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities, denoted by
+$\mathcal{L}_{\text{cal}}=\mathbb{E}_{p_{\mathcal{D}}}\operatorname{KL}[\,p_{\bm{\phi}}(\mathbf{y}|\mathbf{c},\mathbf{x})\,||\,p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})\,].$ (19)
 In this way, the generator loss can be defined as
-$\mathcal{L}_{G}=\mathcal{L}_{adv}+\lambda\cdot\mathcal{L}_{cal},$ (21)
+$\mathcal{L}_{G}=\mathcal{L}_{\text{adv}}+\lambda\cdot\mathcal{L}_{\text{cal}},$ (20)
 with hyperparameter $\lambda\geq 0$.
-The purpose of the calibration network is argued to be three-fold.
-Namely, it (1) sets a calibration target for $\mathcal{L}_{cal}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification.
-The refinement network can be seen as modeling the spatial dependence across the pixels, which enables sampling coherent segmentation maps through latent variable $\mathbf{Z}$.
-IV-B2 Variational Autoencoders
-Techniques such as GANs rely on implicit distributions and are void of any notion of data likelihoods.
-An alternative approach estimates the Bayesian posterior w.r.t.
-the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$ obtained my maximizing conditional Evidence Lower Bound (ELBO)
-$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
-$=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
-$\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
-$\hskip 50.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,].$ (22)
-This is also known as Variational Inference (VI).
-The first term in Equation (22) represents the reconstruction cost of the decoder subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$.
-The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density.
-As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Normal densities and amortized through neural networks parameterized by $\phi$, $\theta$ and $\psi$.
-The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
-[…]
-Implementing the conditional ELBO in Equation (22) can be achieved through a VAE-like architecture [10].
-A few additional design choices specific for uncertainty quantification for segmentation result in the The Probabilistic U-Net (PU-Net) [58].
-Firstly, the latent variable $\mathbf{z}$ is only introduced at the final layers of a U-Net conditioned on $\mathbf{X}$, where the vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolutions.
-When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation is performed in the U-Net, while the information $\mathbf{Z}$ provides is almost exclusively regarding the segmentation variability.
-Therefore, relatively smaller values of $d$ are feasible than what is commonly used in image generation tasks.
-Figure 8: The Probabilistic U-Net [58] based on a conditional Variational Autoencoder [10]. Latent samples are inserted at the final stages of a U-Net through a tiling operation.
-Similar to related research on the VAE [59, 60, 61, 62, 63], much work has been dedicated to improving the PU-Net.
-For instance, investigation on improving VI with novel parameterization of the amortized densities also provide interesting insights into model behaviour [64, 65, 66, 67, 68, 69, 70].
-Furthermore, extending the architecture to multiple decomposition hierarchies [71, 72], three-dimensional modalities [73, 74, 75, 76, 77] and conditioning on the annotator [78] also resulted in substantial performance gains.
-Density parameterization
+The purpose of the calibration network is argued to be threefold.
+Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification.
+The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps through latent variable $\mathbf{Z}$.
+IV-B2 Variational Autoencoders (VAEs)
 Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness [58].
-This phenomena has also been successful for cVAE-like models such as the PU-Net [64, 66].
-NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $\mathbf{f}=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex distributions $\log p(\mathbf{x}|\mathbf{z})$.
+This phenomenon has also been successfully applied to VAE-like models such as the PU-Net [59, 60].
+NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $\mathbf{f}=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex distributions $\log p(\mathbf{x}|\mathbf{z})$, where symbol $\circ$ indicates functional composition.
 These are often denoted as $\log p(\mathbf{x})$ for simplicity and can be determined through
@@
 Bhat et al.
-[67, 68] compare this approach with other parameterizations of the latent space including a mixture of Gaussians and low-rank approximation of the full covariance matrix.
+[61, 62] compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix.
 Valiuddin et al.
 [63] show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse [64, 65].
-Their proposition considers the alternative formulation of the ELBO
+Their proposition considers the alternative formulation of the ELBO, specified as
 $\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (23)
 $\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]$
-$\hskip 70.0pt-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}),$
-which the novel objective maximizes the contribution of the (expected) mutual information between latent and output variables. This can be rewritten to the objective
+$\hskip 70.0pt-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
+The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables. This can be rewritten into the objective
 $\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (24)
@@
 $\hskip 80.0pt+\beta\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$
-with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence [81] and $\alpha,\beta$ being hyperparameters, which results in a more uniform latent space leading to increased model performance.
-Also, modeling the ELBO of the joint density has also been explored [69].
+with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence [66] and $\alpha,\beta$ denoting hyperparameters.
+This adaptation results in a more uniform latent space leading to increased model performance.
+Also, modeling the ELBO of the joint density has been explored [67].
 This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data.
-Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this counters the model collapse phenomena in latent space [70].
+Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon [68].
+Density parameterization
+Techniques such as GANs rely on implicit distributions and are void of any notion of data likelihoods.
+An alternative approach estimates the Bayesian posterior w.r.t.
+the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, obtained by maximizing the conditional Evidence Lower Bound (ELBO)
+$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
+$=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
+$\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
+$\hskip 50.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,].$ (22)
+This approach of estimating $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$ is also known as Variational Inference (VI).
+The first term in Equation (25) represents the reconstruction cost of the decoder subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$.
+The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density.
+As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\phi$, $\theta$ and $\psi$.
+The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
+$p(\mathbf{Y}|\mathbf{x}^{*})=\int p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{x}^{*})q_{\theta}(\mathbf{z}|\mathcal{D})\mathrm{d}\mathbf{z}.$ (26)
+The implementation of the conditional ELBO in Equation (25) can be achieved through a VAE-like architecture [10].
+A few additional design choices specific for uncertainty quantification in segmentation lead to the Probabilistic U-Net (PU-Net) [69].
+Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on $\mathbf{X}$, where the vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolution for classification.
+When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the provided information by $\mathbf{Z}$ is almost exclusively related to the segmentation variability.
+Therefore, relatively smaller values of $d$ are feasible than what is commonly used in conventional image generation tasks.
+Figure 8: Depiction of the Probabilistic U-Net [69] based on a conditional Variational Autoencoder [10]. The latent samples are inserted at the final stages of a U-Net through a tiling operation.
+Similar to related research on the VAE [65, 70, 71, 72, 58], much work has been dedicated to improving the PU-Net.
+For instance, investigation on improving VI with novel parameterization of the amortized densities also provides insights into model behaviour [59, 73, 60, 61, 62, 67, 68] (Section IV-B2).
+Furthermore, extending the architecture to multiple decomposition hierarchies [74, 75] (Section IV-B2), three-dimensional modalities [76, 77, 78, 79, 80] (Section IV-B2), and conditioning on the annotator [81] (Section IV-B2), also result in substantial performance gains.
 Multi-scale approach
 Learning latent features over several hierarchical scales can provide expressive densities and interpretable features across various abstraction levels [82, 83, 84, 85, 86].
-Such models commonly fall under hierarchical VAE umbrella term.
+Such models are commonly categorized under hierarchical VAE (HVAE) modeling.
 Often, an additional Markov assumption of length $T$ is placed on the posterior as
@@
 Simply stated, the architecture learns a latent representation at multiple decomposition levels of the U-Net.
-Furthermore, residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL-divergence rapidly approaching zero.
-For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective was employed, which extends on Equation (29) as
-$\mathcal{L}_{GECO}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (30)
+Furthermore, residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL divergence rapidly approaching zero.
+For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective is employed, which extends Equation (28) to
+$\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (29)
 $-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})\,||\,q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]$
 $\hskip 20.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})\,||\,q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
-Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$ set beforehand to an appropriate value.
+Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value.
 Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick [87, 88].
-PHiSeg [72] takes a similar approach to the HPU-Net.
-However, instead of placing the residual connections in the convolutional layers, PHiSeg uses these between the latent vectors across decomposition.
-Furthermore, auxiliary outputs, or deep supervision, at each decomposition scale is used to enforce disentanglement across latent variables.
+A similar architecture known as PHiSeg [75] takes a similar approach to the HPU-Net.
+However, instead of placing the residual connections in the convolutional layers, PHiSeg uses these between the latent vectors across the decompositions.
+Furthermore, deep supervision at each decomposition scale is used to enforce disentanglement across latent variables.
 Figure 9: Hierarchical Probabilistic U-Net [74] based on a hierarchical Variational Autoencoder [10, 82, 84].
@@
 Extension to 3D
-Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to potential loss of critical spatial information and subtle nuances often necessary for accurate delineation.
+Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to a potential loss of critical spatial information and subtle nuances often necessary for accurate delineation.
 This limitation has spurred extensive research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures.
@@
 [78] additionally augment a Normalizing Flow to the posterior density, as described in Section IV-B2.
-Further enhancements to the architecture include the implementation of the 3D hierarchical PU-Net [74] or an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications [73].
+Further enhancements to the architecture include the implementation of the 3D HPU-Net [77], an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications [76].
 Conditioning on annotator
@@
 This can be achieved by conditioning the learned densities on the annotator itself [81, 89].
-For example, features of a U-Net can be combined with samples from annotator specific Gaussian distributed posterior distributions [89].
+For example, features of a U-Net can be combined with samples from annotator-specific Gaussian-distributed posteriors [89].
 Considering the approach from Gao et al.
-[78], generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ with for the $k$-th annotator.
-Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through image-conditional decoder as $\mathbf{y}=F(\mathbf{x},\mathbf{z}_{k})$.
+[81], generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator.
+Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $\mathbf{y}=F(\mathbf{x},\mathbf{z}_{k})$.
 The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble.
-Additionally, consistency between the model and ground-truth distribution is additionally enforced through an optimal transport loss between the set of predictions and labels.
-IV-B3 Denoising Diffusion Probabilistic Models
-Recent developments in generative modeling have resulted a family of models known as Denoising Diffusion Probabilistic Models (DDPMs) [90, 91, 92].
-Such models are especially renowned for their expressive power by able to encapsulate large and diverse datasets.
-While several perspectives can be used to introduce the the DDPMs, we build upon the earlier discussed HVAE (Section IV-B2) to maintain cohesiveness with the overall manuscript.
-In specific, let us introduce three additional modifications to the HVAE.
-Firstly, the latent dimensionality is set equal to the data dimensions, i.e.
+Additionally, consistency between the model and ground-truth distribution is enforced through an optimal transport loss between the set of predictions and labels.
+IV-B3 Denoising Diffusion Probabilistic Models (DDPMs)
+Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models [90, 91, 92].
+Such models are especially renowned for their expressive power by being able to encapsulate large and diverse datasets.
+While several perspectives can be used to introduce the DDPMs, we build upon the earlier discussed HVAE (Section IV-B2) to maintain cohesiveness with the overall manuscript.
+Specifically, we introduce three additional modifications to the HVAE [93].
+Firstly, the latent dimensionality is set equal to the spatial data dimensions, i.e.
 $d\,$=$\,D$.
@@
 Secondly, the encoding process (or forward process) is predefined as a linear Gaussian model such that
-$q(\mathbf{y}_{T}|\mathbf{y}_{0})=p(\mathbf{y}_{0})\prod^{T}_{t=1}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$ (31)
-and
+$q(\mathbf{y}_{T}|\mathbf{y}_{0})=p(\mathbf{y}_{0})\prod^{T}_{t=1}q(\mathbf{y}_{t}|\mathbf{y}_{t-1}),$ (30)
+where
 $q(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\bm{\mu}=\sqrt{\alpha_{t}}\mathbf{y}_{t-1},\bm{\Sigma}=(1-\alpha_{t})\cdot\mathbf{I}),$ (31)
@@
 $p(\mathbf{y}$ $|\mathbf{x})\geq\mathbb{E}_{q(\mathbf{y}_{1}|\mathbf{y}_{0})}[\,\log p_{\phi}(\mathbf{y}_{0}|\mathbf{y}_{1},\mathbf{x})\,]$ (32)
-$+\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{y}_{t}|\mathbf{Y}_{0})}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})||p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]$
+$+\sum_{t=2}^{T}\mathbb{E}_{q\left(\mathbf{y}_{t}|\mathbf{Y}_{0}\right)}\big{[}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})\,||\,p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]\,\big{]}$
 $\hskip 40.0pt+\underbrace{\mathbb{E}_{q(\mathbf{y}_{T}|\mathbf{y}_{0})}\left[\log\frac{p(\mathbf{y}_{T})}{q(\mathbf{y}_{T}|\mathbf{y}_{0})}\right]}_{\approx 0}.$
-As denoted, the regularization term is assumed to be zero, since we assume that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed.
-With the reparameterization trick, the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$.
-As such, the KL-divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$ depending on the parameterization, and is in almost all instances approximated with a U-Net [2].
-It has also be proposed to model Bernoulli noise instead of Gaussian [93, 94, 95, 96].
+As denoted, the regularization term is assumed to be zero, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed.
+With the reparameterization trick [10], the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$.
+As such, the KL divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$, depending on the parameterization.
+This is in almost all instances approximated with a U-Net [2].
+Also, it has been proposed to model Bernoulli noise instead of Gaussian [94, 95, 96, 97].
 However, most methodologies vary in the conditioning of the reverse process on the input image [98, 99, 100, 101].
@@
 [99] insert encoded image features to the U-Net bottleneck.
-Additionally, information on predictions $\mathbf{Y}_{t}$ at a time step $t$ is also provided in the intermediate layers of the conditioning encoder.
-This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attentive map and the inverse FFT.
-The procedure of applying attention on the spectral domain of the U-Net encoding has also been done with transformers in follow-up work [99].
-Segdiff [100] also encode both current time step and input image, but combine the extracted features by simple summation before applying the U-Net.
-Figure 10: An illustration of the Diffusion Probabilistic Models [90]. The model learns to remove noise that has gradually been added to the input image.
+Additionally, information on predictions $\mathbf{Y}_{t}$ at a time step $t$ is provided in the intermediate layers of the conditioning encoder.
+This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attention map and the inverse FFT.
+The procedure of applying attention on the spectral domain of the U-Net encoding has also been implemented with transformers in follow-up work [100].
+Segdiff [101] also encode both current time step and input image, but combine the extracted features by simple summation prior to inferring it through the U-Net.
+Figure 10: Illustration of the Diffusion Probabilistic Models [90]. The model learns to remove noise that has been gradually added to the input image.
 IV-C Test-time augmentation
 An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest.
-For example, systematic noise, translation or rotation result in many instances realistic variations that approximately retain image semantics.
-Hence, data augmentation [31] has been used at test time (hence the name test-time augmentation, or TTA) for image classification to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function [101].
+For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics.
+Hence, data augmentation [31] has been used at test time (explaining the name test-time augmentation, or TTA) for image classification, to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function [102].
 This technique has been applied to image segmentation as well [103, 104, 105, 106].
-By randomly augmenting input images with invertible transformation $T$ as $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$.
-Repeatedly performing this procedure results in a set segmentations masks, which can serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
+By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$.
+Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
 V Epistemic uncertainty
-The crucial difference between epistemic and aleatoric uncertainty is that the former is related to model ignorance, while the latter reflects statistical ambiguity inherent in the data.
+The crucial difference between epistemic and aleatoric uncertainty is that the former relates to model ignorance, while the latter reflects statistical ambiguity inherent in the data.
 Epistemic uncertainty can be further categorized into two distinct types [12].
 The first type pertains uncertainty related to the capacity of the model.
-For example, under-parameterized models or approximate model posteriors can become too stringent to appropriately resemble the true posterior.
-The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, this is also known as model uncertainty.
+For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior.
+The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty.
 Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible.
@@
 Unfortunately, evaluation of the true Bayesian posterior (formulated in Equation (1)) is inhibited by the intractability of the data-likelihood in the denominator.
-Hence, extensive efforts have been taken to obtain viable approximations such as using Mean-Field Variational Inference [8], Markov Chain Monte Carlo (MCMC) [106], Monte-Carlo Dropout [107, 108], Model Ensembling [109], Laplace approximations [110], Stochastic Gradient MCMCs [111, 112, 113], assumed density filtering [114] and expectation propagation [115, 116].
+Hence, extensive efforts have been taken to obtain viable approximations, such as using Mean-Field Variational Inference [8], Markov Chain Monte Carlo (MCMC) [107], Monte-Carlo Dropout [108, 109], Model Ensembling [110], Laplace approximations [111], Stochastic Gradient MCMCs [112, 113, 114], assumed density filtering [115] and expectation propagation [116, 117].
 We refer to any neural network that approximates the Bayesian posterior of the model parameters as a Bayesian Neural Network (BNN).
-The following sections treat methodologies that have been applied within the context of this paper, which are usually straightforward extensions of BNN networks used for conventional regression and classification tasks.
-Additionally, an illustration of these techniques are presented in Figure 11.
+The following sections treat applied methodologies to segmentation, which are usually straightforward extensions of conventional regression and classification BNN networks.
+These techniques are illustrated in Figure 11.
 V-A Variational Inference
-Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
-Then, we can achieve Variational Inference (VI) w.r.t.
-to the parameters by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior can be written as
+Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
+Then, we can employ Variational Inference (VI) w.r.t.
+to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
 $\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathbf{y},\mathbf{x})\,\right]$
@@
 where the parameter-independent terms are constant and therefore excluded from notation.
-A popular choice for the approximated variational posterior is a the Gaussian distribution, i.e.
+A popular choice for the approximated variational posterior is the Gaussian distribution, i.e.
 a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities.
-However, the priors can also be learned through Empirical Bayes [117].
+However, the priors can be also learned through Empirical Bayes [118].
 Furthermore, backpropagation is possible with the reparameterization trick [10] and within this context, the procedure is referred to as Bayes by Backprop (BBB) [8].
 During testing, a sample-based approach is utilized to approximate the posterior.
-Since using several parameter permutations effectively enriches the hypothesis space due to the model combining effect and can already improve performance [118, 119, 120].
+Using several parameter permutations effectively enriches the hypothesis space due to the model combining effect, and can improve performance [119, 120, 121].
 V-B Monte Carlo Dropout
-Dropout, a common technique used to regularize neural networks [121], can mimic sampling from an implicit parameter distribution, $q(\tilde{\bm{\theta}}|\bm{\theta},p)$, defined as
-$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p)$ (35a)
+Dropout is a common technique used to regularize neural networks [122] and can mimic sampling from an implicit parameter distribution $q(\tilde{\bm{\theta}}|\bm{\theta},p)$.
+This implicit distribution is defined as
+$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p),$ (34a)
 $\tilde{\bm{\theta}}$ $=\bm{\theta}\odot\mathbf{n},$ (34b)
-with probability $p$ and $\mathbf{n}$ operating element-wise on the parameters.
-Using Dropout can also be interpreted as a first-order equivalent $L_{2}$-regularization with additional transforming the input by the inverse diagonal Fisher information matrix [121].
-With Monte-Carlo Dropout (MC Dropout), the random node switching is continued during testing, effectively sampling new sets of parameters.
+with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters.
+Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix [122].
+Using Monte-Carlo Dropout (MC Dropout), the random node switching is continued during test time, effectively sampling new sets of parameters.
 While seemingly arbitrary, it has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process [108].
-In this manner, the authors showcase such method is able to provide multi-modal estimates of the model uncertainty.
+In this manner, such method is able to provide multi-modal estimates of the model uncertainty.
 As noted by Gal et al.
-[122], the model output variance is balanced with the weight magnitudes rather than the dropout rate $p$, which is usually optimized through grid search or simply fixed to $0.5$.
+[123], the model output variance is balanced with the weight magnitudes rather than the dropout rate $p$, which is usually optimized through grid search or simply set to $0.5$.
 Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, such that uncertainty estimates are governed by $p$.
@@
 V-C Model Ensembling
-As mentioned earlier, Monte Carlo dropout effectively optimizes over a set of sparse neural networks.
-This ensemble can also be designed in an explicit manner.
-Let us define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{N}$, with $N$ representing the number of models in the ensemble.
+As mentioned earlier, MC dropout effectively optimizes over a set of sparse neural networks.
+Also, this ensemble can be designed in an explicit manner.
+To this end, we define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{N}$, with $N$ representing the number of models in the ensemble.
 Then, it is relatively simple to obtain $\bm{\Theta}=\{\bm{\theta}_{n}\}_{n=1}^{N}$, which can be interpreted as samples from an approximate posterior.
@@
 Often, the $N$ obtained parameters are from $N$ separate training sessions.
-However, it has also been shown effective to ensemble from single training session by saving the parameters at multiple stages or training with different weight initializations [124, 125, 126].
+However, it has also been proven effective to ensemble from a single training session, by saving the parameters at multiple stages or training with different weight initializations [125, 126, 127].
 (a) Variational Inference
@@
 (c) Ensembling
-Figure 11: Varying techniques to sample parameters for the convolutional kernels.
-An approximation of the parameter density can be made (a), taking samples can be mimicked (b), or an ensemble of $N$ different configurations can explicitly be modeled (c).
+Figure 11: Three types of techniques are visualized to sample parameters of the convolutional kernels, (a) an approximation of the parameter density can be made, (b) taking samples can be mimicked with dropout during test time, or, (c) an ensemble of $N$ different configurations can be explicitly modeled.
 A closely related concept to ensembling is known as Mixture of Experts (MoE), where each model in the ensemble (an ‘expert’) is trained on specific subsets of the data [128].
 In such settings, a gating mechanism is usually applied after combining the expert hypotheses.
-While uncommon, incorporating all predictions can also be regarded as ensembling technique.
+While uncommon, incorporating all predictions can also be regarded as an ensembling technique.
 VI Applications
@@
 These include estimating the segmentation mask distribution subject to observer variability (Section VI-A), model introspection (i.e.
-ability to self-assess, Section VI-B), improved generalization (Section VI-C) and reduced labeling costs using Active Learning (Section VI-D).
+ability to self-assess, Section VI-B), improved generalization (Section VI-C), and reduced labeling costs using Active Learning (Section VI-D).
 VI-A Observer variability
@@
 Additionally, such annotations can also vary due to differences in expertise and experience of annotators.
-This phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator is referred to as the intra-observer variability (see Figure 12).
-To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be treated as models themselves.
+The phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator are referred to as the intra-observer variability (see Figure 12).
+To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be regarded as models themselves.
 For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with $k=1,2,...,K$.
-For a simple segmentation task, it can be expected that $\operatorname{Var}[p(\bm{\phi}_{k})]\rightarrow 0$.
-In other words, each annotator is consistent in their delineation and the intra-observer variability is low.
+For a simple segmentation task, it can be expected that $\text{Var}[\,p(\bm{\phi}_{k})\,]\rightarrow 0$.
+In other words, each annotator is consistent in his delineation and the intra-observer variability is low.
 For cases with consensus across experts, i.e.
-negligible inter-observer variability, the marginal converges to $\operatorname{Var}[p(\bm{\phi})]\rightarrow 0$.
+yielding negligible inter-observer variability, the marginal approaches to $\text{Var}[\,p(\bm{\phi})\,]\rightarrow 0$.
 Asserting these two assumptions, it is valid to simply consider a point estimate of the posterior.
-Yet, this rarely the case in many real-life application and, as such, explicitly modeling the involved distributions becomes imperative.
-Figure 12: A visualization of intra-observer variability in parameter space (left) and inter-observer variability in data space (right).
-For evaluation, a common metric is to minimize the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick.
-This know as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [128], denoted as
-$\operatorname{GED}^{2}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
+Yet, this is rarely the case in many real-life applications and, as such, explicitly modeling the involved distributions becomes necessary.
+Figure 12: Visualization of the intra-observer variability in parameter space (left) and the inter-observer variability in data space (right).
+For evaluation, a commonly used metric minimizes the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick [129].
+This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [130], denoted as
+$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
 $\hskip 10.0pt+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]$
 $\hskip 20.0pt-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (37)
-with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and some kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the 1--IoU or 1--Dice score.
+with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and a distance represented by kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the (1--IoU) or (1--Dice) score (squared by the GED).
 An alternative metric known as Hungarian Matching (HM) compares the predictions against the ground-truth labels through a cost matrix [74].
-Subsequently, the unique optimal coupling between the two sets that minimizes the average cost is determined through combinatorial optimization algorithm.
-This can also be formally denoted as finding the permutation matrix $\mathbf{P}$ subject to the objective
-$\operatorname{HM}(Y,\hat{Y})=\min_{\mathbf{P}}\frac{1}{N^{2}}\operatorname{Tr}(\mathbf{P}\mathbf{M}),$ (37)
-where $M_{i,j}=k(y_{i},\hat{y}_{j})$ and $N^{2}$ represents the number of elements of the matrix.
+Subsequently, the unique optimal coupling between the two sets that minimize the average cost is determined through a combinatorial optimization algorithm.
+This can be also formally denoted as finding the permutation matrix $\mathbf{P}$, subject to the objective
+$\text{HM}(\mathcal{Y},\mathcal{\hat{Y}})=\frac{1}{N^{2}}\min_{\mathbf{P}}\text{Tr}(\mathbf{P}\mathbf{M}),$ (36)
+where the elements of $\mathbf{M}$ are $M_{i,j}=k(\mathcal{Y}_{i},\mathcal{\hat{Y}}_{j})$, and $N^{2}$ represents the number of elements within the matrix.
 For accurate comparison and evaluation of the problem, benchmarking is often constrained to publicly available multi-annotated data.
@@
 Multiple versions have been used, either with 15,096 (LIDCv1) or 12,816 (LIDCv2) patches.
-LIDCv1 is employed with a 60:20:20 split while for LIDCv2 this is 70:15:15 for train:validation:test.
-LIDCv2 has also been used with threefold cross validation (LIDCv2-cv), with either a 90:10 or 80:20 train:test split.
-Other less commonly datasets include the CityScapes [130] dataset, which contains street view images of German cities from the perspective of a driving car, have been used to artificially create class-level label ambiguity [71, 58, 78].
-Some classes are switched to an arbitrary auxiliary class with some probability $p$.
-Since the underlying probabilities are known, the empirical fraction of the classes from the model predictions can directly be compared to the ground-truth values of $p$.
+LIDCv1 is employed with a 60:20:20 split, while for LIDCv2 this is 70:15:15 for train:validation:test.
+LIDCv2 has also been used with threefold cross-validation (LIDCv2-cv), with either a 90:10 or 80:20 train:test split [63, 59, 73].
+Other less commonly employed datasets include the CityScapes [132] dataset, which contains street-view images of German cities from the perspective of a driving car, have been used to artificially create class-level label ambiguity [74, 69, 81].
+Classes are switched to an arbitrary auxiliary class with some probability $p$.
+Since the underlying probabilities are known, the empirical fraction of the classes from the model predictions can be directly compared to the ground-truth values of $p$.
 Furthermore, the QUBIQ Challenge [133] contains MRI and CT data from varying organs.
-Also, the retinal fundus images for glaucoma analysis (RIGA) [132] dataset contains delineations of optic cup and disc boundaries by six experienced ophthalmologists.
+Also, the retinal fundus images for glaucoma analysis (RIGA) [134] dataset contains delineations of optic cup and disc boundaries by 6 experienced ophthalmologists.
 TABLE I: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table adapted from [95].
 LIDCv1 LIDCv2
-Method Year GED16 HM-IoU16 GED16 HM-IoU16
-PU-Net [58] 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
-HPU-Net [71] 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
-PhiSeg [72] 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
-SSN [50] 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
-CAR [56] 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
-JProb. U-Net [133] 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
-PixelSeg [53] 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
-MoSE [78] 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
-AB [134] 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
-CIMD [95] 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
-CCDM [94] 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
-Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotations pairs, provided that annotators have an intrinsically associated expertise [135].
-Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data can become quite burdensome.
-Hence, it can be much simpler to directly model the empirical stochasticity in the annotations and has been extensively explored in modalities such as Lung nodule detection in 2D [58, 72, 71, 136, 137, 70, 66, 68, 65, 50, 64], as well as 3D [76, 138, 139], Brain Tumor [140, 50, 97, 93, 95], White Matter Hyperintensities [141], Pulmonary Tumour Growth [142], Prostate [143, 73] and vascular [144], street scene [145, 71, 94], aerial imaging [100], optic cup [98, 146], abdominal multi-organ [99] and nuclei microscopy [96] segmentation.
+Method GED16 HM-IoU16 GED16 HM-IoU16
+PU-Net [69], 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
+HPU-Net [74], 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
+PhiSeg [75], 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
+SSN [50], 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
+CAR [56], 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
+JProb. U-Net [135], 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
+PixelSeg [53], 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
+MoSE [81], 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
+AB [136], 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
+CIMD [96], 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
+CCDM [95], 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
+Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotation pairs, provided that annotators have an intrinsically associated expertise [137].
+Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome.
+Hence, it can be much simpler to directly model the empirical stochasticity in the annotations, which has been extensively explored in modalities such as Lung nodule detection in 2D [69, 75, 74, 138, 139, 68, 60, 62, 73, 50, 59], as well as 3D [79, 140, 141], Brain Tumor [142, 50, 98, 94, 96], White Matter Hyperintensities [143], Pulmonary Tumour Growth [144], Prostate [145, 76] and vascular [146], street scene [147, 74, 95], aerial imaging [101], optic cup [99, 148], abdominal multi-organ [100] and nuclei microscopy [97] segmentation.
 The overwhelming majority of research for this particular application has been executed with variants of conditional VAEs [69, 60, 59, 73, 68, 62, 74, 75, 139, 144, 149, 142, 150, 141, 146, 79, 76, 145, 140].
-More recently, the growing popularity of DDPMs is also apparent in the field [94, 100, 97, 93, 95].
-Also, GAN-based approaches have been also employed [56].
+More recently, the growing popularity of DDPMs is apparent in the field [95, 101, 98, 94, 96].
+Also, GAN-based approaches have been employed [56].
 VI-B Model introspection
@@
 [151] through two conditional likelihoods.
-Firstly, the accuracy given a certain prediction, $p(\mathrm{A}|\mathrm{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\mathrm{U}|\mathrm{I})$.
-Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{ac}$, $u_{au}$, $u_{ic}$, $u_{iu}$, respectively.
-Consequently, the authors combine $p(\mathrm{A}|\mathrm{C})=\frac{n_{ac}}{n_{ac}+n_{au}}$ and $p(\mathrm{U}|\mathrm{I})=\frac{n_{ic}}{n_{ic}+n_{iu}}$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
-$\mathrm{PAvPU}=\frac{n_{ac}+n_{au}}{n_{ac}+n_{ic}+n_{ic}+n_{iu}}.$ (38)
-This downstream task has been applied to (video) street scene [150, 151], remote sensing [152, 153], instance segmentation of various objects [154], point-cloud indoor scenes [155] brain [156, 157, 158, 17, 159, 160, 161], Multiple Sclerosis [162], cardiac [163, 164], heart ventricle [161], prostate [161], carotid artery [165], lumbosacral [19] MRI, and Optical Coherence Tomography [166, 167], skin imaging [168, 169], lung [169] and liver CT [170] and MRI  [171], and ultrasound [124, 172].
+Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$.
+Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{\text{ac}}$, $u_{\text{au}}$, $u_{\text{ic}}$, $u_{\text{iu}}$, respectively.
+Consequently, the authors combine $p(\text{A}|\text{C})=n_{\text{ac}}/(n_{\text{ac}}+n_{\text{au}})$ and $p(\text{U}|\text{I})=n_{\text{ic}}/(n_{\text{ic}}+n_{\text{iu}})$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
+$\text{PAvPU}=\frac{n_{\text{ac}}+n_{\text{au}}}{n_{\text{ac}}+n_{\text{au}}+n_{\text{ic}}+n_{\text{iu}}}.$ (37)
+This application has been applied in combination with MC dropout to model (video) street scenes [152, 153], remote sensing [154, 155], instance segmentation of various objects [156], point-cloud indoor scenes [157], brain [158, 159, 160, 17, 161, 162, 163], Multiple Sclerosis [164], cardiac [165, 166], heart ventricle [163], prostate [163], carotid artery [167], lumbosacral [19] MRI, and Optical Coherence Tomography [168, 169], skin imaging [170, 171], lung [171] and liver CT [172] and MRI  [173], and ultrasound [125, 174].
 Concrete dropout has been applied to instance segmentation of C.
@@
 To a lesser degree, ensembling [176, 125], Variational Inference both in 2D [177, 178] and 3D [179], M-heads (auxiliary networks) [14, 180, 181], and test-time augmentation [125] have also been used to quantify the uncertainty for quality assessment.
-It can be noted that uncertainty is usually only obtained on a pixel basis, while crucial information can be present in structural statistics.
+It can be noted that uncertainty is usually only obtained on pixel basis, while crucial information can be present in structural statistics.
 Hence, the Coefficient of Variation (CV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples.
 Also, Roy et al.
-[160] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to then determine the pair-wise average overlap between all respective samples as
-$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\left\{\mathrm{Dice}(\mathbf{Y}_{i}=t(\mathbf{y}),\mathbf{Y}_{j}=t(\mathbf{y}))\right\}_{i\neq j}\right].$ (39)
+[162] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to determine the pair-wise average overlap between all respective samples as
+$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\,\text{Dice}(t(\mathbf{y}_{i}),t(\mathbf{y}_{j}))\,\right]\quad\forall i\neq j.$ (38)
 Finally, including information on localized uncertainty to the training objective has shown to improve generalization capabilities [182, 183, 184, 185].
@@
 VI-C Model generalization
-As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance due to the model combining effect.
+As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance, due to the model combining effect.
 For instance, dropout layers at the deeper decomposition levels of the SegNet [4] improves model performance [152].
 Literature also reports improved performance with Concrete Dropout for semantic (instance) segmentation [151, 175].
-This has been observed in multiple domains including out-/indoor scene understanding [150, 149, 184, 185, 179], brain tumor MRI [157, 186, 159, 175], Optical Coherence Tomography [166], low dose Computed Tomography of lung nodules [187], colectoral polyps [188], cardiac MRI [18] and C.elegans roundworm microscopy images [173].
+This has been observed in multiple domains, including in-/outdoor scene understanding [152, 151, 186, 187, 181], brain tumor MRI [159, 188, 161, 177], Optical Coherence Tomography [168], low-dose Computed Tomography of lung nodules [189], colectoral polyps [190], cardiac MRI [18] and C.
+elegans roundworm microscopy images [175].
 Ng et al.
-[18] benchmark multiple techniques for cardiac MRI segmentation and find that ensembling results in best performance improvement, while Bayes By Backprop [8] is more robust to noise distortions.
+[18] have benchmarked multiple techniques for cardiac MRI segmentation and have found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions.
 Furthermore, the improved generalization from ensembling has also shown to produce more calibrated outputs [163].
@@
 Pavlitskaya et al.
-[179] show merits of such approach has been observed in urban outdoor scene segmentation.
-For the optic cup segmentation, Ji et al.
-[135] also condition on a normalized expertness vector, where each each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net.
+[181] show merits of such approach for urban outdoor scene segmentation.
+For optic cup segmentation, Ji et al.
+[137] also condition on a normalized expertness vector, where each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net.
 Gao et al.
-[78] introduced Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach.
+[81] have introduced a Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach.
 Nevertheless, the methodology addresses aleatoric uncertainty quantification and has hence been discussed in Section IV-B2.
@@
 $H[\,\bm{\theta}|\mathcal{D}\,]-\mathbb{E}_{p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})}[\,H[\,\bm{\theta}|\mathbf{y},\mathbf{x}^{*},\mathcal{D}\,]\,],$ (39)
-can provide a notion to describe information gain, and thus the uncertainty, from specific datapoints.
+can provide a notion to describe information gain, and thus the decrease of uncertainty for specific datapoints.
 Notably, minimizing the expected increase in posterior entropy is equivalent to maximizing the mutual information between the data and model parameters.
@@
 $I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D})$ $=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
-$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]].$ (41a)
-Active Learning through uncertainty quantification has been been beneficial for medical [192, 193, 194, 195, 196, 197, 198, 199] multi-view [200], remote sensing [201], street-views [202] and 3D point-cloud data [203].
-These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [183] and novel temporal-ensembling methods [204] have also been used to reduce required human labeling.
+$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,].$ (41)
+Active Learning through uncertainty quantification has been been beneficial for medical [194, 195, 196, 197, 198, 199, 200, 201], multi-view [202], remote sensing [203], street-views [204] and 3D point-cloud data [205].
+These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [185] and novel temporal-ensembling methods [206] have also been used to reduce the required human labeling.
 Some methods extend beyond pixel-level uncertainty, incorporating boundary information in the uncertainty quantification process [204, 200].
@@
 Similarly, Ma et al.
-[198] combine target and boundary-based uncertainty sampling, ensuring diversity, effective and balanced utilization of the available information.
+[200] combine target and boundary-based uncertainty sampling, thereby ensuring diversity, effective and balanced utilization of the available information.
 VII Discussion
-This section builds on the preceding, by highlighting the main trends and discussion points in the field.
+This section highlights the main trends and discussion points in the field.
 The challenges and limitations related to the methods (Section VII-A) and downstream applications (Section VII-B) are discussed.
 Based on these observations, recommendations for future work will be provided in Section VII-C.
-VII-A Methods
-The core challenge with modeling segmentation distributions entails calibrated uncertainty estimates and secondly, producing spatial coherence in its samples.
+VII-A Discussion on methods
+The core challenges with modeling segmentation distributions entails obtaining calibrated uncertainty estimates and secondly, designing spatial coherence between pixels of the samples.
 While the former can be achieved with conventional CNN models, the latter requires modeling correlation across pixels of the segmentation mask.
 As discussed in earlier sections, the desired nature of the uncertainty estimates governs the required methodology.
-Hence, our discussion will follow a similar format as the preceding sections, i.e.
+Similarly, our discussion will follow a comparable format as the preceding sections, i.e.
 the aleatoric and epistemic approaches will initially be treated in a separate manner.
@@
 The literature overview indicates that three distinct routes can be taken to model the correlation.
-The first method is to model the correlation directly in pixels space.
-The second method entails using latent-variable modeling, where the correlation is encoded in latent space.
-Finally, test-time augmentation is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models.
+The first method is to model the correlation directly in the pixel space.
+The second method entails using latent-variable modeling, where the correlation is encoded in the latent space.
+Finally, TTA is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models.
 We will discuss the strengths and weaknesses of each of those approaches.
@@
 Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification.
-Nonetheless, the shortcomings of VAEs are well-known.
-For example, such models suffer from inference suboptimality related to ELBO optimization [205, 59] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [65, 79, 70], which is hypothesized to be caused by excessively strong decoders [80].
-This is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [71], residual connections [71, 72] or deep supervision [72] are required for generalization.
-A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotator styles or the exploration of low-likelihood regions.
-Hence, VAE-based models serve as good choice for the task given its shortcomings are sufficiently addressed.
+Nonetheless, the shortcomings of VAEs are well known.
+For example, such models suffer from inference suboptimality related to ELBO optimization [207, 65] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [73, 63, 68].
+This has been hypothesized to be caused by excessively strong decoders [64] and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [74], residual connections [74, 75], or deep supervision [75] are required for generalization.
+A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions.
+Hence, VAE-based models serve as good choice for the task given their shortcomings are sufficiently addressed and partially solved.
 Compared to the VAE-based models, the adoption of DDPMs is rather limited.
 This is regardless of the fact that it outperforms the VAE-based methods.
-Besides the point that DDPMs emerged much later in the field, its crucial limitation is regarding the tedious sequential inference procedure [92, 206, 207].
-This shortcoming is exacerbated in supervised settings, which often validate through sampling on a separate data split.
-Furthermore, it can be noted that the best performing DDPM models are discrete in nature.
-While it is debatable whether shifting to categorical distributions is required for complex image generation [134], this observation does signals further investigation of the merits of categorical distributions in segmentation settings.
-This can visually be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
-Regardless, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g.
+Besides the point that DDPMs emerged much later in the field, their crucial limitation, similar to thee PixelCNN [53], is the tedious sequential inference procedure [92, 208, 209].
+This shortcoming is exacerbated in supervised settings, which often require validation through sampling on a separate data split.
+Regardless of these shortcomings, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g.
 blurry reconstructions of VAEs).
+Furthermore, it can be noted that the best-performing DDPM models are discrete in nature.
+While it is debatable whether shifting to categorical distributions is required for complex image generation [136], this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings.
+Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
 (a) LIDC-IDRI [131]
@@
 Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
-Implicit methods such as test-time augmentation (TTA) [102, 103, 104, 105] have received some use in literature but have quickly been surpassed by alternative methods.
-However, TTA does lead in the category of simplicity, requiring almost no additional mechanisms or modifications to the employed architecture.
+Implicit methods such as TTA [103, 104, 105, 106] have received some use in literature, but have been quickly surpassed by alternative methods.
+However, TTA does lead in the category of simplicity, requiring no additional mechanisms or modifications to the employed architecture.
 Furthermore, CAR [56] is a lesser used method and this is likely due to well-known training problems of GAN-based models [211].
-Similar to inference suboptimality in VAEs, GANs similarly required additional heuristic terms in the training objective due to instability.
+Due to this, GANs require additional heuristic terms in the training objective due to instability.
 For example, the training objective of CAR is a summation of four separate losses.
-SSNs [50] directly model correlations in pixel space and deems a simple, fast and model-agnostic modeling.
+SSNs [50] directly model correlations in pixel space and deem simple, fast and model-agnostic modeling.
 Notably, SSNs also suffer from training instability due to the invertibility requirement of the covariance matrix.
-This can often simply be circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular.
-In practice, nevertheless, this quick-fix solution was much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
+This can be often simply circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular.
+In practice, nevertheless, this quick-fix solution is much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
 VII-A2 Epistemic methods
-In the context of segmentation, almost all discussed literature employ approximations of Variational Inference.
-In fact, usage of MC Dropout dominates the realm of epistemic uncertainty quantification, mainly due it being a relatively simple, cheap and straightforward approach.
+Almost all discussed literature quantifying epistemic uncertainty employ approximations of VI.
+In fact, usage of MC Dropout dominates, mainly due it being a relatively simple, inexpensive and straightforward approach.
 Nevertheless, MC dropout has been subject to substantial criticism [212, 213, 109, 123].
-For example, it has been shown that MC dropout can assign 0 probability to the true posterior, which can also erroneously possesses multi-modality [210].
-Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate rather than the observed data [211, 108].
-These weaknesses of dropout have been used as a basis for alternative dropout techniques where the dropout rate is learned [108, 122].
-Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of a functional model.
-If accurate uncertainty quantification is critical, MC dropout should be avoided altogether.
+For example, it has been shown that MC dropout can assign a probability of zero to the true posterior, which can also erroneously possess multi-modality [212].
+Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate, rather than the observed data [213, 109].
+These weaknesses have been used as a basis for alternative techniques, where the dropout rate is learned [109, 123].
+Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of an already functional model.
+If accurate uncertainty quantification is critical, MC dropout should be avoided in any case.
 Finally, the optimal method for uncertainty quantification has yet to be determined.
@@
 Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout when evaluating with the PAvPU metric [151].
-All things considered, the preference for a particular methodology is seems to carry a strong data-dependency [124, 14].
-Our recommendation is to experiment with both explicit VI and approximations such as MC Dropout and ensembling.
-VII-B Applications
+All things considered, the preference for a particular methodology seems to carry a strong data-dependency [125, 14].
+Our recommendation is to experiment with both explicit VI and its approximations.
+VII-B Discussion on applications
 Up until now, theoretical insights have mostly been discussed.
-In this section, a deeper dive is taken into the domain-/downstream- level applications of uncertainty quantification.
-See Table III for an overview.
+In this section, a deeper dive is taken into the applications of uncertainty quantification.
+Table III provides an overview of these applications in various domains.
 Notably, most models are employed in healthcare use cases, where ambiguity frequently arises due to the trade-off between accuracy and the incisiveness of medical diagnosis systems.
@@
 TABLE III: Overview of domains using uncertainty quantification for segmentation.
-Observer variability Model introspection Active Learning Model generalization
-Lung [50, 58, 64, 65, 66, 68, 70, 71, 72, 136, 137, 76, 138, 139, 93] [169, 175, 187] [180] [187]
-Brain [50, 97, 95, 140, 93, 98, 137, 142] [161, 157, 159, 158, 175] [195] [157, 186, 159, 175]
-Outdoor scenes [71, 50, 145, 100, 154] [150, 151, 155, 149, 179, 152, 153] [200, 202] [149, 179, 150, 184, 185]
-Cardiovascular [163, 164, 174, 161, 176, 124, 165] [192] [212]
-Prostate [143, 73, 72, 136, 137] [161] [192]
+Observer Model Active Model
+variability Introspection Learning generalization
+Lung [50, 69, 59, 73, 60] [171, 177, 189] [182] [189]
+[62, 68, 74, 75, 138]
+[140, 141, 94, 139, 79]
+Brain [50, 98, 96, 142] [163, 159, 161, 160, 177] [197] [159, 188]
+[99, 139, 144, 94] [161, 177]
+Outdoor scenes [74, 50, 147] [152, 153, 157, 151, 181] [202, 204] [151, 181, 152]
+[101, 156] [186, 187]
+Cardiovascular [165, 166, 176, 163] [194] [214]
+[125, 167, 178]
+Prostate [145, 76, 75] [163] [194]
+Prostate [138, 139]
 Eye [100, 99, 60] [168, 169] [168]
@@
 Microscopy [97, 101, 74] [180, 175] [175]
-Others [139, 65, 98, 64, 95, 137] [170, 171, 175, 19, 172] [192, 198, 95] [188]
+Others [141, 73, 99, 59] [172, 173, 177, 19, 174] [194, 200, 96] [190]
+[96, 139]
 VII-B1 Observer variability
@@
 This observer variability is often a result of viewing limitations and therefore heavily correlate to the input image.
-In turn, this explains the success of learning input-conditional observer variability directly through explicitly modeling the likelihood distribution, rather than to infer the underlying latent parameter distribution.
+In turn, this explains the success of directly learning input-conditional observer variability, through explicitly modeling the likelihood distribution, rather than inferring the underlying latent parameter distribution.
 Nonetheless, this approach does have its limitations.
@@
 Firstly, it is evident from Table III that a wide range of datasets are used in literature, often involving proprietary in-house data.
-This makes replicating the presented results impossible.
-Secondly, consensus on data splitting is also lacking on the publicly available datasets.
-For example, significant incongruity across literature is observed for the LIDC-IDRI dataset regarding data splitting and preprocessing.
-Similarly, there is a lack of agreement on the implementation for the GED and HM-IoU metric.
-Several factors, including the choice of kernel, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
+This inherently makes replicating the presented results impossible.
+Secondly, consensus on data processing is also lacking on the publicly available datasets.
+For example, significant incongruity across literature is observed for the LIDC-IDRI dataset, regarding data splitting and preprocessing.
+Similarly, there is a lack of agreement on the implementation for the GED and HM metric.
+Several factors, including the kernel choice, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
 Regarding the evaluation metrics, we find that the literature often places excessive emphasis on improving them.
@@
 Furthermore, the quality of the GED evaluations have been subject to substantial criticism as well [216, 74, 63].
-Hence, practitioners should be extra vigilant and consider the importance to inquire with domain-level experts for qualitative evaluation.
-Because the GED and HM-IoU evaluations are dependent on on the number of segmentation masks available an alternative is to involve additional annotators per data point for more accurate evaluation, but is likely to be equally expensive.
-This fact additionally calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.
+Hence, practitioners should be extra vigilant and consider the importance to inquire domain-level experts for qualitative evaluation.
+Because the GED and HM are dependent on on the number of segmentation masks available, an alternative is to involve additional annotators per datapoint for more accurate evaluation, but is likely to be equally expensive.
+This aspect calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.
 VII-B2 Model introspection
@@
 This is especially evident by the frequent use of the PAvPU metric.
-Furthermore, a majority of the available research use Monte Carlo Dropout and correlated this to the output entropy to the prediction accuracy.
+Furthermore, a majority of the available research use MC Dropout and correlate the output entropy to the prediction accuracy.
 The disadvantage of relying on MC Dropout when the uncertainty is crucial, has been discussed in Section VII-A2.
-Therefore, we recommend incorporating explicit Variational Inference at specific points in the network to achieve more accurate uncertainty quantification.
-These points could be selected based on the layers that most significantly influence the output segmentation.
+Therefore, we recommend incorporating explicit VI at specific points in the network to achieve more accurate uncertainty quantification.
+These points can be selected based on the layers that influence the output segmentation most significantly.
 VII-B3 Model generalization
-Epistemic uncertainty quantification methods can result in improved performance as it often serves as a surrogate for model combination.
-Therefore, uncertainty quantification is can readily be available in the toolbox of deep learning researchers together with other commonly used regularization techniques.
-In this case, simple methods such as ensembling and MC dropout are despite their criticisms relatively harmless.
-Furthermore, parallels can be made with MC Dropout and placing an L2-norm penalty on the model weights.
-While, this benefits is compelling, improved model performance can also be obtained with more computationally efficient regularizers.
-Therefore, improved performance should not be the end goal of uncertainty quantification, but considered as an ancillary advantage.
+Epistemic uncertainty quantification methods can result in improved performance, as it often serves as a surrogate for model combination.
+Therefore, uncertainty quantification can be readily available in the toolbox of deep learning researchers together with other commonly used regularization techniques.
+In this case, simple methods such as ensembling and MC dropout, are despite their criticisms relatively harmless.
+Furthermore, a parallel characterization can be made with MC Dropout and placing an $L_{2}$-norm penalty on the model weights.
+While this benefit is compelling, improved model performance can be also obtained with more computationally efficient regularizers.
+Therefore, improved performance should not be the ultimate goal of uncertainty quantification, but considered as an ancillary advantage.
 VII-B4 Active Learning
-Active Learning, while being a challenging task, holds significant promise by potentially reducing the need for intensive labeling procedures that often require specialized expertise.
-Furthermore, this kind of approach hints towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate adoption of such models in sectors requiring extensive specialization.
-Furthermore, Active Learning can accelerate privacy-centric collaboration when combined in a federating setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local modals.
-However, active learning requires a well-generalized model on little data.
-This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and often requires datasets of substantial size.
-Furthermore, many uncertainty-based approaches often extend traditional active learning for classification by simply aggregating the pixel’s metrics together.
+The challenging task of Active Learning holds a significant promise and can potentially reduce the need for intensive labeling procedures, which often require specialized expertise.
+Furthermore, this kind of approach points towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate the adoption of such models in sectors requiring extensive specialization.
+Furthermore, Active Learning can accelerate privacy-centric collaboration when combined with a federated setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local models.
+However, Active Learning requires a well-generalized model using little data.
+This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and typically requires datasets of substantial size.
+Furthermore, many uncertainty-based approaches regularly extend conventional Active Learning for classification to segmentation by simply aggregating the pixel metrics together.
 For imbalanced settings, it has been shown that it can perform even worse than random selection [217].
-Therefore, additional modifications such as target, boundary or diversity awareness [198, 203], or region-based annotating [202, 216] are often required to apply Active Learning to segmentation tasks.
+Therefore, additional modifications such as target, boundary or diversity awareness [200, 205], or region-based annotation [204, 218] are often required to apply Active Learning to segmentation tasks.
 VII-C Future work
-In this penultimate section, we will provide point-by-point recommendations for future work.
+In this penultimate section, we will provide recommendations for future work.
 VII-C1 Exploring generative models
@@
 Also, literature on unsupervised VAEs have greatly benefited the PU-Net [73, 63, 68].
-Therefore, we advocate for a deeper contextualization of contemporary research within probabilistic segmentation models.
-In fact, any unsupervised generative model can theoretically be translated to the supervised setting through intricate conditioning and can therefore be used as a probabilistic segmentation model.
-Given this flexibility, it remains unclear why Normalizing Flow (NF) remains underutilized for this task.
-Specifically continuous NFs, which approximate the time-dependent score function, strongly resembles DDPMs which have extensively been used for segmentation problems.
+Therefore, we advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models.
+In fact, any unsupervised generative model can be theoretically translated to the supervised setting through intricate conditioning and can be therefore used as a probabilistic segmentation model.
+Given this flexibility, it remains unclear why Normalizing Flows (NFs) remain underutilized for this task.
+Specifically continuous NFs, which approximate the time-dependent score function, strongly resemble DDPMs which have been extensively used for segmentation problems.
 Furthermore, NFs enable explicit and exact evaluation of the likelihood, which can aid in further interpretation of model predictions.
-Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian and inference times are therefore much faster than that of DDPMs.
-It should be noted, however, that the invertibility constraint greatly hinders expressivity of intermediate functions, leading to memory-intensive architectures.
+Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian, so that inference times are therefore much faster.
+However, it should be noted that the invertibility constraint greatly hinders the expressiveness of intermediate functions, leading to memory-intensive architectures.
 VII-C2 Variational Inference and beyond
@@
 why ensembling or MC Dropout is often preferred over explicit VI).
-Additionally, recent studies have shown the benefits of moving from a few large to many small experts for when using a MoE ensemble (see Section V-C for language modeling [217].
+Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section V-C) for language modeling [219].
 Future work should also experiment with this.
@@
 Especially the Laplace Approximation can be very beneficial, as it is easily applicable to pretrained networks.
-Notably, both Laplace approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are useful when expecting to fit multi-modal parameter distributions.
+Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions.
 To the best of our knowledge, these approaches have not been studied within the context of uncertainty quantification in segmentation.
 VII-C3 Single-pass uncertainty
-The multiple forward passes required in Bayesian uncertainty quantification can incur cumbersome additional costs.
+The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs.
 Hence, considerable efforts have been made towards deterministic uncertainty models [220, 221, 222], which only depend on a single forward pass.
@@
 This approach achieves faster computation, while also providing both epistemic and aleatoric uncertainty.
-Along similar lines, Evidential Deep Learning also possess the advantage of quantifying both uncertainties with a single forward pass.
+Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass.
 This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) [223].
@@
 [225] recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model.
-Unfortunately, there has not been much research beyond this.
+Unfortunately, further research has not been published beyond this.
 VII-C4 Improved CNN architectures
 With the introduction of Vision Transformers [226], CNN-based models have been challenged in the task of semantic segmentation [227, 228, 229].
-Regardless the success of vision transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net remain the preferred backbone [39].
+Regardless the success of Vision Transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture [39].
 This is mainly because CNNs already possess desirable inductive biases, while Transformers conversely require extensive pretraining with large datasets [230].
@@
 For instance, “ConvNext” takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains [231].
-Since many innovations in this field focus on the technique of uncertainty quantification, used backbones receive less attention and might often be outdated.
-Our recommendation is to improve current models with developments in general CNN-based architectures.
+Since many innovations in this field focus on the technique of uncertainty quantification, the applied backbones receive less attention and might gradually become outdated.
+Our recommendation suggests improving current models with developments in generic CNN-based architectures.
 VII-C5 Distributing-free modeling
-A distribution-free framework known as Conformal Deep Learning produces prediction sets that guarantee to contain the ground-truth with a user-defined probability.
+A distribution-free framework known as Conformal Deep Learning, produces prediction sets that guarantee to contain the ground truth with a user-defined probability.
 With the help of an additional calibration set, a heuristic notion of ambiguity (i.e.
-miscalibrated softmax outputs) is transformed to rigorous uncertainty and is especially renowned for being model-agnostic, simple and highly flexible [230].
-Very recently, conformal prediction has been applied to segmentation problems [231, 232, 233, 234], thereby enjoying the aforementioned benefits of this framework and indicating an increased traction.
-We recommend further research in this direction to discover novel applications and to benchmark against current architectures.
+miscalibrated softmax outputs) is transformed to rigorous uncertainty.
+This technique is especially renown for being model-agnostic, simple and highly flexible [232].
+Quite recently, Conformal Prediction has been applied to segmentation problems [233, 234, 235, 236], thereby enjoying the aforementioned benefits and showing an increased traction of this framework.
+We recommend further research in this direction to discover novel applications and benchmark against current architectures.
 VII-C6 Volumetric segmentation
-In many clinical datasets, the volumetric data is in most instances sliced to patches and processed with conventional 2D-based CNN-models.
+In many clinical datasets, the volumetric data are in most instances sliced to patches and processed with conventional 2D-based CNN-models.
 However, such models are often straightforward extensions of existing 2D models (and almost exclusively VAE-based), rarely addressing novel challenges introduced with additional dimensionalities.
-For example, 3D extensions of the PU-Net simply remain to use similar techniques to insert latent samples in the decoding networks.
-Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization.
-Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners looking to implement models for volumetric segmentations problems.
-Also, methodologies need to be developed in order to appropriately compare 2D to 3D models.
+For example, 3D extensions of the PU-Net simply remain to use similar techniques for inserting latent samples in the decoding networks.
+Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization [179].
+Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to implement models for volumetric segmentations problems.
+Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models.
 For example, the GED cannot be estimated volumetrically with 2D models.
-Furthermore, translating already computationally intensive models (e.g.
-based on DDPMs or PixelCNNs) to three dimensions is another huge challenge due to the memory requirement involved volumetric data, which will even increase the training load and inference time even further.
-Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underlines the need of more investigation related to three-dimensional uncertainty quantification, which more closely aligns with real-world clinical practice and therefore encourages faster adoption.
+Moreover, translating the already computation-intensive models (e.g.
+based on DDPMs or PixelCNNs) to 3D is another challenge, due to the memory requirement involved with volumetric data, which will increase the training load and inference time even further.
+Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underline the need of more investigation related to 3D uncertainty quantification, that more closely aligns with real-world clinical practice and therefore encourages faster adoption.
 VIII Conclusion
@@
 With the vast body of literature, encompassing diverse applications and modalities, the need for a comprehensive and systematic overview of the field is addressed.
-We present clear definitions and notation of methodologies that attempt uncertainty modeling by considering the field from a theoretical perspective and relating this to various pertinent applications.
+We present clear definitions and unified notation of methodologies that attempt uncertainty modeling, by considering the field from a theoretical perspective.
 Aleatoric uncertainty can be modeled in pixel or latent space with generative models, or implicitly be expressed with test-time augmentation.
 Epistemic uncertainty is captured with Variational Inference on the parameter distribution, or approximated with Monte Carlo Dropout or model ensembling.
-Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks.
-Which in turn enable us to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and real-world applications.
-Our future recommendation pertain aligning the field with advancements in general generative modeling and deep learning architectures.
-Furthermore, the adoption of deterministic uncertainty quantification methods that do no require multiple forward passes such as Conformal and Evidential Deep Learning is suggested.
-Especially the latter approach is interesting due to its ability to encapsulate and express both uncertainty types.
-Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation is a beneficial contribution to the field.
-Finally, due to the clinical relevancy of uncertainty in semantic segmentation, more attention to models catered to volumetric data is advised.
+Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks, which enable to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and to the real-world applications.
+Our future recommendations pertain aligning the field with advancements in general generative modeling and Deep Learning architectures.
+Furthermore, the adoption of deterministic uncertainty quantification methods that avoid multiple forward passes, such as Conformal and Evidential Deep Learning, are suggested.
+Particularly the latter approach is interesting due to its ability to incorporate both uncertainty types.
+Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques, as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation, is a beneficial contribution to the field.
+Finally, due to the clinical relevance of uncertainty in semantic segmentation, more attention is advised to models catering to volumetric data.
 In this manner, the review paper guides researchers on the topic of probabilistic segmentation and suggests future endeavors within the lightning-fast evolving field of Deep Learning-based Computer Vision.

### v4

@@
 Abstract
-Advancements in image segmentation play an integral role within the broad scope of Deep Learning-based Computer Vision.
-Furthermore, their widespread applicability in critical real-world tasks has resulted in challenges related to the reliability of such algorithms.
-Hence, uncertainty quantification has been extensively studied within this context, enabling the expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision-making.
-Due to the rapid adoption of Convolutional Neural Network (CNN)-based segmentation models in high-stake applications, a substantial body of research has been published on this very topic, causing its swift expansion into a distinct field.
-This work provides a comprehensive overview of probabilistic segmentation, by discussing fundamental concepts of uncertainty quantification, governing advancements in the field as well as the application to various tasks.
-Moreover, literature on both types of uncertainties trace back to four key applications: (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) Active Learning.
-An extensive discussion follows that includes an overview of utilized datasets for each of the applications and evaluation of the available methods.
-We also highlight challenges related to architectures, uncertainty quantification methods, standardization and benchmarking, and finally end with recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
-Index Terms:
-image segmentation, uncertainty quantification, probabilistic modeling, epistemic, aleatoric
-I Introduction
-Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest [1].
-With the rapid development of Convolution Neural Networks (CNNs), Deep Learning-based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7].
-Nevertheless, such methodologies rely on extensive assumptions and relaxations on the Bayesian learning paradigm, omitting crucial information on the uncertainty associated with the model predictions.
-This omission diminishes the reliability and interpretability of such models.
-For example, a difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
-Extensive efforts have been made to align modern neural network optimization with Bayesian Machine Learning [8, 9, 10, 11], such as learning parameter densities, rather than point estimates, to include a notion of epistemic uncertainty.
-Furthermore, explicitly modeling the output likelihood distribution enables expressing the aleatoric uncertainty.
-Notably, literature mention that determining the nature of uncertainty is not often straightforward.
-For example, Hüllermeier et al.
-[12] stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”.
-This sentiment is also shared by Kiureghian and Ditlevsen [13], noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic”.
-Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic.
-Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview.
-Nonetheless, most surveys take the perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19].
-There is a notable absence of a comprehensive overview in this field, relating theoretical foundation to the multitude of applications.
-Furthermore, the abundance of available works can be often overwhelming for both emerging and seasoned researchers.
-This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are discussed through standardized notation and clarified by mutual comparison.
-After reading this work, readers will be able to discern various forms of uncertainty with their pertinent applications to segmentation tasks.
-Additionally, they will have developed comprehensive understanding of challenges and unexplored avenues in the field.
-This review paper is structured as follows.
-Past work with significant impact in general image segmentation is presented in Section II, where closely related surveys are also referenced.
-Then, the theoretical framework and notation that governs the remainder of the paper is introduced in Section III.
-The role of these concepts in image segmentation will be treated in Sections IV and V, which include all architectures and approaches with significant impact on the field.
-We then consider the applications that use these uncertainty estimates in Section VI.
-Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII, which is concluded in Section VIII.
-Figure 1 illustrates a brief overview of the sections.
+Advances in architectural design, data availability, and compute have driven remarkable progress in semantic segmentation.
+Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making.
+The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented, marked by inconsistent terminology, evaluation practices, and theoretical ambiguities.
+In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising.
+By unifying theory, terminology, and applications, we provide a coherent foundation for researchers and identify critical challenges, such as strong assumptions in spatial aggregation, lack of standardized benchmarks, and pitfalls in current uncertainty quantification methods.
+We identify trends such as the adoption of contemporary generative models, driven by advances in the broader field of generative modeling, with segmentation-specific innovation primarily in the conditioning mechanisms.
+Moreover, we observe growing interest in distribution- and sampling-free approaches to uncertainty estimation.
+We further propose directions for advancing uncertainty-aware segmentation in deep learning, including pragmatic strategies for disentangling different sources of uncertainty, novel uncertainty modeling approaches and improved Transformer-based backbones.
+In this way, we aim to support the development of more reliable, efficient, and interpretable segmentation models that effectively incorporate uncertainty into real-world applications.
+1 Introduction
+Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest (Szeliski, 2010).
+The advent of convolutional neural networks (CNNs) has led to major breakthroughs in this domain, with deep learning-based methods achieving state-of-the-art performance on large-scale datasets (Ronneberger et al., 2015; Shelhamer et al., 2014; Badrinarayanan et al., 2015), obtaining impressive scores with large-scale segmentation datasets (Lin et al., 2014; Cordts et al., 2016b; Richter et al., 2016).
+However, these models typically rely on strong assumptions and significant relaxations of the Bayesian learning paradigm, neglecting the uncertainty associated with their predictions.
+This lack of uncertainty modeling reduces both the reliability and interpretability of the predictions.
+In high-stakes applications, such as autonomous driving or medical diagnosis, this can have severe consequences.
+For instance, misclassifying adjacent objects in autonomous driving or overlooking uncertainty in lesion classification can both lead to critical decision-making errors.
+Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications.
+Extensive efforts have been made to align neural network optimization with Bayesian machine learning (Blundell et al., 2015; Guo et al., 2017; Kingma & Welling, 2013; Kendall & Gal, 2017), such as learning parameter distributions rather than point estimates to capture epistemic uncertainty.
+Additionally, explicitly modeling the likelihood of outputs enables estimation of aleatoric uncertainty.
+However, in practice, these two forms of uncertainty are often entangled or even conflicting (Mucsányi et al., 2024; de Jong et al., 2024; Valdenegro-Toro & Mori, 2022; Hüllermeier & Waegeman, 2021), making the distinction difficult to define and apply.
+This conceptual ambiguity presents significant barriers to understanding and implementation (Kahl et al., 2024; Der Kiureghian & Ditlevsen, 2009; Kirchhof et al., 2025).
+At the same time, uncertainty is often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective.
+As a result, many works lack rigorous theoretical grounding, focusing instead on empirical gains without clear definitions or principled formulations of uncertainty.
+Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability.
+Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, without being properly contextualized in segmentation.
+This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions.
+The resulting abundance of literature can be overwhelming, even for experienced researchers.
+This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks.
+Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kahl et al., 2024; Kirchhof et al., 2025), we identify a notable absence of a comprehensive, theory-informed overview in the field.
+Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023).
+However, the connection between theoretical foundations and their diverse applications remains underexplored.
+Additionally, the study by Kahl et al.
+(2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks.
+In contrast, our work provides a comprehensive overview that spans the full range of established uncertainty quantification methods, unifies and contextualizes the underlying theory, and standardizes terminology and notation.
+This broader perspective helps clarify conceptual foundations and highlights cross-cutting insights across segmentation architectures and application domains.
+By the end of this paper, readers will have a clear understanding of the various forms of uncertainty, their relevance to segmentation tasks, and a comprehensive grasp of the key challenges and open research directions in the field.
 […]
 Figure 1: Overview of the sections. The leafs of the presented hierarchical tree are related to subsections of the article.
-II Background
-This section highlights a selection of literature crucial to the development of current probabilistic segmentation models.
-While literature on segmentation itself is vast, this decision has resulted in a limited collection of studies.
+This review paper is structured as follows.
+Past work with significant impact in general image segmentation is presented in Section 2.
+Then, the theoretical framework and notation that govern the remainder of the paper is introduced in Section 3.
+The theory enables us to classify the approaches into two distinct but often related methods.
+That is, aleatoric and epistemic uncertainty modeling methods.
+The role of these concepts in image segmentation will be treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field.
+We then correlate these modeling methods to specific applications in Section 6.
+Here, we also observe for which domains uncertainty quantification has been exploited.
+Then, this overview will be discussed further in Section 7.
+This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task.
+To encourage further research in this field, we provide future recommendations by highlighting key challenges, gaps and pitfalls in literature, concluding our review in Section 8.
+Figure 1 illustrates a brief overview of the sections.
+2 Background
+A common approach in image understanding involves labeling pixels according to semantic categories, which is the core of semantic segmentation.
+This technique is particularly well-suited for amorphous or uncountable subjects.
+In contrast, instance segmentation not only assigns class labels, but also distinguishes and delineates individual object occurrences.
+This makes it more appropriate for scenarios involving countable entities.
+A third variant, panoptic segmentation, unifies both semantic and instance-level labeling, offering a comprehensive view of scene composition.
 As summarized by Minaee et al.
-[20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watershedding [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29].
-After the application of CNNs [30], the domain of image segmentation underwent rapid developments.
-Notably, the Fully Convolutional Network (FCN) [3] adapted the AlexNet [31], VGG16 [32] and GoogLeNet [33] architectures to enable end-to-end semantic segmentation.
-Furthermore, other CNN architectures such as DeepLabv3 [34], and the MobileNetv3 [35] have also been commonly used.
-As the research progressed, increasing success has been observed with encoder-decoder models [36, 4, 37, 2].
-Initially developed for the medical applications, Ronneberger et al.
-[2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps.
+(2020), semantic segmentation has been performed using methods such as thresholding (Otsu, 1979), histogram-based bundling, region-growing (Dhanachandra et al., 2015), k-means clustering (Nock & Nielsen, 2004), watershedding (Najman & Schmitt, 1994), to more advanced algorithms such as active contours (Kass et al., 2004), graph cuts (Boykov et al., 2001), conditional and Markov random fields (Plath et al., 2009), and sparsity-based methods (Starck et al., 2005; Minaee & Wang, 2017).
+While the literature on segmentation is vast and rapidly evolving, a selection of backbone architectures has been particularly influential in shaping current probabilistic segmentation models.
+The focus here is not on exhaustiveness, but on those models most relevant to the development of uncertainty-aware approaches.
+In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures.
+Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan & Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation.
+Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used.
+As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015).
+Initially developed for the medical applications, Ronneberger Ronneberger et al.
+(2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps.
 To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain.
-In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
-Semantic segmentation focuses solely on assigning class labels to each pixel and is particularly suitable for amorphous or uncountable subjects of interest.
-In contrast, instance segmentation not only detects, but also delineates individual objects within the image.
-This form of segmentation is more applicable when identifying and outlining countable instances of objects.
-The third category, panoptic segmentation, combines both class and instance-level classification [20].
-III Probabilistic Image Segmentation
-This section describes the fundamentals of Bayesian inference and its relationship with conventional models for segmentation.
-The provided details introduce standard variables and notations as we proceed throughout the manuscript.
-III-A Bayesian inference
-Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image.
+In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
+3 Probabilistic Image Segmentation
+Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image.
 The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively.
@@
 the variance of the model parameters.
-A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy, which can similarly be decomposed as
+Encapsulating a particular uncertainty can therefore be achieved by incorporating parametrized stochasticity at either feature- or parameter-level.
+A straightforward approach to quantify any of these uncertainties is achieved by obtaining the predictive entropy, $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]$, i.e.
+the entropy of the predictive distribution $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})$.
+Furthermore, disentangling these uncertainties can be achieved through its decomposition
 $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}}.$ (3)
-III-B Conventional segmentation
+In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process.
+Nonetheless, determining the nature of uncertainty is not often straightforward.
+For example, Hüllermeier & Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred".
+This sentiment is also shared by Der Kiureghian & Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic".
+Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty.
+Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
+Figure 2: Aleatoric uncertainty quantification by modeling pixel-level outputs as parameters of a Probability Mass Function.
+3.1 Conventional segmentation
 Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks, which are trained by Maximum Likelihood Estimation (MLE) and is specified as
@@
 This approximation improves as the training data increases and the model parameter variances approach zero.
-As such, MLE does not include any prior knowledge on the structure of the parameter distribution.
-This is typically achieved through Bayesian Maximum A Posteriori (MAP) estimation with
+As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Bayesian Maximum A Posteriori (MAP) estimation with
 $\bm{\theta}_{\text{MAP}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (5)
-For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively [41, 42].
+For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007).
 Additionally, the output of such models can be interpreted as a probability distribution.
@@
 omits information on structural variation in the segmentation masks.
-At the very core, this limitation (See Figure 3) has driven research on probabilistic segmentation, enabling the sampling of spatially coherent segmentation masks.
-This challenge can be addressed either from the aleatoric or epistemic perspective.
-Both approaches have relevant specific use cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
-Figure 2: Aleatoric uncertainty quantification by modeling pixel-level outputs as parameters of a Probability Mass Function.
+In conventional classification, factorizing the categorical distribution is typically regarded as a logical simplification.
+In probabilistic segmentation however, this assumption has caused the emergence of a distinct research direction (See Figure 3).
+The challenge of sampling of spatially coherent segmentation masks can be addressed either from the aleatoric or epistemic perspective, usually indicating reducability of the assumed uncertainty source.
+Both approaches have specific use cases, and each modeling choice comes with distinct advantages and limitations, which will be discussed in the following sections.
 (a) Likelihoods
@@
 (c) However, the coherence of the segmentation suffers when sampling.
-IV Aleatoric uncertainty
-Aleatoric uncertainty quantification reconsiders the non-deterministic relationship between $\mathbf{x}\in\mathcal{X}$ and $\mathbf{y}\in\mathcal{Y}$, which implies that
+4 Aleatoric uncertainty
+[…]
 $p(\mathbf{Y}|\mathbf{X})=\frac{p(\mathbf{Y},\mathbf{X})}{p(\mathbf{X})}\neq\delta(\mathbf{Y}-F(\mathbf{X})),$ (8)
@@
 In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$.
-For example, it is possible to model a distribution parameterized by the output of a CNN (Section IV-A).
-Also, generative models have been used extensively, where the data likelihood is learned through latent variables (Section IV-B).
-Finally, image augmentation during test-time inference can be also applied to obtain a notion of aleatoric uncertainty (Section IV-C).
-IV-A Pixel-level sampling
-A valid approach employs direct pixel-level uncertainty in the annotations.
-In case the independence assumption holds, it is important to ensure proper calibration of the model, which is discussed in Section IV-A1.
-In another method, the spatial correlation is explicitly modeled across the pixels of the segmentation mask and results in spatially coherent samples, which will be treated in Section IV-A2.
-IV-A1 Independence
-The normalized confidence values that result from SoftMax activation can only be interpreted as probabilities after proper validation, which is referred to as model calibration.
-Ideally, the empirical accuracy of a model should be approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e.
+For example, it is possible to model a distribution parameterized by the output of a CNN (Section 4.1).
+Also, conditional generative models are used to learn the data distribution through latent variables.
+For the purpose of taxonomy, we categorize models based on the location of these latent (unobserved) variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2.
+While their theoretical formulations are largely similar, their practical implications can differ significantly.
+4.1 Pixel-level sampling
+Uncertainty in segmentation masks can be modeled directly at the pixel level.
+These approaches can be further categorized into those that assume independence between pixels (Section 4.1.1), and those that explicitly model spatial correlations (Section 4.1.2).
+In the former case, accurate uncertainty estimates rely on well-calibrated models, which is an assumption that often fails in practice without additional tuning on a separate calibration set.
+In the latter case, a stochastic variable is introduced to capture dependencies between neighboring pixels.
+4.1.1 Independence
+(a) Calibrated
+(b) Miscalibrated
+Figure 4: Visualization of a reliability diagram, which illustrates whether the model’s predicted confidences align with the observed empirical accuracies.
+As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution.
+However, these values can only be interpreted as probabilities after proper validation, which is referred to as model calibration.
+Here, it is measured whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e.
 $P(Y=k\,|\,c_{k})=c_{k}$.
 Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4).
-Different methods can be used to measure calibration, but often introduce their own biases.
+Furthermore, different methods can be used to quantify calibration, although each may introduce its own biases.
 A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
@@
 The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence.
-Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
-(a) Calibrated
-(b) Miscalibrated
-Figure 4: Visualization of a reliability diagram indicating if the model confidences accurately reflect the empirical probabilities.
-Contemporary SoftMax-activated neural networks often portray a significantly incongruous reflection of the true data uncertainty because of negative log-likelihood optimization and techniques such as batch normalization, weight decay and other forms of regularization [9].
-Most calibration techniques are post-hoc, i.e.
-they occur after training and thus require a separate validation set.
-For example, Temperature Scaling [9] has been applied in a pixel-wise manner for segmentation problems [43].
-Nonetheless, some methods, such as Label Smoothing [44, 45] or using the Focal Loss [46] can be directly applied on the training data.
-Furthermore, over-fitting has often been considered to be the cause of over-confidence [47, 48] and erroneous pixels can therefore be penalized through regularizing low-entropy outputs [49].
-IV-A2 Spatial correlation
-Figure 5: Depiction of Stochastic Segmentation Networks [50]. Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
-To explicitly model spatial correlation within pixel space of the likelihood distribution, Monteiro et al.
-[50] propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
-$p(\mathbf{a}|\mathbf{x},\bm{\theta})=\mathcal{N}(\mathbf{a};\bm{\mu}=f^{\bm{\mu}}_{\bm{\theta}}(\mathbf{x}),\bm{\Sigma}=f^{\bm{\Sigma}}_{\bm{\theta}}(\mathbf{x})),$ (10)
-where the covariance matrix has a low-rank structure
-$\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}.$ (11)
-Here, $\mathbf{P}$ has dimensionality $((K\times D)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ a diagonal matrix.
-This results in a more structured and expressive distribution, while retaining reasonable efficiency.
-As the SoftMax transform on this low-rank multivariate normal distribution pertains an intractable integral, Monte Carlo sampling is employed.
-SSNs can be augmented as an additional layer to any existing CNN (see Figure 5).
-Another method uses an autoregressive approach to predict pixel $Y_{i}$ based on the preceding pixels. In this case, we can rephrase Equation (7) to
-$p(\mathbf{Y}|\mathbf{X})=\prod_{i}^{K\times D}p(Y_{i}|Y_{1},...,Y_{i-1},\mathbf{X}).$ (12)
-A popular method to model such relationship due to its substantial receptive field is known as the PixelCNN [51, 52].
-Zhang et al.
-[53] suggest to use this network to predict a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$).
+Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.Additionally, when background pixels have a predominant influence, each bin can be weighted equally using the Average Calibration Error (ACE) to mitigate this imbalance Jungo et al.
+(2020); Neumann et al.
+(2018).
+Nonetheless, contemporary neural networks often exhibit poorly calibrated uncertainty estimates.
+This misalignment is hypothesized due to the use of negative log-likelihood as the training objective, along with regularization techniques such as batch normalization, weight decay, and others (Guo et al., 2017).
+As a result, calibration methods are typically required to adjust the predicted confidences.
+Since most of these techniques are post-hoc, i.e.
+applied after training, they necessitate a separate validation set.
+For example, Temperature Scaling (Guo et al., 2017) has been applied in a pixel-wise manner for segmentation problems (Ding et al., 2021).
+Nonetheless, some methods, such as Label Smoothing (Silva & Oliveira, 2021; Liu et al., 2022a) or using the Focal Loss (Mukhoti et al., 2020) can be directly applied on the training data.
+Furthermore, over-fitting has often been considered to be the cause of over-confidence (Szegedy et al., 2016; Pereyra et al., 2017) and erroneous pixels can therefore be penalized through regularizing low-entropy outputs (Larrazabal et al., 2021a).
+4.1.2 Spatial correlation
+Figure 5: Illustration of the PixelCNN-based PixelSeg (Zhang et al., 2022c).
+Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation.
+Dotted elements appear during test-time sampling.
+Assuming pixel independence inhibits spatially coherent uncertainty quantification, which is a key factor for segmentation.
+Introducing spatial correlation can be achieved through an autoregressive approach.
+For instance, we can rephrase Equation (7) to
+$p(\*Y|\*X)=\prod_{i}^{K\times D}p(Y_{i}|Y_{1},...,Y_{i-1},\*X),$ (10)
+where pixel $Y_{i}$ is predicted based on the preceding pixels.
+A popular implementation of this formulation is known as the PixelCNN (Van Den Oord et al., 2016).
+For dense predictions, Zhang et al.
+(2022c) propose PixelSeg, which predicts a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$).
 The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples.
@@
 Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
-Figure 6: Illustration of the PixelCNN-based PixelSeg [53].
-Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation.
-Dotted elements appear during test-time sampling.
-IV-B Latent-level sampling
-Directly learning the conditional data distribution is a challenging task.
-Hence, generative models often employ a simpler latent (unobserved) variable $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{z}\in\mathbb{R}^{d}$, to instead learn the approximate joint density $p(\mathbf{Y}|\mathbf{Z},\mathbf{X})$.
-The marginalized distribution is obtained through integral
+Figure 6: Depiction of Stochastic Segmentation Networks (Monteiro et al., 2020). Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
+Monteiro et al.
+(2020) propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
+$p(\mathbf{a}|\mathbf{x},\bm{\theta})=\mathcal{N}(\mathbf{a};\bm{\mu}=f^{\bm{\mu}}_{\bm{\theta}}(\mathbf{x}),\bm{\Sigma}=f^{\bm{\Sigma}}_{\bm{\theta}}(\mathbf{x})),$ (11)
+where the covariance matrix has a low-rank structure $\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}$, with $\mathbf{P}$ having dimensionality $((K\times D)\times R)$, $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ representing a diagonal matrix.
+The low-rank assumption results in a more structured and expressive distribution, while retaining reasonable efficiency.
+Since, the SoftMax transform pertains an intractable integral, Monte Carlo sampling is employed.
+SSNs can theoretically be augmented as an additional layer to any pretrained CNN (see Figure 6).
+4.2 Latent-level sampling
+The complexity of modeling distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
 $p_{\bm{\theta},\bm{\psi}}(\mathbf{Y}|\mathbf{X})=\int p_{\bm{\theta}}(\mathbf{Y}|\mathbf{z},\mathbf{X})p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\mathrm{d}\mathbf{z},$ (12)
 with parameters $\bm{\theta},\bm{\psi}$.
-Conditioning the latent density on the input images is not a necessity, but usually preferred for smooth optimization trajectories [54].
-As such, the spatial correlation is not explicitly modeled, rather induced through mapping the latent variables to segmentation masks.
-Notable generative model architectures relevant to this section are Generative Adversarial Networks (Section IV-B1), Variational Autoencoders (Section IV-B2) and Denoising Diffusion Probabilistic Models (Section IV-B3).
-IV-B1 Generative Adversarial Networks (GANs)
-A straightforward approach is to simply learn the marginalization in Equation (13) through sampling from an unconditional prior density
-$p_{\mathbf{Z}}=\mathcal{N}(\bm{\mu}=\mathbf{0},\bm{\Sigma}=\mathbf{I}),$ (14)
-and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$.
-Goodfellow et al.
-[55] show that this approach can be notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$.
-In this way, $G_{\bm{\phi}}$ learns to reconstruct realistically looking images through the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient guidance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative.
-We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
-$J_{G_{\bm{\phi}}}=-J_{D_{\bm{\psi}}}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]$ (15)
-$-\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[\,\log(1-D_{\bm{\psi}}(G_{\bm{\phi}}(\mathbf{z},\mathbf{x})))\,].$
-Figure 7: Diagram of the Calibrated Adversarial Refinement network [56], based on the Generative Adversarial Network with additional loss terms.
-While conditional GANs have been used for semantic segmentation earlier [57], Kassapis et al.
-[56] explicitly contextualized the architecture within aleatoric uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7).
-The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\mathbf{c}$, with (cross-entropy) reconstruction loss
+As such, the spatial correlation is induced through mapping the latent variables to segmentation masks.
+Conditioning the latent density, i.e.
+$p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})$, on the input images is not a necessity, but usually preferred for smooth optimization trajectories Zheng et al.
+(2022).
+However, it is also possible to simply employ an unconditional prior $p(\mathbf{z})$.
+As argued by Kahl et al.
+(2024), their predictive uncertainty can be decomposed as
+$H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{z}|\mathbf{x},\bm{\theta}]}_{\text{aleatoric}}+\overbrace{\mathbb{E}_{p(\mathbf{z})}[\,H[\mathbf{Y}|\mathbf{z},\mathbf{x},\bm{\theta}]\,]}^{{\text{epistemic}}}.$ (13)
+In contrast, the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty.
+The key overarching contribution of these methods lies in the conditioning of generative models on the input image.
+This section briefly introduces Generative Adversarial Networks (Section 4.2.1), Variational Autoencoders (Section 4.2.2), and Denoising Diffusion Probabilistic Models (Section 4.2.3), and outlines how these architectures are commonly adapted for probabilistic segmentation.
+4.2.1 Generative Adversarial Networks (GANs)
+[…]
+$J_{G_{\bm{\phi}}}=-J_{D_{\bm{\psi}}}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]-\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[\,\log(1-D_{\bm{\psi}}(G_{\bm{\phi}}(\mathbf{z},\mathbf{x})))\,].$ (14)
+Figure 7: Diagram of the Calibrated Adversarial Refinement network Kassapis et al. (2021), based on the Generative Adversarial Network with additional loss terms.
+While conditional GANs have been used for semantic segmentation earlier (Isola et al., 2017), Kassapis et al.
+(2021) explicitly contextualized the architecture within aleatoric uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7).
+The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\*c$, with (cross-entropy) reconstruction loss
 $\mathcal{L}_{\text{rec}}=-\mathbb{E}_{p_{\mathcal{D}}}[\log p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})].$ (15)
-Then, the conditional refinement network $G_{\bm{\theta}}$ uses $\mathbf{c}$ together with $\mathbf{X}$ and latent samples $\mathbf{Z}_{i}\sim p_{\mathbf{Z}}$ injected at multiple decomposition scales $i$, to predict various segmentation maps.
+Then, the conditional refinement network $G_{\bm{\theta}}$ uses $\mathbf{c}$ together with input image $\mathbf{x}$ and latent samples $\mathbf{z}_{i}\sim p_{\*{Z}}$ injected at multiple decomposition scales $i$, to predict various segmentation maps.
 Furthermore, the refinement network is subject to the adversarial objective
@@
 […]
-$\hskip 5.0pt-\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,].$
 Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities, denoted by
@@
 Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification.
-The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps through latent variable $\mathbf{Z}$.
-IV-B2 Variational Autoencoders (VAEs)
-Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness [58].
-This phenomenon has also been successfully applied to VAE-like models such as the PU-Net [59, 60].
-NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $\mathbf{f}=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex distributions $\log p(\mathbf{x}|\mathbf{z})$, where symbol $\circ$ indicates functional composition.
-These are often denoted as $\log p(\mathbf{x})$ for simplicity and can be determined through
-$\log p(\mathbf{x})=\log p_{\mathbf{Z}}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (21)
-where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are intermediate variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$.
-Equation (21) can be substituted in the conditional ELBO objective in Equation (25) to obtain
-$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
-$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,]$
-$\hskip 20.0pt-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (25)
-where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF.
-Bhat et al.
-[61, 62] compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix.
-Valiuddin et al.
-[63] show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse [64, 65].
-Their proposition considers the alternative formulation of the ELBO, specified as
-$\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (23)
-$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]$
-$\hskip 70.0pt-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
-The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables. This can be rewritten into the objective
-$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (24)
-$\hskip 40.0pt+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]$
-$\hskip 80.0pt+\beta\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$
-with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence [66] and $\alpha,\beta$ denoting hyperparameters.
-This adaptation results in a more uniform latent space leading to increased model performance.
-Also, modeling the ELBO of the joint density has been explored [67].
-This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data.
-Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon [68].
-Density parameterization
+The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps.
+4.2.2 Variational Autoencoders (VAEs)
 Techniques such as GANs rely on implicit distributions and are void of any notion of data likelihoods.
@@
 the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, obtained by maximizing the conditional Evidence Lower Bound (ELBO)
-$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
-$=\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
-$\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
-$\hskip 50.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,].$ (22)
-This approach of estimating $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$ is also known as Variational Inference (VI).
-The first term in Equation (25) represents the reconstruction cost of the decoder subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$.
+$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
+$=\mathbb{E}_{q{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
+$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,].$ (20)
+The first term in Equation (20) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$.
 The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density.
-As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\phi$, $\theta$ and $\psi$.
+As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\bm{\phi}$, $\bm{\theta}$ and $\bm{\psi}$.
 The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
-$p(\mathbf{Y}|\mathbf{x}^{*})=\int p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{x}^{*})q_{\theta}(\mathbf{z}|\mathcal{D})\mathrm{d}\mathbf{z}.$ (26)
-The implementation of the conditional ELBO in Equation (25) can be achieved through a VAE-like architecture [10].
-A few additional design choices specific for uncertainty quantification in segmentation lead to the Probabilistic U-Net (PU-Net) [69].
-Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on $\mathbf{X}$, where the vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolution for classification.
-When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the provided information by $\mathbf{Z}$ is almost exclusively related to the segmentation variability.
+$p(\mathbf{Y}|\mathbf{x}^{*})=\int p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{x}^{*})q_{\bm{\theta}}(\mathbf{z}|\mathbf{x}^{*})\mathrm{d}\mathbf{z}.$ (21)
+The implementation of the conditional ELBO in Equation (20) can be achieved through the well known VAE architecture (Kingma & Welling, 2013).
+Furthermore, some additional design choices lead to the Probabilistic U-Net (PU-Net) (Kohl et al., 2019b).
+Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on the images.
+The latent vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolution for classification.
+When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the variability in the latent samples is almost exclusively related to the segmentation variability.
 Therefore, relatively smaller values of $d$ are feasible than what is commonly used in conventional image generation tasks.
-Figure 8: Depiction of the Probabilistic U-Net [69] based on a conditional Variational Autoencoder [10]. The latent samples are inserted at the final stages of a U-Net through a tiling operation.
-Similar to related research on the VAE [65, 70, 71, 72, 58], much work has been dedicated to improving the PU-Net.
-For instance, investigation on improving VI with novel parameterization of the amortized densities also provides insights into model behaviour [59, 73, 60, 61, 62, 67, 68] (Section IV-B2).
-Furthermore, extending the architecture to multiple decomposition hierarchies [74, 75] (Section IV-B2), three-dimensional modalities [76, 77, 78, 79, 80] (Section IV-B2), and conditioning on the annotator [81] (Section IV-B2), also result in substantial performance gains.
+Similar to related research on the VAE (Zhao et al., 2017; Bousquet et al., 2017; Higgins et al., 2017; Van Den Oord et al., 2017; Rezende & Mohamed, 2015), much work has been dedicated to improving the PU-Net.
+Figure 8: Depiction of the Probabilistic U-Net (Kohl et al., 2019b) based on a conditional Variational Autoencoder (Kingma & Welling, 2013).
+The latent samples are inserted at the final stages of a U-Net through a tiling operation.
+Density reparameterization
+[…]
+$\log p(\mathbf{x})=\log p_{\*Z}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (22)
+[…]
+$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
+$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,]-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (23)
+where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF.
+Bhat et al.
+(2023; 2022a) compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix.
+Valiuddin et al.
+(2024b) show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse (Chen et al., 2016; Zhao et al., 2017).
+Their proposition considers the alternative formulation of the ELBO, specified as
+$\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (24)
+$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
+The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables, i.e.
+the aleatoric uncertainty term of the predictive entropy in Equation (13).
+This can be rewritten into the objective
+$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (25)
+$\hskip 40.0pt+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]+\beta\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$
+with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence Cuturi (2013) and $\alpha,\beta$ denoting hyperparameters.
+This adaptation results in a more uniform latent space leading to increased model performance.
+Also, modeling the ELBO of the joint density has been explored (Zhang et al., 2022b).
+This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data.
+Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon (Qiu & Lui, 2020).
 Multi-scale approach
-Learning latent features over several hierarchical scales can provide expressive densities and interpretable features across various abstraction levels [82, 83, 84, 85, 86].
+Learning latent features over several hierarchical scales can provide expressive densities and interpretable features across various abstraction levels (Sønderby et al., 2016; Kingma et al., 2016; Klushyn et al., 2019; Gregor et al., 2015; Ranganath et al., 2016).
 Such models are commonly categorized under hierarchical VAE (HVAE) modeling.
@@
 $p(\mathbf{\mathbf{Y}|\mathbf{X}})\geq$ $\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$ (27)
-$-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})||q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]$
-$\hskip 20.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
-This objective is implemented in the Hierarchical PU-Net [74] (HPU-Net, depicted in Figure 9).
-Simply stated, the architecture learns a latent representation at multiple decomposition levels of the U-Net.
-Furthermore, residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL divergence rapidly approaching zero.
+$-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})||q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
+This objective is implemented in the Hierarchical PU-Net (Kohl et al., 2019a) (HPU-Net, depicted in Figure 9), which learns a latent representation at multiple decomposition levels of the U-Net.
+Residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL divergence rapidly approaching zero.
 For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective is employed, which extends Equation (27) to
 $\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (28)
-$-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})\,||\,q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]$
-$\hskip 20.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})\,||\,q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
+$-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})\,||\,q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})\,||\,q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
 Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value.
-Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick [87, 88].
-A similar architecture known as PHiSeg [75] takes a similar approach to the HPU-Net.
-However, instead of placing the residual connections in the convolutional layers, PHiSeg uses these between the latent vectors across the decompositions.
-Furthermore, deep supervision at each decomposition scale is used to enforce disentanglement across latent variables.
-Figure 9: Hierarchical Probabilistic U-Net [74] based on a hierarchical Variational Autoencoder [10, 82, 84].
+Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick (Jang et al., 2016; Huijben et al., 2022).
+A similar architecture known as PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net.
+However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers.
+Furthermore, deep supervision at each decomposition scale to enforce the disentanglement between latent variables.
+Figure 9: Hierarchical Probabilistic U-Net (Kohl et al., 2019a) based on a hierarchical Variational Autoencoder (Kingma & Welling, 2013; Sønderby et al., 2016; Klushyn et al., 2019).
 Instead of a single latent code, multiple decomposition scales encode the segmentation variability, depending on the depth of the U-Net structure.
@@
 Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to a potential loss of critical spatial information and subtle nuances often necessary for accurate delineation.
-This limitation has spurred extensive research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures.
-Initial works [79, 80] demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants.
+This limitation has spurred research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures.
+Initial works (Chotzoglou & Kainz, 2019; Long et al., 2021a) demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants.
 Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation.
 Viviers et al.
-[78] additionally augment a Normalizing Flow to the posterior density, as described in Section IV-B2.
-Further enhancements to the architecture include the implementation of the 3D HPU-Net [77], an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications [76].
+(2023c) additionally augment a Normalizing Flow to the posterior density for improved expressiveness, as discussed in Section 4.2.2.
+Further enhancements to the architecture include the implementation of the 3D HPU-Net (Saha et al., 2020), an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications (Saha et al., 2021a).
 Conditioning on annotator
 It can be relevant to model the annotators independently in cases with consistent annotator-segmentation pairs in the dataset.
-This can be achieved by conditioning the learned densities on the annotator itself [81, 89].
-For example, features of a U-Net can be combined with samples from annotator-specific Gaussian-distributed posteriors [89].
+This can be achieved by conditioning the learned densities on the annotator itself (Gao et al., 2022; Schmidt et al., 2023).
+For example, features of a U-Net can be combined with samples from annotator-specific Gaussian-distributed posteriors (Schmidt et al., 2023).
 Considering the approach from Gao et al.
-[81], generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator.
-Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $\mathbf{y}=F(\mathbf{x},\mathbf{z}_{k})$.
+(2022), generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator.
+Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $)\*y=F(\mathbf{x},\mathbf{z}_{k}$ ).
 The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble.
 Additionally, consistency between the model and ground-truth distribution is enforced through an optimal transport loss between the set of predictions and labels.
-IV-B3 Denoising Diffusion Probabilistic Models (DDPMs)
-Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models [90, 91, 92].
+4.2.3 Denoising Diffusion Probabilistic Models (DDPMs)
+Figure 10: Illustration of the Diffusion Probabilistic Models (Ho et al., 2020). The model learns to remove noise that has been gradually added to the input image.
+Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models (Ho et al., 2020; Song et al., 2022; 2020).
 Such models are especially renowned for their expressive power by being able to encapsulate large and diverse datasets.
-While several perspectives can be used to introduce the DDPMs, we build upon the earlier discussed HVAE (Section IV-B2) to maintain cohesiveness with the overall manuscript.
-Specifically, we introduce three additional modifications to the HVAE [93].
+While several perspectives can be used to introduce the DDPMs, we build upon the earlier discussed HVAE (Section 4.2.2) to maintain cohesiveness with the overall manuscript.
+Specifically, we describe three additional modifications to the HVAE (Luo, 2022).
 Firstly, the latent dimensionality is set equal to the spatial data dimensions, i.e.
@@
 $p(\mathbf{Y}|\mathbf{X})\geq$ $\;\mathbb{E}_{q(\mathbf{Y}_{1}|\mathbf{Y}_{0})}[\,\log p_{\phi}(\mathbf{Y}_{0}|\mathbf{y}_{1},\mathbf{X})\,]$ (31)
-$+\sum_{t=2}^{T}\mathbb{E}_{q\left(\mathbf{y}_{t}|\mathbf{Y}_{0}\right)}\big{[}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})\,||\,p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]\,\big{]}$
-$\hskip 40.0pt+\underbrace{\mathbb{E}_{q(\mathbf{y}_{T}|\mathbf{y}_{0})}\left[\log\frac{p(\mathbf{y}_{T})}{q(\mathbf{y}_{T}|\mathbf{y}_{0})}\right]}_{\approx 0}.$
+$+\sum_{t=2}^{T}\mathbb{E}_{q\left(\mathbf{Y}_{t}|\mathbf{Y}_{0}\right)}\big{[}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})\,||\,p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{X})]\,\big{]}+\underbrace{\mathbb{E}_{q(\mathbf{Y}_{T}|\mathbf{Y}_{0})}\left[\log\frac{p(\mathbf{y}_{T})}{q(\mathbf{y}_{T}|\mathbf{Y}_{0})}\right]}_{\approx 0}.$
 As denoted, the regularization term is assumed to be zero, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed.
-With the reparameterization trick [10], the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$.
+With the reparameterization trick (Kingma & Welling, 2013), the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$.
 As such, the KL divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$, depending on the parameterization.
-This is in almost all instances approximated with a U-Net [2].
-Also, it has been proposed to model Bernoulli noise instead of Gaussian [94, 95, 96, 97].
-However, most methodologies vary in the conditioning of the reverse process on the input image [98, 99, 100, 101].
-For instance, Wolleb et al.
-[98] concatenate the input image with the noised segmentation mask.
-Wu et al.
-[99] insert encoded image features to the U-Net bottleneck.
-Additionally, information on predictions $\mathbf{Y}_{t}$ at a time step $t$ is provided in the intermediate layers of the conditioning encoder.
-This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attention map and the inverse FFT.
-The procedure of applying attention on the spectral domain of the U-Net encoding has also been implemented with transformers in follow-up work [100].
-Segdiff [101] also encode both current time step and input image, but combine the extracted features by simple summation prior to inferring it through the U-Net.
-Figure 10: Illustration of the Diffusion Probabilistic Models [90]. The model learns to remove noise that has been gradually added to the input image.
-IV-C Test-time augmentation
-An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest.
-For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics.
-Hence, data augmentation [31] has been used at test time (explaining the name test-time augmentation, or TTA) for image classification, to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function [102].
-This technique has been applied to image segmentation as well [103, 104, 105, 106].
-By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$.
-Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
-V Epistemic uncertainty
-The crucial difference between epistemic and aleatoric uncertainty is that the former relates to model ignorance, while the latter reflects statistical ambiguity inherent in the data.
-Epistemic uncertainty can be further categorized into two distinct types [12].
+This is in almost all instances approximated with a U-Net (Ronneberger et al., 2015).
+[…]
+5 Epistemic uncertainty
+The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance, whereas aleatoric uncertainty reflects inherent statistical variability in the data.
+Consequently, unlike aleatoric uncertainty, epistemic uncertainty should not only be quantified but ideally also minimized.
+Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier & Waegeman, 2021).
 The first type pertains uncertainty related to the capacity of the model.
@@
 A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data.
-Both model and approximation uncertainty contribute to the epistemic uncertainty.
-Unfortunately, evaluation of the true Bayesian posterior (formulated in Equation (1)) is inhibited by the intractability of the data-likelihood in the denominator.
-Hence, extensive efforts have been taken to obtain viable approximations, such as using Mean-Field Variational Inference [8], Markov Chain Monte Carlo (MCMC) [107], Monte-Carlo Dropout [108, 109], Model Ensembling [110], Laplace approximations [111], Stochastic Gradient MCMCs [112, 113, 114], assumed density filtering [115] and expectation propagation [116, 117].
-We refer to any neural network that approximates the Bayesian posterior of the model parameters as a Bayesian Neural Network (BNN).
-The following sections treat applied methodologies to segmentation, which are usually straightforward extensions of conventional regression and classification BNN networks.
-These techniques are illustrated in Figure 11.
-V-A Variational Inference
+As discussed in Section 3, epistemic uncertainty is captured by modeling the posterior distribution defined in Equation (1).
+However, evaluating the true Bayesian posterior is impeded by the intractability of the data likelihood in the denominator.
+Hence, extensive efforts have been taken to obtain viable approximations, such as using Mean-Field Variational Inference Blundell et al.
+(2015), Markov Chain Monte Carlo (MCMC) Neal (2012), Monte-Carlo Dropout Gal & Ghahramani (2016); Kingma et al.
+(2015), Model Ensembling Lakshminarayanan et al.
+(2017), Laplace approximations Mackay (1992), Stochastic Gradient MCMCs Korattikara Balan et al.
+(2015); Springenberg et al.
+(2016); Welling & Teh (2011), assumed density filtering Hernández-Lobato & Adams (2015) and expectation propagation Hasenclever et al.
+(2017); Louizos & Welling (2016).
+We refer to any neural network that approximates the Bayesian posterior over model parameters as a Bayesian Neural Network (BNN).
+Unlike modeling the output distribution directly, sampling parameters from the posterior naturally induces spatial coherence in the output.
+As a result, existing methods are often simple extensions of conventional classification BNNs, rather than leveraging techniques tailored for segmentation tasks.
+This section discusses three methods for modeling epistemic uncertainty.
+The most commonly used approach is Variational Inference (VI) over model parameters (Section 5.1).
+Test-time augmentation, often mistakenly regarded as an aleatoric method, is addressed in Section 5.3.
+Finally, Laplace Approximations are examined in Section 5.2.
+5.1 Variational Inference
 Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
@@
 a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities.
-However, the priors can be also learned through Empirical Bayes [118].
-Furthermore, backpropagation is possible with the reparameterization trick [10] and within this context, the procedure is referred to as Bayes by Backprop (BBB) [8].
-During testing, a sample-based approach is utilized to approximate the posterior.
-Using several parameter permutations effectively enriches the hypothesis space due to the model combining effect, and can improve performance [119, 120, 121].
-V-B Monte Carlo Dropout
-Dropout is a common technique used to regularize neural networks [122] and can mimic sampling from an implicit parameter distribution $q(\tilde{\bm{\theta}}|\bm{\theta},p)$.
-This implicit distribution is defined as
-$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p),$ (34a)
-$\tilde{\bm{\theta}}$ $=\bm{\theta}\odot\mathbf{n},$ (34b)
-with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters.
-Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix [122].
-Using Monte-Carlo Dropout (MC Dropout), the random node switching is continued during test time, effectively sampling new sets of parameters.
-While seemingly arbitrary, it has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process [108].
-In this manner, such method is able to provide multi-modal estimates of the model uncertainty.
-As noted by Gal et al.
-[123], the model output variance is balanced with the weight magnitudes rather than the dropout rate $p$, which is usually optimized through grid search or simply set to $0.5$.
-Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, such that uncertainty estimates are governed by $p$.
-As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution [124, 87], to enable path-wise derivatives through $p$.
-V-C Model Ensembling
-As mentioned earlier, MC dropout effectively optimizes over a set of sparse neural networks.
-Also, this ensemble can be designed in an explicit manner.
-To this end, we define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{N}$, with $N$ representing the number of models in the ensemble.
-Then, it is relatively simple to obtain $\bm{\Theta}=\{\bm{\theta}_{n}\}_{n=1}^{N}$, which can be interpreted as samples from an approximate posterior.
-Ensembling in only the latter parts of a neural network (typically the decoder) is referred to as M-heads, i.e.
-the network has multiple outputs.
-Often, the $N$ obtained parameters are from $N$ separate training sessions.
-However, it has also been proven effective to ensemble from a single training session, by saving the parameters at multiple stages or training with different weight initializations [125, 126, 127].
+However, the priors can be also learned through Empirical Bayes (Bishop, 1995).
+Furthermore, backpropagation is possible with the reparameterization trick (Kingma & Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015).
+In addition to approximating the posterior, sampling from it yields multiple parameter permutations, effectively enriching the model’s hypothesis space, a phenomenon also referred to as model combination (Minka, 2000; Clarke, 2003; Lakshminarayanan, 2016).
+In most cases, VI is not performed explicitly.
+Instead, simpler techniques are employed, with the resulting parameter permutations hypothesized to serve as a proxy.
+We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2).
+See Figure 11 for an illustration of each method.
+See Figure 11 for an illustration of each method.
+All of these methods have been widely applied to segmentation.
+However, unlike aleatoric uncertainty modeling, their implementations are often more generic and not specifically tailored to segmentation tasks.
+Therefore, we provide only a brief overview here, followed by a more in-depth discussion in later sections related to various applications.
 (a) Variational Inference
@@
 Figure 11: Three types of techniques are visualized to sample parameters of the convolutional kernels, (a) an approximation of the parameter density can be made, (b) taking samples can be mimicked with dropout during test time, or, (c) an ensemble of $N$ different configurations can be explicitly modeled.
-A closely related concept to ensembling is known as Mixture of Experts (MoE), where each model in the ensemble (an ‘expert’) is trained on specific subsets of the data [128].
-In such settings, a gating mechanism is usually applied after combining the expert hypotheses.
-While uncommon, incorporating all predictions can also be regarded as an ensembling technique.
-VI Applications
-This section explores literature that employs uncertainty-based downstream tasks on segmentation models.
-These include estimating the segmentation mask distribution subject to observer variability (Section VI-A), model introspection (i.e.
-ability to self-assess, Section VI-B), improved generalization (Section VI-C), and reduced labeling costs using Active Learning (Section VI-D).
-VI-A Observer variability
+5.1.1 Monte Carlo Dropout
+Dropout is a common technique used to regularize neural networks (Wager et al., 2013) by randomly “switching off” nodes of the neural network.
+Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix (Wager et al., 2013).
+Furthermore, with Monte-Carlo Dropout (MC Dropout), the random node switching is continued during test time, effectively sampling new sets of parameters, mimicking samples from an implicit parameter distribution $q(\tilde{\bm{\theta}}|\bm{\theta},p)$, defined as
+$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p),$ (33a)
+$\tilde{\bm{\theta}}$ $=\bm{\theta}\odot\mathbf{n},$ (33b)
+with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters.
+It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal & Ghahramani, 2016).
+In this manner, such a method is able to provide multimodal estimates of the model uncertainty.
+As noted by Gal et al.
+(2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$, which is usually optimized through grid search or simply set to $0.5$.
+Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$.
+As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$.
+Also, a generalization of MC dropout has been proposed, where the weights of the network, rather that its hidden output, are set to zero (Mobiny et al., 2021).
+5.1.2 Ensembling
+As mentioned earlier, MC dropout effectively optimizes over a set of sparse neural networks.
+From this perspective, explicitly ensembling multiple models can also be considered as an approximation of VI (Lakshminarayanan et al., 2017).
+To this end, we define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{M}$, with $M$ representing the number of models in the ensemble.
+Then, it is relatively simple to obtain $\bm{\Theta}=\{\bm{\theta}_{m}\}_{n=1}^{M}$, which can be interpreted as samples from an approximate posterior.
+Ensembling in only the latter parts of a neural network (typically the decoder) is referred to as M-heads, i.e.
+the network has multiple outputs.
+Often, the $M$ obtained parameters are from $M$ separate training sessions.
+However, it has also been proven effective to ensemble from a single training session, by saving the parameters at multiple stages or training with different weight initializations (Dahal et al., 2020; Xie et al., 2013; Huang et al., 2017).
+A closely related concept to ensembling is known as Mixture of Experts (MoE), where each model in the ensemble (an ‘expert’) is trained on specific subsets of the data (Jacobs et al., 1991).
+In such settings, however, a gating mechanism is usually applied after combining the expert hypotheses.
+5.2 Laplace approximation
+It can be shown that the intractable posterior can be estimated as
+$p(\bm{\theta}|\mathcal{D})$ $\approx\frac{(\det\mathbf{H})^{\frac{1}{2}}}{(2\pi)^{\frac{d}{2}}}\exp\left(-\frac{1}{2}(\bm{\theta}-\bm{\theta}_{\text{MAP}})^{T}\mathbf{H}\,(\bm{\theta}-\bm{\theta}_{\text{MAP}})\right)$ (34)
+$=\mathcal{N}(\bm{\theta}|\mu=\bm{\theta}_{\text{MAP}},\Sigma=\mathbf{H}^{-1}),$
+with $\bm{\theta}_{\text{MAP}}$ being the Maximum a Posteriori solution and Hessian matrix $\mathbf{H}=-\nabla^{2}\log h(\bm{\theta})|_{\bm{\theta}_{\text{MAP}}}$, i.e.
+the second derivative of the loss function evaluated at $\bm{\theta}_{\text{MAP}}$.
+This approach, known as the Laplace Approximation (LA) (Laplace, 1774), has the merit that it can be performed post-hoc to .e.g.
+pretrained neural networks.
+However, estimating the Hessian can become infeasible, as computation scales quadratically with the model parameters.
+This is usually circumvented by treating neural network partly probabilistic and/or approximating the Hessian with a more simplified matrix structure (Martens & Grosse, 2015; Botev et al., 2017; Daxberger et al., 2021).
+5.3 Test-time augmentation
+An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest.
+For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics.
+Hence, data augmentation (Krizhevsky et al., 2012) has been used at test time (explaining the name test-time augmentation, or TTA), argued to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function (Ayhan & Berens, 2022).
+By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$.
+Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
+In segmentation literature, it has been hypothesized that TTA encapsualtes the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023).
+However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022).
+Kahl et al.
+(2024) draw parallels with BNNs and demonstrate this through decomposing the predictive uncertainty as
+$H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{t}|\mathbf{x},\bm{\theta}]}_{\text{epistemic}}+\overbrace{\mathbb{E}_{\mathbf{t}}[\,H[\mathbf{Y}|\mathbf{t},\mathbf{x},\bm{\theta}]\,]}^{{\text{aleatoric}}}.$ (35)
+where $t$ are sampled augmentations from the input space $\mathcal{T}$.Aleatoric uncertainty is captured because a perfectly trained model produces consistent predictions across augmentations of known data.
+Epistemic uncertainty arises when the mutual information between the predicted label and augmentation is greater than zero for unseen data, indicating the model has not yet learned invariance for such inputs.
+6 Applications
+This section briefly provides literature that incorporates uncertainty into downstream tasks using segmentation models.
+These tasks can be broadly categorized into four main groups: estimating the segmentation mask distribution subject to observer variability (Section 6.1), reducing labeling costs using Active Learning (Section 6.2), model introspection (i.e.
+ability to self-assess Section 6.3) and improved generalization (Section 6.4).
+For quick reference, Table 5 summarizes the types of dataset domains (e.g., medical, outdoor scenes) associated with various downstream tasks.
+6.1 Observer variability
 After observing sufficient data, the variability in the predictive distribution is often considered to be negligible and is therefore omitted.
@@
 The phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator are referred to as the intra-observer variability (see Figure 12).
+Annotators are models
 To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be regarded as models themselves.
@@
 Figure 12: Visualization of the intra-observer variability in parameter space (left) and the inter-observer variability in data space (right).
-For evaluation, a commonly used metric minimizes the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick [129].
-This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [130], denoted as
-$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
-$\hskip 10.0pt+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]$
-$\hskip 20.0pt-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (37)
-with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and a distance represented by kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the (1--IoU) or (1--Dice) score (squared by the GED).
-An alternative metric known as Hungarian Matching (HM) compares the predictions against the ground-truth labels through a cost matrix [74].
-Subsequently, the unique optimal coupling between the two sets that minimize the average cost is determined through a combinatorial optimization algorithm.
-This can be also formally denoted as finding the permutation matrix $\mathbf{P}$, subject to the objective
+Evaluation
+For evaluation, a commonly used metric minimizes the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick (Shawe-Taylor, 2004).
+This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) (Gretton et al., 2012), denoted as
+$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (36)
+[…]
 $\text{HM}(\mathcal{Y},\mathcal{\hat{Y}})=\frac{1}{N^{2}}\min_{\mathbf{P}}\text{Tr}(\mathbf{P}\mathbf{M}),$ (37)
-where the elements of $\mathbf{M}$ are $M_{i,j}=k(\mathcal{Y}_{i},\mathcal{\hat{Y}}_{j})$, and $N^{2}$ represents the number of elements within the matrix.
-For accurate comparison and evaluation of the problem, benchmarking is often constrained to publicly available multi-annotated data.
-For example, the Lung Image Database Consortium and Image Database Resource Initiative (LIDC-IDRI) [131] contains manually annotated lesions from lung patients and is one of the few datasets that has extensively been evaluated with appropriate metrics (see Table I).
-Multiple versions have been used, either with 15,096 (LIDCv1) or 12,816 (LIDCv2) patches.
-LIDCv1 is employed with a 60:20:20 split, while for LIDCv2 this is 70:15:15 for train:validation:test.
-LIDCv2 has also been used with threefold cross-validation (LIDCv2-cv), with either a 90:10 or 80:20 train:test split [63, 59, 73].
-Other less commonly employed datasets include the CityScapes [132] dataset, which contains street-view images of German cities from the perspective of a driving car, have been used to artificially create class-level label ambiguity [74, 69, 81].
-Classes are switched to an arbitrary auxiliary class with some probability $p$.
-Since the underlying probabilities are known, the empirical fraction of the classes from the model predictions can be directly compared to the ground-truth values of $p$.
-Furthermore, the QUBIQ Challenge [133] contains MRI and CT data from varying organs.
-Also, the retinal fundus images for glaucoma analysis (RIGA) [134] dataset contains delineations of optic cup and disc boundaries by 6 experienced ophthalmologists.
-TABLE I: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table adapted from [95].
-LIDCv1 LIDCv2
-Method GED16 HM-IoU16 GED16 HM-IoU16
-PU-Net [69], 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
-HPU-Net [74], 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
-PhiSeg [75], 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
-SSN [50], 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
-CAR [56], 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
-JProb. U-Net [135], 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
-PixelSeg [53], 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
-MoSE [81], 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
-AB [136], 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
-CIMD [96], 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
-CCDM [95], 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
-Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotation pairs, provided that annotators have an intrinsically associated expertise [137].
-Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome.
-Hence, it can be much simpler to directly model the empirical stochasticity in the annotations, which has been extensively explored in modalities such as Lung nodule detection in 2D [69, 75, 74, 138, 139, 68, 60, 62, 73, 50, 59], as well as 3D [79, 140, 141], Brain Tumor [142, 50, 98, 94, 96], White Matter Hyperintensities [143], Pulmonary Tumour Growth [144], Prostate [145, 76] and vascular [146], street scene [147, 74, 95], aerial imaging [101], optic cup [99, 148], abdominal multi-organ [100] and nuclei microscopy [97] segmentation.
-The overwhelming majority of research for this particular application has been executed with variants of conditional VAEs [69, 60, 59, 73, 68, 62, 74, 75, 139, 144, 149, 142, 150, 141, 146, 79, 76, 145, 140].
-More recently, the growing popularity of DDPMs is apparent in the field [95, 101, 98, 94, 96].
-Also, GAN-based approaches have been employed [56].
-VI-B Model introspection
-The uncertainties obtained from probabilistic models can provide an insight into the reliability of a model when the correlation between uncertainty and model accuracy is strong.
-This relationship has been formalized by Mukhoti et al.
-[151] through two conditional likelihoods.
+[…]
+$\text{NCC}=\frac{1}{n\sigma\hat{\sigma}}\sum_{i}(a_{i}-\mu)\cdot(\hat{b}_{i}-\hat{\mu}),$ (38)
+with $n$ being the number of pixels, and $\mu,\sigma$ and $\hat{\mu},\hat{\sigma}$ the mean and standard deviation of the true and predicted segmentations set, respectively.
+Variables $a$ and $b$ represent the uncertainty map calculated through the pixel-wise variance across the true and predicted segmentations, respectively.
+Approaches
+The most straightforward approach is to directly model the empirical stochasticity in the annotations with models that target the aleatoric uncertainty.
+Given that annotators have an intrinsically associated expertise, it can also be possible to modeling the annotator distribution.
+However, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome.
+This observation is reflected in literature, where we see that most approaches use conditional generative models to encapsulate observer variability (See Table 1).
+While conditional generative models have been successfully applied for the task at hand, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles of annotators Zepf et al.
+(2023a).
+VAE-based Probabilistic U-Net is the most popular approached followed by its hierarchical variants (HPU-Net and PHI-Seg) and SSNs.
+More recently, the growing popularity of DDPMs is apparent in the field.
+Table 1: Methods used for the application of encapsulating Observer Variability.
+Method Literature
+PixelCNN Zhang et al. (2022c)
+SSN Monteiro et al. (2020); Kahl et al. (2024); Zepf et al. (2024; 2023a); Philps et al. (2024)
+GAN Kassapis et al. (2021)
+VAE Gao et al.
+(2022); Kohl et al.
+(2019b); Valiuddin et al.
+(2021); Selvan et al.
+(2020); Bhat et al.
+(2023; 2022a); Valiuddin et al.
+(2024b); Qiu & Lui (2020); Valiuddin et al.
+(2024a); Viviers et al.
+(2023c); Schmidt et al.
+(2023); Long et al.
+(2021b); Zepf et al.
+(2023a; 2024); Hu et al.
+(2022); Viviers et al.
+(2023b); Savadikar et al.
+(2021); Philps et al.
+(2024)
+HVAE Zhang et al. (2022b); Rafael-Palou et al. (2021); Kohl et al. (2019a); Baumgartner et al. (2019); Gantenbein et al. (2020)
+DDPM Chen et al. (2023); Zbinden et al. (2023); Rahman et al. (2023)
+MC dropout Kahl et al. (2024); Philps et al. (2024)
+Ensembling Zepf et al. (2024); Hu et al. (2023)
+TTA Kahl et al. (2024)
+6.2 Active Learning
+The field of active learning (Settles, 2009; Ren et al., 2021) aims to reduce the costly annotation procedure by careful selection of the most informative samples for training (Settles, 2009).
+In this way, the required labeled data is minimized and training convergence can be accelerated due to reduced redundancy.
+Active Learning is an extensively researched topic in both traditional (Cohn et al., 1996) and deep Machine Learning (Gal et al., 2017b), including uncertainty-based approaches applied to image-segmentation pairs.
+In such methods, it is assumed that predictions with high uncertainty provide the most informative samples for training.
+Evaluation
+Selecting the most informative samples for training contains two elements.
+Firstly, the uncertainty needs to be quantified.
+This is usually the predictive entropy Kasarla et al.
+(2019); Shen et al.
+(2021); Xie et al.
+(2022) or variance Ozdemir et al.
+(2021), but in some instances the epistemic uncertainty is explicitly estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021).
+Simply stated, the reduction in posterior entropy indicates the informative of a data sample.
+In turn, this can also be formulated by the reduction in predictive entropy in the dual formulation.
+We can formally denote this as
+BALD $\equiv H[\,\bm{\theta}|\mathcal{D}\,]-\mathbb{E}_{p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})}[\,H[\,\bm{\theta}|\mathbf{y},\mathbf{x}^{*},\mathcal{D}\,]\,],$
+$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
+$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]=I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}).$ (39)
+Note the similarity between the mutual information obtained from the predictive entropy decomposition in Equation (3).
+Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e.
+a percentage of the original dataset) Yang et al.
+(2017); Sourati et al.
+(2018); Kasarla et al.
+(2019); Mahapatra et al.
+(2018); Sinha et al.
+(2019); Siddiqui et al.
+(2020); Li & Alstrøm (2020); Kim et al.
+(2021); Shen et al.
+(2021); Burmeister et al.
+(2022); Ma et al.
+(2024b).
+Kahl et al.
+(2024) argue that Active Learning should be evaluated on image-level as humans dont annotate on pxiel-level.
+Assuming a saturated performance on in-distribution data (training cycle $t_{1}$), the author propose to measure the improved relative performance, $C$, subject to a second training sample ($t_{2}$) from selected samples
+$C_{\text{total}}=C_{\text{method}}-C_{\text{random}},$ (40)
+with
+$C=\frac{P_{t_{2}}-P_{t_{1}}}{P_{t_{1}}},$ (41)
+where $C_{\text{method}}$ denotes the relative improvement achieved by the employed method, and $C_{\text{random}}$ is a correction term accounting for gains due to random sampling, and $P$ suggested to be the Dice-score, but can theoretically be any segmentation metric.
+Table 2: Methods used for the application of Active Learning
+Method Literature
+SoftMax Kasarla et al. (2019); García Rodríguez et al. (2020); Xie et al. (2022); Wu et al. (2021); Burmeister et al. (2022); Gaillochet et al. (2023)
+SSN Kahl et al. (2024)
+Adversarial VAE/ GAN Mahapatra et al. (2018); Kim et al. (2021); Sinha et al. (2019)
+Ensembling Yang et al. (2017); Nath et al. (2020); Khalili et al. (2024)
+MC dropout Kahl et al.
+(2024); Gorriz et al.
+(2017); Ozdemir et al.
+(2021); Hiasa et al.
+(2019); Li & Alstrøm (2020); Shen et al.
+(2021); Ma et al.
+(2024b); Gaillochet et al.
+(2023); Siddiqui et al.
+(2020); Sadafi et al.
+(2019)
+TTA Gaillochet et al. (2023)
+Approaches
+Since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate.
+Hence, as can be seen in Table 2, most of the literature use MC dropout of ensembles.
+Furhtermore, using the plain SoftMax activation has also been commonly used.
+Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness Ozdemir et al.
+(2021); Shen et al.
+(2021); Wu et al.
+(2021); Mahapatra et al.
+(2018); Sinha et al.
+(2019).
+Another crucial modeling element pertains the aggregation of the pixel-level uncertainties.
+Therefore, closeness to boundary Gorriz et al.
+(2017); Kasarla et al.
+(2019); Ma et al.
+(2024b) or specific regions Gaillochet et al.
+(2023); Wu et al.
+(2021); Kasarla et al.
+(2019) are additionally incorporated when combining the uncertainty values across pixels.
+Siddiqui et al.
+(2020) determine the entropy through estimating the entropy across viewpoints in multi-view datasets.
+Gaillochet et al.
+(2023) argue the superiority of selecting samples based on batch-level uncertainty.
+6.3 Model introspection
+Erroneous predictions should correlate with the uncertainty in cases the source is assumed to be epistemic.
+In turn, this implies that the provided uncertainty can be used to to gauge prediction quality of a model.
+High uncertainty predictions can be discarded to ensure a level of model reliability.
+Hence, the described problem also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021).
+Including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021).
+Note that using uncertainty to guide model training is closely related to Active Learning, which has been discussed in Section 6.2.
+Table 3: Methods used for the application of Model Introspection
+Method Literature
+VI Ng et al. (2022); LaBonte et al. (2019)
+LA Zepf et al. (2023b)
+TTA Kahl et al. (2024); Wang et al. (2019b); Whitbread & Jenkinson (2022); Roshanzamir et al. (2023); Dahal et al. (2020)
+Ensembling Jungo et al.
+(2020); Mehrtash et al.
+(2020); Czolbe et al.
+(2021); Holder & Shafique (2021); Jungo & Reyes (2019); Ng et al.
+(2022); Hann et al.
+(2021); Jungo & Reyes (2019); Linmans et al.
+(2020); Pavlitskaya et al.
+(2020); Valada et al.
+(2017)
+MC Dropout Jungo et al.
+(2020); Kahl et al.
+(2024); Whitbread & Jenkinson (2022); Kendall et al.
+(2016); Kampffmeyer et al.
+(2016); Dechesne et al.
+(2021); Morrison et al.
+(2019); Qi et al.
+(2023); Eaton-Rosen et al.
+(2018); Jungo et al.
+(2018a; b); Roy et al.
+(2018a; b); Mehrtash et al.
+(2020); Nair et al.
+(2020); Sander et al.
+(2019); Hasan & Linte (2022b); Camarasa et al.
+(2021); Roshanzamir et al.
+(2023); Sedai et al.
+(2018); Seeböck et al.
+(2019); DeVries & Taylor (2018); Czolbe et al.
+(2021); Hoebel et al.
+(2019); Bhat et al.
+(2021); Dahal et al.
+(2020); Antico et al.
+(2020); Lambert et al.
+(2022); Hasan & Linte (2022a); Jungo & Reyes (2019); Ng et al.
+(2022); Iwamoto et al.
+(2021); Hoebel et al.
+(2020)
+SSN Kahl et al. (2024); Ng et al. (2022)
+VAE Chotzoglou & Kainz (2019); Viviers et al. (2023a); Czolbe et al. (2021); Bian et al. (2020)
+Evaluation
+The relationship between uncertainty and model accuracy has been formalized by Mukhoti & Gal (2018) through the Patch Accuracy vs Patch Uncertainty (PAvPU) metric.
 Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$.
@@
 $\text{PAvPU}=\frac{n_{\text{ac}}+n_{\text{au}}}{n_{\text{ac}}+n_{\text{au}}+n_{\text{ic}}+n_{\text{iu}}}.$ (42)
-This application has been applied in combination with MC dropout to model (video) street scenes [152, 153], remote sensing [154, 155], instance segmentation of various objects [156], point-cloud indoor scenes [157], brain [158, 159, 160, 17, 161, 162, 163], Multiple Sclerosis [164], cardiac [165, 166], heart ventricle [163], prostate [163], carotid artery [167], lumbosacral [19] MRI, and Optical Coherence Tomography [168, 169], skin imaging [170, 171], lung [171] and liver CT [172] and MRI  [173], and ultrasound [125, 174].
-Concrete dropout has been applied to instance segmentation of C.
-elegans assays [175] and street-scenes [151].
-To a lesser degree, ensembling [176, 125], Variational Inference both in 2D [177, 178] and 3D [179], M-heads (auxiliary networks) [14, 180, 181], and test-time augmentation [125] have also been used to quantify the uncertainty for quality assessment.
 It can be noted that uncertainty is usually only obtained on pixel basis, while crucial information can be present in structural statistics.
-Hence, the Coefficient of Variation (CV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples.
-Also, Roy et al.
-[162] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to determine the pair-wise average overlap between all respective samples as
-$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\,\text{Dice}(t(\mathbf{y}_{i}),t(\mathbf{y}_{j}))\,\right]\quad\forall i\neq j.$ (38)
-Finally, including information on localized uncertainty to the training objective has shown to improve generalization capabilities [182, 183, 184, 185].
-Note that using uncertainty to guide model training is closely related to Active Learning, which is discussed in Section VI-D.
-VI-C Model generalization
-As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance, due to the model combining effect.
-For instance, dropout layers at the deeper decomposition levels of the SegNet [4] improves model performance [152].
-Literature also reports improved performance with Concrete Dropout for semantic (instance) segmentation [151, 175].
-This has been observed in multiple domains, including in-/outdoor scene understanding [152, 151, 186, 187, 181], brain tumor MRI [159, 188, 161, 177], Optical Coherence Tomography [168], low-dose Computed Tomography of lung nodules [189], colectoral polyps [190], cardiac MRI [18] and C.
-elegans roundworm microscopy images [175].
+Hence, Roy et al.
+(2018b) propose to use the Coefficient of Variation (CoV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples.
+The authors also evaluate structural uncertainties by assuming predictions with the pair-wise average overlap between all respective samples.
+$\overline{\mathrm{D}}=\mathbb{E}_{p_{\*Y|\*X^{*}}}\left[\,\text{Dice}(t(\*y_{i}),t(\*y_{j}))\,\right]\quad\forall i\neq j.$ (43)
+Approaches
+Considering Table 3, it can be obseverd that epistemic methods such as MC dropout and Ensemblign are by far the most popul;ar choice of uncertainty encapsulation method.
+However, Concrete dropout (Rumberger et al., 2020b; Mukhoti & Gal, 2018), M-heads (auxiliary networks) (Jungo & Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
+6.4 Model generalization
+As mentioned earlier in Section 5.1, sampling new parameter permutations often improves segmentation performance, due to the model combining effect.
+In contrast to other downstream tasks, evaluation does not require specialized metrics besides quantifying the relative change in model performance.
+For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) improves model performance (Kendall et al., 2016).In Table 4, we can see improved performance with MC Dropout and also in specific with Concrete Dropout (Mukhoti & Gal, 2018; Rumberger et al., 2020b).
+The improved generalization from ensembling has also shown to produce more calibrated outputs (Mehrtash et al., 2020).
 Ng et al.
-[18] have benchmarked multiple techniques for cardiac MRI segmentation and have found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions.
-Furthermore, the improved generalization from ensembling has also shown to produce more calibrated outputs [163].
-In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits [49].
+(2022) found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions.
+In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits (Larrazabal et al., 2021b).
+Ensembling has also been performed using M-Heads (Hu et al., 2023; Jungo & Reyes, 2019).
 Nonetheless, individual models in conventional ensembles receive data in an unstructured manner.
-However, it is also possible to assign specific subsets of the data to particular models (so-called ‘experts’) in the ensemble [128], commonly referred to as a Mixture of Experts (MoE).
-While MoE resembles ensembling in many ways, the approach additionally relies on a learnable gate that inherits the decision-taking logic.
-Pavlitskaya et al.
-[181] show merits of such approach for urban outdoor scene segmentation.
-For optic cup segmentation, Ji et al.
-[137] also condition on a normalized expertness vector, where each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net.
-Gao et al.
-[81] have introduced a Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach.
-Nevertheless, the methodology addresses aleatoric uncertainty quantification and has hence been discussed in Section IV-B2.
-VI-D Active Learning
-The field of active learning [191, 192] aims to reduce the costly annotation procedure by careful selection of unlabeled training samples.
-A wide range of methodologies exist, but since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate.
-In terms of metrics, the expected increase in posterior entropy, defined by
-$H[\,\bm{\theta}|\mathcal{D}\,]-\mathbb{E}_{p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})}[\,H[\,\bm{\theta}|\mathbf{y},\mathbf{x}^{*},\mathcal{D}\,]\,],$ (39)
-can provide a notion to describe information gain, and thus the decrease of uncertainty for specific datapoints.
-Notably, minimizing the expected increase in posterior entropy is equivalent to maximizing the mutual information between the data and model parameters.
-This enables reformulation in terms of the output space, rather than the complex parameter space and is known as Bayesian Active Learning by Disagreement (BALD) [193], denoted as
-$I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D})$ $=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
-$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,].$ (41)
-Active Learning through uncertainty quantification has been been beneficial for medical [194, 195, 196, 197, 198, 199, 200, 201], multi-view [202], remote sensing [203], street-views [204] and 3D point-cloud data [205].
-These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [185] and novel temporal-ensembling methods [206] have also been used to reduce the required human labeling.
-Some methods extend beyond pixel-level uncertainty, incorporating boundary information in the uncertainty quantification process [204, 200].
-For example, Kasarla et al.
-[204] make use of both image and pixel-level uncertainty.
-Furthermore, pixels are weighed by their closeness to edges, as boundary cases are more likely to be uncertain.
-Similarly, Ma et al.
-[200] combine target and boundary-based uncertainty sampling, thereby ensuring diversity, effective and balanced utilization of the available information.
-VII Discussion
-This section highlights the main trends and discussion points in the field.
-The challenges and limitations related to the methods (Section VII-A) and downstream applications (Section VII-B) are discussed.
-Based on these observations, recommendations for future work will be provided in Section VII-C.
-VII-A Discussion on methods
-The core challenges with modeling segmentation distributions entails obtaining calibrated uncertainty estimates and secondly, designing spatial coherence between pixels of the samples.
-While the former can be achieved with conventional CNN models, the latter requires modeling correlation across pixels of the segmentation mask.
-As discussed in earlier sections, the desired nature of the uncertainty estimates governs the required methodology.
-Similarly, our discussion will follow a comparable format as the preceding sections, i.e.
-the aleatoric and epistemic approaches will initially be treated in a separate manner.
-VII-A1 Aleatoric methods
-The literature overview indicates that three distinct routes can be taken to model the correlation.
-The first method is to model the correlation directly in the pixel space.
-The second method entails using latent-variable modeling, where the correlation is encoded in the latent space.
-Finally, TTA is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models.
-We will discuss the strengths and weaknesses of each of those approaches.
-A summary of this is presented in Table II.
-TABLE II: Comparison between models that quantify aleatoric uncertainty.
-Method Advantages Disadvantages Examples
-TTA Model agnostic, no additional training parameters Implicit likelihoods [103, 104, 105, 106]
-SSN Model agnostic, explicit likelihoods, fast sampling Unstable training [50]
-PixelCNN Explicit, exact likelihoods Sequential sampling, memory intensive [53]
-GAN Fast sampling, flexible Unstable training, poorly defined objective, implicit likelihoods [56, 57]
-VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap [69, 74, 75, 59, 60, 73, 62, 61, 141, 89, 81, 135, 68]
-DDPM Flexible, expressive Sequential sampling [98, 99, 100, 101, 94, 95, 96, 97]
+Hence, specific subsets of the data can also be assigned to particular models with MoEs (Pavlitskaya et al., 2020; Ji et al., 2021; Gao et al., 2022).
+Generative models, though typically used for other applications, have also led to performance improvements across a wide range of studies Saha et al.
+(2021a); Wolleb et al.
+(2021); Wu et al.
+(2023b); Amit et al.
+(2021); Bogensperger et al.
+(2023); Zbinden et al.
+(2023); Viviers et al.
+(2023a); Zepf et al.
+(2023a); Hu et al.
+(2022).
+Table 4: Methods used for the application of improved Model Generalization.
+Method Literature
+SoftMax Kasarla et al. (2019); García Rodríguez et al. (2020); Xie et al. (2022); Wu et al. (2021); Burmeister et al. (2022); Gaillochet et al. (2023)
+SSN Zepf et al. (2023a); Ng et al. (2022)
+VAE Viviers et al. (2023a); Zepf et al. (2023a); Hu et al. (2022)
+DDPM Wolleb et al. (2021); Wu et al. (2023b; a); Amit et al. (2021); Zbinden et al. (2023); Bogensperger et al. (2023); Hoogeboom et al. (2021)
+MC dropout Mukhoti & Gal (2018); Rumberger et al.
+(2020b); Saha et al.
+(2021a); Zhang et al.
+(2022a); Whitbread & Jenkinson (2022); Huang et al.
+(2018); Mukhoti et al.
+(2020); Ng et al.
+(2022); Wickstrøm et al.
+(2020)
+Ensembling Ng et al. (2022); Larrazabal et al. (2021b); Kamnitsas et al. (2018); Saha et al. (2021b); Hu et al. (2023); Ji et al. (2021)
+TTA Wang et al. (2019a); Rakic et al. (2024); Whitbread & Jenkinson (2022)
+6.5 Application domains
+In this Table 5, we link the discussed methodologies and applications to their respective domains.
+It is evident that the medical domain, followed by the automotive field, is the most prominent for probabilistic segmentation tasks.
+The high-stakes nature of both domains contributes to their prominence, as uncertainty-aware models are especially vital in safety-critical applications.
+Specifically, the delineation of lung nodules in CT images are ambiguous.
+The LIDC-IDRI dataset is a multi-annotated Lung CT dataset, which serves as a good benchmark for quantifying observer variability and is reflected in the table.
+Table 5: Overview that links literature to various domains and applications.
+Observer Variability Active Learning Model Introspection Model Generalization
+Outdoor scenes Kahl et al.
+(2024); Kohl et al.
+(2019b); Kassapis et al.
+(2021); Gao et al.
+(2022) Sinha et al.
+(2019); Kahl et al.
+(2024); Kasarla et al.
+(2019); Kim et al.
+(2021); Xie et al.
+(2022); Wu et al.
+(2021) Pavlitskaya et al.
+(2020); Hasan & Linte (2022b); Sander et al.
+(2019); Ng et al.
+(2022); Hasan & Linte (2022a); Mehrtash et al.
+(2020); Kendall et al.
+(2016); Kahl et al.
+(2024); Holder & Shafique (2021); Qi et al.
+(2023) Ng et al.
+(2022); Huang et al.
+(2018); Zbinden et al.
+(2023); Mukhoti & Gal (2018); Amit et al.
+(2021); Hoogeboom et al.
+(2021); Valada et al.
+(2017); Rakic et al.
+(2024)
+Indoor scenes Siddiqui et al. (2020); Wu et al. (2021) Qi et al. (2023)
+Various objects Philps et al. (2024) Kendall et al. (2016); Qi et al. (2023); Morrison et al. (2019)
+Remote sensing García Rodríguez et al. (2020) Dechesne et al. (2021; 2021); Kampffmeyer et al. (2016); Dechesne et al. (2021)
+Microscopy Kohl et al.
+(2019a); Zepf et al.
+(2023a); Schmidt et al.
+(2023); Philps et al.
+(2024) Khalili et al.
+(2024); Li & Alstrøm (2020); Yang et al.
+(2017); Sadafi et al.
+(2019) Linmans et al.
+(2020); Iwamoto et al.
+(2021) Bogensperger et al.
+(2023); Amit et al.
+(2021)
+Brain MRI Philps et al.
+(2024); Bhat et al.
+(2023); Hu et al.
+(2023); Philps et al.
+(2024); Savadikar et al.
+(2021); Zhang et al.
+(2022c) Zepf et al.
+(2023b); Ma et al.
+(2024b); Shen et al.
+(2021) Roy et al.
+(2018b); Mehrtash et al.
+(2020); Jungo & Reyes (2019); Eaton-Rosen et al.
+(2018); Wang et al.
+(2019b); Jungo et al.
+(2020); Whitbread & Jenkinson (2022); Monteiro et al.
+(2020); Lambert et al.
+(2022); Jungo et al.
+(2018b; a) Wolleb et al.
+(2021); Wu et al.
+(2023b; a); Chen et al.
+(2023); Wang et al.
+(2019a); Kamnitsas et al.
+(2018); Larrazabal et al.
+(2021b); Wang et al.
+(2019b); Whitbread & Jenkinson (2022)
+Prostate MRI Baumgartner et al.
+(2019); Hu et al.
+(2022; 2023); Zepf et al.
+(2024) Burmeister et al.
+(2022); Gaillochet et al.
+(2023) Mehrtash et al.
+(2020); Zepf et al.
+(2023b) Hu et al.
+(2022); Saha et al.
+(2021a); Rakic et al.
+(2024); Saha et al.
+(2021b)
+Various MRI Rahman et al.
+(2023) Ozdemir et al.
+(2021); Hiasa et al.
+(2019); Ng et al.
+(2022); Burmeister et al.
+(2022); Ma et al.
+(2024b); Nath et al.
+(2020); Gaillochet et al.
+(2023) Bhat et al.
+(2021); Camarasa et al.
+(2021); Nair et al.
+(2020); Roshanzamir et al.
+(2023); Hann et al.
+(2021) Ji et al.
+(2021); Ng et al.
+(2022)
+OCT imaging Selvan et al. (2020) Seeböck et al. (2019); Bian et al. (2020); Sedai et al. (2018) Rakic et al. (2024); Ji et al. (2021); Wu et al. (2023a; b)
+Dermoscopy Li & Alstrøm (2020); Gorriz et al. (2017) Zepf et al. (2023b); Jungo & Reyes (2019); DeVries & Taylor (2018); Czolbe et al. (2021) Wu et al. (2023a); Zepf et al. (2023a)
+Lung CT Rafael-Palou et al.
+(2021); Gao et al.
+(2022); Valiuddin et al.
+(2024b); Chen et al.
+(2023); Zhang et al.
+(2022b); Monteiro et al.
+(2020); Qiu & Lui (2020); Kohl et al.
+(2019a); Hu et al.
+(2022); Zhang et al.
+(2022c); Gantenbein et al.
+(2020); Viviers et al.
+(2023b; c); Valiuddin et al.
+(2024a); Zbinden et al.
+(2023); Zepf et al.
+(2024); Kahl et al.
+(2024); Bhat et al.
+(2022a); Long et al.
+(2021b); Valiuddin et al.
+(2021); Rahman et al.
+(2023); Kohl et al.
+(2019b); Hu et al.
+(2023); Bhat et al.
+(2023); Kassapis et al.
+(2021); Long et al.
+(2021a); Selvan et al.
+(2020); Baumgartner et al.
+(2019) Kahl et al.
+(2024); Hoebel et al.
+(2020) Kahl et al.
+(2024); Czolbe et al.
+(2021); Chotzoglou & Kainz (2019) Hu et al.
+(2022); Rakic et al.
+(2024); Zhang et al.
+(2022a)
+Various CT Valiuddin et al. (2024b); Hu et al. (2023; 2023) Hiasa et al. (2019) Hoebel et al. (2019); LaBonte et al. (2019; 2019) Viviers et al. (2023a)
+Various US Rahman et al. (2023) Yang et al. (2017) Antico et al. (2020); Dahal et al. (2020); Bian et al. (2020) Wu et al. (2023b; a); Rumberger et al. (2020b)
+Others Valiuddin et al. (2021) Mahapatra et al. (2018) Wickstrøm et al. (2020)
+7 Discussion
+This section builds on the reviewed theory and literature to guide researchers in selecting methods suited to their goals.
+Furthermore, it highlights key gaps and limitations, offering insights for addressing current challenges in the field and identifying promising directions for future research.
+Our main discussion points can be summarized as follows:
+  • The interpretation and quantification of epistemic and aleatoric uncertainty remain complex and warrant careful reconsideration, despite the literature often indicating a clear distinction between them.
+(Section 7.1).
+  • Spatial uncertainty introduces additional challenges, including the accurate estimation of structural entropy, the aggregation of pixel-wise uncertainties, and increased complexity in volumetric segmentation tasks (Section 7.2).
+  • Lack of standardization across the downstream tasks remains a major issue in the field, hindering fair benchmarking and limiting further progress (Section 7.3).
+  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4)..
+  • Segmentation backbones commonly used in the field tend to be outdated, with Transformer-based architectures having surpassed them in performance yet remaining underutilized in probabilistic dense predictions (Section 7.5).
+  • Aleatoric uncertainty is often modeled using conditional generative approaches; as this area advances rapidly, its developments should be promptly integrated into probabilistic segmentation research (Section 7.6).
+  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout and its ability to capture true model uncertainty (Section 7.7).
+7.1 Disentangling uncertainties
+Proper disentanglement of the uncertainties allows to accredit high-entropy outputs to either model ignorance about the data (epistemic), or inherent noise in the data generation process (aleatoric).
+Importantly, this distinction relates to the suggestion whether the uncertainty is reducible or not.
+Furthermore, Equation( 2) indicates this can be achieved by introducing stochasticity into the model parameters or the features.
+Seemingly straightforward, sufficient literature and discussion in the scientific community exist the applicability of this taxonomy in practical cases .
+As paraphrased from Kirchhof et al.
+(2025), definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another.
+We further explore this concept and show how confusion and contradiction have led to common pitfalls in related studies.
+To address these issues, we highlight key nuances in uncertainty modeling and contextualize this to segmentation problems by providing recommendations for future research.
+7.1.1 Nuances in uncertainty modeling
+Interpretation of uncertainty is task dependent
+It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context.
+For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Section 6.3 and 6.4.
+Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2).
+Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty Depeweg et al.
+(2018), such as commonly done in Active Learning (Section 6.2).
+Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk Apostolakis (1990); Helton (1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1).
+However, it was highlighted that the decision to model a specific type of uncertainty was largely influenced by how the problem was framed.
+To avoid confusion.
+this task-dependency should be noted when interpreting uncertainty.
+Interpretation of uncertainty is model dependent
+Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian & Ditlevsen, 2009).
+For instance, the expected entropy w.r.t.
+a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3).
+However, when employing generative models, this behavior changes.
+The latent (feature) stochasticity variable is employed to model the aleatoric uncertainty.
+Hence, the remaining uncertainty, i.e.
+epistemic, is encapsulated in the expected entropy w.r.t.
+this latent variable (Equation (13).
+Give this, the calibration network of the GAN-based CAR model of Kassapis et al.
+(2021) (discussed in Section 4.2.1) actually quantified the epistemic rather than aleatoric uncertainty.
+This is dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Kahl et al., 2024; Hu et al., 2022; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023).
+The key message thus far is that, after selecting the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty type.
+Uncertainties are intertwined
+Despite the seemingly rigorous and well-defined distinction, it can be seen that implementation rapidly blurs this notion.
+A systematic study has found that the ability to separate the uncertainties can strongly depend on the data Kahl et al.
+(2024).
+These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods de Jong et al.
+(2024); Mucsányi et al.
+(2024), indicating that they represent uncertainty of similar kind.
+Kendall & Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other.
+In probabilistic models, we can also find that the epistemic encapsulation methods have been used to model observer variability.
+Furthermore, using a combination of aleatoric end epistemic modeling techniques can also beneficial (Gao et al., 2022).
+These finding further highlight the notion of the uncertainties being intertwined.
+On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions Kingma et al.
+(2015), further supporting this argument.
+Encapsulating vs. quantifying
+A common misconception pertains that modeling a specific uncertainty also implies that the predicted output represents that same uncertainty.
+This undermines the entropy decomposition in Equation (3) distinguishing between encapsulating, i.e.
+targeting an intended uncertainty type, versus quantifying, i.e.
+measuring the modeled uncertainty in isolation.
+For example, the estimation of mutual information is required to express the epistemic uncertainty in isolation.
+Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties(Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018).
+At the very most, the output correlates to an increase in epistemic uncertainty.
+However, given the entanglement of both uncertainties, part of the predictive variance can be accredited to noise in the data.
+This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance Zhang et al.
+(2022c); Monteiro et al.
+(2020); Valiuddin et al.
+(2024b); Gao et al.
+(2022).
+However, epistemic uncertainty is often assumed to be explained away or assumed to be comparatively negligible with large datasets.
+Finally, we have found it to be common to interchangeably use the total and model uncertainty, especially in the context of Monte Carlo Dropout methods Kendall et al.
+(2016); Roy et al.
+(2018b); DeVries & Taylor (2018); Mehrtash et al.
+(2020); Burmeister et al.
+(2022); Iwamoto et al.
+(2021).
+To prevent confusion in already established terminology, the term model uncertainty should remain reserved for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
+7.1.2 Pragmatic use of uncertainty
+Unless being the core contribution, there seems to be no compelling need to investigate uncertainty disentanglement for the applications of uncertainty in dense prediction.
+If performance on the downstream application is improved, it can be assumed that modeled uncertainty correlates with changes in variance of the predictive distribution.
+In such cases, no claims need to be made on the uncertainty type and neither is it required to isolate the two.
+Hence, as described in previous work, aleatoric and epistemic uncertainty should be used as to communicate whether the uncertainty is reducible or not (Der Kiureghian & Ditlevsen, 2009; Faber, 2005), to then select appropriate modeling methods.
+As argued by Kahl et al.
+(2024), claiming to quantify and uncertainty type would ideally require validation across all relevant downstream tasks.
+For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but is found in literature that often either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022).
+As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often found to be evaluated on a single annotator or without the use of distribution-level statistics Wang et al.
+(2019a); Whitbread & Jenkinson (2022); Kendall & Gal (2017); Liu et al.
+(2022b); Schmidt et al.
+(2023); Savadikar et al.
+(2021).
+Hence, simply avoiding any serious claims related to either type alleviates this burden, and allows one to focus solely on the task at hand.
+In contrast to the work of Kahl et al.
+(2024), however, we do not consider calibration of models as a downstream task, but rather as a tool to quantify model reliability.
+As a consequence, the pixel-wise uncertainties can be used for possible downstream applications.
+7.2 Spatial coherence
+The inter-pixel dependency of segmentation masks often introduces additional complexity in proper uncertainty quantification, questioning the commonly repurposed techniques from conventional image-level classification.
+In this section, we further dive into import details and challenges arising due to the added dimension(s) in both 2D and 3D data.
+Aggregation methods
+Quantifying the epistemic uncertainty requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (39)).
+In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values Mukhoti & Gal (2018); Camarasa et al.
+(2021); Bhat et al.
+(2022b); Mukhoti et al.
+(2020); Ma et al.
+(2024b); Shen et al.
+(2021); Li & Alstrøm (2020); Dechesne et al.
+(2021); Huang et al.
+(2018); Nair et al.
+(2020); Zepf et al.
+(2023b).
+However, this asserts the assumption of a factorized categorical distribution, i.e.
+uncorrelated pixels, which is a strong assumption that ignores the core rationale for the field’s existence.
+Since this could explain failure cases of entropy-based uncertainty evaluation methods, it is important to shift towards contemporary generative models that enable exact likelihoods, improving the theoretical soundness of the evaluated entropies.
+Moreover, elements such as object class dependency, size, shape, proximity, contextual information etc.
+can all influence the decision-making of the aggregation method.
+Notably, Kahl et al.
+(2024) argue that pixel aggregation should be considered as a separate, distinct modeling choice in probabilistic segmentation.
+The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score.
+Fortunately, some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021).
+However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
+Volumetric uncertainty modeling
+In clinical settings, data is often acquired volumetrically but sliced into 2D patches due to memory limits.
+With improved compute, 3D segmentation models have gained traction for preserving spatial continuity and reducing inter-slice inconsistencies.
+As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue 3D segmentation to introduce or exacerbate previously mentioned issues related to entropy estimation and aggregation of pixels.
+Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023b; Chotzoglou & Kainz, 2019; Long et al., 2021b; Saha et al., 2021a; 2020; Hoebel et al., 2019).
+Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D patches are often highly redundant Shen et al.
+(2021); Burmeister et al.
+(2022).
+Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors.
+Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models.
+For example, the GED cannot be estimated volumetrically with 2D models.
+Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data.
+Fortunately, research dedicated to probabilstic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al.
+(2025); Burmeister et al.
+(2022); Viviers et al.
+(2023b).
+7.3 Standardization
+The widespread success of contemporary machine learning models can be largely attributed to open-source initiatives and rigorous benchmarking practices.
+The field of probabilistic falls short on this.
+A key factor contributing to these diverging standards is that the literature is predominantly application-driven, uncertainty often considered merely an auxiliary tool, leading to a fragmented landscape where studies rely on varied datasets and methodologies without a shared evaluation protocol.
+This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024).
+We have incorporated the suggestions of the authors in the suggested evaluation of the applications in Section 6.
+Despite the scattered nature of related works, efforts to quantify observer variability show some implicit consensus on data preprocessing.
+However, these practices occasionally diverge from the recommendations of Kahl et al.
+(2024).
+For instance, for the widely used LIDC-IDRI dataset (Armato, 2011), we recommend following the preprocessing protocol of Baumgartner et al.
+(2019), which is supported by publicly available code and has been adopted in several subsequent studies.
+Similarly, the Cityscapes dataset (Cordts et al., 2016a), comprising street-view images from German cities, has been used to introduce artificial class-level label ambiguity (Kassapis et al., 2021; Kohl et al., 2019b; Gao et al., 2022).
+We advise using the label-switching parameters proposed by Kohl et al.
+(2019b), whose preprocessing scripts are also available.
+Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), though they follow less standardized preprocessing.
+The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al.
+(2022).
+The evaluation metrics used across studies also diverge substantially.
+While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of empty segmentations (which often cause NaNs)—vary widely.
+This inconsistency is particularly evident in the work of Zbinden et al.
+(2023), who attempt to identify common ground across benchmarks but find over 10 (!) different evaluation methods.
+Additionally, many reported results are not taken from original papers but from follow-up studies using them as baselines.
+See Table 6 for a comparison of benchmarks on the LIDC-IDRI dataset.
+We recommend using 16 predictive samples as the standard for evaluating the LIDC-IDRI (v2, see experiments section of Zbinden et al.
+(2023)) dataset.
+Additionally, correctly predicted empty segmentations should be assigned full score.
+Notably, the literature often overemphasizes metric improvement, which can incentivize models to simply replicate ground-truth masks rather than capture the underlying distribution.
+This risks undermining the original goal of predicting plausible, unseen segmentations, a textbook case of Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” and is further compounded by known biases in commonly used metrics (Kohl et al., 2019a; Zepf et al., 2023a; Valiuddin et al., 2024b; a).
+Hence, practitioners should remain cautious and complement quantitative evaluations with qualitative assessments from domain experts.
+Table 6: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table extended from (Zbinden et al., 2023). Indicates evaluation results not obtained from the original work.
+LIDCv1 LIDCv2
+Method GED16 HM-IoU16 GED16 HM-IoU16
+PU-Net (Kohl et al., 2019b) $0.310\pm 0.010^{*}$ $0.552\pm 0.000^{*}$ $0.320\pm 0.030^{*}$ $0.500\pm 0.030^{*}$
+HPU-Net (Kohl et al., 2019a) $0.270\pm 0.010^{*}$ $0.530\pm 0.010^{*}$ $0.270\pm 0.010^{*}$ $0.530\pm 0.010^{*}$
+PhiSeg (Baumgartner et al., 2019) $0.262\pm 0.000^{*}$ $0.586\pm 0.000^{*}$
+SSN (Monteiro et al., 2020) $0.259\pm 0.000^{*}$ $0.558\pm 0.000^{*}$
+CAR (Kassapis et al., 2021) - - $0.264\pm 0.002^{\phantom{*}}$ $0.592\pm 0.005^{\phantom{*}}$
+JProb. U-Net (Zhang et al., 2022d) - - $0.262\pm 0.000^{\phantom{*}}$ $0.585\pm 0.000^{\phantom{*}}$
+PixelSeg (Zhang et al., 2022c) $0.243\pm 0.010^{\phantom{*}}$ $0.614\pm 0.000^{\phantom{*}}$ $0.260\pm 0.000^{\phantom{*}}$ $0.587\pm 0.010^{\phantom{*}}$
+MoSE (Gao et al., 2022) $0.218\pm 0.001^{\phantom{*}}$ $0.624\pm 0.004^{\phantom{*}}$
+AB (Chen et al., 2022) $0.213\pm 0.001^{*}$ $0.614\pm 0.001^{*}$
+CIMD (Rahman et al., 2023) $0.234\pm 0.005^{*}$ $0.587\pm 0.001^{*}$
+CCDM (Zbinden et al., 2023) $0.212\pm 0.002^{\phantom{*}}$ $0.623\pm 0.002^{\phantom{*}}$ $0.239\pm 0.003^{\phantom{*}}$ $0.598\pm 0.001^{\phantom{*}}$
+BerDiff (Chen et al., 2023) - - $0.238\pm 0.010^{\phantom{*}}$ $0.596\pm 0.000^{\phantom{*}}$
+MedSegDiff (Wu et al., 2023b) - - $0.420\pm 0.030^{*}$ $0.413\pm 0.030^{*}$
+SPU-Net (Valiuddin et al., 2024b) - - $0.327\pm 0.003^{\phantom{*}}$ $0.560\pm 0.005^{\phantom{*}}$
+7.4 Future of Active Learning
+Among the applications considered, Active Learning emerges as a promising approach to uncertainty quantification, showing true potential of uncertainty quantification by potentially reducing annotation costs while preserving or even improving model generalization.
+Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains.
+Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos.
+However, key limitations remain and require further research.
+Model sensitivity and baselines
+While deep learning-based segmentation typically involves high-dimensional data and large datasets, active learning requires a well-generalized model trained on limited data.
+This mismatch renders performance highly sensitive to budget constraints, model architecture, hyperparameters, and regularization Mittal et al.
+(2019); Kirsch (2024); Munjal et al.
+(2022); Kirsch et al.
+(2019); Gaillochet et al.
+(2023).
+Furthermore, random sampling, particularly when combined with MC Dropout, often proves to be a surprisingly strong baseline that is difficult to outperform Kahl et al.
+(2024); Mittal et al.
+(2019); Lüth et al.
+(2023); Burmeister et al.
+(2022); Siddiqui et al.
+(2020); Kim et al.
+(2021); Sinha et al.
+(2019); Nath et al.
+(2020), especially with highly imbalanced data Ma et al.
+(2024a).
+This observation can partly be attributed to the fact that MC Dropout likely fails to capture true Bayesian uncertainty (see Section 7.7), inviting exploration of other techniques.
+Sample informativeness
+Besides uncertainty, the diversity of selected samples is equally critical.
+For instance, a set of highly uncertain samples may share similar semantic content, leading the model to overfit a narrow and homogeneous region of the data distribution during early training.
+Hence, sample selection should be based on both uncertainty and diversity (Ozdemir et al., 2021; Wu et al., 2021; Jensen et al., 2019; Burmeister et al., 2022).
+Ozdemir et al.
+(2021) leverage the latent space of a VAE to select representative samples, whereas other studies use feature vectors from prediction models (Wu et al., 2021; Burmeister et al., 2022).
+Owing to their significantly greater representational capacity, modern generative models, as discussed in Section 7.6, merit experimental investigation together with segmentation specific conditioning.
+Furthermore, the sample diversity offered by such models can be effectively combined with adversarial training (Mahapatra et al., 2018; Sinha et al., 2019).
+Uncertainty calibration and aggregation
+As discussed in Section 4.1, using the predictive entropy of SoftMax-based uncertainty requires proper calibration.
+Unfortunately, this is rarely verified in practice Kasarla et al.
+(2019); García Rodríguez et al.
+(2020); Xie et al.
+(2022); Wu et al.
+(2021); Burmeister et al.
+(2022); Gaillochet et al.
+(2023).
+Furhtermore, most works rely on heuristics to aggregate the pixel-level information such as closeness to edges, boundaries or the use of super-pixels Gorriz et al.
+(2017); Kasarla et al.
+(2019); Ma et al.
+(2024b); Gaillochet et al.
+(2023); Wu et al.
+(2021); Kasarla et al.
+(2019); Bengar et al.
+(2021).
+Moreover, the mutual information is often quantified by plain summation (i.e.
+assuming pixel independence) Shen et al.
+(2021); Ma et al.
+(2024b).
+We hypothesize that these ad-hoc aggregation techniques, as mentioned in Section 7.2, are a key limiting factor to uncertainty-based Active Learning.
+The Cold-Start Problem
+Finally, the “cold start” problem Nath et al.
+(2022); Houlsby et al.
+(2014); Yuan et al.
+(2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed in most work.
+Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty.
+Addressing this gap, post-hoc techniques such as that of Sourati et al.
+(2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.
+7.5 Segmentation backbone
+Due to emphasis on the uncertainty methods, the backbone feature extractor is often an overlooked element in uncertainty modeling.
+For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023).
+Nevertheless, Vision Transformers (ViTs) (Vaswani et al., 2017) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Hümmer et al., 2023).
+With the exception of experiments in Kahl et al.
+(2024) (combinding SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone.
+The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021).
+Therefore, we recommend integrating the discussed methods into similar architectures to achieve both strong classification performance and effective ambiguity modeling.
+Also, CNNs have benefited from the recent developments in Transformers.
+For instance, “ConvNext" takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c).
+7.6 Modeling feature distributions
+In many instances, latent-variable models are employed to model the observer variability.
+To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches.
+A summary of this is presented in Table 7.
+Note that while classified PixelCNNs and SSNs are classified as ‘pixel-level’ methods, both can be phrased as a latent-variable model as well.
+ELBO suboptimality
 Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification.
 Nonetheless, the shortcomings of VAEs are well known.
-For example, such models suffer from inference suboptimality related to ELBO optimization [207, 65] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [73, 63, 68].
-This has been hypothesized to be caused by excessively strong decoders [64] and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [74], residual connections [74, 75], or deep supervision [75] are required for generalization.
+For example, such models suffer from inference suboptimality related to ELBO optimization Cremer (2018); Zhao et al.
+(2017) and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse Valiuddin et al.
+(2024a; b); Qiu & Lui (2020).
+This has been hypothesized to be caused by excessively strong decoders Chen et al.
+(2016) and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective Kohl et al.
+(2019a), residual connections Kohl et al.
+(2019a); Baumgartner et al.
+(2019), or deep supervision Baumgartner et al.
+(2019) are required for generalization.
 A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions.
-Hence, VAE-based models serve as good choice for the task given their shortcomings are sufficiently addressed and partially solved.
-Compared to the VAE-based models, the adoption of DDPMs is rather limited.
-This is regardless of the fact that it outperforms the VAE-based methods.
-Besides the point that DDPMs emerged much later in the field, their crucial limitation, similar to thee PixelCNN [53], is the tedious sequential inference procedure [92, 208, 209].
-This shortcoming is exacerbated in supervised settings, which often require validation through sampling on a separate data split.
-Regardless of these shortcomings, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g.
-blurry reconstructions of VAEs).
-Furthermore, it can be noted that the best-performing DDPM models are discrete in nature.
-While it is debatable whether shifting to categorical distributions is required for complex image generation [136], this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings.
+Limitation of sequantial modeling
+The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models.
+This is unexpected, as both models outperforms the latter.
+This is likly due to their tedious sequential inference procedure Song et al.
+(2020); Zheng et al.
+(2023); Meng et al.
+(2023); Zhang et al.
+(2022c), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split.
+Despite these shortcomings, the strengths of DDPMs should not be overlooked.
+Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions.
+Moreover, with existing work on faster sampling methods such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al.
+(2022), these models warrant further investigation.
+Discrete vs. continuous
+It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al.
+(2023); Zbinden et al.
+(2023).
+While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al.
+(2022), this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings.
 Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
-(a) LIDC-IDRI [131]
-(b) CityScapes [132].
+(a) LIDC-IDRI Armato (2011)
+(b) CityScapes Cordts et al. (2016a).
 Figure 13: Continuous vs.
-categorical forward diffusion process with cosine noise scheduling [210].
+categorical forward diffusion process with cosine noise scheduling Nichol & Dhariwal (2021).
 Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
-Implicit methods such as TTA [103, 104, 105, 106] have received some use in literature, but have been quickly surpassed by alternative methods.
-However, TTA does lead in the category of simplicity, requiring no additional mechanisms or modifications to the employed architecture.
-Furthermore, CAR [56] is a lesser used method and this is likely due to well-known training problems of GAN-based models [211].
-Due to this, GANs require additional heuristic terms in the training objective due to instability.
-For example, the training objective of CAR is a summation of four separate losses.
-SSNs [50] directly model correlations in pixel space and deem simple, fast and model-agnostic modeling.
-Notably, SSNs also suffer from training instability due to the invertibility requirement of the covariance matrix.
-This can be often simply circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular.
-In practice, nevertheless, this quick-fix solution is much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
-VII-A2 Epistemic methods
-Almost all discussed literature quantifying epistemic uncertainty employ approximations of VI.
-In fact, usage of MC Dropout dominates, mainly due it being a relatively simple, inexpensive and straightforward approach.
-Nevertheless, MC dropout has been subject to substantial criticism [212, 213, 109, 123].
-For example, it has been shown that MC dropout can assign a probability of zero to the true posterior, which can also erroneously possess multi-modality [212].
-Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate, rather than the observed data [213, 109].
-These weaknesses have been used as a basis for alternative techniques, where the dropout rate is learned [109, 123].
-Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of an already functional model.
-If accurate uncertainty quantification is critical, MC dropout should be avoided in any case.
+Others
+SSNs are a simple, fast, and model-agnostic approach for modeling uncertainty.
+However, they can suffer from training instability due to the requirement that the covariance matrix be invertible.
+To address this, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular.
+Despite these fixes, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space.
+CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017).
+GANs often require additional heuristic terms in the loss function to ensure stable training.
+CAR is no exception to this, as its objective function combines four separate losses.
+Recommendation
+We advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models.
+Generative modeling is a rapidly evolving field and any such model can be modified for segmentation through intricate conditioning.
+For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation.
+Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling Lipman et al.
+(2022).
+In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022)
+Table 7: Comparison between generative models that model aleatoric uncertainty.
+Method Advantages Disadvantages Examples
+SSN Model agnostic, explicit likelihoods, fast sampling possibly unstable, memory intensive (Monteiro et al., 2020; Kahl et al., 2024; Zepf et al., 2023a; Philps et al., 2024; Kahl et al., 2024)
+PixelCNN Explicit, exact likelihoods Sequential sampling, memory intensive Zhang et al. (2022c)
+GAN Fast sampling, flexible Unstable training, poorly defined objective, implicit likelihoods (Kassapis et al., 2021; Isola et al., 2017)
+VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap (Kohl et al., 2019b; Selvan et al., 2020; Valiuddin et al., 2021; 2024a; Qiu & Lui, 2020; Bhat et al., 2022a; 2023; Valiuddin et al., 2024b; Zhang et al., 2022b; Kohl et al., 2019a; Baumgartner et al., 2019; Hu et al., 2023; Liu et al., 2022b; Fischer et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Viviers et al., 2023b; a; Chotzoglou & Kainz, 2019; Saha et al., 2021a; b; Long et al., 2021a; Schmidt et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Zepf et al., 2024; 2023a; Hu et al., 2022)
+DDPM Flexible, expressive Sequential sampling (Chen et al., 2023; Zbinden et al., 2023; Rahman et al., 2023)
+7.7 Modeling parameter distributions
+In the previous sections, we reviewed feature modeling methods, typically used to quantify observer variability.
+Other downstream applications usually rely on epistemic uncertainty modeling, i.e.
+modeling parameter distributions, which will be discussed in this section.
+Criticism of MC Dropout
+Nearly all literature on epistemic uncertainty quantification relies on approximations of variational inference (VI).
+Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation.
+However, MC Dropout has faced substantial criticism Folgoc et al.
+(2021); Osband (2016); Kingma et al.
+(2015); Gal et al.
+(2017a).
+For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality Folgoc et al.
+(2021).
+It is also sensitive to model size and dropout rate rather than the observed data Osband (2016); Kingma et al.
+(2015).
+To address these issues, some alternatives attempt to learn the dropout rate Kingma et al.
+(2015); Gal et al.
+(2017a).
+Still, uncertainty estimates from MC Dropout should be treated as a convenient byproduct—not a reliable core feature.
+When precise uncertainty quantification is critical, such methods are generally insufficient and best avoided.
+When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary.
+In such cases, methods like ensembling and MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation.
+However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors.
+For instance, both ensembling and MC Dropout introduce a model combination effect—effectively averaging multiple networks.
+MC Dropout also resembles placing an $L_{2}$ penalty on weights.
+While these effects can improve generalization, similar gains might be achieved with simpler, more efficient regularizers.
+Perhaps it is the stochasticity itself that helps.
+Still, if performance is the main goal, uncertainty modeling is likely not the most efficient route due to its added computational cost.
+Instead, improved performance may simply be a side effect of these techniques.
+Comparative studies
 Finally, the optimal method for uncertainty quantification has yet to be determined.
-In some works, MC dropout was found to perform better than ensembling [189, 162], while in other works ensembling excels [163].
-Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout when evaluating with the PAvPU metric [151].
-All things considered, the preference for a particular methodology seems to carry a strong data-dependency [125, 14].
-Our recommendation is to experiment with both explicit VI and its approximations.
-VII-B Discussion on applications
-Up until now, theoretical insights have mostly been discussed.
-In this section, a deeper dive is taken into the applications of uncertainty quantification.
-Table III provides an overview of these applications in various domains.
-Notably, most models are employed in healthcare use cases, where ambiguity frequently arises due to the trade-off between accuracy and the incisiveness of medical diagnosis systems.
-Furthermore, in the automotive industry, sensors often operate under constraints, and objects of interest are typically at a considerable distance which induces ambiguity in the acquired images.
-It is evident that in these fields, uncertainty is primarily utilized to quantify observer variability and/or to correlate uncertainty with prediction accuracy.
-Research on improved generalization and uncertainty-based active learning is much sparser.
-In the following sections, each respective downstream task is further discussed.
-TABLE III: Overview of domains using uncertainty quantification for segmentation.
-Observer Model Active Model
-variability Introspection Learning generalization
-Lung [50, 69, 59, 73, 60] [171, 177, 189] [182] [189]
-[62, 68, 74, 75, 138]
-[140, 141, 94, 139, 79]
-Brain [50, 98, 96, 142] [163, 159, 161, 160, 177] [197] [159, 188]
-[99, 139, 144, 94] [161, 177]
-Outdoor scenes [74, 50, 147] [152, 153, 157, 151, 181] [202, 204] [151, 181, 152]
-[101, 156] [186, 187]
-Cardiovascular [165, 166, 176, 163] [194] [214]
-[125, 167, 178]
-Prostate [145, 76, 75] [163] [194]
-Prostate [138, 139]
-Eye [100, 99, 60] [168, 169] [168]
-Skin [100] [171, 172] [195, 198]
-Indoor scenes [101, 156] [152, 157] [202, 205] [152]
-Microscopy [97, 101, 74] [180, 175] [175]
-Others [141, 73, 99, 59] [172, 173, 177, 19, 174] [194, 200, 96] [190]
-[96, 139]
-VII-B1 Observer variability
-Varying delineation hypotheses present themselves as “noise” in the ground-truth of the data.
-This observer variability is often a result of viewing limitations and therefore heavily correlate to the input image.
-In turn, this explains the success of directly learning input-conditional observer variability, through explicitly modeling the likelihood distribution, rather than inferring the underlying latent parameter distribution.
-Nonetheless, this approach does have its limitations.
-For example, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles [215].
-A significant challenge in this domain is the lack of standardization, making benchmarking extremely difficult.
-The reasoning for this is twofold.
-Firstly, it is evident from Table III that a wide range of datasets are used in literature, often involving proprietary in-house data.
-This inherently makes replicating the presented results impossible.
-Secondly, consensus on data processing is also lacking on the publicly available datasets.
-For example, significant incongruity across literature is observed for the LIDC-IDRI dataset, regarding data splitting and preprocessing.
-Similarly, there is a lack of agreement on the implementation for the GED and HM metric.
-Several factors, including the kernel choice, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
-Regarding the evaluation metrics, we find that the literature often places excessive emphasis on improving them.
-This essentially forces the algorithm to predict segmentations identical to the available ground-truth masks.
-This can be problematic, as it defeats the original goal to predict plausible unseen segmentations, becoming a textbook case of Goodhart’s law: ”When a measure becomes a target, it ceases to be a good measure”.
-Furthermore, the quality of the GED evaluations have been subject to substantial criticism as well [216, 74, 63].
-Hence, practitioners should be extra vigilant and consider the importance to inquire domain-level experts for qualitative evaluation.
-Because the GED and HM are dependent on on the number of segmentation masks available, an alternative is to involve additional annotators per datapoint for more accurate evaluation, but is likely to be equally expensive.
-This aspect calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.
-VII-B2 Model introspection
-Literature that evaluates models by correlating model uncertainty with error is, comparatively, much more thorough and standardized.
-This is especially evident by the frequent use of the PAvPU metric.
-Furthermore, a majority of the available research use MC Dropout and correlate the output entropy to the prediction accuracy.
-The disadvantage of relying on MC Dropout when the uncertainty is crucial, has been discussed in Section VII-A2.
-Therefore, we recommend incorporating explicit VI at specific points in the network to achieve more accurate uncertainty quantification.
-These points can be selected based on the layers that influence the output segmentation most significantly.
-VII-B3 Model generalization
-Epistemic uncertainty quantification methods can result in improved performance, as it often serves as a surrogate for model combination.
-Therefore, uncertainty quantification can be readily available in the toolbox of deep learning researchers together with other commonly used regularization techniques.
-In this case, simple methods such as ensembling and MC dropout, are despite their criticisms relatively harmless.
-Furthermore, a parallel characterization can be made with MC Dropout and placing an $L_{2}$-norm penalty on the model weights.
-While this benefit is compelling, improved model performance can be also obtained with more computationally efficient regularizers.
-Therefore, improved performance should not be the ultimate goal of uncertainty quantification, but considered as an ancillary advantage.
-VII-B4 Active Learning
-The challenging task of Active Learning holds a significant promise and can potentially reduce the need for intensive labeling procedures, which often require specialized expertise.
-Furthermore, this kind of approach points towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate the adoption of such models in sectors requiring extensive specialization.
-Furthermore, Active Learning can accelerate privacy-centric collaboration when combined with a federated setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local models.
-However, Active Learning requires a well-generalized model using little data.
-This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and typically requires datasets of substantial size.
-Furthermore, many uncertainty-based approaches regularly extend conventional Active Learning for classification to segmentation by simply aggregating the pixel metrics together.
-For imbalanced settings, it has been shown that it can perform even worse than random selection [217].
-Therefore, additional modifications such as target, boundary or diversity awareness [200, 205], or region-based annotation [204, 218] are often required to apply Active Learning to segmentation tasks.
-VII-C Future work
-In this penultimate section, we will provide recommendations for future work.
-VII-C1 Exploring generative models
-Generative modeling, a rapidly evolving field, has successfully been employed for quantifying observer variability.
-The benefits of quickly applying its developments to segmentation problems have been evident with state-of-the-art DDPMs [90], which were initially proposed for unsupervised image generation.
-Also, literature on unsupervised VAEs have greatly benefited the PU-Net [73, 63, 68].
-Therefore, we advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models.
-In fact, any unsupervised generative model can be theoretically translated to the supervised setting through intricate conditioning and can be therefore used as a probabilistic segmentation model.
-Given this flexibility, it remains unclear why Normalizing Flows (NFs) remain underutilized for this task.
-Specifically continuous NFs, which approximate the time-dependent score function, strongly resemble DDPMs which have been extensively used for segmentation problems.
-Furthermore, NFs enable explicit and exact evaluation of the likelihood, which can aid in further interpretation of model predictions.
-Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian, so that inference times are therefore much faster.
-However, it should be noted that the invertibility constraint greatly hinders the expressiveness of intermediate functions, leading to memory-intensive architectures.
-VII-C2 Variational Inference and beyond
+In some works, MC dropout was found to perform better than ensembling Hoebel et al.
+(2020); Roy et al.
+(2018b), while in other works ensembling excels Mehrtash et al.
+(2020); Ng et al.
+(2022).
+Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout Mukhoti & Gal (2018).
+All things considered, the preference for a particular methodology seems to carry a strong data-dependency Dahal et al.
+(2020); Jungo & Reyes (2019).
 A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets.
 In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e.
-why ensembling or MC Dropout is often preferred over explicit VI).
-Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section V-C) for language modeling [219].
+why ensembling or MC Dropout is often preferred).
+Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section LABEL:subsec:ME) for language modeling (He, 2024).
 Future work should also experiment with this.
+Variational Inference Alternatives
 Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior.
@@
 Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions.
-To the best of our knowledge, these approaches have not been studied within the context of uncertainty quantification in segmentation.
-VII-C3 Single-pass uncertainty
+Zepf et al.
+(2023b) explored the impact of the Laplace Approximation on segmentation networks; however, beyond this work, the method has received limited attention in the field.
+Sampling-free uncertainty
 The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs.
-Hence, considerable efforts have been made towards deterministic uncertainty models [220, 221, 222], which only depend on a single forward pass.
+Hence, considerable efforts have been made towards sampling-free uncertainty models (Mukhoti et al., 2023; Liu et al., 2020; Van Amersfoort et al., 2020; Postels et al., 2019), only depending on a single forward pass.
 Mukhoti et al.
-[220] show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling.
-This approach achieves faster computation, while also providing both epistemic and aleatoric uncertainty.
+(2023) show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling.
 Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass.
-This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) [223].
+This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) (Dempster, 1967).
 While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e.
 the set of all possible outcomes.
-The use of Evidential Deep Learning has seen success in conventional classification problems [224], and Ancha et al.
-[225] recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model.
-Unfortunately, further research has not been published beyond this.
-VII-C4 Improved CNN architectures
-With the introduction of Vision Transformers [226], CNN-based models have been challenged in the task of semantic segmentation [227, 228, 229].
-Regardless the success of Vision Transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture [39].
-This is mainly because CNNs already possess desirable inductive biases, while Transformers conversely require extensive pretraining with large datasets [230].
-Nonetheless, CNNs have benefited from the recent developments in Transformers.
-For instance, “ConvNext” takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains [231].
-Since many innovations in this field focus on the technique of uncertainty quantification, the applied backbones receive less attention and might gradually become outdated.
-Our recommendation suggests improving current models with developments in generic CNN-based architectures.
-VII-C5 Distributing-free modeling
-A distribution-free framework known as Conformal Deep Learning, produces prediction sets that guarantee to contain the ground truth with a user-defined probability.
-With the help of an additional calibration set, a heuristic notion of ambiguity (i.e.
-miscalibrated softmax outputs) is transformed to rigorous uncertainty.
-This technique is especially renown for being model-agnostic, simple and highly flexible [232].
-Quite recently, Conformal Prediction has been applied to segmentation problems [233, 234, 235, 236], thereby enjoying the aforementioned benefits and showing an increased traction of this framework.
-We recommend further research in this direction to discover novel applications and benchmark against current architectures.
-VII-C6 Volumetric segmentation
-In many clinical datasets, the volumetric data are in most instances sliced to patches and processed with conventional 2D-based CNN-models.
-However, such models are often straightforward extensions of existing 2D models (and almost exclusively VAE-based), rarely addressing novel challenges introduced with additional dimensionalities.
-For example, 3D extensions of the PU-Net simply remain to use similar techniques for inserting latent samples in the decoding networks.
-Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization [179].
-Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to implement models for volumetric segmentations problems.
-Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models.
-For example, the GED cannot be estimated volumetrically with 2D models.
-Moreover, translating the already computation-intensive models (e.g.
-based on DDPMs or PixelCNNs) to 3D is another challenge, due to the memory requirement involved with volumetric data, which will increase the training load and inference time even further.
-Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underline the need of more investigation related to 3D uncertainty quantification, that more closely aligns with real-world clinical practice and therefore encourages faster adoption.
-VIII Conclusion
-Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability in their predictions.
-With the vast body of literature, encompassing diverse applications and modalities, the need for a comprehensive and systematic overview of the field is addressed.
-We present clear definitions and unified notation of methodologies that attempt uncertainty modeling, by considering the field from a theoretical perspective.
-Aleatoric uncertainty can be modeled in pixel or latent space with generative models, or implicitly be expressed with test-time augmentation.
-Epistemic uncertainty is captured with Variational Inference on the parameter distribution, or approximated with Monte Carlo Dropout or model ensembling.
-Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks, which enable to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and to the real-world applications.
-Our future recommendations pertain aligning the field with advancements in general generative modeling and Deep Learning architectures.
-Furthermore, the adoption of deterministic uncertainty quantification methods that avoid multiple forward passes, such as Conformal and Evidential Deep Learning, are suggested.
-Particularly the latter approach is interesting due to its ability to incorporate both uncertainty types.
-Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques, as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation, is a beneficial contribution to the field.
-Finally, due to the clinical relevance of uncertainty in semantic segmentation, more attention is advised to models catering to volumetric data.
-In this manner, the review paper guides researchers on the topic of probabilistic segmentation and suggests future endeavors within the lightning-fast evolving field of Deep Learning-based Computer Vision.
+The use of Evidential Deep Learning has seen success in conventional classification problems (Sensoy et al., 2018), and Ancha et al.
+(2024) recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model.
+However, follow-up research in this direction remains limited to date.
+Reliable uncertainty estimates
+A distribution-free framework known as Conformal Prediction (CP) produces prediction sets that guarantee inclusion of the ground truth with a user-specified probability.
+By using an additional calibration set, CP transforms heuristic model confidence scores into rigorous uncertainty estimates.
+This technique is particularly renowned for being model-agnostic, simple, and highly flexible (Angelopoulos & Bates, 2021).
+Recently, CP has been applied to segmentation tasks (Wieslander et al., 2020; Brunekreef et al., 2024; Mossina et al., 2024; Wundram et al., 2024a), benefiting from these advantages and reflecting the growing interest in the framework for structured prediction.
+We encourage further research to explore novel applications and benchmark CP against existing architectures.
+8 Conclusion
+Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability of their predictions.
+Given the vast and diverse body of related literature spanning various applications and modalities, often scattered and marked by inconsistent evaluation and benchmarking practices, there is a clear need for a comprehensive and systematic overview.
+This work addresses that need by presenting clear definitions and a unified notation for methodologies in uncertainty modeling.
+Furthermore, confusion and misunderstanding around uncertainty modeling and quantification have led to diverging beliefs within the field.
+By structuring the methodologies by targeted uncertainty type and followed by downstream pipeline into four distinct tasks, it becomes easier for researchers to navigate the field and build upon existing work.
+We place particular emphasis on addressing key conceptual gaps to support both pragmatic and theoretically sound research.
+We also highlight the potential and broader impact of advancing uncertainty-based active learning.
+A noticeable trend is the adaptation of generative modeling advances to segmentation through novel, task-specific conditioning strategies for modeling ambiguity in segmentation masks.
+While such efforts are already underway, it remains important to closely follow progress in general generative machine learning and continue exploring its effective application to segmentation.
+Finally, we recommend future work to focus on improving spatial aggregation methods, standardization in benchmarking, backbone architectures, as well as the adoption of novel uncertainty estimation techniques, as many commonly used methods lack rigor.
+In particular, there is growing interest in distribution- and sampling-free approaches.
+In this manner, this review paper aims to guide researchers toward building more reliable, efficient, and uncertainty-aware segmentation models within the lightning-fast evolving field of Deep Learning-based Computer Vision.

### v5

@@
 Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making.
-The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented, marked by inconsistent terminology, evaluation practices, and theoretical ambiguities.
+The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented.
 In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising.
@@
 Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability.
-Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, without being properly contextualized in segmentation.
+Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, and can be insufficiently contextualized within the domain of segmentation.
 This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions.
@@
 This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks.
-Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kahl et al., 2024; Kirchhof et al., 2025), we identify a notable absence of a comprehensive, theory-informed overview in the field.
+Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kirchhof et al., 2025), we note that a comprehensive overview in the field remains limited.
 Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023).
@@
 Past work with significant impact in general image segmentation is presented in Section 2.
-Then, the theoretical framework and notation that govern the remainder of the paper is introduced in Section 3.
+Then, the theoretical framework and notation that govern the remainder of the paper are introduced in Section 3.
 The theory enables us to classify the approaches into two distinct but often related methods.
-That is, aleatoric and epistemic uncertainty modeling methods.
-The role of these concepts in image segmentation will be treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field.
+That is, methods targeting either aleatoric or epistemic uncertainty, which can be distinguished by either modeling the feature or parameter distribution(s).
+The role of these concepts in image segmentation are treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field.
 We then correlate these modeling methods to specific applications in Section 6.
 Here, we also observe for which domains uncertainty quantification has been exploited.
-Then, this overview will be discussed further in Section 7.
+Then, this overview is discussed further in Section 7.
 This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task.
@@
 (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps.
-To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain.
-In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
+To this day, the U-Net is still often utilized as the default backbone model for many semantic segmentation and even general image generation architectures (Ho et al., 2020; Song et al., 2020), particularly in the medical domain.
+In fact, reports of recent research indicate that the relatively simple U-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
 3 Probabilistic Image Segmentation
-Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image.
+Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{Z}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then instance $\mathbf{y}$ can be considered as the ground-truth of a $K$-class segmentation task and instance $\mathbf{x}$ as the query image.
 The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively.
@@
 $p(\bm{\theta}|\mathbf{y},\mathbf{x})=\frac{p(\mathbf{y}|\mathbf{x},\bm{\theta})p(\bm{\theta})}{p(\mathbf{y|x})},$ (1)
-where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence).
+assuming independence of $\mathbf{x}$ from $\bm{\theta}$, and where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence).
 After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
@@
 Furthermore, disentangling these uncertainties can be achieved through its decomposition
-$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}}.$ (3)
+$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}},$ (3)
+where $I$ represents the mutual information.
 In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process.
@@
 This approximation improves as the training data increases and the model parameter variances approach zero.
-As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Bayesian Maximum A Posteriori (MAP) estimation with
+As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Maximum A Posteriori (MAP) estimation with
 $\bm{\theta}_{\text{MAP}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (5)
-For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007).
-Additionally, the output of such models can be interpreted as a probability distribution.
-For instance, consider spatial image dimensions $D$ and a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$.
-Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
+For example, assuming Gaussian or Laplacian priors leads to regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007).
+To model $p(\mathbf{y}|\mathbf{x},\bm{\theta})$, we make use of function $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ that infers the parameters of a Probability Mass Function (PMF).
+For instance, consider spatial image dimensions $D$ and a CNN with $\mathbf{a}=f_{\theta}(\mathbf{x})$.
+Then, we can write
 $p(\mathbf{Y}=k\,|\,\mathbf{x}^{*},\bm{\theta})=\frac{e^{\mathbf{a}_{k}}}{\sum_{k}e^{\mathbf{a}_{k}}},$ (6)
@@
 (c) However, the coherence of the segmentation suffers when sampling.
-4 Aleatoric uncertainty
+4 Feature modeling
 […]
@@
 The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
-In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$.
-For example, it is possible to model a distribution parameterized by the output of a CNN (Section 4.1).
-Also, conditional generative models are used to learn the data distribution through latent variables.
-For the purpose of taxonomy, we categorize models based on the location of these latent (unobserved) variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2.
+In most practical scenarios, aleatoric uncertainty modeling methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$.
+For example, it is possible to model a distribution parameterized by the output of a CNN.
+Also, conditional generative models are used to learn the data distribution through so-called “latent” (i.e.
+unobserved) variables.
+For the purpose of taxonomy, we categorize models based on the location of these latent variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2.
 While their theoretical formulations are largely similar, their practical implications can differ significantly.
@@
 Figure 4: Visualization of a reliability diagram, which illustrates whether the model’s predicted confidences align with the observed empirical accuracies.
-As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution.
-However, these values can only be interpreted as probabilities after proper validation, which is referred to as model calibration.
+As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution, but rerely reflect the true probabilities in modern neural networks Guo et al.
+(2017).
+Thus, interpreting the confidences as true probabilities is often only justified after proper validation, which is referred to as model calibration.
 Here, it is measured whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e.
@@
 Dotted elements appear during test-time sampling.
-Assuming pixel independence inhibits spatially coherent uncertainty quantification, which is a key factor for segmentation.
+Assuming pixel independence inhibits spatial coherence in uncertainty quantification, which is a key factor for segmentation.
 Introducing spatial correlation can be achieved through an autoregressive approach.
@@
 Monteiro et al.
-(2020) propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
+(2020) propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$.
+Given the output features $f_{\bm{\theta}}(\mathbf{x})=\mathbf{a}$ of a deterministic model, we can denote the logits distribution as
 $p(\mathbf{a}|\mathbf{x},\bm{\theta})=\mathcal{N}(\mathbf{a};\bm{\mu}=f^{\bm{\mu}}_{\bm{\theta}}(\mathbf{x}),\bm{\Sigma}=f^{\bm{\Sigma}}_{\bm{\theta}}(\mathbf{x})),$ (11)
-where the covariance matrix has a low-rank structure $\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}$, with $\mathbf{P}$ having dimensionality $((K\times D)\times R)$, $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ representing a diagonal matrix.
-The low-rank assumption results in a more structured and expressive distribution, while retaining reasonable efficiency.
-Since, the SoftMax transform pertains an intractable integral, Monte Carlo sampling is employed.
-SSNs can theoretically be augmented as an additional layer to any pretrained CNN (see Figure 6).
+where the covariance matrix has a low-rank structure $\bm{\Sigma}=\mathbf{P}\mathbf{P}^{T}+\bm{\Lambda}$, with $\mathbf{P}$ having dimensionality $((K\times D)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ representing a diagonal matrix.
+The low-rank assumption results in a more structured distribution, while retaining reasonable efficiency.
+Monte Carlo sampling is used to generate predictions, which are then mapped to categorical values using the SoftMax activation.
+SSNs can theoretically be augmented to any pretrained CNNs as an additional layer (see Figure 6).
 4.2 Latent-level sampling
-The complexity of modeling distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
+The limitation of modeling intricate, complex distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
 $p_{\bm{\theta},\bm{\psi}}(\mathbf{Y}|\mathbf{X})=\int p_{\bm{\theta}}(\mathbf{Y}|\mathbf{z},\mathbf{X})p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\mathrm{d}\mathbf{z},$ (12)
@@
 $H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{z}|\mathbf{x},\bm{\theta}]}_{\text{aleatoric}}+\overbrace{\mathbb{E}_{p(\mathbf{z})}[\,H[\mathbf{Y}|\mathbf{z},\mathbf{x},\bm{\theta}]\,]}^{{\text{epistemic}}}.$ (13)
-In contrast, the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty.
-The key overarching contribution of these methods lies in the conditioning of generative models on the input image.
+In contrast to Equation (3), the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty.
+Also, note that the mutual information is between the output and latent variables, rather than the parameters of the model.
+The key overarching contribution of latent-level sampling methods lies in the conditioning of generative models on the input image.
 This section briefly introduces Generative Adversarial Networks (Section 4.2.1), Variational Autoencoders (Section 4.2.2), and Denoising Diffusion Probabilistic Models (Section 4.2.3), and outlines how these architectures are commonly adapted for probabilistic segmentation.
@@
 Furthermore, the refinement network is subject to the adversarial objective
-$\mathcal{L}_{\text{adv}}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (16)
+$\mathcal{L}_{\text{adv}}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x})],$ (16)
 which is argued to elicit superior structural qualities compared to relying solely on cross-entropy loss. At the same time, the discriminator opposes the optimization with
@@
 the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, obtained by maximizing the conditional Evidence Lower Bound (ELBO)
-$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p(\mathbf{Y}|\mathbf{X})\,]$
-$=\mathbb{E}_{q{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{p(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$
-$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,].$ (20)
+$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\log\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$ (20a)
+$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$ (20b)
+$=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,],$ (20c)
+where Jensen’s inequality justifies moving the expectation outside the logarithm.
 The first term in Equation (20) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$.
@@
 […]
-$\log p(\mathbf{x})=\log p_{\*Z}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (22)
-[…]
+$\log p(\mathbf{x})=\log p_{\mathbf{Z}}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (22)
+where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are intermediate variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$.
+Equation (22) can be substituted in the conditional ELBO objective in Equation (20) to obtain
 $p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $\geq\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]$
-$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}|\mathbf{X})\,]-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (23)
+$\hskip 10.0pt-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})||q_{\psi}(\mathbf{z}_{k}|\mathbf{X})\,]-\mathbb{E}_{q_{\theta}(\mathbf{z}_{0}|\mathbf{Y},\mathbf{X})}\left[\,\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|\,\right],$ (23)
 where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF.
@@
 $\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (24)
-$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
+$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
 The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables, i.e.
 the aleatoric uncertainty term of the predictive entropy in Equation (13).
-This can be rewritten into the objective
+This enables the introduction of the updated objective
 $\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (25)
@@
 Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick (Jang et al., 2016; Huijben et al., 2022).
-A similar architecture known as PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net.
+Furthermore, PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net.
 However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers.
@@
 […]
-5 Epistemic uncertainty
-The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance, whereas aleatoric uncertainty reflects inherent statistical variability in the data.
+5 Parameter modeling
+The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance (i.e.
+related to the parameters), whereas aleatoric uncertainty reflects inherent statistical variability in the data.
 Consequently, unlike aleatoric uncertainty, epistemic uncertainty should not only be quantified but ideally also minimized.
@@
 5.1 Variational Inference
-Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
+Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$ (e.g.
+a gaussian with $\bm{\eta}=\{\mu,\sigma\}$), to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
 Then, we can employ Variational Inference (VI) w.r.t.
 to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
-$\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathbf{y},\mathbf{x})\,\right]$
+$\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathcal{D})\,\right]$
 $=\operatorname*{arg\,min}_{\bm{\theta}}\int q(\bm{\theta}|\bm{\eta})\log\frac{q(\bm{\theta}|\bm{\eta})}{p(\mathbf{Y}|\mathbf{x},\bm{\theta})p(\mathbf{x},\bm{\theta})}d\bm{\theta}$
@@
 where the parameter-independent terms are constant and therefore excluded from notation.
+In the case of a deterministic encoder, i.e.
+$q(\bm{\theta}|\bm{\eta})=\delta(\bm{\theta}-\bm{\theta}^{*})$, the formulation collapses to the MAP estimate in Equation (5).
 A popular choice for the approximated variational posterior is the Gaussian distribution, i.e.
@@
 We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2).
-See Figure 11 for an illustration of each method.
 See Figure 11 for an illustration of each method.
@@
 As noted by Gal et al.
-(2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$, which is usually optimized through grid search or simply set to $0.5$.
-Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$.
+(2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$.
+The dropout rate is usually optimized through grid search or simply set to $p=0.5$.
+The relationship between $p$ and the magnitude of the model weights has also be exploited to probabilistically prune neural networks Gonzalez-Carabarin et al.
+(2022).
+Gal et al.
+(2017a) propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$.
 As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$.
@@
 5.2 Laplace approximation
-It can be shown that the intractable posterior can be estimated as
+It can be shown that the intractable posterior in Equation (2) can be estimated as
 $p(\bm{\theta}|\mathcal{D})$ $\approx\frac{(\det\mathbf{H})^{\frac{1}{2}}}{(2\pi)^{\frac{d}{2}}}\exp\left(-\frac{1}{2}(\bm{\theta}-\bm{\theta}_{\text{MAP}})^{T}\mathbf{H}\,(\bm{\theta}-\bm{\theta}_{\text{MAP}})\right)$ (34)
@@
 Kahl et al.
-(2024) draw parallels with BNNs and demonstrate this through decomposing the predictive uncertainty as
+(2024) draw parallels with BNNs and demonstrate this through the (by now) familiar decomposition of the predictive uncertainty
 $H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{t}|\mathbf{x},\bm{\theta}]}_{\text{epistemic}}+\overbrace{\mathbb{E}_{\mathbf{t}}[\,H[\mathbf{Y}|\mathbf{t},\mathbf{x},\bm{\theta}]\,]}^{{\text{aleatoric}}}.$ (35)
-where $t$ are sampled augmentations from the input space $\mathcal{T}$.Aleatoric uncertainty is captured because a perfectly trained model produces consistent predictions across augmentations of known data.
+where $t$ are sampled augmentations from the input space $\mathcal{T}$.
+Aleatoric uncertainty is captured because a perfectly trained model produces consistent predictions across augmentations of known data.
 Epistemic uncertainty arises when the mutual information between the predicted label and augmentation is greater than zero for unseen data, indicating the model has not yet learned invariance for such inputs.
@@
 HVAE Zhang et al. (2022b); Rafael-Palou et al. (2021); Kohl et al. (2019a); Baumgartner et al. (2019); Gantenbein et al. (2020)
-DDPM Chen et al. (2023); Zbinden et al. (2023); Rahman et al. (2023)
+DDPM Chen et al. (2023a); Zbinden et al. (2023); Rahman et al. (2023)
 MC dropout Kahl et al. (2024); Philps et al. (2024)
@@
 (2021), but in some instances the epistemic uncertainty is explicitly estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021).
-Simply stated, the reduction in posterior entropy indicates the informative of a data sample.
+Simply stated, the reduction in posterior entropy indicates the informativeness of a data sample.
 In turn, this can also be formulated by the reduction in predictive entropy in the dual formulation.
@@
 BALD $\equiv H[\,\bm{\theta}|\mathcal{D}\,]-\mathbb{E}_{p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})}[\,H[\,\bm{\theta}|\mathbf{y},\mathbf{x}^{*},\mathcal{D}\,]\,],$
-$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
+$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}],$
 $=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]=I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}).$ (39)
-Note the similarity between the mutual information obtained from the predictive entropy decomposition in Equation (3).
+Note the similarity between the mutual information $I$ obtained from the predictive entropy decomposition in Equation (3), where both are the objective of epistemic uncertainty modeling.
 Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e.
@@
 Hence, as can be seen in Table 2, most of the literature use MC dropout of ensembles.
-Furhtermore, using the plain SoftMax activation has also been commonly used.
+Furthermore, using the plain SoftMax activation has also been commonly used.
 Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness Ozdemir et al.
@@
 High uncertainty predictions can be discarded to ensure a level of model reliability.
-Hence, the described problem also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021).
+Hence, the described problem can also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021).
 Including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021).
@@
 Approaches
-Considering Table 3, it can be obseverd that epistemic methods such as MC dropout and Ensemblign are by far the most popul;ar choice of uncertainty encapsulation method.
+Considering Table 3, it can be observed that epistemic methods such as MC dropout and Ensembles are by far the most popular choice of uncertainty encapsulation method.
 However, Concrete dropout (Rumberger et al., 2020b; Mukhoti & Gal, 2018), M-heads (auxiliary networks) (Jungo & Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
@@
 In contrast to other downstream tasks, evaluation does not require specialized metrics besides quantifying the relative change in model performance.
-For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) improves model performance (Kendall et al., 2016).In Table 4, we can see improved performance with MC Dropout and also in specific with Concrete Dropout (Mukhoti & Gal, 2018; Rumberger et al., 2020b).
+For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) improves model performance (Kendall et al., 2016).
+In Table 4, we can see improved performance with MC Dropout and also in specific with Concrete Dropout (Mukhoti & Gal, 2018; Rumberger et al., 2020b).
 The improved generalization from ensembling has also shown to produce more calibrated outputs (Mehrtash et al., 2020).
@@
 (2023b; a); Chen et al.
-(2023); Wang et al.
+(2023a); Wang et al.
 (2019a); Kamnitsas et al.
@@
 (2024b); Chen et al.
-(2023); Zhang et al.
+(2023a); Zhang et al.
 (2022b); Monteiro et al.
@@
   • Lack of standardization across the downstream tasks remains a major issue in the field, hindering fair benchmarking and limiting further progress (Section 7.3).
-  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4)..
+  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4).
   • Segmentation backbones commonly used in the field tend to be outdated, with Transformer-based architectures having surpassed them in performance yet remaining underutilized in probabilistic dense predictions (Section 7.5).
   • Aleatoric uncertainty is often modeled using conditional generative approaches; as this area advances rapidly, its developments should be promptly integrated into probabilistic segmentation research (Section 7.6).
-  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout and its ability to capture true model uncertainty (Section 7.7).
+  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout, questioning its ability to capture true model uncertainty (Section 7.7).
 7.1 Disentangling uncertainties
@@
 Furthermore, Equation( 2) indicates this can be achieved by introducing stochasticity into the model parameters or the features.
-Seemingly straightforward, sufficient literature and discussion in the scientific community exist the applicability of this taxonomy in practical cases .
+Seemingly straightforward, sufficient literature and discussion in the scientific community exist related to the applicability of this taxonomy in practical cases.
 As paraphrased from Kirchhof et al.
-(2025), definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another.
+(2025), “definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another”.
 We further explore this concept and show how confusion and contradiction have led to common pitfalls in related studies.
@@
 It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context.
-For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Section 6.3 and 6.4.
+For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Sections 6.3 and 6.4).
 Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2).
@@
 However, it was highlighted that the decision to model a specific type of uncertainty was largely influenced by how the problem was framed.
-To avoid confusion.
-this task-dependency should be noted when interpreting uncertainty.
+To avoid confusion, this task-dependency should be noted when interpreting uncertainty.
 Interpretation of uncertainty is model dependent
@@
 For instance, the expected entropy w.r.t.
-a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3).
+a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3)).
 However, when employing generative models, this behavior changes.
@@
 epistemic, is encapsulated in the expected entropy w.r.t.
-this latent variable (Equation (13).
-Give this, the calibration network of the GAN-based CAR model of Kassapis et al.
+this latent variable (Equation (13)).
+Hence, the calibration network of the GAN-based CAR model of Kassapis et al.
 (2021) (discussed in Section 4.2.1) actually quantified the epistemic rather than aleatoric uncertainty.
@@
 Furthermore, using a combination of aleatoric end epistemic modeling techniques can also beneficial (Gao et al., 2022).
-These finding further highlight the notion of the uncertainties being intertwined.
+These findings further highlight the notion of the uncertainties being intertwined.
 On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions Kingma et al.
@@
 Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties (Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018).
-At the very most, the output correlates to an increase in epistemic uncertainty.
+At the very most, the output correlates with an increase in epistemic uncertainty.
 However, given the entanglement of both uncertainties, part of the predictive variance can be accredited to noise in the data.
@@
 (2021).
-To prevent confusion in already established terminology, the term model uncertainty should remain reserved for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
+To prevent confusion in already established terminology, the term model uncertainty should be exclusively used for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
 7.1.2 Pragmatic use of uncertainty
@@
 As argued by Kahl et al.
-(2024), claiming to quantify and uncertainty type would ideally require validation across all relevant downstream tasks.
-For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but is found in literature that often either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022).
+(2024), claiming to quantify an uncertainty type would ideally require validation across all relevant applications.
+For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but it is often found in literature that either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022).
 As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often found to be evaluated on a single annotator or without the use of distribution-level statistics Wang et al.
@@
 In contrast to the work of Kahl et al.
-(2024), however, we do not consider calibration of models as a downstream task, but rather as a tool to quantify model reliability.
-As a consequence, the pixel-wise uncertainties can be used for possible downstream applications.
+(2024), however, we do not consider calibration of models itself as a downstream task, but rather as a tool to gauge model reliability to be later used for other tasks.
+As a consequence, the pixel-wise uncertainties can be later used for possible downstream applications.
 7.2 Spatial coherence
@@
 The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score.
-Fortunately, some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021).
+Some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021).
 However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
 Volumetric uncertainty modeling
-In clinical settings, data is often acquired volumetrically but sliced into 2D patches due to memory limits.
+In clinical settings, data is often acquired volumetrically but sliced into patches due to memory limits.
 With improved compute, 3D segmentation models have gained traction for preserving spatial continuity and reducing inter-slice inconsistencies.
-As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue 3D segmentation to introduce or exacerbate previously mentioned issues related to entropy estimation and aggregation of pixels.
+As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue that 3D segmentation introduces or exacerbates previously mentioned issues related to entropy estimation and aggregation of pixels.
 Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023b; Chotzoglou & Kainz, 2019; Long et al., 2021b; Saha et al., 2021a; 2020; Hoebel et al., 2019).
-Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D patches are often highly redundant Shen et al.
+Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D volumes are often highly redundant Shen et al.
 (2021); Burmeister et al.
@@
 Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors.
-Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models.
+Furthermore, methodologies need to be developed in order to appropriately compare 2D- to 3D models.
 For example, the GED cannot be estimated volumetrically with 2D models.
 Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data.
-Fortunately, research dedicated to probabilstic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al.
+Fortunately, research dedicated to probabilistic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al.
 (2025); Burmeister et al.
@@
 The widespread success of contemporary machine learning models can be largely attributed to open-source initiatives and rigorous benchmarking practices.
-The field of probabilistic falls short on this.
+The field of probabilistic segmentation falls short on this.
 A key factor contributing to these diverging standards is that the literature is predominantly application-driven, uncertainty often considered merely an auxiliary tool, leading to a fragmented landscape where studies rely on varied datasets and methodologies without a shared evaluation protocol.
 This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024).
-We have incorporated the suggestions of the authors in the suggested evaluation of the applications in Section 6.
-Despite the scattered nature of related works, efforts to quantify observer variability show some implicit consensus on data preprocessing.
-However, these practices occasionally diverge from the recommendations of Kahl et al.
-(2024).
+These practices are included in our recommended practices, as discussed in Section 6 per application.
+Despite the aforementioned points, efforts to quantify observer variability show some implicit consensus.
+Notably, the recommendations of Kahl et al.
+(2024) diverge from these practices, suggesting a more rigorous examination of this application.
 For instance, for the widely used LIDC-IDRI dataset (Armato, 2011), we recommend following the preprocessing protocol of Baumgartner et al.
@@
 (2019b), whose preprocessing scripts are also available.
-Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), though they follow less standardized preprocessing.
+Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), although they follow less standardized preprocessing.
 The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al.
@@
 The evaluation metrics used across studies also diverge substantially.
-While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of empty segmentations (which often cause NaNs)—vary widely.
+While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of correct empty segmentations (which cause NaNs) vary widely.
 This inconsistency is particularly evident in the work of Zbinden et al.
@@
 CCDM (Zbinden et al., 2023) $0.212\pm 0.002^{\phantom{*}}$ $0.623\pm 0.002^{\phantom{*}}$ $0.239\pm 0.003^{\phantom{*}}$ $0.598\pm 0.001^{\phantom{*}}$
-BerDiff (Chen et al., 2023) - - $0.238\pm 0.010^{\phantom{*}}$ $0.596\pm 0.000^{\phantom{*}}$
+BerDiff (Chen et al., 2023a) - - $0.238\pm 0.010^{\phantom{*}}$ $0.596\pm 0.000^{\phantom{*}}$
 MedSegDiff (Wu et al., 2023b) - - $0.420\pm 0.030^{*}$ $0.413\pm 0.030^{*}$
@@
 7.4 Future of Active Learning
-Among the applications considered, Active Learning emerges as a promising approach to uncertainty quantification, showing true potential of uncertainty quantification by potentially reducing annotation costs while preserving or even improving model generalization.
+Among the applications considered, uncertainty quantification emerges as a promising approach to Active Learning, showing true potential of potentially reducing annotation costs while preserving or even improving model generalization.
 Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains.
@@
 (2014); Yuan et al.
-(2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed in most work.
+(2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed.
 Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty.
@@
 For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023).
-Nevertheless, Vision Transformers (ViTs) (Vaswani et al., 2017) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Hümmer et al., 2023).
+Nevertheless, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Cheng et al., 2022).
 With the exception of experiments in Kahl et al.
-(2024) (combinding SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone.
+(2024) (combining SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone.
 The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021).
@@
 Also, CNNs have benefited from the recent developments in Transformers.
-For instance, “ConvNext" takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c).
+For instance, ConvNext takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c).
+These same inductive biases have been added to ViTs through CNN adapters to also enhance performance Chen et al.
+(2023b).
 7.6 Modeling feature distributions
@@
 To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches.
-A summary of this is presented in Table 7.
-Note that while classified PixelCNNs and SSNs are classified as ‘pixel-level’ methods, both can be phrased as a latent-variable model as well.
+A summary of this discussion is presented in Table 7.
+Note that, while classified PixelCNNs and SSNs are classified as pixel-level methods, both can also be phrased as a latent-variable model.
 ELBO suboptimality
@@
 A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions.
-Limitation of sequantial modeling
+Limitation of sequential modeling
 The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models.
 This is unexpected, as both models outperforms the latter.
-This is likly due to their tedious sequential inference procedure Song et al.
+This is likely due to their tedious sequential inference procedure Song et al.
 (2020); Zheng et al.
@@
 Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions.
-Moreover, with existing work on faster sampling methods such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al.
-(2022), these models warrant further investigation.
+Furthermore, recent advances in faster sampling methods, such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al.
+(2022), make these models promising candidates for further exploration.
 Discrete vs. continuous
 It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al.
-(2023); Zbinden et al.
+(2023a); Zbinden et al.
 (2023).
 While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al.
-(2022), this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings.
+(2022).
 Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
-(a) LIDC-IDRI Armato (2011)
+However, Table 1 shows AB Chen et al.
+(2022) performs almost identically to CCDM Zbinden et al.
+(2023), suggesting that straightforward thresholding of continuous models have little-to-no performance sacrifice compared with explicitly modeling discrete distributions.
+These observation call for further investigation of the merits of categorical distributions in segmentation settings.
+(a) LIDC-IDRI Armato (2011).
 (b) CityScapes Cordts et al. (2016a).
@@
 Others
-SSNs are a simple, fast, and model-agnostic approach for modeling uncertainty.
+Stochastic Segmentation Networks (SSNs) are a simple, fast, and model-agnostic approach for modeling uncertainty.
 However, they can suffer from training instability due to the requirement that the covariance matrix be invertible.
-To address this, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular.
-Despite these fixes, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space.
+To address this limitation, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular.
+Despite these efforts, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space.
 CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017).
@@
 (2022).
-In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022)
+In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022).
 Table 7: Comparison between generative models that model aleatoric uncertainty.
@@
 VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap (Kohl et al., 2019b; Selvan et al., 2020; Valiuddin et al., 2021; 2024a; Qiu & Lui, 2020; Bhat et al., 2022a; 2023; Valiuddin et al., 2024b; Zhang et al., 2022b; Kohl et al., 2019a; Baumgartner et al., 2019; Hu et al., 2023; Liu et al., 2022b; Fischer et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Viviers et al., 2023b; a; Chotzoglou & Kainz, 2019; Saha et al., 2021a; b; Long et al., 2021a; Schmidt et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Zepf et al., 2024; 2023a; Hu et al., 2022)
-DDPM Flexible, expressive Sequential sampling (Chen et al., 2023; Zbinden et al., 2023; Rahman et al., 2023)
+DDPM Flexible, expressive Sequential sampling (Chen et al., 2023a; Zbinden et al., 2023; Rahman et al., 2023)
 7.7 Modeling parameter distributions
@@
 When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary.
-In such cases, methods like ensembling and MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation.
+In such cases, methods like MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation.
 However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors.
@@
 why ensembling or MC Dropout is often preferred).
-Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section LABEL:subsec:ME) for language modeling (He, 2024).
+Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble for language modeling (He, 2024).
 Future work should also experiment with this.
-Variational Inference Alternatives
+Alternatives to Variational Inference
 Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior.

### v8

@@
 Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making.
-The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented.
-In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising.
-By unifying theory, terminology, and applications, we provide a coherent foundation for researchers and identify critical challenges, such as strong assumptions in spatial aggregation, lack of standardized benchmarks, and pitfalls in current uncertainty quantification methods.
-We identify trends such as the adoption of contemporary generative models, driven by advances in the broader field of generative modeling, with segmentation-specific innovation primarily in the conditioning mechanisms.
-Moreover, we observe growing interest in distribution- and sampling-free approaches to uncertainty estimation.
-We further propose directions for advancing uncertainty-aware segmentation in deep learning, including pragmatic strategies for disentangling different sources of uncertainty, novel uncertainty modeling approaches and improved Transformer-based backbones.
-In this way, we aim to support the development of more reliable, efficient, and interpretable segmentation models that effectively incorporate uncertainty into real-world applications.
+Despite growing interest in probabilistic segmentation to address point-estimate limitations, the research landscape remains fragmented.
+In response, this review synthesizes foundational concepts in uncertainty modeling, analyzing how feature- and parameter-distribution modeling impact four key segmentation tasks: Observer Variability, Active Learning, Model Introspection, and Model Generalization.
+Our work establishes a common framework by standardizing theory, notation, and terminology, thereby bridging the gap between method developers, task specialists, and applied researchers.
+We then discuss critical challenges, including the nuanced distinction between uncertainty types, strong assumptions in spatial aggregation, the lack of standardized benchmarks, and pitfalls in current quantification methods.
+We identify promising avenues for future research, such as uncertainty-aware active learning, data-driven benchmarks, transformer-based models, and novel techniques to move from simple segmentation problems to uncertainty in holistic scene understanding.
+Based on our analysis, we offer practical guidelines for researchers on method selection, evaluation, reproducibility, and meaningful uncertainty estimation.
+Ultimately, our goal is to facilitate the development of more reliable, efficient, and interpretable segmentation models that can be confidently deployed in real-world applications.
 1 Introduction
 Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest (Szeliski, 2010).
-The advent of convolutional neural networks (CNNs) has led to major breakthroughs in this domain, with deep learning-based methods achieving state-of-the-art performance on large-scale datasets (Ronneberger et al., 2015; Shelhamer et al., 2014; Badrinarayanan et al., 2015), obtaining impressive scores with large-scale segmentation datasets (Lin et al., 2014; Cordts et al., 2016b; Richter et al., 2016).
-However, these models typically rely on strong assumptions and significant relaxations of the Bayesian learning paradigm, neglecting the uncertainty associated with their predictions.
-This lack of uncertainty modeling reduces both the reliability and interpretability of the predictions.
-In high-stakes applications, such as autonomous driving or medical diagnosis, this can have severe consequences.
-For instance, misclassifying adjacent objects in autonomous driving or overlooking uncertainty in lesion classification can both lead to critical decision-making errors.
+The advent of convolutional neural networks (CNNs) has led to major breakthroughs in this domain (Ronneberger et al., 2015; Shelhamer et al., 2014; Badrinarayanan et al., 2015), obtaining impressive scores with large-scale segmentation datasets (Lin et al., 2014; Cordts et al., 2016b; Richter et al., 2016).
+However, these models typically rely on strong assumptions and significant relaxations of the Bayesian learning paradigm, neglecting the uncertainty associated within their predictions.
+This lack of uncertainty modeling reduces both the reliability and interpretability of these predictions.
+In high-stake applications, such as autonomous driving or medical diagnosis, this can have severe consequences.
+For instance, misclassifying objects in autonomous driving or overlooking uncertainty in lesion classification can both lead to critical decision-making errors.
 Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications.
-Extensive efforts have been made to align neural network optimization with Bayesian machine learning (Blundell et al., 2015; Guo et al., 2017; Kingma & Welling, 2013; Kendall & Gal, 2017), such as learning parameter distributions rather than point estimates to capture epistemic uncertainty.
-Additionally, explicitly modeling the likelihood of outputs enables estimation of aleatoric uncertainty.
-However, in practice, these two forms of uncertainty are often entangled or even conflicting (Mucsányi et al., 2024; de Jong et al., 2024; Valdenegro-Toro & Mori, 2022; Hüllermeier & Waegeman, 2021), making the distinction difficult to define and apply.
-This conceptual ambiguity presents significant barriers to understanding and implementation (Kahl et al., 2024; Der Kiureghian & Ditlevsen, 2009; Kirchhof et al., 2025).
-At the same time, uncertainty is often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective.
-As a result, many works lack rigorous theoretical grounding, focusing instead on empirical gains without clear definitions or principled formulations of uncertainty.
-Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability.
+Hence, extensive efforts have been made to align neural network optimization with Bayesian machine learning, such as learning parameter distributions rather than point estimates or to explicitly model the likelihood distribution of model features (Blundell et al., 2015; Guo et al., 2017; Kingma and Welling, 2013; Kendall and Gal, 2017).
+At the same time, uncertainty in image segmentation often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective.
+Consequently, many studies lack a rigorous theoretical foundation, prioritizing empirical gains over principled formulations and clear definitions of uncertainty.
+This narrow focus frequently results in methods tailored to specific datasets or modalities, which limits their generalizability.
+Due to this application-driven nature, most existing surveys adopt a medical perspective (Jungo and Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023).
+Additionally, the study by Kahl et al.
+(2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks.
 Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, and can be insufficiently contextualized within the domain of segmentation.
@@
 The resulting abundance of literature can be overwhelming, even for experienced researchers.
-This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks.
-Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kirchhof et al., 2025), we note that a comprehensive overview in the field remains limited.
-Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023).
-However, the connection between theoretical foundations and their diverse applications remains underexplored.
-Additionally, the study by Kahl et al.
-(2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks.
-In contrast, our work provides a comprehensive overview that spans the full range of established uncertainty quantification methods, unifies and contextualizes the underlying theory, and standardizes terminology and notation.
-This broader perspective helps clarify conceptual foundations and highlights cross-cutting insights across segmentation architectures and application domains.
-By the end of this paper, readers will have a clear understanding of the various forms of uncertainty, their relevance to segmentation tasks, and a comprehensive grasp of the key challenges and open research directions in the field.
-[…]
+Given the fragmented nature of the existing literature and the renewed interest in uncertainty (Papamarkou et al., 2024; Kirchhof et al., April 28, 2025), we note that a comprehensive overview in the field remains limited.
+Our central contribution is to unify core uncertainty modeling concepts across methods, tasks, and applications, thereby offering substantial theoretical insights and demonstrating their relevance across diverse domains.
+We unify and contextualize the underlying theory, and standardize terminology and notation.
+We systematically compare methodologies, link them to specific tasks, and connect those tasks to practical applications.
+Hence, this review is not only a conventional "zero-to-hero" survey; rather, it additionally serves as a cross-domain synthesis that bridges method developers, task specialists, and application-focused researchers.
+Our aim is to provide a cohesive framework for navigating and integrating the complex uncertainty modeling landscape.
+This broader perspective helps clarify conceptual foundations, highlight cross-cutting insights and provide a comprehensive grasp of the key challenges and open research directions in the field.
+In the absence of broad quantitative consensus, we adopt a robust, evidence-based methodology for making recommendations.
+By leveraging a deep qualitative analysis of the literature, we derive reliable method comparisons and guidelines that offer essential and practical guidance.
+However, this reliance on qualitative synthesis rather than quantitative evidence is an important limitation and our recommendations should be interpreted accordingly.
+We begin by establishing the background on image segmentation in Section 2.
+Then, Section 3 introduces the notation and general theoretical principles of probabilistic image segmentation.
+Subsequently, we organize the review based on the source of introduced stochasticity: the feature level (Section 4) and the parameter level (Section 5).
+This structure aligns with the underlying theoretical decomposition of uncertainty and the utility of the modeled uncertainty for the downstream tasks and application domains discussed in Section 6.
+The review concludes by moving to the discussion (Section 7), where we highlight important pitfalls, challenges and opportunities in the field.
+This is followed by recommendations and guidelines for researchers (Section 8), before concluding in Section 9.
+The core structure and organizational flow of this paper is visualized in Figure 1, which serves as a top-level roadmap for the reader.
+{forest}
 Figure 1: Overview of the sections. The leafs of the presented hierarchical tree are related to subsections of the article.
-This review paper is structured as follows.
-Past work with significant impact in general image segmentation is presented in Section 2.
-Then, the theoretical framework and notation that govern the remainder of the paper are introduced in Section 3.
-The theory enables us to classify the approaches into two distinct but often related methods.
-That is, methods targeting either aleatoric or epistemic uncertainty, which can be distinguished by either modeling the feature or parameter distribution(s).
-The role of these concepts in image segmentation are treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field.
-We then correlate these modeling methods to specific applications in Section 6.
-Here, we also observe for which domains uncertainty quantification has been exploited.
-Then, this overview is discussed further in Section 7.
-This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task.
-To encourage further research in this field, we provide future recommendations by highlighting key challenges, gaps and pitfalls in literature, concluding our review in Section 8.
-Figure 1 illustrates a brief overview of the sections.
 2 Background
@@
 As summarized by Minaee et al.
-(2020), semantic segmentation has been performed using methods such as thresholding (Otsu, 1979), histogram-based bundling, region-growing (Dhanachandra et al., 2015), k-means clustering (Nock & Nielsen, 2004), watershedding (Najman & Schmitt, 1994), to more advanced algorithms such as active contours (Kass et al., 2004), graph cuts (Boykov et al., 2001), conditional and Markov random fields (Plath et al., 2009), and sparsity-based methods (Starck et al., 2005; Minaee & Wang, 2017).
+(2020), semantic segmentation has been performed using methods such as thresholding (Otsu, 1979), histogram-based bundling, region-growing (Dhanachandra et al., 2015), k-means clustering (Nock and Nielsen, 2004), watershedding (Najman and Schmitt, 1994), to more advanced algorithms such as active contours (Kass et al., 2004), graph cuts (Boykov et al., 2001), conditional and Markov random fields (Plath et al., 2009), and sparsity-based methods (Starck et al., 2005; Minaee and Wang, 2017).
 While the literature on segmentation is vast and rapidly evolving, a selection of backbone architectures has been particularly influential in shaping current probabilistic segmentation models.
@@
 In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures.
-Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan & Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation.
+Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan and Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation.
 Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used.
 As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015).
-Initially developed for the medical applications, Ronneberger Ronneberger et al.
+Initially developed for the medical applications, Ronneberger et al.
 (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps.
@@
 3 Probabilistic Image Segmentation
-Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{Z}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then instance $\mathbf{y}$ can be considered as the ground-truth of a $K$-class segmentation task and instance $\mathbf{x}$ as the query image.
+Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ that take values in $\mathcal{Y}\in\mathbb{Z}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then instance $\mathbf{y}$ can be considered as the ground-truth of a $K$-class segmentation task and instance $\mathbf{x}$ as the query image.
 The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively.
@@
 assuming independence of $\mathbf{x}$ from $\bm{\theta}$, and where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence).
-After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
+After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution for a new datapoint $\mathbf{x}^{*}$ can be denoted as
 $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})=\int\underbrace{p(\mathbf{Y}|\mathbf{x}^{*},\bm{\theta})}_{\text{Data}}\overbrace{p(\bm{\theta}|\mathcal{D})}^{\text{Model}}\mathrm{d}\bm{\theta}.$ (2)
 As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution.
-Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e.
-the variance of the model parameters.
-Encapsulating a particular uncertainty can therefore be achieved by incorporating parametrized stochasticity at either feature- or parameter-level.
-A straightforward approach to quantify any of these uncertainties is achieved by obtaining the predictive entropy, $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]$, i.e.
-the entropy of the predictive distribution $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})$.
-Furthermore, disentangling these uncertainties can be achieved through its decomposition
+A straightforward approach to quantify the total uncertainty is achieved by obtaining the predictive entropy, $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]$, i.e.
+the entropy of the predictive distribution $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})$, as
 $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{Y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{Y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}},$ (3)
 where $I$ represents the mutual information.
-In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process.
-Nonetheless, determining the nature of uncertainty is not often straightforward.
-For example, Hüllermeier & Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred".
-This sentiment is also shared by Der Kiureghian & Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic".
+The first term represents the expected information gain about the model parameters $\bm{\theta}$ given the true label $\mathbf{Y}$.
+A high value indicates the model is very uncertain about its prediction, meaning observing the true label would provide a large information gain.
+The second term represents the inherent randomness in the data, found by averaging the predictive uncertainty across the posterior of plausible models from Equation (1).
+This reflects the noise that cannot be reduced, regardless of how much more training data is observed.
+This distinction allows us to determine if a model’s uncertainty stems from its own ignorance or from ambiguity inherent in the data.
+Therefore, uncertainty is often split into two types: epistemic uncertainty, which reflects the model’s limited knowledge of its parameters, and aleatoric uncertainty, which captures the irreducible statistical randomness in the data itself.
+Nonetheless, determining the nature of uncertainty is rarely straightforward.
+For example, Hüllermeier and Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred".
+This sentiment is also shared by Der Kiureghian and Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic".
 Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty.
 Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
-Figure 2: Aleatoric uncertainty quantification by modeling pixel-level outputs as parameters of a Probability Mass Function.
+Figure 2: Uncertainty modeling by treating pixel-level confidence values as parameters of a Probability Mass Function.
 3.1 Conventional segmentation
-Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks, which are trained by Maximum Likelihood Estimation (MLE) and is specified as
+Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks.
+Such models are trained by Maximum Likelihood Estimation (MLE), specified as
 $\bm{\theta}_{\text{MLE}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta}),$ (4)
@@
 This approximation improves as the training data increases and the model parameter variances approach zero.
-As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Maximum A Posteriori (MAP) estimation with
+As such, MLE does not include any prior knowledge on the structure of the parameter distribution.
+This, however, can be achieved through Maximum A Posteriori (MAP) estimation with
 $\bm{\theta}_{\text{MAP}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (5)
 For example, assuming Gaussian or Laplacian priors leads to regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007).
-To model $p(\mathbf{y}|\mathbf{x},\bm{\theta})$, we make use of function $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ that infers the parameters of a Probability Mass Function (PMF).
+To model $p(\mathbf{y}|\mathbf{x},\bm{\theta})$, we make use of a function $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ that infers the parameters of a Probability Mass Function (PMF).
 For instance, consider spatial image dimensions $D$ and a CNN with $\mathbf{a}=f_{\theta}(\mathbf{x})$.
@@
 with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2).
-This approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature.
-In fact, the approximated distribution can represent and localize uncertain regions.
+Hence, this approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature.
+As such, the approximated distribution can represent and localize uncertain regions.
 However, the implicit pixel-independence assumption
@@
 In probabilistic segmentation however, this assumption has caused the emergence of a distinct research direction (See Figure 3).
-The challenge of sampling of spatially coherent segmentation masks can be addressed either from the aleatoric or epistemic perspective, usually indicating reducability of the assumed uncertainty source.
-Both approaches have specific use cases, and each modeling choice comes with distinct advantages and limitations, which will be discussed in the following sections.
+Key takeaways: Conventional models yield independent uncertainty estimates for each pixel, which omits spatial coherence.
+This failure to model spatial correlation precludes realistic inference and inhibits discerning common underlying causes for uncertainty in neighboring regions.
 (a) Likelihoods
@@
 (b) Maximum likelihood thresholding can be applied.
-(c) However, the coherence of the segmentation suffers when sampling.
+(c) However, the coherence of the segmentation suffers when sampling because information on neighboring pixels is disregarded.
+3.2 Review organization and structure
+The body of this review is structured around three interconnected concepts that serve as the fundamental organizing framework for the field of uncertainty in image segmentation.
+This structure is intended to facilitate cross-domain synthesis and enable researchers from diverse backgrounds to navigate the literature effectively.
+Consistent with the decomposition in Equation 2, we structure the methods based on feature-level (Section 4) and parameter-level (Section 5) uncertainty modeling.
+We then detail the utility of these uncertainties for four key downstream tasks in Section 6, linking them to methods via Tables 14, where the connection to practical application domains is summarized in Table 5.
+Following this structured overview, we provide a detailed discussion (Section 7) to synthesize key insights, culminating in practical recommendations for the field (Section 8).
+Core concepts: Methods: How uncertainty is mathematically modeled and quantified.
+Tasks: What specific goals are accomplished using these uncertainty estimates.
+Applications: Where these methods and tasks are applied in real-world domains.
 4 Feature modeling
-[…]
-$p(\mathbf{Y}|\mathbf{X})=\frac{p(\mathbf{Y},\mathbf{X})}{p(\mathbf{X})}\neq\delta(\mathbf{Y}-F(\mathbf{X})),$ (8)
-with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$.
-This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g.
-not all cats have tails).
-Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality.
-The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
-In most practical scenarios, aleatoric uncertainty modeling methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$.
-For example, it is possible to model a distribution parameterized by the output of a CNN.
-Also, conditional generative models are used to learn the data distribution through so-called “latent” (i.e.
-unobserved) variables.
-For the purpose of taxonomy, we categorize models based on the location of these latent variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2.
-While their theoretical formulations are largely similar, their practical implications can differ significantly.
+Following Equation (2), we focus in this section on approaches that place uncertainty at the level of features.
+In most practical scenarios, methods assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$, where parameters are point estimates.
+Accordingly, the central design decision is where to introduce auxiliary stochastic variables.
+Hence, we categorize models based on the location of these latent variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional or latents embedded deeper within the architecture are covered in Section 4.2.
+Although their theoretical formulations are similar, the practical consequences can differ substantially.
 4.1 Pixel-level sampling
@@
 In the former case, accurate uncertainty estimates rely on well-calibrated models, which is an assumption that often fails in practice without additional tuning on a separate calibration set.
-In the latter case, a stochastic variable is introduced to capture dependencies between neighboring pixels.
+Furthermore, such methods do not address the challenge of spatial coherence.
+In contrast, the latter approach introduces a stochastic variable to capture dependencies between neighboring pixels.
 4.1.1 Independence
-(a) Calibrated
-(b) Miscalibrated
-Figure 4: Visualization of a reliability diagram, which illustrates whether the model’s predicted confidences align with the observed empirical accuracies.
-As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution, but rerely reflect the true probabilities in modern neural networks Guo et al.
-(2017).
+While it is well established that assuming independence across pixels is problematic, it was a standard approach in much of the early literature.
+As discussed demonstrated in Equation (6), neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution.
+However, these values rarely reflect the true probabilities in modern neural networks, hypothesized due to the use of negative log-likelihood as the training objective, along with regularization techniques such as batch normalization, weight decay, and others (Guo et al., 2017).
 Thus, interpreting the confidences as true probabilities is often only justified after proper validation, which is referred to as model calibration.
-Here, it is measured whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e.
+Calibration
+Different methods can be used to quantify whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e.
 $P(Y=k\,|\,c_{k})=c_{k}$.
-Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4).
-Furthermore, different methods can be used to quantify calibration, although each may introduce its own biases.
 A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
@@
 The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence.
-Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.Additionally, when background pixels have a predominant influence, each bin can be weighted equally using the Average Calibration Error (ACE) to mitigate this imbalance Jungo et al.
-(2020); Neumann et al.
-(2018).
-Nonetheless, contemporary neural networks often exhibit poorly calibrated uncertainty estimates.
-This misalignment is hypothesized due to the use of negative log-likelihood as the training objective, along with regularization techniques such as batch normalization, weight decay, and others (Guo et al., 2017).
-As a result, calibration methods are typically required to adjust the predicted confidences.
-Since most of these techniques are post-hoc, i.e.
+Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
+Additionally, when background pixels have a predominant influence, each bin can be weighted equally using the Average Calibration Error (ACE) to mitigate this imbalance (Jungo et al., 2020; Neumann et al., 2018).
+Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4).
+(a) Calibrated
+(b) Miscalibrated
+Figure 4: Visualization of model calibration with a reliability diagram, which demonstrates whether the predicted model confidences align with the observed empirical accuracies.
+Methods
+Most of the calibration techniques are post-hoc, i.e.
 applied after training, they necessitate a separate validation set.
-For example, Temperature Scaling (Guo et al., 2017) has been applied in a pixel-wise manner for segmentation problems (Ding et al., 2021).
-Nonetheless, some methods, such as Label Smoothing (Silva & Oliveira, 2021; Liu et al., 2022a) or using the Focal Loss (Mukhoti et al., 2020) can be directly applied on the training data.
-Furthermore, over-fitting has often been considered to be the cause of over-confidence (Szegedy et al., 2016; Pereyra et al., 2017) and erroneous pixels can therefore be penalized through regularizing low-entropy outputs (Larrazabal et al., 2021a).
+For example.
+over-fitting has often been considered to be the cause of over-confidence (Szegedy et al., 2016; Pereyra et al., 2017) and erroneous pixels can therefore be penalized through regularizing low-entropy outputs (Larrazabal et al., 2021b).
+Another prominent method is Temperature Scaling (Guo et al., 2017), which has been successfully applied in a pixel-wise manner to segmentation problems (Ding et al., 2021).
+The standard softmax probability in Equation (6) is modified as
+$p(\mathbf{Y}=k\,|\,\mathbf{x}^{*},\bm{\theta})=\frac{e^{\mathbf{a}_{k}/T}}{\sum_{k}e^{\mathbf{a}_{k}/T}},$ (9)
+with some scalar $T$ optimized on a validation set.
+For overconfident models, $T>1$ softens the estimated distribution, pushing output probabilities away from the extremes of 0 and 1 toward more moderate values.
+In contrast, label Smoothing (Silva and Oliveira, 2021; Liu et al., 2022a) require only modifying the training data.
+Instead of one-hot encoding $\mathbf{Y}$ in to $[0,1]$, a small positive random variable $\epsilon$ is introduced, which converts the targets into $\epsilon,1-\epsilon]$, discouraging the model from producing overly confident logit outputs.
+The Focal Loss (Mukhoti et al., 2020) adds a modulating factor to down-weight the loss from well-classified examples.
+This forces the model to focus more on hard-to-classify pixels, which can implicitly improve calibration by preventing it from becoming overly certain about easy examples.
+Ultimately, relying on raw model outputs as probabilities typically yields sample-level spatial incoherence.
+Because modern neural networks are often uncalibrated by default, they require post-hoc calibration.
+Crucially, resolving spatial incoherence does not preclude good calibration, i.e., spatially correlated models can still be well calibrated.
+Conversely, calibration alone cannot impose spatial structure.
+Well-calibrated models may remain spatially incoherent.
+This asymmetry motivates placing greater emphasis on modeling spatial coherence rather than calibration alone.
 4.1.2 Spatial correlation
-Figure 5: Illustration of the PixelCNN-based PixelSeg (Zhang et al., 2022c).
+Figure 5: Illustration of the PixelCNN-based PixelSeg (Zhang et al., 2022d).
 Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation.
 Dotted elements appear during test-time sampling.
-Assuming pixel independence inhibits spatial coherence in uncertainty quantification, which is a key factor for segmentation.
-Introducing spatial correlation can be achieved through an autoregressive approach.
-For instance, we can rephrase Equation (7) to
+Introducing spatial correlation can be achieved through an autoregressive approach. For instance, we can rephrase Equation (7) to
 $p(\*Y|\*X)=\prod_{i}^{K\times D}p(Y_{i}|Y_{1},...,Y_{i-1},\*X),$ (10)
@@
 For dense predictions, Zhang et al.
-(2022c) propose PixelSeg, which predicts a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$).
-The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples.
-Figure 5 illustrates this concept.
-Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
-Figure 6: Depiction of Stochastic Segmentation Networks (Monteiro et al., 2020). Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
+(2022d) propose PixelSeg, which predicts a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$).
+The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples (Figure 5).
+Notably, the recursive sampling process also enables completion/inpainting of user-given inputs.
 Monteiro et al.
@@
 SSNs can theoretically be augmented to any pretrained CNNs as an additional layer (see Figure 6).
+Figure 6: Depiction of a Stochastic Segmentation Network (Monteiro et al., 2020). Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
+Key takeaways: Calibrating standard segmentation models can produce valid confidence values.
+However, it still preserves the incorrect pixel-independence assumption and requires a separate validation set.
+A possible solution is to model the correlation directly in pixel space through autoregressive modeling or a low-rank approximation of the output distribution.
 4.2 Latent-level sampling
-The limitation of modeling intricate, complex distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
+The limitation of modeling intricate, complex distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{Z}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximation through
 $p_{\bm{\theta},\bm{\psi}}(\mathbf{Y}|\mathbf{X})=\int p_{\bm{\theta}}(\mathbf{Y}|\mathbf{z},\mathbf{X})p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\mathrm{d}\mathbf{z},$ (12)
@@
 As such, the spatial correlation is induced through mapping the latent variables to segmentation masks.
-Conditioning the latent density, i.e.
-$p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})$, on the input images is not a necessity, but usually preferred for smooth optimization trajectories Zheng et al.
-(2022).
-However, it is also possible to simply employ an unconditional prior $p(\mathbf{z})$.
-As argued by Kahl et al.
-(2024), their predictive uncertainty can be decomposed as
-$H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{z}|\mathbf{x},\bm{\theta}]}_{\text{aleatoric}}+\overbrace{\mathbb{E}_{p(\mathbf{z})}[\,H[\mathbf{Y}|\mathbf{z},\mathbf{x},\bm{\theta}]\,]}^{{\text{epistemic}}}.$ (13)
-In contrast to Equation (3), the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty.
-Also, note that the mutual information is between the output and latent variables, rather than the parameters of the model.
+On a related note, an unconditional prior $p(\mathbf{z})$ can also be employed.
+However, conditioning the latent density, i.e.
+$p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})$, is usually preferred for smooth optimization trajectories (Zheng et al., 2022).
 The key overarching contribution of latent-level sampling methods lies in the conditioning of generative models on the input image.
@@
 4.2.1 Generative Adversarial Networks (GANs)
-[…]
+A straightforward approach is to simply learn the marginalization in Equation (12) through sampling from an unconditional prior density, $p_{Z}$, and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$.
+Goodfellow et al.
+(2014) show that this approach can be notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$.
+In this way, $G_{\bm{\phi}}$ learns to reconstruct realistic images using the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient guidance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative.
+We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
 $J_{G_{\bm{\phi}}}=-J_{D_{\bm{\psi}}}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]-\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[\,\log(1-D_{\bm{\psi}}(G_{\bm{\phi}}(\mathbf{z},\mathbf{x})))\,].$ (13)
-Figure 7: Diagram of the Calibrated Adversarial Refinement network Kassapis et al. (2021), based on the Generative Adversarial Network with additional loss terms.
-While conditional GANs have been used for semantic segmentation earlier (Isola et al., 2017), Kassapis et al.
-(2021) explicitly contextualized the architecture within aleatoric uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7).
+Conditional GANs have been used for semantic segmentation in early literature (Isola et al., 2017).
+However, Kassapis et al.
+(2021) explicitly contextualized the architecture within uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7).
 The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\*c$, with (cross-entropy) reconstruction loss
@@
 which is argued to elicit superior structural qualities compared to relying solely on cross-entropy loss. At the same time, the discriminator opposes the optimization with
-[…]
+$\mathcal{L}_{D}=$- $\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[1-\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x})\,]-\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,].$ (16)
 Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities, denoted by
@@
 The purpose of the calibration network is argued to be threefold.
-Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification.
+Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free uncertainty quantification.
 The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps.
+Figure 7: Diagram of the Calibrated Adversarial Refinement network (Kassapis et al., 2021), based on the Generative Adversarial Network (Goodfellow et al., 2014) with additional mechanisms and loss terms.
 4.2.2 Variational Autoencoders (VAEs)
@@
 the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, obtained by maximizing the conditional Evidence Lower Bound (ELBO)
-$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\log\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$ (20a)
-$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$ (20b)
-$=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,],$ (20c)
-where Jensen’s inequality justifies moving the expectation outside the logarithm.
+$\log p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\log\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$
+$\geq\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$
+$=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,],$ (19)
+where Jensen’s inequality justifies moving the logarithm inside the integral.
 The first term in Equation (19) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$.
@@
 $p(\mathbf{Y}|\mathbf{x}^{*})=\int p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{x}^{*})q_{\bm{\theta}}(\mathbf{z}|\mathbf{x}^{*})\mathrm{d}\mathbf{z}.$ (20)
-The implementation of the conditional ELBO in Equation (20) can be achieved through the well known VAE architecture (Kingma & Welling, 2013).
-Furthermore, some additional design choices lead to the Probabilistic U-Net (PU-Net) (Kohl et al., 2019b).
+The implementation of the conditional ELBO in Equation (19) can be achieved through the well known VAE architecture (Kingma and Welling, 2013).
+Furthermore, some additional design choices lead to the segmentation-specific Probabilistic U-Net (PU-Net) (Kohl et al., 2019b).
 Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on the images.
-The latent vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolution for classification.
+The latent vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of $\times1$ convolution for classification.
 When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the variability in the latent samples is almost exclusively related to the segmentation variability.
 Therefore, relatively smaller values of $d$ are feasible than what is commonly used in conventional image generation tasks.
-Similar to related research on the VAE (Zhao et al., 2017; Bousquet et al., 2017; Higgins et al., 2017; Van Den Oord et al., 2017; Rezende & Mohamed, 2015), much work has been dedicated to improving the PU-Net.
-Figure 8: Depiction of the Probabilistic U-Net (Kohl et al., 2019b) based on a conditional Variational Autoencoder (Kingma & Welling, 2013).
+Similar to related research on the VAE (Zhao et al., 2017; Bousquet et al., 2017; Higgins et al., 2017; Van Den Oord et al., 2017; Rezende and Mohamed, 2015), much work has been dedicated to improving the PU-Net.
+Figure 8: Depiction of the Probabilistic U-Net (Kohl et al., 2019b) based on a conditional Variational Autoencoder (Kingma and Welling, 2013).
 The latent samples are inserted at the final stages of a U-Net through a tiling operation.
 Density reparameterization
-[…]
+Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness (Rezende and Mohamed, 2015).
+This phenomenon has also been successfully applied to VAE-like models such as the PU-Net (Valiuddin et al., 2021; Selvan et al., 2020).
+NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $f=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex approximations of the posterior $p(\mathbf{z}|\mathbf{x})$, where symbol $\circ$ indicates functional composition.
+The approximated posteriors in turn induce an approximate of $p(\mathbf{x}|\mathbf{z})=p(\mathbf{z}|\mathbf{x})\frac{p(\mathbf{x})}{p(\mathbf{z})}$ and are often denoted as $\log p(\mathbf{x})$ for simplicity.
+The fractional term accounts for the change in probability space and corresponds to the log-Jacobian determinant, as given by the Change of Variables theorem.
+Thus, we can denote
 $\log p(\mathbf{x})=\log p_{\mathbf{Z}}(\mathbf{z}_{0})-\sum_{k=1}^{K}\log\left|\det\frac{\mathrm{d}f_{k}(\mathbf{z}_{k-1})}{\mathrm{d}\mathbf{z}_{k-1}}\right|,$ (21)
-where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are intermediate variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$.
+where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$.
 Equation (21) can be substituted in the conditional ELBO objective in Equation (19) to obtain
@@
 (2024b) show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse (Chen et al., 2016; Zhao et al., 2017).
-Their proposition considers the alternative formulation of the ELBO, specified as
-$\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (24)
-$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
-The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables, i.e.
-the aleatoric uncertainty term of the predictive entropy in Equation (13).
-This enables the introduction of the updated objective
-$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (25)
-$\hskip 40.0pt+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]+\beta\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$
-with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence Cuturi (2013) and $\alpha,\beta$ denoting hyperparameters.
+Their proposition considers a modified objective
+$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]+(\beta-\alpha)\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$ (23)
+with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence (Cuturi, 2013) and $\alpha,\beta$ denoting hyperparameters.
 This adaptation results in a more uniform latent space leading to increased model performance.
-Also, modeling the ELBO of the joint density has been explored (Zhang et al., 2022b).
+Also, modeling the ELBO of the joint density has been explored (Zhang et al., 2022c).
 This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data.
-Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon (Qiu & Lui, 2020).
+Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon (Qiu and Lui, 2020).
 Multi-scale approach
@@
 Often, an additional Markov assumption of length $T$ is placed on the posterior as
-$q_{\theta}(\mathbf{Z}_{1:T}|\mathbf{Y},\mathbf{X})=q_{\theta}(\mathbf{Z}_{1}|\mathbf{Y},\mathbf{X})\prod^{T}_{t=2}q_{\theta}(\mathbf{Z}_{t}|\mathbf{Z}_{t-1}).$ (26)
+$q_{\theta}(\mathbf{Z}_{1:T}|\mathbf{Y},\mathbf{X})=q_{\theta}(\mathbf{Z}_{1}|\mathbf{Y},\mathbf{X})\prod^{T}_{t=2}q_{\theta}(\mathbf{Z}_{t}|\mathbf{Z}_{t-1},\mathbf{X}).$ (24)
 Consequently, the conditional ELBO is denoted as
@@
 For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective is employed, which extends Equation (25) to
-$\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (28)
+$\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot||\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa||_{1}$ (26)
 $-\textstyle\sum\nolimits^{T}_{t=2}\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{t}|\mathbf{Y},\mathbf{X},\mathbf{z}_{1:t-1})\,||\,q_{\psi}(\mathbf{z}_{t}|\mathbf{z}_{1:t-1})\,]-\operatorname{KL}[\,q_{\theta}(\mathbf{z}_{1}|\mathbf{Y},\mathbf{X})\,||\,q_{\psi}(\mathbf{z}_{1}|\mathbf{X})\,].$
@@
 However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers.
-Furthermore, deep supervision at each decomposition scale to enforce the disentanglement between latent variables.
-Figure 9: Hierarchical Probabilistic U-Net (Kohl et al., 2019a) based on a hierarchical Variational Autoencoder (Kingma & Welling, 2013; Sønderby et al., 2016; Klushyn et al., 2019).
-Instead of a single latent code, multiple decomposition scales encode the segmentation variability, depending on the depth of the U-Net structure.
+Furthermore, deep supervision is placed at each decomposition scale to enforce disentanglement between the latent variables.
+Figure 9: Hierarchical Probabilistic U-Net (Kohl et al., 2019a) based on a hierarchical Variational Autoencoder (Kingma and Welling, 2013; Sønderby et al., 2016; Klushyn et al., 2019).
+Instead of a single latent code, multiple decomposition scales encode the segmentation variability.
 Extension to 3D
@@
 This limitation has spurred research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures.
-Initial works (Chotzoglou & Kainz, 2019; Long et al., 2021a) demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants.
+Initial works (Chotzoglou and Kainz, 2019; Long et al., 2021b) demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants.
 Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation.
 Viviers et al.
-(2023c) additionally augment a Normalizing Flow to the posterior density for improved expressiveness, as discussed in Section 4.2.2.
+(2023a) additionally augment a Normalizing Flow to the posterior density for improved expressiveness.
 Further enhancements to the architecture include the implementation of the 3D HPU-Net (Saha et al., 2020), an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications (Saha et al., 2021a).
@@
 (2022), generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator.
-Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $)\*y=F(\mathbf{x},\mathbf{z}_{k}$ ).
+Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $y=F(\mathbf{x},\mathbf{z}_{k})$.
 The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble.
@@
 4.2.3 Denoising Diffusion Probabilistic Models (DDPMs)
-Figure 10: Illustration of the Diffusion Probabilistic Models (Ho et al., 2020). The model learns to remove noise that has been gradually added to the input image.
 Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models (Ho et al., 2020; Song et al., 2022; 2020).
@@
 Firstly, the latent dimensionality is set equal to the spatial data dimensions, i.e.
-$d\,$=$\,D$.
+dd\,=D\,D.
 As a consequence, redundant notation of $\mathbf{Z}$ is removed and $\mathbf{Y}$ is instead subscripted with $t\in\{1,...,T\}$, indicating the encoding depth, where $\mathbf{Y}_{0}$ is the initial segmentation mask.
@@
 $q(\mathbf{Y}_{T}|\mathbf{Y}_{0})=p(\mathbf{Y}_{0})\prod^{T}_{t=1}q(\mathbf{Y}_{t}|\mathbf{Y}_{t-1}),$ (27)
-where
-$q(\mathbf{y}_{t}|\mathbf{Y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\bm{\mu}=\sqrt{\alpha_{t}}\mathbf{Y}_{t-1},\bm{\Sigma}=(1-\alpha_{t})\cdot\mathbf{I}),$ (30)
-with noise schedule $\bm{\alpha}=\{\alpha_{t}\}^{T}_{t=1}$.
+with $q(\mathbf{y}_{t}|\mathbf{Y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\bm{\mu}=\sqrt{\alpha_{t}}\,\mathbf{Y}_{t-1},\bm{\Sigma}=(1-\alpha_{t})\cdot\mathbf{I}),$ and noise schedule $\bm{\alpha}=\{\alpha_{t}\}^{T}_{t=1}$.
 Then, the decoding or reverse process can be learned through $p_{\phi}(\mathbf{Y}_{t-1}|\mathbf{Y}_{t},\mathbf{x})$.
@@
 $+\sum_{t=2}^{T}\mathbb{E}_{q\left(\mathbf{Y}_{t}|\mathbf{Y}_{0}\right)}\big[\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})\,||\,p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{X})]\,\big]+\underbrace{\mathbb{E}_{q(\mathbf{Y}_{T}|\mathbf{Y}_{0})}\left[\log\frac{p(\mathbf{y}_{T})}{q(\mathbf{y}_{T}|\mathbf{Y}_{0})}\right]}_{\approx 0}.$
-As denoted, the regularization term is assumed to be zero, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed.
-With the reparameterization trick (Kingma & Welling, 2013), the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$.
+As denoted, the regularization term is ignored, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed.
+With the reparameterization trick (Kingma and Welling, 2013), the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$.
 As such, the KL divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$, depending on the parameterization.
-This is in almost all instances approximated with a U-Net (Ronneberger et al., 2015).
-[…]
+This is in almost all instances approximated with U-Net variants (Ronneberger et al., 2015; Song et al., 2020).
+The proposed methodologies for segmentation vary in the conditioning of the reverse process on the input image.
+For instance, Wolleb et al.
+(2021) concatenate the input image with the noised segmentation mask.
+Wu et al.
+(2023b) insert encoded image features to the U-Net bottleneck.
+Additionally, information on predictions $Y_{t}$ at a time step $t$ is provided in the intermediate layers of the conditioning encoder.
+This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attention map and the inverse FFT.
+The procedure of applying attention on the spectral domain of the U-Net encoding has also been implemented with transformers in follow-up work (Wu et al., 2023a).
+Amit et al.
+(2021) also encode both current time step and input image, but combine the extracted features by simple summation prior to inferring it through the U-Net.
+It has also been proposed to model Bernoulli instead of Gaussian noise (Chen et al., 2023a; Zbinden et al., 2023; Rahman et al., 2023; Bogensperger et al., 2023).
+Figure 10: Illustration of conditional Diffusion Probabilistic Models (Ho et al., 2020). The model learns to remove noise that has been gradually added to the input image.
+Key takeaways: There is a large focus on improving amortized inference in VAE-based models (richer posteriors, hierarchies, refined objectives).
+The GAN-based CAR model has received limited follow-up.
+More contemporary work incline towards using DDPMs, where innovations largely center on input-conditioning strategies and employing categorical distributions in the noising process.
 5 Parameter modeling
-The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance (i.e.
-related to the parameters), whereas aleatoric uncertainty reflects inherent statistical variability in the data.
-Consequently, unlike aleatoric uncertainty, epistemic uncertainty should not only be quantified but ideally also minimized.
-Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier & Waegeman, 2021).
-The first type pertains uncertainty related to the capacity of the model.
-For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior.
-The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty.
-Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible.
-A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data.
-As discussed in Section 3, epistemic uncertainty is captured by modeling the posterior distribution defined in Equation (1).
+A significant advantage of this parameter-sampling approach, when compared to modeling the output distribution directly, is that it naturally induces spatial coherence across the output segmentations.
+As a result, existing methods are often simple extensions of uncertainty modeling in conventional classification, rather than leveraging techniques tailored for segmentation tasks.
+As established in Section 3, the posterior distribution from Equation (1) is used to generate a predictive distribution via Equation (2), suggesting parameter distribution modeling as an alternative to the modeling in feature space.
+We refer to any neural network that approximates the Bayesian posterior over model parameters as a Bayesian Neural Network (BNN).
 However, evaluating the true Bayesian posterior is impeded by the intractability of the data likelihood in the denominator.
-Hence, extensive efforts have been taken to obtain viable approximations, such as using Mean-Field Variational Inference Blundell et al.
-(2015), Markov Chain Monte Carlo (MCMC) Neal (2012), Monte-Carlo Dropout Gal & Ghahramani (2016); Kingma et al.
-(2015), Model Ensembling Lakshminarayanan et al.
-(2017), Laplace approximations Mackay (1992), Stochastic Gradient MCMCs Korattikara Balan et al.
-(2015); Springenberg et al.
-(2016); Welling & Teh (2011), assumed density filtering Hernández-Lobato & Adams (2015) and expectation propagation Hasenclever et al.
-(2017); Louizos & Welling (2016).
-We refer to any neural network that approximates the Bayesian posterior over model parameters as a Bayesian Neural Network (BNN).
-Unlike modeling the output distribution directly, sampling parameters from the posterior naturally induces spatial coherence in the output.
-As a result, existing methods are often simple extensions of conventional classification BNNs, rather than leveraging techniques tailored for segmentation tasks.
-This section discusses three methods for modeling epistemic uncertainty.
+This section discusses the methods used in image segmentation for approximating the parameter distribution.
 The most commonly used approach is Variational Inference (VI) over model parameters (Section 5.1).
-Test-time augmentation, often mistakenly regarded as an aleatoric method, is addressed in Section 5.3.
-Finally, Laplace Approximations are examined in Section 5.2.
+Laplace Approximations are examined in Section 5.2.
+Test-Time Augmentation (TTA) is covered in Section 5.3 (we justify this non-trivial choice later in Section 7.1).
 5.1 Variational Inference
-Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$ (e.g.
-a gaussian with $\bm{\eta}=\{\mu,\sigma\}$), to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
-Then, we can employ Variational Inference (VI) w.r.t.
-to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
+Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$.
+Variational Inference (VI) w.r.t.
+to the parameters can be employed, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
 $\bm{\eta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\operatorname{KL}\left[\,q(\bm{\theta}|\bm{\eta})\,||\,p(\bm{\theta}|\mathcal{D})\,\right]$
@@
 $q(\bm{\theta}|\bm{\eta})=\delta(\bm{\theta}-\bm{\theta}^{*})$, the formulation collapses to the MAP estimate in Equation (5).
-A popular choice for the approximated variational posterior is the Gaussian distribution, i.e.
-a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities.
+A popular choice for the approximated variational posterior is the gaussian distribution, i.e., a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary gaussian prior densities.
 However, the priors can be also learned through Empirical Bayes (Bishop, 1995).
-Furthermore, backpropagation is possible with the reparameterization trick (Kingma & Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015).
+Furthermore, backpropagation is possible with the reparameterization trick (Kingma and Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015).
 In addition to approximating the posterior, sampling from it yields multiple parameter permutations, effectively enriching the model’s hypothesis space, a phenomenon also referred to as model combination (Minka, 2000; Clarke, 2003; Lakshminarayanan, 2016).
@@
 We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2).
-See Figure 11 for an illustration of each method.
+See Figure 11 for an illustration how each method approximates the convolutional kernel distribution.
 All of these methods have been widely applied to segmentation.
-However, unlike aleatoric uncertainty modeling, their implementations are often more generic and not specifically tailored to segmentation tasks.
+However, unlike feature-level distribution modeling, their implementations are often more generic and not specifically tailored to segmentation tasks.
 Therefore, we provide only a brief overview here, followed by a more in-depth discussion in later sections related to various applications.
@@
 (c) Ensembling
-Figure 11: Three types of techniques are visualized to sample parameters of the convolutional kernels, (a) an approximation of the parameter density can be made, (b) taking samples can be mimicked with dropout during test time, or, (c) an ensemble of $N$ different configurations can be explicitly modeled.
+Figure 11: Three techniques for sampling convolutional kernels, with grayscale values indicating kernel weight magnitudes.
+These include: (a) Sampling an gaussian approximation parameterized by a mean and standard deviation kernel.
+(b) Applying dropout at inference time.
+(c) Training NN independently trained models, providing a discrete set of parameter configurations.
 5.1.1 Monte Carlo Dropout
-Dropout is a common technique used to regularize neural networks (Wager et al., 2013) by randomly “switching off” nodes of the neural network.
+Dropout is a common technique used to regularize neural networks by randomly “switching off” nodes of the neural network.
 Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix (Wager et al., 2013).
@@
 with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters.
-It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal & Ghahramani, 2016).
-In this manner, such a method is able to provide multimodal estimates of the model uncertainty.
+It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal and Ghahramani, 2016).
+In this manner, such a method is able to provide multimodal estimates of the uncertainty.
+The dropout rate is usually optimized through grid search or simply set to $p=0.5$.
+The relationship between $p$ and the magnitude of the model weights can also be exploited to probabilistically prune neural networks (Gonzalez-Carabarin et al., 2022).
 As noted by Gal et al.
 (2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$.
-The dropout rate is usually optimized through grid search or simply set to $p=0.5$.
-The relationship between $p$ and the magnitude of the model weights has also be exploited to probabilistically prune neural networks Gonzalez-Carabarin et al.
-(2022).
-Gal et al.
-(2017a) propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$.
+Hence, they propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$.
 As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$.
-Also, a generalization of MC dropout has been proposed, where the weights of the network, rather that its hidden output, are set to zero (Mobiny et al., 2021).
+Also, a generalization of MC dropout has been proposed, where the weights of the network, rather than its hidden output, are set to zero (Mobiny et al., 2021).
 5.1.2 Ensembling
@@
 the second derivative of the loss function evaluated at $\bm{\theta}_{\text{MAP}}$.
-This approach, known as the Laplace Approximation (LA) (Laplace, 1774), has the merit that it can be performed post-hoc to .e.g.
-pretrained neural networks.
+This approach, known as the Laplace Approximation (LA) (Laplace, 1774), has the merit that it can be performed post-hoc on pretrained neural networks.
 However, estimating the Hessian can become infeasible, as computation scales quadratically with the model parameters.
-This is usually circumvented by treating neural network partly probabilistic and/or approximating the Hessian with a more simplified matrix structure (Martens & Grosse, 2015; Botev et al., 2017; Daxberger et al., 2021).
+This is usually circumvented by treating neural networks partly probabilistic and/or approximating the Hessian with a more simplified matrix structure (Martens and Grosse, 2015; Botev et al., 2017; Daxberger et al., 2021).
 5.3 Test-time augmentation
-An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest.
-For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics.
-Hence, data augmentation (Krizhevsky et al., 2012) has been used at test time (explaining the name test-time augmentation, or TTA), argued to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function (Ayhan & Berens, 2022).
-By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$.
-Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
-In segmentation literature, it has been hypothesized that TTA encapsualtes the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023).
-However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022).
-Kahl et al.
-(2024) draw parallels with BNNs and demonstrate this through the (by now) familiar decomposition of the predictive uncertainty
-$H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{t}|\mathbf{x},\bm{\theta}]}_{\text{epistemic}}+\overbrace{\mathbb{E}_{\mathbf{t}}[\,H[\mathbf{Y}|\mathbf{t},\mathbf{x},\bm{\theta}]\,]}^{{\text{aleatoric}}}.$ (35)
-where $t$ are sampled augmentations from the input space $\mathcal{T}$.
-Aleatoric uncertainty is captured because a perfectly trained model produces consistent predictions across augmentations of known data.
-Epistemic uncertainty arises when the mutual information between the predicted label and augmentation is greater than zero for unseen data, indicating the model has not yet learned invariance for such inputs.
-6 Applications
-This section briefly provides literature that incorporates uncertainty into downstream tasks using segmentation models.
-These tasks can be broadly categorized into four main groups: estimating the segmentation mask distribution subject to observer variability (Section 6.1), reducing labeling costs using Active Learning (Section 6.2), model introspection (i.e.
-ability to self-assess Section 6.3) and improved generalization (Section 6.4).
-For quick reference, Table 5 summarizes the types of dataset domains (e.g., medical, outdoor scenes) associated with various downstream tasks.
+[…]
+Key takeaways: Compared with earlier methods, parameter distribution modeling is largely model agnostic and rarely incorporates segmentation-specific architectures.
+In practice, many implementations reduce to approximate Variational Inference techniques, such as MC Dropout or Ensembling, followed by methods like Test-Time Augmentation and Laplace Approximations.
+Echoing their origins in classical classification, these concepts apply broadly to other domains.
+6 Tasks
+This section reviews literature on the downstream utilization of uncertainty in segmentation models, classifying the corresponding tasks into four main groups.
+Furthermore, Table 5 in Section 6.5 explicitly connects the downstream tasks to the specific application domains where they have been successfully employed.
+Tasks: Estimating Observer Variability, the range of plausible segmentation masks (Section 6.1) Reducing labeling costs using Active Learning (Section 6.2) Giving the model the ability to self-assess with Model Introspection (Section 6.3) Improving Model Generalization (Section 6.4)
 6.1 Observer variability
@@
 The phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator are referred to as the intra-observer variability (see Figure 12).
+For example, in critical applications like radiation therapy planning, the precise delineation of tumor contours is essential.
+However, inherent ambiguity in the image acquisition process often leads to significant inter-observer variability.
+Multiple expert radiologists might annotate the same tumor with different, yet equally plausible, boundaries.
+In this context, no single annotation can be considered definitively correct, and this irreducible uncertainty in the ground truth must be accounted for to mitigate potentially serious impacts on patient care.
+Similarly, the perception of an autonomous driving system can suffer from label-level uncertainty.
+Consider an object partially obscured at a distance.
+The sensor data may be insufficient to distinguish whether it is a pedestrian about to cross the road or a stationary object like a mailbox due to a genuine ambiguity in the scene itself.
+The decision-making process must handle this uncertainty, as the consequences of misclassifying a vulnerable road user are severe.
+Figure 12: Visualization of the intra-observer variability in parameter space (left) and the inter-observer variability in data space (right).
 Annotators are models
 To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be regarded as models themselves.
-For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with $k=1,2,...,K$.
+For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with k=1,2,,Kk=1,2,...,K.
 For a simple segmentation task, it can be expected that $\text{Var}[\,p(\bm{\phi}_{k})\,]\rightarrow 0$.
-In other words, each annotator is consistent in his delineation and the intra-observer variability is low.
+In other words, each annotator is consistent in their delineation and the intra-observer variability is low.
 For cases with consensus across experts, i.e.
@@
 Yet, this is rarely the case in many real-life applications and, as such, explicitly modeling the involved distributions becomes necessary.
-Figure 12: Visualization of the intra-observer variability in parameter space (left) and the inter-observer variability in data space (right).
 Evaluation
@@
 This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) (Gretton et al., 2012), denoted as
-$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (36)
 […]
+with marginals $P_{\*Y}$ and $P_{\hat{\*Y}}$, representing the true and predictive segmentation distribution, and a distance represented by kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the (1-IoU) or (1-Dice) score (squared by the GED).
+However, solely relying on the GED has been criticized due to undesirable inductive biases (Kohl et al., 2019a; Zepf et al., 2024).
+Hence, Kohl et al.
+(2019a) introduce an alternative metric known as Hungarian Matching (HM), which compares the predictions against the ground-truth labels through a cost matrix.
+This can be also formally denoted as finding the permutation matrix P\*P, subject to the objective
 $\text{HM}(\mathcal{Y},\mathcal{\hat{Y}})=\frac{1}{N^{2}}\min_{\mathbf{P}}\text{Tr}(\mathbf{P}\mathbf{M}),$ (33)
+where the elements of M\*M are $M_{i,j}=k(\mathcal{Y}_{i},\mathcal{\hat{Y}}_{j})$, and $N^{2}$ represents the number of elements within the matrix.
+Subsequently, the unique optimal coupling between the two sets that minimize the average cost is determined through a combinatorial optimization algorithm.
+Hu et al.
+(2022) use the Normalized Cross Correlation (NCC) to measure the similarity between the true and predictive set, which can formally be defined as
 […]
-$\text{NCC}=\frac{1}{n\sigma\hat{\sigma}}\sum_{i}(a_{i}-\mu)\cdot(\hat{b}_{i}-\hat{\mu}),$ (38)
 with $n$ being the number of pixels, and $\mu,\sigma$ and $\hat{\mu},\hat{\sigma}$ the mean and standard deviation of the true and predicted segmentations set, respectively.
@@
 Approaches
-The most straightforward approach is to directly model the empirical stochasticity in the annotations with models that target the aleatoric uncertainty.
-Given that annotators have an intrinsically associated expertise, it can also be possible to modeling the annotator distribution.
+The most straightforward approach is to directly model the empirical stochasticity in the annotations with stochasticity in the features.
+Given that annotators have an intrinsically associated expertise, it can also be possible to model the annotator distribution in the parameters.
 However, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome.
 This observation is reflected in literature, where we see that most approaches use conditional generative models to encapsulate observer variability (See Table 1).
-While conditional generative models have been successfully applied for the task at hand, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles of annotators Zepf et al.
-(2023a).
+While conditional generative models have been successfully applied for the task at hand, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles of annotators (Zepf et al., 2023a).
 VAE-based Probabilistic U-Net is the most popular approached followed by its hierarchical variants (HPU-Net and PHI-Seg) and SSNs.
@@
 Method Literature
-PixelCNN Zhang et al. (2022c)
+PixelCNN Zhang et al. (2022d)
 SSN Monteiro et al. (2020); Kahl et al. (2024); Zepf et al. (2024; 2023a); Philps et al. (2024)
@@
 (2023; 2022a); Valiuddin et al.
-(2024b); Qiu & Lui (2020); Valiuddin et al.
+(2024b); Qiu and Lui (2020); Valiuddin et al.
 (2024a); Viviers et al.
-(2023c); Schmidt et al.
+(2023a); Schmidt et al.
 (2023); Long et al.
-(2021b); Zepf et al.
+(2021a); Zepf et al.
 (2023a; 2024); Hu et al.
 (2022); Viviers et al.
-(2023b); Savadikar et al.
+(2023c); Savadikar et al.
 (2021); Philps et al.
 (2024)
-HVAE Zhang et al. (2022b); Rafael-Palou et al. (2021); Kohl et al. (2019a); Baumgartner et al. (2019); Gantenbein et al. (2020)
+HVAE Zhang et al. (2022c); Rafael-Palou et al. (2021); Kohl et al. (2019a); Baumgartner et al. (2019); Gantenbein et al. (2020)
 DDPM Chen et al. (2023a); Zbinden et al. (2023); Rahman et al. (2023)
@@
 6.2 Active Learning
-The field of active learning (Settles, 2009; Ren et al., 2021) aims to reduce the costly annotation procedure by careful selection of the most informative samples for training (Settles, 2009).
-In this way, the required labeled data is minimized and training convergence can be accelerated due to reduced redundancy.
+The field of Active Learning aims to reduce the costly annotation procedure by careful selection of the most informative samples for training (Settles, 2009; Ren et al., 2021).
+Instead of randomly labeling data, the model directs expert time toward the most challenging examples, leading to more efficient learning.
 Active Learning is an extensively researched topic in both traditional (Cohn et al., 1996) and deep Machine Learning (Gal et al., 2017b), including uncertainty-based approaches applied to image-segmentation pairs.
 In such methods, it is assumed that predictions with high uncertainty provide the most informative samples for training.
+Moreover, this core idea extends to active perception, where an autonomous agent, like a robot or a drone, physically acts to resolve its own environmental uncertainty.
+For instance, instead of guessing what an obscured object is, it might move its camera to get a better view.
+In both cases, the system actively seeks the most valuable information, either from a human annotator or by interacting with the world, to improve its understanding and performance.
 Evaluation
@@
 Firstly, the uncertainty needs to be quantified.
-This is usually the predictive entropy Kasarla et al.
-(2019); Shen et al.
-(2021); Xie et al.
-(2022) or variance Ozdemir et al.
-(2021), but in some instances the epistemic uncertainty is explicitly estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021).
+This is usually the predictive entropy (Kasarla et al., 2019; Shen et al., 2021; Xie et al., 2022) or variance (Ozdemir et al., 2021), but in some instances the uncertainty is estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021).
 Simply stated, the reduction in posterior entropy indicates the informativeness of a data sample.
@@
 $=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]=I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}).$ (36)
-Note the similarity between the mutual information $I$ obtained from the predictive entropy decomposition in Equation (3), where both are the objective of epistemic uncertainty modeling.
+Note the similarity between the mutual information $I$ obtained from the predictive entropy decomposition in Equation (3).
 Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e.
-a percentage of the original dataset) Yang et al.
-(2017); Sourati et al.
-(2018); Kasarla et al.
-(2019); Mahapatra et al.
-(2018); Sinha et al.
-(2019); Siddiqui et al.
-(2020); Li & Alstrøm (2020); Kim et al.
-(2021); Shen et al.
-(2021); Burmeister et al.
-(2022); Ma et al.
-(2024b).
+a percentage of the original dataset) (Yang et al., 2017; Sourati et al., 2018; Kasarla et al., 2019; Mahapatra et al., 2018; Sinha et al., 2019; Siddiqui et al., 2020; Li and Alstrøm, 2020; Kim et al., 2021; Shen et al., 2021; Burmeister et al., 2022; Ma et al., 2024b).
 Kahl et al.
-(2024) argue that Active Learning should be evaluated on image-level as humans dont annotate on pxiel-level.
+(2024) argue that Active Learning should be evaluated on image-level as humans do not annotate on pixel-level.
 Assuming a saturated performance on in-distribution data (training cycle $t_{1}$), the author propose to measure the improved relative performance, $C$, subject to a second training sample ($t_{2}$) from selected samples
@@
 where $C_{\text{method}}$ denotes the relative improvement achieved by the employed method, and $C_{\text{random}}$ is a correction term accounting for gains due to random sampling, and $P$ suggested to be the Dice-score, but can theoretically be any segmentation metric.
-Table 2: Methods used for the application of Active Learning
+Table 2: Methods used for the application of Active Learning.
 Method Literature
@@
 (2021); Hiasa et al.
-(2019); Li & Alstrøm (2020); Shen et al.
+(2019); Li and Alstrøm (2020); Shen et al.
 (2021); Ma et al.
@@
 Approaches
-Since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate.
-Hence, as can be seen in Table 2, most of the literature use MC dropout of ensembles.
+As can be seen in Table 2, most of the literature use MC dropout or ensembles.
 Furthermore, using the plain SoftMax activation has also been commonly used.
-Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness Ozdemir et al.
-(2021); Shen et al.
-(2021); Wu et al.
-(2021); Mahapatra et al.
-(2018); Sinha et al.
-(2019).
+Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness (Ozdemir et al., 2021; Shen et al., 2021; Wu et al., 2021; Mahapatra et al., 2018; Sinha et al., 2019).
 Another crucial modeling element pertains the aggregation of the pixel-level uncertainties.
-Therefore, closeness to boundary Gorriz et al.
-(2017); Kasarla et al.
-(2019); Ma et al.
-(2024b) or specific regions Gaillochet et al.
-(2023); Wu et al.
-(2021); Kasarla et al.
-(2019) are additionally incorporated when combining the uncertainty values across pixels.
+Therefore, closeness to boundary (Gorriz et al., 2017; Kasarla et al., 2019; Ma et al., 2024b) or specific regions (Gaillochet et al., 2023; Wu et al., 2021; Kasarla et al., 2019) are additionally incorporated when combining the uncertainty values across pixels.
 Siddiqui et al.
@@
 6.3 Model introspection
-Erroneous predictions should correlate with the uncertainty in cases the source is assumed to be epistemic.
-In turn, this implies that the provided uncertainty can be used to to gauge prediction quality of a model.
-High uncertainty predictions can be discarded to ensure a level of model reliability.
-Hence, the described problem can also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021).
-Including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021).
-Note that using uncertainty to guide model training is closely related to Active Learning, which has been discussed in Section 6.2.
-Table 3: Methods used for the application of Model Introspection
+The provided uncertainty can be used to gauge prediction quality of a model.
+For example, an agricultural drone’s segmentation model provides a clear example.
+When distinguishing "crop" vs.
+"weed", it will segment healthy fields with low uncertainty, signaling a reliable output.
+However, for an anomalous patch of stressed plants, it will produce a segmentation with high uncertainty.
+This high score directly flags the prediction as low-quality and likely erroneous, preventing a costly automated action like misapplying herbicide.
+As evident, the described problem can also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder and Shafique, 2021).
+Table 3: Methods used for the application of Model Introspection.
 Method Literature
@@
 LA Zepf et al. (2023b)
-TTA Kahl et al. (2024); Wang et al. (2019b); Whitbread & Jenkinson (2022); Roshanzamir et al. (2023); Dahal et al. (2020)
+TTA Kahl et al. (2024); Wang et al. (2019b); Whitbread and Jenkinson (2022); Roshanzamir et al. (2023); Dahal et al. (2020)
 Ensembling Jungo et al.
@@
 (2020); Czolbe et al.
-(2021); Holder & Shafique (2021); Jungo & Reyes (2019); Ng et al.
+(2021); Holder and Shafique (2021); Jungo and Reyes (2019); Ng et al.
 (2022); Hann et al.
-(2021); Jungo & Reyes (2019); Linmans et al.
+(2021); Jungo and Reyes (2019); Linmans et al.
 (2020); Pavlitskaya et al.
@@
 (2020); Kahl et al.
-(2024); Whitbread & Jenkinson (2022); Kendall et al.
+(2024); Whitbread and Jenkinson (2022); Kendall et al.
 (2016); Kampffmeyer et al.
@@
 (2020); Sander et al.
-(2019); Hasan & Linte (2022b); Camarasa et al.
+(2019); Hasan and Linte (2022b); Camarasa et al.
 (2021); Roshanzamir et al.
@@
 (2018); Seeböck et al.
-(2019); DeVries & Taylor (2018); Czolbe et al.
+(2019); DeVries and Taylor (2018); Czolbe et al.
 (2021); Hoebel et al.
@@
 (2020); Lambert et al.
-(2022); Hasan & Linte (2022a); Jungo & Reyes (2019); Ng et al.
+(2022); Hasan and Linte (2022a); Jungo and Reyes (2019); Ng et al.
 (2022); Iwamoto et al.
@@
 SSN Kahl et al. (2024); Ng et al. (2022)
-VAE Chotzoglou & Kainz (2019); Viviers et al. (2023a); Czolbe et al. (2021); Bian et al. (2020)
+VAE Chotzoglou and Kainz (2019); Viviers et al. (2023b); Czolbe et al. (2021); Bian et al. (2020)
 Evaluation
-The relationship between uncertainty and model accuracy has been formalized by Mukhoti & Gal (2018) through the Patch Accuracy vs Patch Uncertainty (PAvPU) metric.
+The relationship between uncertainty and model accuracy has been formalized by Mukhoti and Gal (2018) through the Patch Accuracy vs Patch Uncertainty (PAvPU) metric.
 Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$.
@@
 Hence, Roy et al.
-(2018b) propose to use the Coefficient of Variation (CoV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples.
+(2018b) propose to use the Coefficient of Variation (CoV), which addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples.
 The authors also evaluate structural uncertainties by assuming predictions with the pair-wise average overlap between all respective samples.
@@
 Approaches
-Considering Table 3, it can be observed that epistemic methods such as MC dropout and Ensembles are by far the most popular choice of uncertainty encapsulation method.
-However, Concrete dropout (Rumberger et al., 2020b; Mukhoti & Gal, 2018), M-heads (auxiliary networks) (Jungo & Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
+Considering Table 3, it can be observed that parameter modeling methods such as MC Dropout and Ensembles are by far the most popular choice of uncertainty encapsulation method.
+However, Concrete dropout (Rumberger et al., 2020b; Mukhoti and Gal, 2018), M-heads (auxiliary networks) (Jungo and Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
+Moreover, including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021).
 6.4 Model generalization
@@
 In contrast to other downstream tasks, evaluation does not require specialized metrics besides quantifying the relative change in model performance.
-For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) improves model performance (Kendall et al., 2016).
-In Table 4, we can see improved performance with MC Dropout and also in specific with Concrete Dropout (Mukhoti & Gal, 2018; Rumberger et al., 2020b).
+For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) has shown to improve model performance (Kendall et al., 2016).
+MC Dropout, and specifically Concrete Dropout (Mukhoti and Gal, 2018; Rumberger et al., 2020b)), has also led to improved performance.
 The improved generalization from ensembling has also shown to produce more calibrated outputs (Mehrtash et al., 2020).
@@
 (2022) found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions.
-In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits (Larrazabal et al., 2021b).
-Ensembling has also been performed using M-Heads (Hu et al., 2023; Jungo & Reyes, 2019).
+In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits (Larrazabal et al., 2021a).
+Ensembling has also been performed using M-Heads (Hu et al., 2023; Jungo and Reyes, 2019).
 Nonetheless, individual models in conventional ensembles receive data in an unstructured manner.
 Hence, specific subsets of the data can also be assigned to particular models with MoEs (Pavlitskaya et al., 2020; Ji et al., 2021; Gao et al., 2022).
-Generative models, though typically used for other applications, have also led to performance improvements across a wide range of studies Saha et al.
-(2021a); Wolleb et al.
-(2021); Wu et al.
-(2023b); Amit et al.
-(2021); Bogensperger et al.
-(2023); Zbinden et al.
-(2023); Viviers et al.
-(2023a); Zepf et al.
-(2023a); Hu et al.
-(2022).
+Generative models, though typically used for other applications, have also led to performance improvements across a wide range of studies (Saha et al., 2021a; Wolleb et al., 2021; Wu et al., 2023b; Amit et al., 2021; Bogensperger et al., 2023; Zbinden et al., 2023; Viviers et al., 2023b; Zepf et al., 2023a; Hu et al., 2022).
+Key takeaways: Parameter distribution methods are commonly utilized for Active Learning, Model Introspection, and Model Generalization, while feature distribution modeling is employed for Observer Variability.
+This observation suggests the irreducibility of uncertainty tied to parameter-level modeling, while reducibility aligns with feature-level modeling.
 Table 4: Methods used for the application of improved Model Generalization.
@@
 SSN Zepf et al. (2023a); Ng et al. (2022)
-VAE Viviers et al. (2023a); Zepf et al. (2023a); Hu et al. (2022)
+VAE Viviers et al. (2023b); Zepf et al. (2023a); Hu et al. (2022)
 DDPM Wolleb et al. (2021); Wu et al. (2023b; a); Amit et al. (2021); Zbinden et al. (2023); Bogensperger et al. (2023); Hoogeboom et al. (2021)
-MC dropout Mukhoti & Gal (2018); Rumberger et al.
+MC dropout Mukhoti and Gal (2018); Rumberger et al.
 (2020b); Saha et al.
 (2021a); Zhang et al.
-(2022a); Whitbread & Jenkinson (2022); Huang et al.
+(2022a); Whitbread and Jenkinson (2022); Huang et al.
 (2018); Mukhoti et al.
@@
 (2020)
-Ensembling Ng et al. (2022); Larrazabal et al. (2021b); Kamnitsas et al. (2018); Saha et al. (2021b); Hu et al. (2023); Ji et al. (2021)
-TTA Wang et al. (2019a); Rakic et al. (2024); Whitbread & Jenkinson (2022)
+Ensembling Ng et al. (2022); Larrazabal et al. (2021a); Kamnitsas et al. (2018); Saha et al. (2021b); Hu et al. (2023); Ji et al. (2021)
+TTA Wang et al. (2019a); Rakic et al. (2024); Whitbread and Jenkinson (2022)
 6.5 Application domains
-In this Table 5, we link the discussed methodologies and applications to their respective domains.
+In Table 5, we link the discussed methodologies and tasks to their respective application domains.
 It is evident that the medical domain, followed by the automotive field, is the most prominent for probabilistic segmentation tasks.
 The high-stakes nature of both domains contributes to their prominence, as uncertainty-aware models are especially vital in safety-critical applications.
-Specifically, the delineation of lung nodules in CT images are ambiguous.
-The LIDC-IDRI dataset is a multi-annotated Lung CT dataset, which serves as a good benchmark for quantifying observer variability and is reflected in the table.
+Several benchmark datasets are central to these applications.
+For instance, the LIDC-IDRI dataset (Armato, 2011), with its multiple expert delineations of lung nodules in CT scans, is a key benchmark for quantifying observer variability.
+In the autonomous driving domain, outdoor scene understanding applications widely utilize the Cityscapes dataset.
+Similarly, MRI-based research is often advanced by datasets like BraTS or challenges such as QUBIQ (Menze et al., 2021).
 Table 5: Overview that links literature to various domains and applications.
@@
 (2021) Pavlitskaya et al.
-(2020); Hasan & Linte (2022b); Sander et al.
+(2020); Hasan and Linte (2022b); Sander et al.
 (2019); Ng et al.
-(2022); Hasan & Linte (2022a); Mehrtash et al.
+(2022); Hasan and Linte (2022a); Mehrtash et al.
 (2020); Kendall et al.
 (2016); Kahl et al.
-(2024); Holder & Shafique (2021); Qi et al.
+(2024); Holder and Shafique (2021); Qi et al.
 (2023) Ng et al.
@@
 (2018); Zbinden et al.
-(2023); Mukhoti & Gal (2018); Amit et al.
+(2023); Mukhoti and Gal (2018); Amit et al.
 (2021); Hoogeboom et al.
@@
 Various objects Philps et al. (2024) Kendall et al. (2016); Qi et al. (2023); Morrison et al. (2019)
-Remote sensing García Rodríguez et al. (2020) Dechesne et al. (2021; 2021); Kampffmeyer et al. (2016); Dechesne et al. (2021)
+Remote sensing García Rodríguez et al. (2020) Dechesne et al. (2021); Kampffmeyer et al. (2016); Dechesne et al. (2021)
 Microscopy Kohl et al.
@@
 (2024) Khalili et al.
-(2024); Li & Alstrøm (2020); Yang et al.
+(2024); Li and Alstrøm (2020); Yang et al.
 (2017); Sadafi et al.
@@
 (2021); Zhang et al.
-(2022c) Zepf et al.
+(2022d) Zepf et al.
 (2023b); Ma et al.
@@
 (2018b); Mehrtash et al.
-(2020); Jungo & Reyes (2019); Eaton-Rosen et al.
+(2020); Jungo and Reyes (2019); Eaton-Rosen et al.
 (2018); Wang et al.
 (2019b); Jungo et al.
-(2020); Whitbread & Jenkinson (2022); Monteiro et al.
+(2020); Whitbread and Jenkinson (2022); Monteiro et al.
 (2020); Lambert et al.
@@
 (2018); Larrazabal et al.
-(2021b); Wang et al.
-(2019b); Whitbread & Jenkinson (2022)
+(2021a); Wang et al.
+(2019b); Whitbread and Jenkinson (2022)
 Prostate MRI Baumgartner et al.
@@
 OCT imaging Selvan et al. (2020) Seeböck et al. (2019); Bian et al. (2020); Sedai et al. (2018) Rakic et al. (2024); Ji et al. (2021); Wu et al. (2023a; b)
-Dermoscopy Li & Alstrøm (2020); Gorriz et al. (2017) Zepf et al. (2023b); Jungo & Reyes (2019); DeVries & Taylor (2018); Czolbe et al. (2021) Wu et al. (2023a); Zepf et al. (2023a)
+Dermoscopy Li and Alstrøm (2020); Gorriz et al. (2017) Zepf et al. (2023b); Jungo and Reyes (2019); DeVries and Taylor (2018); Czolbe et al. (2021) Wu et al. (2023a); Zepf et al. (2023a)
 Lung CT Rafael-Palou et al.
@@
 (2023a); Zhang et al.
-(2022b); Monteiro et al.
-(2020); Qiu & Lui (2020); Kohl et al.
+(2022c); Monteiro et al.
+(2020); Qiu and Lui (2020); Kohl et al.
 (2019a); Hu et al.
 (2022); Zhang et al.
-(2022c); Gantenbein et al.
+(2022d); Gantenbein et al.
 (2020); Viviers et al.
-(2023b; c); Valiuddin et al.
+(2023c; a); Valiuddin et al.
 (2024a); Zbinden et al.
@@
 (2022a); Long et al.
-(2021b); Valiuddin et al.
+(2021a); Valiuddin et al.
 (2021); Rahman et al.
@@
 (2021); Long et al.
-(2021a); Selvan et al.
+(2021b); Selvan et al.
 (2020); Baumgartner et al.
@@
 (2024); Czolbe et al.
-(2021); Chotzoglou & Kainz (2019) Hu et al.
+(2021); Chotzoglou and Kainz (2019) Hu et al.
 (2022); Rakic et al.
@@
 (2022a)
-Various CT Valiuddin et al. (2024b); Hu et al. (2023; 2023) Hiasa et al. (2019) Hoebel et al. (2019); LaBonte et al. (2019; 2019) Viviers et al. (2023a)
+Various CT Valiuddin et al. (2024b); Hu et al. (2023) Hiasa et al. (2019) Hoebel et al. (2019); LaBonte et al. (2019) Viviers et al. (2023b)
 Various US Rahman et al. (2023) Yang et al. (2017) Antico et al. (2020); Dahal et al. (2020); Bian et al. (2020) Wu et al. (2023b; a); Rumberger et al. (2020b)
@@
 This section builds on the reviewed theory and literature to guide researchers in selecting methods suited to their goals.
-Furthermore, it highlights key gaps and limitations, offering insights for addressing current challenges in the field and identifying promising directions for future research.
-Our main discussion points can be summarized as follows:
-  • The interpretation and quantification of epistemic and aleatoric uncertainty remain complex and warrant careful reconsideration, despite the literature often indicating a clear distinction between them.
-(Section 7.1).
-  • Spatial uncertainty introduces additional challenges, including the accurate estimation of structural entropy, the aggregation of pixel-wise uncertainties, and increased complexity in volumetric segmentation tasks (Section 7.2).
-  • Lack of standardization across the downstream tasks remains a major issue in the field, hindering fair benchmarking and limiting further progress (Section 7.3).
-  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4).
-  • Segmentation backbones commonly used in the field tend to be outdated, with Transformer-based architectures having surpassed them in performance yet remaining underutilized in probabilistic dense predictions (Section 7.5).
-  • Aleatoric uncertainty is often modeled using conditional generative approaches; as this area advances rapidly, its developments should be promptly integrated into probabilistic segmentation research (Section 7.6).
-  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout, questioning its ability to capture true model uncertainty (Section 7.7).
+Furthermore, it highlights key gaps and limitations, offering insights to address current challenges and identify promising directions for future research.
+Main discussion points: The interpretation of epistemic and aleatoric uncertainty remain complex and warrant careful reconsideration, despite the literature often suggesting a clear distinction (Section 7.1).
+Challenges related to spatial coherence remain unaddressed (Section 7.2).
+The application-driven field lacks standardization across the downstream tasks, hindering benchmarking and further progress (Section 7.3).
+Uncertainty-based Active Learning is promising and has many open challenges (Section 7.4).
+There exist various limitations and trade-offs across methods (Section 7.5).
+Understanding the critical influence of data characteristics on uncertainty models remains a relatively under-explored area of research (Section 7.6).
+The field of segmentation could advance significantly by treating the backbone as a critical modeling choice and embracing contemporary architectures (Section 7.7).
+The segmentation community should actively explore and integrate novel uncertainty quantification methodologies (Section 7.8).
+Uncertainty modeling should address the growing importance of multiclass, instance, and panoptic segmentation problems (Section 7.9).
 7.1 Disentangling uncertainties
 Proper disentanglement of the uncertainties allows to accredit high-entropy outputs to either model ignorance about the data (epistemic), or inherent noise in the data generation process (aleatoric).
-Importantly, this distinction relates to the suggestion whether the uncertainty is reducible or not.
-Furthermore, Equation( 2) indicates this can be achieved by introducing stochasticity into the model parameters or the features.
+As Equation (2) implies, this can be achieved by introducing stochasticity into either the model’s parameters or its features.
+This fundamental distinction not only provides the organizing principle for our subsequent review but also aligns with the approaches empirically observed in the literature.
 Seemingly straightforward, sufficient literature and discussion in the scientific community exist related to the applicability of this taxonomy in practical cases.
 As paraphrased from Kirchhof et al.
-(2025), “definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another”.
-We further explore this concept and show how confusion and contradiction have led to common pitfalls in related studies.
-To address these issues, we highlight key nuances in uncertainty modeling and contextualize this to segmentation problems by providing recommendations for future research.
-7.1.1 Nuances in uncertainty modeling
-Interpretation of uncertainty is task dependent
+(April 28, 2025), “definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another”.
+We further explore this, highlighting key nuances in uncertainty modeling and contextualize this to segmentation problems.
+To begin, we first establish the original definitions of these concepts to better understand why these concepts have lead to incorrect definitions, claims and general confusion.
+Aleatoric uncertainty
+reconsiders the non-deterministic relationship between $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, which implies that
+$p(\mathbf{Y}|\mathbf{X})=\frac{p(\mathbf{Y},\mathbf{X})}{p(\mathbf{X})}\neq\delta(\mathbf{Y}-F(\mathbf{X})),$ (40)
+with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$.
+This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g.
+not all cats have tails).
+Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality.
+The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (input dependent) or homoscedastic (input independent) aleatoric uncertainty.
+Epistemic uncertainty
+arises from model ignorance (i.e.
+related to the parameters).
+Consequently, epistemic uncertainty should not only be quantified but ideally also minimized.
+Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier and Waegeman, 2021).
+The first type pertains uncertainty related to the capacity of the model.
+For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior.
+The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty.
+Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible.
+A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data.
+Additionally, we have found it to be common to interchangeably use the total epistemic and solely model uncertainty, especially in the context of Monte Carlo Dropout methods (Kendall et al., 2016; Roy et al., 2018b; DeVries and Taylor, 2018; Mehrtash et al., 2020; Burmeister et al., 2022; Iwamoto et al., 2021).
+To prevent confusion in already established terminology, the term model uncertainty should be exclusively used for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
+Task dependency
 It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context.
@@
 Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2).
-Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty Depeweg et al.
-(2018), such as commonly done in Active Learning (Section 6.2).
-Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk Apostolakis (1990); Helton (1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1).
-However, it was highlighted that the decision to model a specific type of uncertainty was largely influenced by how the problem was framed.
-To avoid confusion, this task-dependency should be noted when interpreting uncertainty.
-Interpretation of uncertainty is model dependent
-Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian & Ditlevsen, 2009).
+Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty (Depeweg et al., 2018), such as commonly done in Active Learning (Section 6.2).
+Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk (Apostolakis, 1990; Helton, 1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1).
+To avoid confusion, this task-dependency should be considered when interpreting uncertainty.
+Model dependency
+Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian and Ditlevsen, 2009).
 For instance, the expected entropy w.r.t.
-a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3)).
+a stochastic parameter variable is often assumed to represent the epistemic uncertainty as shown in Equation (3).
 However, when employing generative models, this behavior changes.
-The latent (feature) stochasticity variable is employed to model the aleatoric uncertainty.
-Hence, the remaining uncertainty, i.e.
-epistemic, is encapsulated in the expected entropy w.r.t.
-this latent variable (Equation (13)).
+As argued by Kahl et al.
+(2024), the predictive uncertainty can be decomposed as
+$H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{Z}|\mathbf{x},\bm{\theta}]}_{\text{aleatoric}}+\overbrace{\mathbb{E}_{p(\mathbf{z})}[\,H[\mathbf{Y}|\mathbf{Z},\mathbf{x},\bm{\theta}]\,]}^{{\text{epistemic}}}.$ (41)
+In contrast to Equation (3), the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty.
+This is because generative models operate under the assumption that the latent variable $\mathbf{Z}$ captures the inherent data ambiguity, thus representing the aleatoric uncertainty.
+The mutual information describes how much about this aleatoric uncertainty can be learned by observing the output $\mathbf{Y}$.
+If one were to know the optimal latent code $\mathbf{Z}$, the mutual information would be zero.
+In this case, all stochasticity in the output $\mathbf{Y}$ is perfectly represented by $\mathbf{Z}$, meaning the aleatoric uncertainty has been fully captured.
+Under this view, if all data uncertainty can be fully encapsulated by the latent variable $\mathbf{Z}$, there is no need to model this irreducible stochasticity in the predictive model $p(\mathbf{Y}|\mathbf{Z,\mathbf{x},\bm{\theta}})$, implying it corresponds to the epistemic uncertainty.
+This is consistent with the principle that epistemic uncertainty is the residual component isolated by subtracting the aleatoric uncertainty from the total predictive entropy.
 Hence, the calibration network of the GAN-based CAR model of Kassapis et al.
-(2021) (discussed in Section 4.2.1) actually quantified the epistemic rather than aleatoric uncertainty.
-This is dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Kahl et al., 2024; Hu et al., 2022; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023).
-The key message thus far is that, after selecting the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty type.
-Uncertainties are intertwined
+(2021) (discussed in Section 4.2.1) is actually more inclined to model the epistemic rather than aleatoric uncertainty.
+This dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA.
+In some works, it is argued that this approach models the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan and Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread and Jenkinson, 2022; Roshanzamir et al., 2023).
+However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022).
+Kahl et al.
+(2024) draw parallels with BNNs and demonstrate that
+$H[\mathbf{Y}|\mathbf{x},\bm{\theta}]=\underbrace{I[\mathbf{Y},\mathbf{t}|\mathbf{x},\bm{\theta}]}_{\text{epistemic}}+\overbrace{\mathbb{E}_{\mathbf{t}}[\,H[\mathbf{Y}|\mathbf{t},\mathbf{x},\bm{\theta}]\,]}^{{\text{aleatoric}}}.$ (42)
+where $t$ are sampled augmentations from the input space $\mathcal{T}$.
+A well-trained model should be invariant to data augmentations.
+Consequently, predictive variance across augmentations is treated as irreducible aleatoric uncertainty, which is quantified by a non-zero expected entropy over these predictions.
+Non-zero mutual information indicates the model has not yet learned invariance to the augmentations and can be reduced with further training, deeming it to be epistemic.
+This decomposition is analogous to the one for conventional BNNs (Equation 3), while it contrasts with the decomposition used for generative models (Equation 41).
+The key message thus far is that, after considering the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty.
+Intertwinedness
 Despite the seemingly rigorous and well-defined distinction, it can be seen that implementation rapidly blurs this notion.
-A systematic study has found that the ability to separate the uncertainties can strongly depend on the data Kahl et al.
-(2024).
-These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods de Jong et al.
-(2024); Mucsányi et al.
-(2024), indicating that they represent uncertainty of similar kind.
-Kendall & Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other.
-In probabilistic models, we can also find that the epistemic encapsulation methods have been used to model observer variability.
-Furthermore, using a combination of aleatoric end epistemic modeling techniques can also beneficial (Gao et al., 2022).
+A systematic study has found that the ability to separate the uncertainties can strongly depend on the data (Kahl et al., 2024).
+These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods (de Jong et al., 2024; Mucsányi et al., 2024), indicating that they represent uncertainty of similar kind.
+Kendall and Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other.
+In probabilistic models, we can also find that the parameter modeling methods, traditionally used for epistemic uncertainty, have been used to model observer variability, by for example using MC-Dropout or ensembling (see Table 1).
+Furthermore, using a combination of distribution and paramter modeling techniques can also beneficial (Gao et al., 2022).
 These findings further highlight the notion of the uncertainties being intertwined.
-On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions Kingma et al.
-(2015), further supporting this argument.
-Encapsulating vs. quantifying
-A common misconception pertains that modeling a specific uncertainty also implies that the predicted output represents that same uncertainty.
+On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions (Kingma et al., 2015), further supporting this argument.
+Modeling vs. quantifying
+A common misconception pertains that modeling a specific uncertainty also implies that the predicted output distribution represents that same uncertainty.
 This undermines the entropy decomposition in Equation (3) distinguishing between encapsulating, i.e.
@@
 Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties (Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018).
-At the very most, the output correlates with an increase in epistemic uncertainty.
-However, given the entanglement of both uncertainties, part of the predictive variance can be accredited to noise in the data.
-This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance Zhang et al.
-(2022c); Monteiro et al.
-(2020); Valiuddin et al.
-(2024b); Gao et al.
-(2022).
-However, epistemic uncertainty is often assumed to be explained away or assumed to be comparatively negligible with large datasets.
-Finally, we have found it to be common to interchangeably use the total and model uncertainty, especially in the context of Monte Carlo Dropout methods Kendall et al.
-(2016); Roy et al.
-(2018b); DeVries & Taylor (2018); Mehrtash et al.
-(2020); Burmeister et al.
-(2022); Iwamoto et al.
-(2021).
-To prevent confusion in already established terminology, the term model uncertainty should be exclusively used for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
-7.1.2 Pragmatic use of uncertainty
-Unless being the core contribution, there seems to be no compelling need to investigate uncertainty disentanglement for the applications of uncertainty in dense prediction.
-If performance on the downstream application is improved, it can be assumed that modeled uncertainty correlates with changes in variance of the predictive distribution.
-In such cases, no claims need to be made on the uncertainty type and neither is it required to isolate the two.
-Hence, as described in previous work, aleatoric and epistemic uncertainty should be used as to communicate whether the uncertainty is reducible or not (Der Kiureghian & Ditlevsen, 2009; Faber, 2005), to then select appropriate modeling methods.
-As argued by Kahl et al.
-(2024), claiming to quantify an uncertainty type would ideally require validation across all relevant applications.
-For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but it is often found in literature that either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022).
-As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often found to be evaluated on a single annotator or without the use of distribution-level statistics Wang et al.
-(2019a); Whitbread & Jenkinson (2022); Kendall & Gal (2017); Liu et al.
-(2022b); Schmidt et al.
-(2023); Savadikar et al.
-(2021).
-Hence, simply avoiding any serious claims related to either type alleviates this burden, and allows one to focus solely on the task at hand.
-In contrast to the work of Kahl et al.
-(2024), however, we do not consider calibration of models itself as a downstream task, but rather as a tool to gauge model reliability to be later used for other tasks.
-As a consequence, the pixel-wise uncertainties can be later used for possible downstream applications.
+This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance (Zhang et al., 2022d; Monteiro et al., 2020; Valiuddin et al., 2024b; Gao et al., 2022).
+Since predictive variance can encompass both uncertainties, a crucial distinction must be made.
+Targeting a specific uncertainty for a task through either parameter or feature modeling is different from truly quantifying a type in isolation, which may require separate validation.
 7.2 Spatial coherence
 The inter-pixel dependency of segmentation masks often introduces additional complexity in proper uncertainty quantification, questioning the commonly repurposed techniques from conventional image-level classification.
-In this section, we further dive into import details and challenges arising due to the added dimension(s) in both 2D and 3D data.
+In this section, we further dive into important details and challenges arising due to the added dimension(s) in both 2D and 3D data.
 Aggregation methods
-Quantifying the epistemic uncertainty requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (39)).
-In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values Mukhoti & Gal (2018); Camarasa et al.
-(2021); Bhat et al.
-(2022b); Mukhoti et al.
-(2020); Ma et al.
-(2024b); Shen et al.
-(2021); Li & Alstrøm (2020); Dechesne et al.
-(2021); Huang et al.
-(2018); Nair et al.
-(2020); Zepf et al.
-(2023b).
+The mutual information is often utilized to express the reducible, epistemic uncertainty.
+This requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (35)).
+In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values (Mukhoti and Gal, 2018; Camarasa et al., 2021; Bhat et al., 2022b; Mukhoti et al., 2020; Ma et al., 2024b; Shen et al., 2021; Li and Alstrøm, 2020; Dechesne et al., 2021; Huang et al., 2018; Nair et al., 2020; Zepf et al., 2023b).
 However, this asserts the assumption of a factorized categorical distribution, i.e.
 uncorrelated pixels, which is a strong assumption that ignores the core rationale for the field’s existence.
-Since this could explain failure cases of entropy-based uncertainty evaluation methods, it is important to shift towards contemporary generative models that enable exact likelihoods, improving the theoretical soundness of the evaluated entropies.
-Moreover, elements such as object class dependency, size, shape, proximity, contextual information etc.
+Hence, spatial coherence is a fundamental consideration and is frequently neglected in the uncertainty estimation frameworks used for Active Learning.
+The problem with factorized entropy can be demonstrated with a simple example: a binary segmentation of a circle on a $\times32$ grid.
+If we assume that the ’latent’ uncertainty is purely translational, Figure 13 visualizes the example distributions of possible segmentation maps.
+The true entropy of the generating source scales with the number of possible states $N$ as $\log_{2}(N)$.
+However, using a factorized approach severely overestimates the latent, true uncertainty.
+This behavior is a direct consequence of the subadditivity property of entropy, which states that
+$H[X_{1},X_{2},...,X_{D}]\leq\sum^{D}_{i}H[X_{i}].$ (43)
+(a) 22 states, true entropy $log_{2}(2)=1$ bit, factorized entropy $p_{i}\sum\log_{2}p_{i}\approx 112$ bits
+(b) 55 states, true entropy $log_{2}(2)\approx 2.32$ bits, factorized entropy $p_{i}\sum\log_{2}p_{i}\approx 114$ bits
+(c) 88 states, true entropy $log_{2}(8)=3$ bits, factorized entropy $p_{i}\sum\log_{2}p_{i}\approx 190$ bits
+Figure 13: This figure illustrates why factorized entropy is a poor proxy for the true uncertainty in structured images.
+The common method to estimate this entropy is by averaging multiple sampled binary segmentation masks, where the resulting pixel intensity indicates the average confidence value.
+The shapes shown possess only translational uncertainty.
+The factorized model fails to capture the object’s spatial coherence and leads to a significant overestimation of the true entropy.
+Importantly, this fallacy could explain failure cases of entropy-based uncertainty evaluation methods in tasks such as Active Learning where sample informativeness/diversity is crucial.
+Elements such as object class dependency, size, shape, proximity, contextual information etc.
 can all influence the decision-making of the aggregation method.
@@
 The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score.
-Some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021).
+Some empirical studies have explored other types of aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021).
 However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
 Volumetric uncertainty modeling
-In clinical settings, data is often acquired volumetrically but sliced into patches due to memory limits.
-With improved compute, 3D segmentation models have gained traction for preserving spatial continuity and reducing inter-slice inconsistencies.
+Processing volumetric clinical data knows two dominant strategies.
+The first, 2D slice processing, analyzes the volume as a stack of individual images, preserving full context within each slice but severing spatial continuity along the third dimension.
+The second, 3D patch processing, divides the volume into smaller ones, which maintains local 3D information but loses the global anatomical structure.
+As compute power increases, end-to-end 3D segmentation models are gaining traction because they can analyze increasingly more volume, preserving spatial continuity and reducing the inconsistencies inherent in these slice- or patch-based methods (Liu et al., 2025; Burmeister et al., 2022; Viviers et al., 2023c; Monteiro et al., 2020).
 As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue that 3D segmentation introduces or exacerbates previously mentioned issues related to entropy estimation and aggregation of pixels.
-Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023b; Chotzoglou & Kainz, 2019; Long et al., 2021b; Saha et al., 2021a; 2020; Hoebel et al., 2019).
-Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D volumes are often highly redundant Shen et al.
-(2021); Burmeister et al.
-(2022).
+Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023c; Chotzoglou and Kainz, 2019; Long et al., 2021a; Saha et al., 2021a; 2020; Hoebel et al., 2019).
+Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D volumes are often highly redundant (Shen et al., 2021; Burmeister et al., 2022).
 Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors.
-Furthermore, methodologies need to be developed in order to appropriately compare 2D- to 3D models.
-For example, the GED cannot be estimated volumetrically with 2D models.
+Moreover, methodologies need to be developed in order to appropriately compare 2D to 3D models.
+For example, a valid volumetric GED cannot be estimated with 2D models.
 Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data.
-Fortunately, research dedicated to probabilistic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al.
-(2025); Burmeister et al.
-(2022); Viviers et al.
-(2023b).
 7.3 Standardization
@@
 This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024).
-These practices are included in our recommended practices, as discussed in Section 6 per application.
+Our manuscript further contributes by introducing coherence and a unifying synthesis across the distinct disciplines of applications, tasks, and methodological developments, with these concepts being further detailed in our recommendations in Section 8.
 Despite the aforementioned points, efforts to quantify observer variability show some implicit consensus.
-Notably, the recommendations of Kahl et al.
-(2024) diverge from these practices, suggesting a more rigorous examination of this application.
-For instance, for the widely used LIDC-IDRI dataset (Armato, 2011), we recommend following the preprocessing protocol of Baumgartner et al.
-(2019), which is supported by publicly available code and has been adopted in several subsequent studies.
-Similarly, the Cityscapes dataset (Cordts et al., 2016a), comprising street-view images from German cities, has been used to introduce artificial class-level label ambiguity (Kassapis et al., 2021; Kohl et al., 2019b; Gao et al., 2022).
-We advise using the label-switching parameters proposed by Kohl et al.
-(2019b), whose preprocessing scripts are also available.
-Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), although they follow less standardized preprocessing.
-The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al.
-(2022).
-The evaluation metrics used across studies also diverge substantially.
-While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of correct empty segmentations (which cause NaNs) vary widely.
+Notably, some of the recommendations of Kahl et al.
+(2024) diverge from these practices, such as how to employ the widely used LIDC-IDRI dataset (Armato, 2011) or the Cityscapes Cordts et al.
+(2016a) dataset, suggesting a more rigorous examination of this application.
+Furthermore, the evaluation metrics used across various studies also diverge substantially.
+While GED and Hungarian Matching are commonly applied to assess model performance, key details, such as the choice of distance kernel, the number of predictive samples, and the handling of correct empty segmentations (which cause NaNs), vary widely.
 This inconsistency is particularly evident in the work of Zbinden et al.
@@
 See Table 6 for a comparison of benchmarks on the LIDC-IDRI dataset.
-We recommend using 16 predictive samples as the standard for evaluating the LIDC-IDRI (v2, see experiments section of Zbinden et al.
-(2023)) dataset.
-Additionally, correctly predicted empty segmentations should be assigned full score.
-Notably, the literature often overemphasizes metric improvement, which can incentivize models to simply replicate ground-truth masks rather than capture the underlying distribution.
-This risks undermining the original goal of predicting plausible, unseen segmentations, a textbook case of Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” and is further compounded by known biases in commonly used metrics (Kohl et al., 2019a; Zepf et al., 2023a; Valiuddin et al., 2024b; a).
-Hence, practitioners should remain cautious and complement quantitative evaluations with qualitative assessments from domain experts.
 Table 6: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table extended from (Zbinden et al., 2023). Indicates evaluation results not obtained from the original work.
@@
 HPU-Net (Kohl et al., 2019a) $0.270\pm 0.010^{*}$ $0.530\pm 0.010^{*}$ $0.270\pm 0.010^{*}$ $0.530\pm 0.010^{*}$
-PhiSeg (Baumgartner et al., 2019) $0.262\pm 0.000^{*}$ $0.586\pm 0.000^{*}$
-SSN (Monteiro et al., 2020) $0.259\pm 0.000^{*}$ $0.558\pm 0.000^{*}$
+PhiSeg (Baumgartner et al., 2019) $0.262\pm 0.000^{*}$ $0.586\pm 0.000^{*}$ - -
+SSN (Monteiro et al., 2020) $0.259\pm 0.000^{*}$ $0.558\pm 0.000^{*}$ - -
 CAR (Kassapis et al., 2021) - - $0.264\pm 0.002^{\phantom{*}}$ $0.592\pm 0.005^{\phantom{*}}$
-JProb. U-Net (Zhang et al., 2022d) - - $0.262\pm 0.000^{\phantom{*}}$ $0.585\pm 0.000^{\phantom{*}}$
-PixelSeg (Zhang et al., 2022c) $0.243\pm 0.010^{\phantom{*}}$ $0.614\pm 0.000^{\phantom{*}}$ $0.260\pm 0.000^{\phantom{*}}$ $0.587\pm 0.010^{\phantom{*}}$
-MoSE (Gao et al., 2022) $0.218\pm 0.001^{\phantom{*}}$ $0.624\pm 0.004^{\phantom{*}}$
-AB (Chen et al., 2022) $0.213\pm 0.001^{*}$ $0.614\pm 0.001^{*}$
-CIMD (Rahman et al., 2023) $0.234\pm 0.005^{*}$ $0.587\pm 0.001^{*}$
+JProb. U-Net (Zhang et al., 2022b) - - $0.262\pm 0.000^{\phantom{*}}$ $0.585\pm 0.000^{\phantom{*}}$
+PixelSeg (Zhang et al., 2022d) $0.243\pm 0.010^{\phantom{*}}$ $0.614\pm 0.000^{\phantom{*}}$ $0.260\pm 0.000^{\phantom{*}}$ $0.587\pm 0.010^{\phantom{*}}$
+MoSE (Gao et al., 2022) $0.218\pm 0.001^{\phantom{*}}$ $0.624\pm 0.004^{\phantom{*}}$ - -
+AB (Chen et al., 2022) $0.213\pm 0.001^{*}$ $0.614\pm 0.001^{*}$ - -
+CIMD (Rahman et al., 2023) $0.234\pm 0.005^{*}$ $0.587\pm 0.001^{*}$ - -
 CCDM (Zbinden et al., 2023) $0.212\pm 0.002^{\phantom{*}}$ $0.623\pm 0.002^{\phantom{*}}$ $0.239\pm 0.003^{\phantom{*}}$ $0.598\pm 0.001^{\phantom{*}}$
@@
 Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos.
+Similarly, these principles can also translate tp Reinforcement Learning, where uncertainty guides an agent’s exploration policy to accelerate learning in complex environments.
 However, key limitations remain and require further research.
 Model sensitivity and baselines
-While deep learning-based segmentation typically involves high-dimensional data and large datasets, active learning requires a well-generalized model trained on limited data.
-This mismatch renders performance highly sensitive to budget constraints, model architecture, hyperparameters, and regularization Mittal et al.
-(2019); Kirsch (2024); Munjal et al.
-(2022); Kirsch et al.
-(2019); Gaillochet et al.
-(2023).
-Furthermore, random sampling, particularly when combined with MC Dropout, often proves to be a surprisingly strong baseline that is difficult to outperform Kahl et al.
-(2024); Mittal et al.
-(2019); Lüth et al.
-(2023); Burmeister et al.
-(2022); Siddiqui et al.
-(2020); Kim et al.
-(2021); Sinha et al.
-(2019); Nath et al.
-(2020), especially with highly imbalanced data Ma et al.
-(2024a).
+Active Learning requires a well-generalized model trained on limited data, rendering performance highly sensitive to budget constraints, model architecture, hyperparameters, and regularization (Mittal et al., 2019; Kirsch, 2024; Munjal et al., 2022; Kirsch et al., 2019; Gaillochet et al., 2023).
+Furthermore, random sampling, particularly when combined with MC Dropout, often proves to be a surprisingly strong baseline that is difficult to outperform (Kahl et al., 2024; Mittal et al., 2019; Lüth et al., 2023; Burmeister et al., 2022; Siddiqui et al., 2020; Kim et al., 2021; Sinha et al., 2019; Nath et al., 2020), especially with highly imbalanced data (Ma et al., 2024a).
 This observation can partly be attributed to the fact that MC Dropout likely fails to capture true Bayesian uncertainty (see Section 7.5), inviting exploration of other techniques.
 Sample informativeness
-Besides uncertainty, the diversity of selected samples is equally critical.
-For instance, a set of highly uncertain samples may share similar semantic content, leading the model to overfit a narrow and homogeneous region of the data distribution during early training.
+A set of highly uncertain samples may share similar semantic content, leading the model to overfit a narrow and homogeneous region of the data distribution during early training.
 Hence, sample selection should be based on both uncertainty and diversity (Ozdemir et al., 2021; Wu et al., 2021; Jensen et al., 2019; Burmeister et al., 2022).
@@
 (2021) leverage the latent space of a VAE to select representative samples, whereas other studies use feature vectors from prediction models (Wu et al., 2021; Burmeister et al., 2022).
-Owing to their significantly greater representational capacity, modern generative models, as discussed in Section 7.6, merit experimental investigation together with segmentation specific conditioning.
 Furthermore, the sample diversity offered by such models can be effectively combined with adversarial training (Mahapatra et al., 2018; Sinha et al., 2019).
@@
 As discussed in Section 4.1, using the predictive entropy of SoftMax-based uncertainty requires proper calibration.
-Unfortunately, this is rarely verified in practice Kasarla et al.
-(2019); García Rodríguez et al.
-(2020); Xie et al.
-(2022); Wu et al.
-(2021); Burmeister et al.
-(2022); Gaillochet et al.
-(2023).
-Furhtermore, most works rely on heuristics to aggregate the pixel-level information such as closeness to edges, boundaries or the use of super-pixels Gorriz et al.
-(2017); Kasarla et al.
-(2019); Ma et al.
-(2024b); Gaillochet et al.
-(2023); Wu et al.
-(2021); Kasarla et al.
-(2019); Bengar et al.
-(2021).
+Unfortunately, this is rarely verified in practice (Kasarla et al., 2019; García Rodríguez et al., 2020; Xie et al., 2022; Wu et al., 2021; Burmeister et al., 2022; Gaillochet et al., 2023).
+Furhtermore, most works rely on heuristics to aggregate the pixel-level information such as closeness to edges, boundaries or the use of super-pixels (Gorriz et al., 2017; Kasarla et al., 2019; Ma et al., 2024b; Gaillochet et al., 2023; Wu et al., 2021; Kasarla et al., 2019; Bengar et al., 2021).
 Moreover, the mutual information is often quantified by plain summation (i.e.
-assuming pixel independence) Shen et al.
-(2021); Ma et al.
-(2024b).
+assuming pixel independence) (Shen et al., 2021; Ma et al., 2024b).
 We hypothesize that these ad-hoc aggregation techniques, as mentioned in Section 7.2, are a key limiting factor to uncertainty-based Active Learning.
 The Cold-Start Problem
-Finally, the “cold start” problem Nath et al.
-(2022); Houlsby et al.
-(2014); Yuan et al.
-(2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed.
+Finally, the “cold start” problem (Nath et al., 2022; Houlsby et al., 2014; Yuan et al., 2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed.
 Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty.
@@
 (2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.
+7.5 Method suboptimality
+To provide a necessary contrast to the benefits of many models, this section discusses key methodological limitations.
+These limitations will serve as guidance for readers selecting the most optimal methods and will be utilized when we present our recommendations in Section 8.
+Invertible covariance of SSNs
+Stochastic Segmentation Networks (SSNs), operate by placing a multivariate gaussian over the classification head of an existing model.
+However, they can suffer from training instability primarily due to the requirement that the covariance matrix remain invertible.
+To address this critical limitation, the authors propose two distinct remedial strategies: masking out the background to prevent exploding variances and employing uncorrelated gaussians when the covariance matrix becomes singular, which may impact model performance and theoretical purity.
+GAN training
+The Calibrated Adversarial Refinement network (CAR, as treated in Section 4.2.1), on the other hand, is a less commonly used method in literature.
+This can perhaps be accredited due to the well-known instability of GAN-based models (Arjovsky and Bottou, 2017).
+GANs often require additional heuristic terms in the loss function to ensure stable training.
+CAR is no exception to this, as its objective function combines four separate losses.
+The multiplicity of individual losses is critical to the model’s success.
+Nonetheless, this complex objective function suggests the training process may pose substantial fine-tuning challenges, potentially explaining the model’s limited practical adoption.
+Mode collapse with ELBOs
+Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, Variational Autoencoders (VAEs) seem to be the most popular choice for the task of encapsulating Observer Variability.
+(See Table 1 in Section 6.1).
+Nonetheless, the shortcomings of VAEs are well known.
+For example, such models suffer from inference suboptimality related to ELBO optimization (Cremer, 2018; Zhao et al., 2017) and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of mode collapse (Valiuddin et al., 2024a; b; Qiu and Lui, 2020).
+This has been hypothesized to be caused by excessively strong decoders (Chen et al., 2016) and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective (Kohl et al., 2019a), residual connections (Kohl et al., 2019a; Baumgartner et al., 2019), or deep supervision (Baumgartner et al., 2019) are required for generalization.
+Limitation of sequential modeling
+The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models.
+This is unexpected, as both models outperforms the latter.
+This is likely due to their tedious sequential inference procedure (Song et al., 2020; Zheng et al., 2023; Meng et al., 2023; Zhang et al., 2022d), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split.
+Despite these shortcomings, the strengths of DDPMs should not be overlooked.
+Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions.
+Discrete vs. continuous
+It can be noted in Table 6 that the best-performing Denoising Diffusion Probabilistic Models (DDPMs) are discrete in nature (Chen et al., 2023a; Zbinden et al., 2023).
+While it is debatable whether shifting to categorical distributions is required for complex image generation (Chen et al., 2022).
+This can visually be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 14).
+However, Table 6 shows AB (Chen et al., 2022) performs almost identically to CCDM (Zbinden et al., 2023), suggesting that straightforward thresholding of continuous models have little-to-no performance sacrifice compared with explicitly modeling discrete distributions.
+These observations suggests that the modeling intuition of a categorical latent space, despite its fundamental soundness, lacks demonstrable merit.
+(a) LIDC-IDRI (Armato, 2011).
+(b) CityScapes (Cordts et al., 2016a).
+Figure 14: Continuous vs.
+categorical forward diffusion process with cosine noise scheduling (Nichol and Dhariwal, 2021).
+Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
+Criticism of MC Dropout
+The majority of literature on parameter distribution modeling appears to rely on approximations of variational inference (VI).
+Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation.
+However, MC Dropout has faced substantial criticism (Folgoc et al., 2021; Osband, 2016; Kingma et al., 2015; Gal et al., 2017a).
+For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality (Folgoc et al., 2021).
+It is also sensitive to model size and dropout rate rather than the observed data (Osband, 2016; Kingma et al., 2015).
+To address these issues, some alternatives attempt to learn the dropout rate (Kingma et al., 2015; Gal et al., 2017a).
+In cases of improved Model Generalization (Section 6.4), it is also debatable whether the performance gains stem from better uncertainty modeling or from other factors.
+For instance, both ensembling and MC Dropout introduce a model combination effect, effectively averaging multiple networks.
+MC Dropout also resembles placing an $L_{2}$ penalty on weights (Gal and Ghahramani, 2016).
+While these effects may improve generalization, similar gains might be achieved with simpler, more efficient regularizers.
+This suggests that perhaps it is the inherent stochasticity of the methods that is the key beneficial factor.
+Still, if achieving peak performance is the primary goal, uncertainty modeling is likely not the most efficient route due to its added computational cost.
+Instead, the improved performance may simply be an unintended side effect of techniques chosen for a different purpose.
+7.6 Data dependency
+In contrast to many other domains within computer vision, this field exhibits a notably greater absence of a single state-of-the-art method.
+Performance is heavily contingent upon multiple factors such as modeling context and application.
+However, data dependency often remains unexamined.
+Especially in cases where parameter distributions are modeled.
+For instance, MC dropout was found to perform better than ensembling (Hoebel et al., 2020; Roy et al., 2018b), while in other works ensembling excels (Mehrtash et al., 2020; Ng et al., 2022).
+Also, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout (Mukhoti and Gal, 2018).
+All things considered, the preference for a particular methodology seems to carry a strong data-dependency (Dahal et al., 2020; Jungo and Reyes, 2019; Burmeister et al., 2022; Ng et al., 2022; Kahl et al., 2024).
+For feature modeling, some models in Table 6 may appear to be overall best performers.
+However, their validation is often limited to binary, multi-annotated data, thereby compromising any claims of generalization to other data types, such as multi-class or single-annotated scenarios.
+Current research often focuses on incorporating uncertainty into already high-performing systems rather than comparing distinct methods.
+This trend necessitates a more rigorous, data-driven perspective to validate and understand the optimal uncertainty modeling strategy.
+Hence, a valuable benchmark study should comprehensively compare all available methods (including both feature and parameter modeling approaches) across all relevant task with special focus on the data dependencies.
 7.7 Segmentation backbone
-Due to emphasis on the uncertainty methods, the backbone feature extractor is often an overlooked element in uncertainty modeling.
+Due to emphasis on the method, task and application, the backbone feature extractor is often an overlooked element in uncertainty modeling.
 For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023).
-Nevertheless, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Cheng et al., 2022).
+Nevertheless, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang and Yang, 2021; Cheng et al., 2022).
 With the exception of experiments in Kahl et al.
@@
 The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021).
-Therefore, we recommend integrating the discussed methods into similar architectures to achieve both strong classification performance and effective ambiguity modeling.
-Also, CNNs have benefited from the recent developments in Transformers.
+Nonetheless, CNNs have benefited from the recent developments in Transformers.
 For instance, ConvNext takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c).
-These same inductive biases have been added to ViTs through CNN adapters to also enhance performance Chen et al.
-(2023b).
-7.6 Modeling feature distributions
-In many instances, latent-variable models are employed to model the observer variability.
-To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches.
-A summary of this discussion is presented in Table 7.
-Note that, while classified PixelCNNs and SSNs are classified as pixel-level methods, both can also be phrased as a latent-variable model.
-ELBO suboptimality
-Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification.
-Nonetheless, the shortcomings of VAEs are well known.
-For example, such models suffer from inference suboptimality related to ELBO optimization Cremer (2018); Zhao et al.
-(2017) and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse Valiuddin et al.
-(2024a; b); Qiu & Lui (2020).
-This has been hypothesized to be caused by excessively strong decoders Chen et al.
-(2016) and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective Kohl et al.
-(2019a), residual connections Kohl et al.
-(2019a); Baumgartner et al.
-(2019), or deep supervision Baumgartner et al.
-(2019) are required for generalization.
-A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions.
-Limitation of sequential modeling
-The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models.
-This is unexpected, as both models outperforms the latter.
-This is likely due to their tedious sequential inference procedure Song et al.
-(2020); Zheng et al.
-(2023); Meng et al.
-(2023); Zhang et al.
-(2022c), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split.
-Despite these shortcomings, the strengths of DDPMs should not be overlooked.
-Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions.
-Furthermore, recent advances in faster sampling methods, such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al.
-(2022), make these models promising candidates for further exploration.
-Discrete vs. continuous
-It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al.
-(2023a); Zbinden et al.
-(2023).
-While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al.
-(2022).
-Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
-However, Table 1 shows AB Chen et al.
-(2022) performs almost identically to CCDM Zbinden et al.
-(2023), suggesting that straightforward thresholding of continuous models have little-to-no performance sacrifice compared with explicitly modeling discrete distributions.
-These observation call for further investigation of the merits of categorical distributions in segmentation settings.
-(a) LIDC-IDRI Armato (2011).
-(b) CityScapes Cordts et al. (2016a).
-Figure 13: Continuous vs.
-categorical forward diffusion process with cosine noise scheduling Nichol & Dhariwal (2021).
-Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
-Others
-Stochastic Segmentation Networks (SSNs) are a simple, fast, and model-agnostic approach for modeling uncertainty.
-However, they can suffer from training instability due to the requirement that the covariance matrix be invertible.
-To address this limitation, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular.
-Despite these efforts, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space.
-CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017).
-GANs often require additional heuristic terms in the loss function to ensure stable training.
-CAR is no exception to this, as its objective function combines four separate losses.
-Recommendation
-We advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models.
-Generative modeling is a rapidly evolving field and any such model can be modified for segmentation through intricate conditioning.
-For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation.
-Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling Lipman et al.
-(2022).
-In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022).
-Table 7: Comparison between generative models that model aleatoric uncertainty.
-Method Advantages Disadvantages Examples
-SSN Model agnostic, explicit likelihoods, fast sampling possibly unstable, memory intensive (Monteiro et al., 2020; Kahl et al., 2024; Zepf et al., 2023a; Philps et al., 2024; Kahl et al., 2024)
-PixelCNN Explicit, exact likelihoods Sequential sampling, memory intensive Zhang et al. (2022c)
-GAN Fast sampling, flexible Unstable training, poorly defined objective, implicit likelihoods (Kassapis et al., 2021; Isola et al., 2017)
-VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap (Kohl et al., 2019b; Selvan et al., 2020; Valiuddin et al., 2021; 2024a; Qiu & Lui, 2020; Bhat et al., 2022a; 2023; Valiuddin et al., 2024b; Zhang et al., 2022b; Kohl et al., 2019a; Baumgartner et al., 2019; Hu et al., 2023; Liu et al., 2022b; Fischer et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Viviers et al., 2023b; a; Chotzoglou & Kainz, 2019; Saha et al., 2021a; b; Long et al., 2021a; Schmidt et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Zepf et al., 2024; 2023a; Hu et al., 2022)
-DDPM Flexible, expressive Sequential sampling (Chen et al., 2023a; Zbinden et al., 2023; Rahman et al., 2023)
-7.7 Modeling parameter distributions
-In the previous sections, we reviewed feature modeling methods, typically used to quantify observer variability.
-Other downstream applications usually rely on epistemic uncertainty modeling, i.e.
-modeling parameter distributions, which will be discussed in this section.
-Criticism of MC Dropout
-Nearly all literature on epistemic uncertainty quantification relies on approximations of variational inference (VI).
-Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation.
-However, MC Dropout has faced substantial criticism Folgoc et al.
-(2021); Osband (2016); Kingma et al.
-(2015); Gal et al.
-(2017a).
-For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality Folgoc et al.
-(2021).
-It is also sensitive to model size and dropout rate rather than the observed data Osband (2016); Kingma et al.
-(2015).
-To address these issues, some alternatives attempt to learn the dropout rate Kingma et al.
-(2015); Gal et al.
-(2017a).
-Still, uncertainty estimates from MC Dropout should be treated as a convenient byproduct—not a reliable core feature.
-When precise uncertainty quantification is critical, such methods are generally insufficient and best avoided.
-When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary.
-In such cases, methods like MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation.
-However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors.
-For instance, both ensembling and MC Dropout introduce a model combination effect—effectively averaging multiple networks.
-MC Dropout also resembles placing an $L_{2}$ penalty on weights.
-While these effects can improve generalization, similar gains might be achieved with simpler, more efficient regularizers.
-Perhaps it is the stochasticity itself that helps.
-Still, if performance is the main goal, uncertainty modeling is likely not the most efficient route due to its added computational cost.
-Instead, improved performance may simply be a side effect of these techniques.
-Comparative studies
-Finally, the optimal method for uncertainty quantification has yet to be determined.
-In some works, MC dropout was found to perform better than ensembling Hoebel et al.
-(2020); Roy et al.
-(2018b), while in other works ensembling excels Mehrtash et al.
-(2020); Ng et al.
-(2022).
-Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout Mukhoti & Gal (2018).
-All things considered, the preference for a particular methodology seems to carry a strong data-dependency Dahal et al.
-(2020); Jungo & Reyes (2019).
-A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets.
-In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e.
-why ensembling or MC Dropout is often preferred).
-Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble for language modeling (He, 2024).
-Future work should also experiment with this.
-Alternatives to Variational Inference
+These same inductive biases have been added to ViTs through CNN adapters to also enhance performance (Chen et al., 2023b).
+Building upon these observations, we conclude that future research should actively explore the integration of state-of-the-art Vision Transformer (ViT) and hybrid CNN-Transformer backbones to advance the field.
+7.8 Contemporary uncertainty
+While current efforts frequently leverage established uncertainty modeling techniques for segmentation, the practical integration involves significant challenges related to ensuring architectural and task compatibility.
+To accelerate progress, the segmentation community should actively explore and integrate newly developed uncertainty quantification methodologies, rather than relying solely on standard, established practices.
+VI alternatives
 Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior.
@@
 Zepf et al.
-(2023b) explored the impact of the Laplace Approximation on segmentation networks; however, beyond this work, the method has received limited attention in the field.
+(2023b) explored the impact of the Laplace Approximation on segmentation networks.
+However, beyond this work, the method has received limited attention in the field.
 Sampling-free uncertainty
@@
 Mukhoti et al.
-(2023) show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling.
+(2023) show that Gaussian Discriminant Analysis after training with a SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling.
 Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass.
@@
 The use of Evidential Deep Learning has seen success in conventional classification problems (Sensoy et al., 2018), and Ancha et al.
-(2024) recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model.
+(2024) recently applied this concept to segmentation attempt to decouple aleatoric and epistemic uncertainty within a single model.
 However, follow-up research in this direction remains limited to date.
@@
 By using an additional calibration set, CP transforms heuristic model confidence scores into rigorous uncertainty estimates.
-This technique is particularly renowned for being model-agnostic, simple, and highly flexible (Angelopoulos & Bates, 2021).
+This technique is particularly renowned for being model-agnostic, simple, and highly flexible (Angelopoulos and Bates, 2021).
 Recently, CP has been applied to segmentation tasks (Wieslander et al., 2020; Brunekreef et al., 2024; Mossina et al., 2024; Wundram et al., 2024a), benefiting from these advantages and reflecting the growing interest in the framework for structured prediction.
+However, the validity of conformal prediction requires exchangeable draws for train, calibration, and test data.
+In practical settings test data often comes with distribution shift, which can coverage to fall below the nominal level and predictions can become overconfident.
+Such overconfidence yields prediction sets that are too tight, suppressing deferrals and increasing missed or undersegmented regions which can be especially detrimental in safety-critical tasks.
 We encourage further research to explore novel applications and benchmark CP against existing architectures.
+Generative models
+Given the established time-lag in the adoption of SOTA techniques from the broader generative modeling field, there remains a persistent opportunity for research to effectively bridge the gap by translating the latest architectural and methodological improvements from image generation into more performant and robust segmentation solutions.
+For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation.
+Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling compared to DDPMs (Lipman et al., 2022).
+7.9 Towards complex scene understanding
+A clear trend shows that most uncertainty applications remain narrowly focused on binary semantic segmentation.
+While limited work exists in probabilistic instance segmentation (Kohl et al., 2019a; Morrison et al., 2019), there is a notable absence of research addressing uncertainty in this domain, as well as in panoptic segmentation.
+Furthermore, work involving uncertainty across multiple classes is also scarce.
+This represents a critical gap, given that contemporary segmentation research is increasingly centered on multiclass problems across all segmentation types (Kerssies et al., 2025), often even incorporating class hierarchies (Atigh et al., 2022).
+Consequently, the significance of research focused solely on binary-class semantic segmentation is diminishing.
+8 Guidelines and recommendations
+This section synthesizes the preceding discussions into practical guidelines and recommendations.
+We will first critically examine specific methods and offer criteria for selecting the optimal approach for different tasks.
+Subsequently, we will discuss the essential topic of evaluating uncertainty and the potential biases that can arise.
+The section concludes with a more philosophical look at what constitutes truly meaningful uncertainty in machine learning.
+8.1 Method selection
+The proposed framework is designed to simplify the complex landscape of available techniques by deconstructing the decision into a series of manageable questions.
+By systematically addressing these questions, practitioners can navigate the available options more effectively and clearly justify their final choice of method.
+Decision framework: 1.
+Baseline: Does the vanilla backbone reach sufficient performance?
+(Section 8.1.1) 2.
+Reducible: Is the target uncertainty reducible?
+(Section 8.1.2) 3.
+Purpose: Is the uncertainty auxiliary or crucial to the main task?
+(Section 8.1.3) 4.
+Data regime: Are labels multi-annotated or multi-class?
+(Section 8.1.4) 5.
+Budget: Are there budgetary constraints that limit compute?
+(Table 7) For each element, evaluate potential method with criteria: Prioritizing methods with stronger evidence in literature.
+Considering the theoretical soundness.
+Favoring architecture-agnostic and post-hoc designs.
+Evaluating computational efficiency, prioritizing methods that offer fast sampling time.
+The final two points address the modern paradigm of using large, pre-trained models.
+These models are typically fine-tuned once for a specific task but are then deployed for continuous use.
+While the initial training is a significant one-time cost, deployment involves ongoing inference.
+Therefore, optimizing for inference efficiency is paramount and will lead to substantial long-term savings in computational resources.
+A comprehensive comparison of available methods is provided in Table 7, which will be used as a key reference when making recommendations.
+The complete selection process for our proposed framework is outlined in the flowchart in Figure 15.
+We now proceed to detail each component of this framework in the subsequent sections.
+{forest}
+Figure 15: Flowchart illustrating the complete framework for selecting the appropriate uncertainty modeling method.
+8.1.1 Baseline
+Probabilistic segmentation models perform best when they are built on a robust, capable backbone.
+Hence, it is crucial to establish a strong deterministic baseline first.
+If the underlying model underfits the data or fails to learn the task, adding a probabilistic layer will not solve this issue.
+Only when the baseline is solid can probabilistic modeling effectively provide calibrated uncertainty and support decision-making.
+Besides model backbone, improvements through data collection, parameter tuning, and different training schemes can also be explored.
+8.1.2 Reducibility of uncertainty
+Unless it is the core scientific contribution, we argue that disentangling uncertainty sources is often unnecessary for dense prediction applications.
+The primary benchmark for success is improved performance on a downstream task.
+If a model’s total uncertainty score achieves this, its practical value is already established, regardless of our ability to precisely attribute that uncertainty to its underlying sources.
+This pragmatic view avoids the need for complex validation methods designed solely to prove disentanglement.
+Instead, the distinction between aleatoric and epistemic uncertainty is most valuable as a conceptual guide to communicate whether the uncertainty is reducible and, consequently, to select the most appropriate modeling strategy, as noted in prior work (Der Kiureghian and Ditlevsen, 2009; Faber, 2005).
+Disentanglement shouldn’t be mistaken for the primary objective, but might be a useful diagnostic tool.
+For example, it makes sense to assume disentanglement of the uncertainty to choose the appropriate method.
+In a theoretical view, a model that perfectly captures all learnable structure would have its remaining prediction error driven solely by irreducible noise.
+This suggests that the fitted parameters can be treated as near-point estimates, and consequently, most residual uncertainty would be expected to arise from the predictive distribution rather than from parameter variability.
+In contrast, if we were to model the reducible noise within the parameter distribution by markedly narrowing the predictive distribution while the parameter variance remains materially above zero, we would be suggesting that a single sample from the parameter distribution could yield predictions with near-zero variance.
+This seems in tension with the premise of inherent data noise, since some uncertainty should still appear in predictions when the data-generating process is noisy.
+Thus, although the disentanglement of uncertainty is a theoretical debate (see Section 7.1), both our analyses and literature overview conclude that it remains a useful tool for method selection.
+Specifically, by modeling parameter distributions for reducible uncertainty and feature distributions for irreducible uncertainty.
+While a conventional neural network, by contrast, should theoretically be able to model both uncertainties within the feature-based output distribution, the empirical failure of this simple approach to provide robust uncertainty quantification is, in fact, the central driving force behind the entire field of study.
+8.1.3 Purpose of modeling
+Once a decision has been made on the uncertainty reducibility, the next step is to determine its purpose.
+For cases where the uncertainty is irreducible, our literature overview has identified a single use case: estimating Observer Variability (Section 6.1).
+For reducible uncertainty, this overview identified three key tasks.
+For the task of Model Generalization (Section 6.4), uncertainty is not the primary objective, but a beneficial byproduct of the modeling approach.
+In contrast, for Active Learning and Model Introspection (Section 6.2 and 6.3, respectively), uncertainty quantification is the core objective and must therefore be critically sound and theoretically justified.
+Observer Variability
+As demonstrated in Table 6, the top-performing methods for this task are SSN and DDPM-based approaches.
+Given its model-agnostic nature, fast sampling capabilities, compatibility with 3D data, and strong recent benchmarks (Kahl et al., 2024; Monteiro et al., 2020), we recommend prioritizing SSNs.
+Should the performance prove insufficient, further exploration based on specific data characteristics and computational constraints is warranted.
+As discussed in Section 7.6, the existing literature on the data dependency of these specific methods is limited.
+However, we will provide an overview and discussion in the following section based on the available studies.
+Model Generalization
+Though their theoretical grounding are limited (see discussion in Section 7.5, a pragmatic first step is to use either MC Dropout or TTA.
+Both methods are simple, inexpensive, and have been shown to be effective.
+Their limited theoretical support is not a major concern here, as the uncertainty is auxiliary to the main task of improving generalization.
+Nevertheless, MC dropout has more evidence in literature to perform well (see Table 4) and is generally better understood (i.e.
+the connection to variational inference made in Section 5.1) than TTA (its connection to Bayesian learning is under discussion as mentioned in Section 7.1).
+Active Learning & Model Introspection
+In contrast to Model Generalization, the reducible uncertainty is a crucial factor for both tasks.
+MC Dropout remains overwhelmingly the most popular uncertainty estimation method (see Tables 2 and 3) due primarily to its minimal implementation overhead, making it an easy baseline for numerous papers.
+Despite this prevalence, studies have reported that the simple random sampling baseline in Active Learning remains a highly competitive approach (Sinha et al., 2019).
+Recent benchmarks consistently show that ensemble methods are superior (Kahl et al., 2024), and Variational Inference (VI) closely trails ensembling while clearly outperforming MC Dropout(Ng et al., 2022).
+Since VI offers a more theoretically sound Bayesian approach that is also more cost-efficient than full ensembling, we recommend researchers adopt rigorous methods like VI for uncertainty characterization.
+Furthermore, caution is warranted concerning the aggregation methods detailed in Section 7.2.
+Namely, entropy calculation, often used to quantify uncertainty severity, is frequently severely overestimated.
+Future research must therefore pursue fundamentally sound methodologies that avoid these aggregation pitfalls.
+8.1.4 Nature of the data
+Our analysis in Section 7.6 indicates that uncertainty modeling is strongly data-dependent and remains largely inconclusive for modeling parameter distributions.
+Nonetheless, a limited but practical ordering can be established for modeling irreducible uncertainty via feature distribution methods, using criteria like class count, labels, and data dimensions.
+Multi-label, binary class
+The LIDC-IDRI benchmark heavily influences current literature, leading to a focus on (H)VAE-based methods.
+However, as shown in Table 6, DDPMs achieve superior scores on this key dataset.
+While (H)VAE approaches remain more popular for other datasets, the lack of robust benchmarking makes broader conclusions difficult.
+We therefore recommend DDPMs for superior performance, but recognize that (H)VAEs remain a viable alternative when constrained by computational limits.
+Single-label, binary class
+When evaluating single-label problems, accounting for ground-truth uncertainty (i.e., observer variability) is critical.
+The only rigorous verification method involves ablation studies on multi-annotated datasets where the model is trained only on a single mask.
+Since these crucial studies are performed almost exclusively on (H)VAE-based approaches (e.g., (Zhang et al., 2022b; Gao et al., 2022; Zepf et al., 2023a; Selvan et al., 2020), plus one instance with PixelSeg (Zhang et al., 2022d)), our recommendation favors (H)VAE models due to this strong accumulated evidence.
+However, given their superior mode coverage (Xiao et al., 2021), there is no indication that DDPMs would perform poorly in this specific ablation.
+Single label, multi-class
+Crucially, we have not found a multi-annotated, multi-class dataset used for ablation by training only on a single mask.
+In the absence of this rigorous evidence, we must draw inferences from related applications, such as Model Generalization.
+Within this context, Table 4 suggests that DDPMs are a popular choice.
+Multi-label, multi-class
+In the rare instances of multi-class, multi-annotated problems, such as the Cityscapes dataset when combined with artificial labels, (H)VAEs currently have the most supporting evidence (Kohl et al., 2019b; Gao et al., 2022).
+Volumetric data
+For problems involving 3D data, the primary evidence strongly supports the use of (H)VAE-based methods (Viviers et al., 2023a; b; Long et al., 2021b).
+This body of work encompasses both multi-class single-annotated and binary-class multi-annotated datasets.
+Given that this is the most substantial lead, outside of SSNs, we recommend the adoption of the (H)VAE model architecture for 3D uncertainty estimation.
+Table 7: Comparison between methods.
+NN represents number of parallel or sequential models in ensembles and DDPM denoising, respectively.
+Params.
+indicate the number of parameters required for feature-based methods.
+For parameter-based modeling, this indicates how the parameter count scales with the uncertainty method.
+Samples indicate how many evaluations are needed for a single forward pass for feature-based modeling.
+For parameter-based modeling, this indicates if the uncertainty evaluation quality scales with the number of forward passes of parameter-based approaches.
+Method Advantages Limitations Params. Samples Likelihoods Extra notes
+——–FEATURE MODELING——–
+SSN Model agnostic, fast sampling possibly unstable $\sim$41M 1 explicit, appr. Occasional diagonal covariance during training. No prior, likelihoods impractical.
+PixelCNN High sample quality Sequential sampling, weak global structure $\sim$3.4M $H\times W$ Explicit, exact
+CAR Fast sampling, flexible Unstable training, poorly defined objective, extensive tuning $\sim$43M 1 implicit Additional losses for segmentation
+(H)VAE Fast sampling, flexible, interpretable latent space Mode/posterior collapse, amortization gap $\sim$(19)5M 1 appr. lower bound additional deep supervision, GECO objective
+DDPM Flexible, expressive Sequential sampling $\sim$30M $\times N$ appr. lower bound Compute can be reduced with DDIMs at cost of sample quality
+——–PARAMETER MODELING——–
+LA Model agnostic, no extra parameters Local approx., full posterior infeasible $\times 1$ No Explicit exact Usually on last layer
+Explicit VI Full weight posterior; principled uncertainty slow convergence $\times$2 No appr. lower bound Usually on selected layers
+Ensemble simple, parallelization expensive, member correlation risk $\times N$ Yes Implicit variants: M-Heads, MoE
+MC dropout Minimal code change; regularizes; model-agnostic poor uncertainty, not principled $\times 1$ Yes Implicit variants: concrete, drop-weights, theoretical groundedness to be disputed
+TTA Zero training changes; boosts robustness boundary errors, aggregation sensitivity $\times 1$ Yes Implicit
+8.2 Evaluation and reproducibility
+Based on our literature overview and in-depth discussions, this section outlines general and task-specific recommendations aimed at improving evaluation and reproducibility.
+Compliance with these guidelines is essential for validating experimental findings and ensuring comparability across the broader research landscape.
+8.2.1 General
+Reporting
+Achieving robust reproducibility and generalizability hinges on detailed reporting, an area where current research practice often falls short, as discussed in Section 7.3.
+Their inconsistent application and reporting inhibit the validation and critical assessment of published methods, eroding confidence in the field’s reported progress.
+Furthermore, the routine omission of source code and model checkpoints exacerbates this issue, making result verification impractical.
+This practice stands in sharp contradiction to the standards of modern machine learning, which mandates standardized data splits, clear evaluation metrics, and the highly encouraged release of code and model weights.
+Structural splitting
+Furthermore, to prevent information leakage and ensure a truly generalizable model, data must be split strategically based on its underlying structure.
+For instance, splitting by patient, time point, or acquisition condition (e.g., device, institution, or environment).
+A notable failure to enforce the critical step of patient wise splitting is frequently observed (see, for example, (Baumgartner et al., 2019; Valiuddin et al., 2024b; Monteiro et al., 2020)).
+Inference details
+To ensure fair and accurate comparisons, the number of forward passes used to generate uncertainty maps and evaluation scores must be explicitly reported for every method.
+This is a prerequisite because the quality of the uncertainty estimate is often dependent on this hyperparameter (see Table 7).
+Such transparency, exemplified by the standardization work of Zbinden et al.
+(2023), allows the community to determine optimal validation strategies.
+Similarly, the specific parameters for any data augmentation must be detailed, particularly when using TTA (Section 5.3).
+For latent variable models, it is crucial to specify the dimensionality of the latent space and detail the precise mechanism for incorporating these variables into the network architecture (e.g., through tiling in Kohl et al.
+(2019b) or by using scale and translation parameters as done in Kassapis et al.
+(2021)).
+Quantifying uncertainty types
+As argued by Kahl et al.
+(2024), claiming to quantify an uncertainty type would ideally require validation across all relevant applications.
+For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but it is often found in literature that either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti and Gal, 2018; Mobiny et al., 2021; Whitbread and Jenkinson, 2022).
+As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often be found to be evaluated on a single annotator or without the use of distribution-level statistics (Wang et al., 2019a; Whitbread and Jenkinson, 2022; Kendall and Gal, 2017; Liu et al., 2022b; Schmidt et al., 2023; Savadikar et al., 2021).
+Hence, simply avoiding any serious claims related to either type alleviates this burden, and allows one to focus solely on the task at hand.
+In contrast to the work of Kahl et al.
+(2024), however, we do not consider calibration of models itself as a downstream task, but rather as a tool to gauge model reliability to be later used for other tasks.
+As a consequence, the pixel-wise uncertainties can be later used for possible downstream applications.
+Aggregation
+The aggregation method should be considered as a standalone design choice (discussed in Section 7.2) and researchers must explicitly state how per-pixel probabilities are combined.
+Ideally, multiple aggregation techniques should be evaluated to rule out potential biases, such as those related to object size.
+When working with volumetric data, it is crucial to evaluate both 3D volumetric and 2D slice-based approaches.
+Furthermore, entropy must be calculated in one of two ways: either correctly, using the true likelihood under the model or by using a factorized assumption.
+If the latter approach is taken, this significant methodological limitation must be made explicit.
+Feature extractors
+The feature extracting backbone is a key component in uncertainty modeling (Section 7.7).
+Thus, any proposed method or improvement must be ablated across diverse feature extractors to confirm its general functionality.
+For instance, latent variable methods should demonstrate performance stability when interchanging CNN or ViT encoders, and parameter modeling approaches should similarly be tested across various architectures.
+Dataset
+Data dependency is a crucial element in uncertainty modeling (Section  7.6).
+For new methodological contributions, ablation across diverse datasets is essential.
+These datasets should encompass variations in: number of classes, modality, input channels, dimensionality (2D vs.
+3D), and annotation types (single vs.
+multi-annotated).
+This scope should also extend to complex tasks beyond simple semantic segmentation, notably instance and panoptic segmentation.
+8.2.2 Task-specific guidelines
+Observer Variability
+We recommend the standard of 16 predictive samples for evaluating the GED and HM-IoU on the LIDC-IDRI dataset, a convention established in literature (Table 6).
+Another commonly used dataset is the CityScapes benchmark.
+for this, it is advisable to use the label-switching parameters proposed by Kohl et al.
+(2019b).
+Other multi-annotated datasets, such as those from the QUBIQ Challenge (Menze et al., 2021), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), although they follow less standardized preprocessing.
+The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al.
+(2022).
+Furthermore, it is crucial to explicitly report the strategy for handling correct empty segmentation maps; we recommend assigning a maximum score (Valiuddin et al., 2024b) rather than removing NaNs (Kohl et al., 2019a).
+Finally, evaluations should include an ablation study on the number of annotations to fully assess the generalization capabilities of a model.
+Active Learning
+Researchers should report the relative change in performance (corrected for random sampling as shown in Equation (36)) and evaluate sample diversity using metrics like FID.
+Furthermore, an ablation study should be conducted where varying percentages of the data are withheld from the initial training pool to assess performance under different data availability scenarios.
+Model Introspection
+It is essential to report PAvPU and CoV (see Section 6.3) based on rigorous uncertainty calculations.
+Furthermore, studies must critically analyze the relationship between the characteristics of the ground truth segmentation mask and the model’s predicted error.
+This analysis is crucial for ensuring that the model’s uncertainty estimates are reliable and not biased by specific data attributes (i.e.
+size, complexity, location, contrast).
+Model Generalization
+Report the relative change in performance as discussed in Section 6.4.
+8.3 Limitations in the recommendations
+Selection bias
+A limitation of this review stems from the prevalence of specific methods in the current literature such as VAE-based models.
+While dominance of such methods may reflect the technique’s genuine efficacy or versatility, it may also be attributed to factors such as ease of implementation, readily available codebases, or simplicity of initial adoption, thereby potentially introducing a selection bias into the overall analysis.
+Furthermore, their extensive application may not inherently confirm their superiority.
+In fact, it could simply reflect the community momentum and historical popularity the technique.
+Literature gaps
+The heterogeneity and lack of rigor in the current literature necessitated a degree of generalization in method classification, which is a key constraint on our comparative analysis.
+To maintain clarity, we primarily refer to methods by their underlying conceptual class (e.g., VAE, DDPM) rather than specific implementations, as many variants (including fine-grained hierarchical or categorical versions) offer only incremental changes.
+Exceptions are made solely for highly task-specific methods, such as SSNs and GAN-based CAR, which lack a broader architectural category.
+This constraint was amplified by the widespread absence of standardized benchmarks and comprehensive failure analyses in literature.
+Consequently, to ensure analytical coherence across all surveyed tasks, we sometimes infer the properties of a specific method from its parent class or rely on qualitative observations based on general architectural principles, prioritizing broad insight across fragmented literature over implementation-specific empirical precision.
+Soundness vs. prevalence
+Finally, a notable tension emerged between theoretical soundness and empirical prevalence.
+Methods built upon stronger theoretical justifications, such as those utilizing principled Bayesian inference, are often less prevalent in the literature than more heuristic, theoretically weaker, approaches such as MC dropout or ensembling.
+This disparity suggests a practical barrier to adoption (e.g., complexity or computational cost).
+In several instances, our analytical preference for theoretical rigor superseded the volume of available literature on more widely adopted alternatives.
+This choice risks under-representing the pragmatic, empirically dominant approaches that currently drive much of the applied progress in the field.
+8.4 Criteria for good uncertainty
+Following the review of the fundamental need and methodological approaches for uncertainty modeling, this analysis culminates in the essential assessment of utility.
+This section transcends evaluations based on algorithmic complexity, theoretical soundness, and standard metrics to articulate the defining characteristics of "good" uncertainty.
+We will therefore establish the practical criteria required to ensure the resulting estimates possess genuine relevance in real-world scenarios.
+8.4.1 Reliable
+Trustworthy systems fundamentally require high accuracy in the point prediction, as being correct most of the time is the prerequisite for the model’s uncertainty to be low (sharp).
+Secondly, and critically, the model must be well-calibrated.
+In cases of potential inaccuracy, the model must reliably report high uncertainty to maintain trustworthiness.
+A miscalibrated model, often manifesting as dangerous overconfidence, is fundamentally unreliable.
+High confidence misleads users to assume low risk, which is particularly problematic in safety-critical fields like medicine or autonomous driving, thus compromising safe decision-making.
+8.4.2 Explainable
+To be effective, uncertainty must be well-communicated and easily understood.
+For example, a raw entropy map is often insufficient because it provides only a relative comparison between pixels while lacking absolute, actionable meaning.
+Therefore, a better approach is to process this raw data into an interpretable signal for the end-user.
+This can be achieved by displaying distinct, plausible alternative boundaries or by using a bright color to highlight only the regions of highest uncertainty, which effectively guides an expert’s attention where it is needed most.
+8.4.3 Actionable
+The key insight is that uncertainty should not be seen merely as a measure of model doubt, but as a crucial actionable signal.
+In practical applications, this signal is used to trigger specific, risk-mitigating behaviors, such as altering a plan, halting a procedure, or deferring control to a human expert, to ensure safety and prevent errors.
+For effective utility, users must be clearly informed about the reducibility of the uncertainty.
+High uncertainty should signal whether there is a need to collect more training data or refine the model or if it simply represents irreducible, inherent data noise.
+In this latter case of irreducible uncertainty, the system must still recommend a specific risk-mitigating strategy based on the bounds of the prediction.
+8.4.4 Unbiased
+A fundamental tenet of robust and equitable modeling is that the system’s predictions must remain unbiased with respect to the underlying data distribution.
+To address this core challenge, we will categorize and discuss several common type of biases, providing examples of how they can manifest.
+Distributional bias
+If the validation data does not reflect the real-world deployment environment, the model’s uncertainty estimates will not be reliable in practice.
+This bias can stem from subtle shifts in acquisition conditions or population characteristics.
+Therefore, to ensure robustness, datasets must represent a balanced collection encompassing diverse acquisition and demographic variations.
+Examples include an object detection model validated exclusively on footage from sunny highways deployed in a snowy snowstorm.
+Or, a diagnostic model validated on MRI from one hospital and specific demographic deployed at a new institution with different scanners and a diverse patient population.
+Annotation bias
+Relying on a single ground-truth mask introduces a significant source of bias, as the model learns to be certain about one annotator’s idiosyncratic choices, even where boundaries are ambiguous to other experts.
+This results in a false sense of certainty, where the model’s confidence reflects its ability to replicate a single opinion rather than the inherent ambiguity of the task.
+Furthermore, the selection of annotators is a critical consideration.
+To mitigate biases from individual annotators and their tools, researchers should collect multiple annotations from a diverse group of experts to build a ground truth that more accurately captures the inherent ambiguity of the task.
+Examples of this can be delineations provided by highly trained medical experts often differing systematically from those of trainees or multiple experts, which may have their own valid interpretation of the boundary.
+Another example of this can be that a polygon tool for annotating encourages simplified masks with straight edges, while a freehand brush tool allows for more detailed curves but can introduce minor inconsistencies.
+Modeling bias
+Different uncertainty quantification methods may have their own intrinsic biases.
+For instance, Gaussian densities assume single-moded uncertainty as well as a light-tail bias.
+Consider segmenting a brain tumor with ambiguous boundaries, a single-moded Bayesian model produces only one average segmentation.
+This misrepresents the true, multimodal uncertainty to a clinician.
+For autonomous vehicles segmenting a road, the light-tail bias, often inherent in models relying on assumptions like the Gaussian distribution, causes the system to view critical safety hazards, such as massive potholes, as statistically near-impossible events because training focused primarily on smooth surfaces.
+Sampling bias
+A system trained to reduce uncertainty often focuses its labeling budget on rare or complex outlier samples.
+This wastes resources by developing expertise in statistically insignificant cases, ultimately resulting in a model that fails to reliably generalize or improve performance in the common, high-value scenarios it was designed to analyze.
+For example, in flawed active perception, a soccer robot aiming to segment the ball focuses its attention where segmentation uncertainty is highest, i.e.
+the visually complex sidelines.
+This draws focus away from the simple green field, causing the robot to become an expert at tracking the ball in the least important part of the game.
+Another example can be a model used in satellite imaging may focus its uncertainty-driven budget on unique, complex buildings instead of common houses resulting in a model expert in architectural oddities but failing to reliably segment critical residential areas.
+Validation bias
+Overemphasizing metric improvement can incentivize models to undermining the original goal.
+A textbook case of Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure”.
+Examples of this are optimization of metrics that minimize the discrepancy between the prediction and ground-truth distribution can incentivize the model to simply replicate the observed ground-truth masks rather than capture the full range of plausible, unseen segmentations.
+Furthermore, a model can achieve a perfect calibration (ECE score) while remaining dangerously useless for quantification; this occurs when the model only expresses binary confidence (100% or 0%).
+While technically calibrated, this failure mode is useless because the system fails to communicate the necessary nuanced doubt about ambiguous regions.
+Another example can be entropy-guided uncertainty can perversely incentivize models to suppress doubt and minimize uncertainty in genuinely ambiguous regions to achieve better metric scores.
+Key takeaways: Good predictive uncertainty is defined by: Reliability: The uncertainty should accurately reflect the model’s true error rate.
+Explainability: The source and meaning of the uncertainty should be clear to the end-user.
+Actionability: The uncertainty estimate must be useful for downstream decision-making.
+Unbiasedness: The estimates should not be systematically skewed by factors like subgroup characteristics.
 9 Conclusion
-Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability of their predictions.
-Given the vast and diverse body of related literature spanning various applications and modalities, often scattered and marked by inconsistent evaluation and benchmarking practices, there is a clear need for a comprehensive and systematic overview.
-This work addresses that need by presenting clear definitions and a unified notation for methodologies in uncertainty modeling.
-Furthermore, confusion and misunderstanding around uncertainty modeling and quantification have led to diverging beliefs within the field.
-By structuring the methodologies by targeted uncertainty type and followed by downstream pipeline into four distinct tasks, it becomes easier for researchers to navigate the field and build upon existing work.
-We place particular emphasis on addressing key conceptual gaps to support both pragmatic and theoretically sound research.
-We also highlight the potential and broader impact of advancing uncertainty-based active learning.
-A noticeable trend is the adaptation of generative modeling advances to segmentation through novel, task-specific conditioning strategies for modeling ambiguity in segmentation masks.
-While such efforts are already underway, it remains important to closely follow progress in general generative machine learning and continue exploring its effective application to segmentation.
-Finally, we recommend future work to focus on improving spatial aggregation methods, standardization in benchmarking, backbone architectures, as well as the adoption of novel uncertainty estimation techniques, as many commonly used methods lack rigor.
-In particular, there is growing interest in distribution- and sampling-free approaches.
-In this manner, this review paper aims to guide researchers toward building more reliable, efficient, and uncertainty-aware segmentation models within the lightning-fast evolving field of Deep Learning-based Computer Vision.
+This review synthesizes the fragmented literature on uncertainty in semantic segmentation, establishing a unified framework to bridge the gap between method developers, task specialists and applied researchers.
+Our analysis highlights critical challenges, including the nuanced distinction between uncertainty types, flawed assumptions in spatial aggregation, and a lack of standardization.
+We identify active learning, data-driven benchmarks, and transformer-based models as the most promising avenues for future research, particularly for extending novel uncertainty methods to holistic tasks like instance/panoptic segmentation.
+Our conclusion provides a clear method selection guide and places a strong emphasis on the need for future work to produce uncertainty that is reliable, explainable, actionable, and unbiased.